A Complete Bibliography of Statistics in Medicine (2020–2029)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

12 October 2024 Version 1.11

Title word cross-reference

1 [296, 431]. 2 × 2 [47, 1278, 8, 18, 784]. 2 × 2 × S [971]. 5 [309]. _{0} [59]. α [502]. B [809]. C [1170]. ϵ [1342]. F_{1} [1322]. G [1562, 1306, 1326, 1102]. k [665, 772]. κ [871]. N [1147, 1210, 296, 431]. P [1492, 1301]. R [1010]. R^{2} [377]. t [968]. Z [470].

-accelerated [1342]. -diversity [502]. -estimation [1306]. -estimators [1326]. -formula [1562]. -indices [1170]. -Inflated [1468]. -modeling [1102]. -of- [296, 431]. -of-1 [1147, 1210]. -Resolution [665]. -scores [470, 1322]. -squared [1010]. -step [309]. -type [871]. -value [809]. -values [1492, 1301].

067 [818].

1 [76, 153]. 1.0 [934]. 101 [1060]. 12 [1090]. 19 [578, 1049, 1378, 655, 887, 935, 1153, 1467, 477, 556, 764, 856, 529, 643, 480, 1528, 478, 483, 587, 914, 1016, 1373, 975].

2 [1225, 1467, 420, 1514, 154]. 2-in-1 [1328, 486]. 2.0 [866]. 2/3 [162]. 2019 [797]. 2021 [559]. 2331 [115].

3 [1517]. 38 [47, 854, 115].

40th [795]. 4269 [47].

5 [593]. 5-STAR [309, 593]. 5-step [593]. 50th [797].

A. [328, 330]. abilities [598]. absence [1644]. Absolute [1167, 584, 978]. absorbing [897]. abundance [1348]. abundant [313]. accelerated [1234, 830, 1361, 602, 1397, 1609, 357, 1316, 807, 1627, 1063, 1342]. accelerometer [1128, 1374]. accelerometer-measured [1374]. accelerometry [199]. access [268, 569]. accommodate [982, 122]. Accommodating [1439, 1026]. account [1168, 1344, 575, 176]. accountable [997]. Accounting [698, 954, 1327, 489, 835, 536, 562, 958, 1346, 776, 1377]. accounts [1404, 1296]. accrual [1404]. accumulation [708]. accumulative [1582]. accuracy [619, 853, 1318, 567, 1346, 9, 581, 1379, 674, 1376, 625, 44, 811, 804, 923, 166, 388, 1314, 53, 1142]. accurate [1225]. aCFO [1582]. acquisition [122]. across [997, 1106, 232, 851, 85, 1631, 1249]. actigraphy [917]. acting [298]. activation [447]. Active [236, 1607]. activity [451, 823, 1019, 199, 1374, 396, 307]. acute [1460]. acyclic [931]. adaptation [1405, 530, 1526, 1025]. adaptations [624]. Adaptive [368, 946, 84, 1560, 1007, 53, 1328, 506, 1591, 1041, 1429, 1503, 533, 816, 146, 672, 1385, 863, 1416, 174, 582, 769, 443, 486, 501, 1084, 821, 869, 1387, 1492, 351, 520, 653, 1185, 1104, 1233, 1228, 296, 431, 858, 761, 992, 1107, 1282, 97, 41, 1473, 1137, 1043, 632, 17]. adaptive-decision [1137]. Adaptively [746]. Adding [1596, 111]. addition [1605]. additive [1032, 1324, 1499, 1148, 978, 980, 710, 1001, 1244, 1000, 333, 492, 1196, 277, 1111, 1009, 213, 717, 354]. address [1428, 365]. Addressing [1478, 1369, 1621, 1099, 43]. adherence [656, 1532, 818, 881, 544, 1317, 616]. adjust [1179, 1641, 459, 1521]. Adjusted [7, 1122, 638, 113, 649, 80, 361, 667, 1312, 948, 1318, 70, 353, 1243, 1628, 505, 1387, 844, 1074, 1076, 1075, 608, 701, 1289, 944, 78, 271]. Adjusting [1216, 1105, 1132, 1241, 464, 301, 858]. adjustment [1309, 1511, 1565, 533, 153, 1252, 317, 496, 1419, 495, 154, 140, 376, 1613, 1380]. adjustments [906, 1401, 685]. Admin [420]. Admin-2 [420]. administered [32]. administration [887]. adults [1332, 1019, 1460, 79]. advanced [154]. Adverse [78, 101, 1207, 588, 1274, 1182, 107]. advising [801]. Africa [799, 1496]. after [411, 932, 467, 1157, 145]. against [388, 1314]. age [367, 1200, 505, 874, 1011, 395, 193]. age-adjusted [505]. age-period-cohort [367, 1200, 395]. age-related [1011]. age-standardized [874]. agent [1647]. agents [607]. aggregate [845, 903, 1400, 85, 1037, 1426]. aggregated [13, 630, 1271]. aggregation [106, 1544]. aging [1349]. agnostic [1524]. agree [286]. agreement [457, 453, 138, 875, 62, 1191, 303, 871, 32]. AI [1471]. aid [331]. aided [1188]. AIDS [476, 474, 1223, 480, 478, 483, 719, 596]. air [405]. airway [656]. al. [620]. Alan [1069]. Albert [559]. ALERT [87]. Alfred [321, 323]. algorithm [87, 1658, 1023, 315, 374, 1564, 1365, 715, 671, 278, 868, 708, 1342, 1479, 1635]. algorithms [978, 807, 1399]. allocating [525]. Allocation [1249, 661, 764, 816, 769, 821, 72, 574, 1149, 761, 1469]. allowing [661, 1192]. allows [1605]. also [1412]. alternative [47, 927, 8, 18, 419, 1012]. Altman [453]. always [1392]. always-survivors [1392]. Alzheimer [996, 767, 523, 1460, 213, 543, 741, 1181, 951]. amalgamation [309, 593]. ambulatory [1204]. among [367, 1332, 470, 771, 550, 363]. amyotrophic [510]. analyses [24, 450, 1127, 1532, 1014, 30, 731, 610, 1577, 1254, 1261, 1490, 897, 206, 1048, 198, 79, 668, 555]. Analysis [339, 1431, 105, 847, 54, 1003, 617, 1210, 332, 1355, 1193, 508, 1625, 1224, 1208, 1372, 1310, 1183, 415, 735, 1266, 442, 598, 1032, 714, 471, 970, 1512, 355, 1655, 1166, 268, 569, 779, 1028, 1563, 1571, 1065, 1190, 1531, 1203, 25, 194, 438, 1472, 1209, 876, 1139, 165, 1554, 802, 1461, 1121, 903, 512, 738, 1050, 1550, 1551, 1223, 1579, 10, 580, 740, 1421, 1332, 401, 1147, 643, 250, 978, 1346, 498, 1530, 1278, 347, 648, 1358, 316, 538, 991, 444, 248, 1222, 371, 229, 201, 812, 225, 111, 721, 1448, 644, 1280, 1618, 12, 130, 1211, 232, 551, 9]. analysis [581, 68, 1218, 374, 502, 1122, 1659, 234, 579, 104, 468, 924, 42, 265, 197, 443, 566, 710, 851, 1252, 1240, 1351, 1437, 155, 67, 352, 976, 1498, 380, 929, 1520, 1201, 954, 92, 571, 1084, 143, 338, 1172, 94, 883, 309, 593, 449, 1490, 1442, 916, 719, 696, 1608, 1615, 766, 1267, 603, 604, 297, 219, 402, 1021, 1037, 1016, 995, 387, 783, 253, 671, 728, 148, 172, 1030, 979, 1230, 1643, 386, 204, 544, 1446, 1359, 1012, 418, 776, 366, 369, 959, 1607, 510, 43, 1426, 212, 36, 1545, 441, 1654, 554, 693, 1123, 941, 1627, 841, 606, 565, 993, 2, 781, 56, 200]. analysis [49, 362, 1363, 944, 1326, 1221, 340, 1589, 226, 171, 123, 670, 549, 775, 1056, 1373, 1345, 1569, 1506, 393, 918, 1297, 1120, 1009, 503, 1340, 41, 1619, 11, 785, 23, 666, 590, 745, 842, 1598, 1146, 1287, 1329, 756, 651, 1390, 39, 121, 563, 884, 1192, 1597, 983, 685, 717, 514, 596, 1142]. analytic [139, 531, 721, 623, 705, 85, 632]. Analytical [1058, 1096]. analytics [1051]. analyze [1576, 1000]. analyzers [1273]. Analyzing [955, 1140, 345, 1460, 614, 658, 789, 1245, 729, 1082, 242, 1415]. anchored [317, 496, 495, 844]. Andersen [988]. annealing [1171]. Anniversary [795, 797]. ANOVA [1000]. answers [319]. Anthony [1129]. antibody [550]. anticancer [607]. anticoagulants [298]. antimicrobial [1544]. antipsychotic [1487]. antiretroviral [1496]. Antonio [1076]. apnea [656]. Application [1463, 361, 536, 583, 157, 268, 569, 1121, 1536, 1332, 996, 444, 818, 374, 687, 1572, 844, 419, 1367, 1234, 217, 872, 836, 974, 1205, 1265, 1324, 791, 1133, 692, 1466, 912, 1311, 1223, 28, 856, 529, 734, 210, 894, 834, 538, 991, 1246, 1293, 1239, 977, 1618, 12, 551, 1562, 854, 527, 1254, 6, 470, 754, 1292, 380, 771, 264, 1556, 1244, 810, 883, 916, 716, 1401, 1273, 297, 715, 1037, 1016, 1406, 653, 1575, 577, 448, 1317, 1426, 758, 554, 169, 1107, 609, 122, 944, 226, 224, 491, 1373, 1349, 1494, 942, 213, 1087, 1060, 216, 218, 75, 983, 1182, 596]. application [741, 1181, 772]. Applications [90, 1204, 131, 502, 337, 1544, 1626, 958, 1338, 1202, 367, 1421, 59, 222, 535, 284, 1268, 1505, 421, 1462, 949, 1154, 831, 1308, 204, 1288, 612, 1048, 435, 89, 123, 283, 1444, 1141, 756, 1102, 1018]. applied [1620, 1389, 526]. applying [239]. approach [1463, 1636, 1570, 1193, 1310, 712, 655, 442, 1241, 636, 305, 157, 95, 179, 355, 1366, 692, 1220, 1145, 740, 1158, 518, 466, 737, 920, 248, 1222, 1038, 898, 114, 26, 721, 1033, 622, 505, 12, 1211, 1013, 98, 854, 1356, 1275, 6, 385, 269, 710, 926, 20, 427, 862, 143, 392, 1, 1315, 1457, 395, 1037, 387, 1593, 683, 728, 1039, 979, 1304, 1117, 1135, 1022, 1109, 399, 1214, 1316, 941, 631, 860, 112, 888, 49, 1363, 775, 1408, 1144, 375, 901, 1619, 1186, 363, 1367, 732, 668, 181, 244, 1621]. approaches [1216, 220, 69, 290, 1576, 1213, 1649, 902, 358, 348, 359, 1391, 1639, 214, 1446, 43, 675, 1160, 1085, 1599, 1006, 1444, 651, 1380]. appropriateness [1109]. Approximate [1533, 1265, 805, 1346, 810, 636, 22, 372, 1361, 1377, 1115]. approximation [189, 1421, 1540]. approximations [1245]. April [797]. arbitrarily [984, 549]. arbitrary [1257]. area [312, 1052, 110, 258, 389]. areas [178, 1635]. arise [1557]. arising [1058]. arm [1041, 1596, 1434, 487, 1605, 1263, 1252, 467, 1126, 727, 1446, 848, 1251, 198, 393, 1340, 1180, 1469, 864]. arm-based [1446, 198]. Armitage [797]. arms [1596, 661, 1605]. array [509, 949]. arrayed [1100]. arrest [1131]. arthritis [527, 894]. article [475]. artificial [1609]. ascertained [49]. Ascertaining [235]. ascertainment [404, 1167, 169]. ask [17]. aspects [269, 1]. assay [1531, 1106]. assays [1198, 509, 1106, 1029]. assess [1158, 783]. Assessable [1390]. Assessing [1532, 690, 299, 894, 646, 879, 1581, 317, 496, 495, 865, 405, 56, 32, 44, 305, 251, 1161, 898, 1657, 1300, 1109, 1170, 1281, 616, 363, 517]. Assessment [567, 52, 1463, 125, 453, 1614, 1652, 341, 875, 1094, 1273, 507, 1226, 1080, 1077]. assessments [1159, 1018, 859]. assignment [384, 347, 1630, 375]. assisted [1175, 287, 897, 1616, 999]. associated [1276, 997, 1551, 3, 1290, 1212, 1544]. Association [1442, 40, 901, 139, 1546, 1466, 738, 96, 1245, 1093, 197, 470, 780, 1360, 254, 1037, 337, 1108, 776, 302, 550, 1619, 436, 245]. associations [1207, 1362]. assuming [1106]. assumption [833, 99]. assumptions [120, 930]. Assurance [1604, 1340]. asthma [217]. asthma-related [217]. asymptomatic [935]. asymptotic [1008, 1258]. atherosclerosis [526]. Attack [1421]. attitude [1049]. attributable [1456, 860]. attribute [211]. attrition [1189, 1525, 1420]. AUC [283, 749]. audiology [609]. augmentation [242]. Augmented [1001, 905, 1252, 1638, 390]. Aurélien [183]. AUROC [377]. Austin [323, 324, 325, 326, 183, 329, 322]. Austria [529]. Author [184, 580, 434]. autism [1584]. autoantibodies [716]. autocorrelated [1210]. autocorrelation [1168, 773, 1412]. automatic [237]. autoregressive [217, 931, 776]. auxiliary [779, 1148, 1330, 1422, 39]. available [1554]. average [1452, 398, 894, 462, 1486, 1608, 1184, 1432, 1288, 718, 1545, 1150, 1408, 1125, 216]. averaged [735]. averaging [1449, 1645, 1522, 1593, 817, 1502, 1367]. Avoiding [703]. Azocar [1076].

B [328, 330, 707]. Backward [597]. bacteriophages [1544]. balance [1650, 863, 170, 1240, 665, 999]. balance-improving [999]. Balancing [220, 1579, 1171, 1533, 591]. Ball [1386]. band [1114]. bandit [431]. BAREB [149]. basal [1451]. based [1463, 192, 1570, 1224, 1216, 66, 887, 1205, 471, 953, 640, 147, 1051, 840, 830, 846, 1461, 1044, 1014, 724, 843, 1551, 1068, 1633, 1421, 838, 398, 1256, 250, 824, 920, 304, 444, 1652, 248, 1222, 201, 1052, 1106, 26, 812, 1033, 955, 12, 863, 1538, 1155, 133, 199, 203, 911, 1254, 144, 108, 647, 1572, 197, 739, 926, 1422, 623, 421, 810, 1302, 282, 831, 1509, 674, 357, 1016, 1308, 783, 1212, 1593, 683, 600, 1640, 576, 1117, 1446, 574, 1149, 1135, 1025, 366, 541, 490, 718, 88, 1248, 302, 1411, 631, 860, 609, 701, 633, 49, 1054, 456, 1085, 532]. based [15, 198, 195, 649, 1347, 1502, 1301, 918, 1297, 1539, 1231, 335, 1541, 63, 363, 1573, 1583, 57, 1342, 422, 1141, 668, 181, 1018, 999, 1142, 599, 522, 354, 1286]. baseline [170, 727, 1289, 575, 140]. Basic [153, 1060]. basket [1334, 1358, 232, 1416, 1005, 763, 1250]. batched [964]. Bayes [1612, 673, 888]. Bayesian [26, 1069, 269, 1067, 1234, 1193, 1611, 508, 585, 1625, 1090, 1372, 852, 712, 102, 415, 973, 735, 1595, 1637, 1178, 1241, 1405, 1265, 1591, 1086, 1484, 1429, 1096, 95, 640, 935, 1512, 1503, 179, 125, 310, 1098, 186, 126, 885, 1366, 1057, 692, 948, 937, 832, 1467, 1334, 1220, 802, 1565, 1449, 1434, 903, 1014, 384, 1260, 786, 533, 641, 637, 518, 904, 704, 261, 292, 466, 1560, 1207, 493, 1587, 531, 812, 977, 622, 505, 1020, 293, 232, 774, 1385, 551, 1538, 610, 98, 1364, 504, 406, 1140, 1333, 315, 1275, 521, 627, 1279, 1081, 582, 385, 647, 149]. Bayesian [542, 443, 1437, 155, 1210, 748, 1005, 752, 1438, 1357, 311, 1360, 626, 1459, 338, 961, 314, 351, 768, 1, 359, 1624, 604, 1305, 1021, 907, 85, 1037, 1103, 896, 683, 337, 1575, 728, 815, 55, 601, 889, 1601, 1230, 77, 1022, 688, 1251, 817, 1607, 431, 212, 693, 492, 1248, 1382, 1627, 1555, 1590, 763, 781, 1377, 1194, 56, 1160, 1111, 226, 171, 198, 611, 381, 827, 1373, 1569, 1539, 1414, 1250, 1009, 503, 923, 1340, 41, 543, 785, 1473, 814, 63, 1445, 1598, 1367, 390, 413, 1018, 884, 1066, 999, 1070, 1192, 1399, 1415, 1417, 983, 717, 790, 867, 1629, 741, 296]. Bayesian-bandit [431]. BBI [1573, 1583]. be [1557, 678, 258, 933]. beat [1573, 1583]. beat-to-beat [1573, 1583]. BEATS [1405]. bed [1562]. behavior [1018]. Belgium [253]. below [1531]. benefit [981, 1187, 360, 1161, 879, 348, 877, 1150, 1147]. benefits [368, 1082]. benfluorex [1551]. Bernoulli [1351]. Bernoulli-gated [1351]. Bernstein [372]. Best [694, 852, 1506, 1542]. beta [1468, 1414]. beta-binomial [1414]. between [1586, 1372, 457, 661, 138, 1187, 548, 1052, 1448, 1093, 780, 254, 773, 1412, 1082, 866, 148, 1230, 1079, 226, 178]. between-person [773, 1412]. between-study [1372, 1448, 1230]. Beyond [1000, 764]. BHAFT [1627]. Bias [289, 682, 43, 1054, 853, 1071, 1429, 450, 948, 1209, 1565, 703, 398, 1641, 1243, 404, 130, 568, 1515, 954, 94, 365, 167, 301, 971, 1432, 858, 489, 56, 456, 383, 1031, 936, 1517, 685, 1142, 1298]. bias-adjusted [948, 1243]. biased [1503, 1223, 312, 1353]. biases [139]. biclustering [149]. big [244]. BILITE [292]. Binary [374, 1463, 1208, 852, 102, 667, 1008, 1595, 35, 885, 1247, 25, 843, 191, 1243, 299, 432, 731, 689, 1333, 109, 1577, 1623, 1489, 748, 902, 467, 62, 1371, 434, 358, 1387, 254, 729, 913, 172, 586, 377, 1030, 576, 433, 1117, 688, 510, 43, 629, 631, 992, 1322, 2, 435, 1299, 1603, 200, 1160, 32, 283, 198, 383, 901, 657, 1541, 422, 1296, 436, 242, 1192, 1255]. binomial [562, 916, 891, 715, 1037, 36, 231, 1375, 1414, 140, 1621]. biobanks [66]. bioequivalence [1486, 1593]. biological [1261, 867, 951]. Biomarker [1248, 570, 563, 428, 958, 1612, 824, 498, 1478, 504, 174, 1151, 284, 902, 427, 230, 605, 1016, 727, 660, 1114, 388, 1314, 1558, 124]. Biomarker-based [1248, 824]. biomarker-driven [174]. Biomarker-guided [563, 1558]. biomarkers [1440, 1133, 1511, 1466, 939, 1480, 1145, 290, 444, 73, 545, 1175, 1384, 446, 1212, 1170, 888, 1115, 400, 749, 75, 639, 244]. biomedical [421, 883, 831, 1000, 1060]. biometrics [1461]. biosimilar [1057]. biosimilarity [251]. biostatistician [801]. BIPSE [824]. bisulfite [1621]. bivariate [217, 836, 641, 252, 622, 203, 954, 233, 1037, 896, 277, 1393, 122, 245]. BivRegBLS [286]. Bland [453]. bleeding [298]. Blinded [1083, 260, 140]. block [1658, 632]. block-coordinate [1658]. BMI [1188]. BOB [1057]. Bock [175]. body [1451]. BOIN12 [867]. Bonferroni [1433]. Book [309, 296]. boost [1438]. Boosting [1196, 851, 274]. Bootstrap [1008, 1564, 674, 123, 150]. bootstrap-based [674]. borne [1305]. Borrowing [642, 779, 1334, 531, 232, 672, 1385, 1494, 1250]. both [1183, 874, 254, 1270, 1132, 1569, 257]. bound [721, 809]. bounded [438, 1347]. bounds [899]. Box [1238]. Bradford [322, 323, 324, 325, 326, 329]. brain [1234, 1311, 1350, 991, 1618, 1349, 983]. branching [901]. Brazil [1153]. breast [1265, 1567, 715, 1303, 1107]. breath [75]. Bridging [1586, 866, 1450]. broad [138]. BUGS [508]. Building [1407, 799]. built [868]. bulk [573]. burden [13, 387, 877]. burdensome [877].

C [183, 377]. cadmium [551]. Calculating [1030, 945]. calculation [852, 985, 731, 1197, 905, 240, 227, 988, 784, 1566, 880, 1272]. calculations [700, 132, 410, 377, 828]. calibrated [1352]. Calibrating [1391, 1639]. Calibration [833, 1561, 187, 1368, 1582, 1652, 1013, 365, 1600, 1132, 344, 702, 29, 150]. calibration-free [1582]. can [832, 528, 1651]. Canada [423]. Cancer [61, 792, 415, 411, 958, 972, 553, 974, 1265, 1065, 308, 250, 1530, 444, 812, 232, 774, 315, 374, 521, 133, 269, 13, 1, 1567, 715, 815, 1407, 1156, 238, 1303, 1382, 1107, 1221, 224, 611, 79, 57, 75]. cancers [190, 57]. candidates [210]. capacity [799, 100]. CAPITAL [998]. capture [1256]. capture-recapture [1256]. Capturing [437, 1022]. carcinoma [888]. cardiac [1551, 1131]. cardiovascular [1285]. cardioverter [86]. care [1378, 997, 1332, 1657, 1624, 337, 1496, 919, 772]. caregiver [1401]. Carlo [126, 315, 1423]. carriage [810]. carry [219]. Case [1071, 1586, 1312, 1153, 1467, 703, 1256, 856, 117, 854, 6, 1515, 197, 1175, 1384, 709, 346, 995, 682, 921, 1135, 865, 169, 1131, 340, 813, 15, 570, 1642, 375, 514]. case-cohort [514]. case-control [117, 197, 1175, 1384, 709, 346, 682, 921, 1135, 169, 813, 15, 570, 1642, 375]. case-study [340]. cases [709]. cataract [204]. Categorical [1597, 33, 527, 1389, 1123, 614]. Categorisation [1521]. categorized [696]. Category [573, 300, 1140, 497]. Causal [1276, 1035, 1563, 1531, 1203, 3, 27, 1218, 82, 566, 1226, 43, 670, 842, 1271, 1463, 642, 669, 1098, 1571, 802, 267, 894, 646, 1038, 266, 870, 1211, 930, 1564, 205, 765, 1648, 1556, 106, 1608, 1315, 359, 167, 995, 50, 1152, 758, 770, 1654, 405, 158, 1409, 606, 860, 1077, 1408, 1024, 280, 1335, 590, 1294, 91, 422, 1296, 837, 1099, 1015]. Cause [430, 163, 1547, 723, 375]. cause-of-disease [375]. Cause-specific [430, 163, 1547, 723]. causes [222, 1547, 1170]. cautionary [1458, 899]. celebration [798, 322]. cell [956, 573]. Censored [500, 76, 1090, 1626, 228, 1591, 981, 1235, 1546, 138, 1527, 1550, 1223, 1421, 1119, 372, 1361, 188, 1106, 163, 955, 873, 1623, 414, 1398, 1520, 862, 262, 968, 662, 638, 345, 1581, 1406, 135, 928, 663, 1109, 1306, 550, 1375, 984, 171, 549, 827, 1468, 166, 1231, 745, 1444, 960, 1342, 1540, 218, 274, 255, 93, 516, 514, 596, 373]. Censoring [743, 189, 847, 1045, 403, 1112, 20, 1108, 1206, 1317, 575, 44, 942, 1144, 1296]. Censoring-robust [743]. center [1588, 1572, 1295]. central [1588]. cervical [536, 238]. cessation [912]. chain [355, 1204, 126]. chained [88]. chains [491]. challenge [603]. challenges [66, 577, 1507]. Change [451, 1349, 1133, 186, 1470, 108, 535, 89, 1297]. Change-point [451, 108, 89]. changepoint [16]. changepoints [1414]. changes [1510]. Channing [1421]. characteristic [428, 228, 281, 1145, 312, 646, 1052, 679, 110, 168, 743, 258, 389, 1142, 1635]. characteristics [929]. CHARMS [219]. charting [70]. chemical [259]. Chiang [1069]. child [420, 169]. childhood [1466]. China [791]. choice [1164, 1339]. choices [957]. Cholesky [385]. cholesterol [366]. Choosing [705]. chronic [1273, 532, 1558]. chronobiology [31]. cigarette [1583]. circadian [1275, 1272]. circular [192]. civil [896]. claims [1551, 1302, 1460]. class [619, 1166, 620, 843, 1145, 353, 518, 26, 1275, 468, 1229, 1313, 618, 621, 499, 1322, 628, 1517, 373]. classes [27, 1379, 85, 1413, 1376]. classical [91]. Classification [1023, 199, 1365, 1042, 278, 972, 1344, 774, 826, 534, 1623, 967, 1459, 883, 395, 1103, 283, 1100, 573]. Classified [1477, 1362]. classifier [488, 1461]. classifiers [763]. Clayton [1317]. Clinical [434, 358, 433, 336, 712, 974, 714, 1591, 1174, 1503, 152, 1309, 35, 819, 223, 472, 1383, 25, 194, 1047, 1565, 1465, 843, 1582, 84, 404, 299, 1199, 905, 839, 252, 531, 293, 98, 1263, 1105, 521, 527, 1659, 1577, 1588, 463, 582, 170, 205, 34, 501, 1437, 976, 1498, 1649, 1500, 427, 1370, 469, 1474, 705, 1634, 309, 593, 449, 459, 260, 1307, 1404, 17, 1313, 577, 913, 601, 586, 828, 77, 1604, 547, 1025, 1078, 1496, 848, 1251, 1607, 431, 629, 492, 940, 761, 1248, 628, 112, 1054, 1513, 1072, 871, 532, 123, 195, 46, 611, 711, 699, 962, 1345]. clinical [1647, 1569, 1494, 1617, 1227, 751, 257, 1367, 376, 1613, 1517, 1380, 1417, 1469, 685, 717, 596, 1089, 1453]. clinically [705, 1638]. closed [1122, 249]. closed-form [1122]. cLRT [471]. cLRT-Mod [471]. Cluster [1410, 1224, 1216, 1585, 1283, 1032, 678, 661, 105, 191, 900, 816, 1560, 731, 132, 426, 964, 1155, 68, 109, 1003, 468, 924, 42, 265, 240, 1455, 227, 753, 1371, 1242, 1419, 600, 1485, 1300, 574, 1149, 895, 608, 784, 835, 1495, 1408, 393, 1301, 1006, 1340, 657, 279, 275, 864, 880, 522]. cluster-based [574, 1149]. cluster-memberships [1006]. cluster-period [426]. cluster-randomized [1419, 1495, 1408]. Clustered [356, 1216, 1585, 513, 1533, 1040, 963, 35, 906, 1059, 1038, 1140, 719, 723, 137, 1432, 40, 1299, 1331, 196, 745, 540]. Clustering [656, 1344, 1324, 937, 1121, 12, 504, 534, 1295, 752, 1394, 1606, 1357, 338, 1185, 778, 925, 218]. clusters [1585, 953, 661, 924, 754, 1419, 393]. clusterwise [673]. co [1057, 1246, 886, 768, 1305, 804, 1399, 686]. co-data [686]. co-developing [768]. co-expression [1246, 1399]. co-infection [1305]. co-primary [1057, 886, 804]. coarsening [868]. Cochran [264]. code [770]. coding [172]. coefficient [977, 914, 706, 356, 1549, 1297]. coefficients [1308, 871, 965]. cognition [489]. cognitive [598, 186, 996, 523, 1244, 491]. cohort [1183, 1265, 177, 692, 707, 1223, 367, 1200, 111, 955, 132, 42, 221, 395, 345, 1024, 514]. cohorts [1614, 1303, 649, 1420]. coin [1503]. collaboration [800]. Collaborative [1529]. collected [953, 179, 432, 1583]. combination [1205, 1480, 1173, 902, 949, 1092, 768, 913, 1352]. combination-schedule [1092]. combinations [230, 815]. combine [1475, 868]. Combined [507, 1267]. Combining [737, 818, 13, 1447, 1106, 470, 1175, 1495]. combo [107]. commensurate [41]. Comment [481, 480, 560, 183, 831, 482, 331]. Commentary [798, 479, 620, 323, 326, 1075, 1067]. Comments [47, 477, 580, 908, 854, 579, 269, 17, 495, 844, 329, 206, 1079, 496, 18]. commercial [1018]. Common [863, 1106, 121]. commonality [851]. commonly [62]. communicating [1307]. Communication [1044]. Communication-efficient [1044]. communities [538]. Community [1133]. comparability [1249]. Comparative [472, 1055, 705, 1369, 1227]. comparators [1607]. compare [384, 718, 1089]. Comparing [102, 1106, 1657, 298, 418, 303, 811, 1602, 742, 1047, 1219, 131, 180, 1083, 424, 421, 831, 988, 1031, 1566, 1635]. Comparison [1591, 1475, 172, 1352, 1043, 1650, 1091, 1283, 1519, 1526, 1334, 1179, 1465, 286, 367, 1243, 755, 290, 234, 1213, 879, 1394, 613, 358, 927, 1391, 1270, 729, 913, 844, 525, 758, 1447, 88, 858, 51, 675, 936, 124, 422, 651]. Comparisons [1189, 390, 361, 1527, 1565, 1187, 360, 1147, 1628, 410, 317, 496, 495, 844, 1074, 1076, 1075, 38, 1450, 215, 231, 412, 1100]. compatibility [541]. compatible [394, 408, 1039]. compensation [1551]. Competing [980, 1090, 19, 607, 974, 1368, 268, 569, 906, 759, 454, 438, 1536, 1059, 229, 222, 1167, 1201, 1644, 719, 183, 1053, 715, 723, 1143, 1638, 1170, 45, 40, 1036, 1227, 1599, 91, 951]. competing-risks [723]. competition [1411]. competition-based [1411]. complete [960]. completeness [565]. completion [521]. Complex [793, 391, 1198, 1174, 1133, 1189, 27, 967, 1313, 287, 154, 702, 1656, 15, 97, 503]. Complexity [383]. compliance [1241, 1409]. complications [1462]. complier [894, 1608]. component [847, 903, 1633, 444, 644, 1644, 338, 366, 369, 959, 78, 1009, 539]. component-based [1633]. component-wise [847]. components [1557, 970, 1546, 622, 729, 945]. Composite [684, 847, 54, 949, 886, 348, 729, 1025, 675, 1513, 1589, 1473, 945]. compositional [1260, 538, 780, 340, 379, 1415]. comprehensive [1391]. computation [1265, 1377, 1573]. computational [427, 770, 1517]. computationally [236]. compute [1145]. computer [999]. Computing [546, 90, 694]. concept [301, 1417]. concerning [875]. conclusions [832]. concordance [1217, 648, 1175]. concordance-assisted [1175]. Conditional [1577, 803, 769, 443, 1175, 1085, 813, 1599, 777, 442, 146, 646, 873, 687, 209, 1330, 902, 495, 776, 1079, 1054]. Conditions [1078]. conduct [148]. Confidence [142, 1198, 1534, 394, 548, 16, 1474, 674, 1341, 1464, 178, 522, 1276, 457, 1440, 3, 1122, 822, 467, 1492, 605, 1379, 1653, 231, 633, 757, 1114, 1113, 1454, 1301, 651]. confirmatory [803, 301]. confirming [628]. Conflating [495]. confounded [1648]. Confounder [364, 289, 1157]. confounders [1203, 1179, 167, 1125]. Confounding [1252, 1650, 229, 1544, 552, 81, 38, 1456, 993, 1124, 1120]. congruence [518]. Connecting [669]. connectivity [127, 1350, 991, 1618]. connectome [983]. consecutive [83, 1279]. consensus [752, 328]. consider [1068, 52]. consideration [948]. considerations [1247, 838, 347, 42, 1080, 1233, 211, 1281, 657, 1541]. considering [348]. considers [773, 1412]. Consistency [448, 299, 783]. Consistent [1125, 1400, 1651]. constant [974, 727]. constituent [997]. Constrained [763, 864, 21, 998, 1627]. constraint [334, 311]. constraints [90, 635, 1131]. construct [952]. Constructing [215, 93]. Construction [852, 1440, 1505, 605]. consultant [328]. contact [1290]. contagion [578]. contaminated [1316]. Contamination [528]. contemplated [216]. content [426]. context [1585, 420, 315]. contingency [47, 8, 18, 340]. contingent [859]. continual [305, 72, 1092, 982]. Continued [619]. Continuous [656, 1463, 336, 1008, 973, 1440, 397, 1204, 126, 1190, 677, 194, 138, 1127, 524, 347, 129, 527, 144, 82, 1151, 1292, 260, 696, 889, 172, 1228, 1117, 547, 1521, 1289, 455, 1347, 388, 1314, 390, 1287, 436, 1255]. continuous-time [1204]. Contrast [1061, 780, 65, 1509, 1446]. contrast-based [1446]. Contribution [478]. control [1193, 667, 472, 1068, 533, 755, 531, 1069, 293, 1385, 117, 545, 197, 1252, 1175, 1292, 1384, 709, 1486, 346, 103, 682, 727, 214, 1232, 77, 921, 1135, 38, 169, 1482, 813, 15, 570, 1642, 1539, 1617, 375, 1137, 1067, 1066, 1070]. Controlled [1208, 1004, 1162, 2, 194, 788, 703, 856, 276, 1333, 1515, 1252, 1644, 729, 660, 860, 1632, 722, 864]. controlled-direct-effect-based [860]. Controlling [1423, 1605, 170]. controls [334, 105]. controversy [619]. convalescent [643]. conversion [767]. convolutional [666]. Conway [406]. coordinate [1658, 199]. coordinated [693]. Copula [281, 316, 201, 1201, 1037, 1317, 631, 49, 181]. copula-based [201, 631, 49, 181]. corner [605]. coronary [366]. coronavirus [703, 1396]. correct [344, 868]. Corrected [493, 1316, 388]. Correcting [853, 1353, 659]. Correction [910, 1276, 1208, 1328, 1372, 185, 271, 26, 581, 855, 1498, 806, 532, 115, 134, 1613, 175, 161, 1429, 289, 755, 674, 682, 1054, 685]. corrections [1452, 1301]. correctly [1500]. correlate [1277]. Correlated [772, 1163, 360, 1497, 259, 771, 1268, 1462, 233, 168, 365, 344, 1058, 1654, 693, 124, 436, 242, 1635]. Correlation [638, 1229, 42, 947, 362, 1331]. correlations [1557, 1354, 1187, 1246, 385, 1281]. Corrigendum [1138]. cosinor [1272]. Cost [1310, 547, 981, 1332, 111, 721, 92, 1575]. Cost-effectiveness [1310, 981, 111, 721]. Cost-efficient [547]. cost-penalized [1575]. costs [958, 525]. could [1528]. count [655, 95, 640, 938, 912, 401, 1083, 265, 1398, 233, 243, 1348, 1545, 295, 880]. countries [420, 896]. counts [640, 105, 1256, 1128, 1254, 1522]. Coupling [1137]. course [278]. CoV [1225, 1467, 1514]. CoV2 [1655]. covariance [915, 538, 1448, 634, 617, 1279, 1267, 907, 208, 1257, 198]. Covariate [1309, 1628, 1151, 1650, 782, 805, 1235, 1312, 450, 677, 1318, 16, 1222, 114, 679, 445, 873, 1385, 863, 170, 108, 542, 501, 156, 1084, 168, 1387, 346, 1457, 365, 1021, 676, 1419, 1030, 631, 370, 1347, 376, 1613]. covariate-adaptive [1385, 863, 501, 1084]. Covariate-adjusted [1628, 1312, 1318]. covariate-dependent [542]. covariate-driven [677]. Covariate-specific [1151, 679, 168]. covariate-treatment [16]. covariates [1171, 19, 1595, 1204, 938, 1028, 1139, 1421, 37, 130, 1645, 1268, 665, 1520, 1397, 1191, 183, 1401, 148, 419, 727, 100, 386, 1316, 247, 1521, 1027, 575, 1036, 1373, 1506, 1144, 288, 1231, 1314, 989, 128]. coverage [455, 1085, 150]. COVID [476, 578, 1378, 655, 887, 935, 1153, 1467, 477, 556, 764, 474, 856, 529, 643, 480, 1528, 483, 587, 914, 1016, 1373, 975, 1049, 478]. COVID-19 [476, 578, 1378, 655, 887, 935, 1153, 1467, 477, 556, 764, 856, 529, 643, 480, 1528, 483, 587, 914, 1016, 1373, 975, 1049, 478]. Cox [782, 1585, 1266, 805, 1470, 371, 59, 108, 1547, 1622, 1295, 113, 1238, 1021, 377, 1407, 386, 1653, 630, 1341, 1027, 1424, 1580, 942, 1144, 213, 382]. Coxian [419, 57]. create [1352]. creation [1183]. criterion [1155, 572, 1018]. criterion-based [1018]. Critical [875, 269, 1, 1163, 1624]. critique [409]. CRM [1090, 1090, 448, 1228]. Crohn [131]. cross [415, 1578, 838, 731, 1458, 221, 1053, 1341, 1362, 383, 39]. cross-classified [1362]. cross-sectional [1578, 838, 731, 221, 383, 39]. cross-validation [1053, 1341]. crossed [1289]. crossing [1386]. crossover [594, 690, 332]. crude [599]. crude-rate-based [599]. cubic [782]. Cumulative [584, 1551, 1636, 177, 1017, 1086, 70, 163, 99, 45, 409, 413]. cure [1319, 268, 569, 1004, 1119, 734, 904, 250, 726, 468, 222, 156, 990, 715, 1238, 1110, 1109, 1375, 893, 942, 61, 792, 539, 1087, 1215]. cured [745]. current [1336, 468, 710, 1165, 122, 1502]. curvature [1200]. curve [428, 228, 312, 1293, 679, 168, 743, 448, 204, 409, 258, 389]. curves [187, 428, 407, 83, 1574, 1179, 281, 1052, 110, 303, 124, 1386, 178, 884, 1602, 1635]. Cutoff [1440, 428, 605]. cutoff-points [605]. cutoffs [958]. cystic [64]. cytokine [1016]. cytokine-based [1016].

daily [994, 307]. Data [314, 1610, 960, 1636, 192, 391, 845, 1234, 313, 872, 1097, 585, 423, 1428, 1626, 1216, 361, 562, 1585, 513, 655, 1163, 1071, 1637, 442, 291, 228, 553, 642, 1265, 90, 1519, 21, 1418, 87, 953, 83, 95, 640, 791, 1133, 1035, 970, 1095, 151, 179, 125, 310, 1040, 963, 1309, 1451, 35, 268, 569, 938, 779, 1051, 981, 1235, 1366, 1571, 1065, 759, 1203, 1153, 25, 677, 194, 849, 1554, 1121, 903, 512, 724, 1336, 1551, 524, 1223, 28, 740, 1004, 1260, 1421, 401, 533, 1119, 643, 120, 734, 1243, 1290, 372, 1361, 300, 978, 432, 312, 1197, 316]. data [538, 1059, 996, 248, 615, 1529, 114, 1106, 163, 789, 1083, 812, 1402, 225, 110, 622, 644, 1280, 955, 943, 12, 1020, 774, 1064, 551, 1538, 1013, 610, 27, 767, 1140, 59, 1218, 1105, 854, 1275, 174, 634, 133, 199, 1564, 787, 234, 957, 468, 1254, 222, 6, 385, 822, 1094, 1489, 149, 82, 209, 537, 564, 494, 710, 967, 765, 829, 947, 1330, 1112, 424, 497, 922, 1292, 1649, 437, 623, 380, 820, 1398, 1488, 421, 1520, 752, 862, 1201, 1606, 233, 1001, 1244, 613, 1154, 956, 1389, 810, 709, 1165, 338, 968, 883, 1525, 662, 168, 449, 1302, 13, 638, 916, 719, 831, 1000, 1435, 1229, 1395]. data [1401, 473, 52, 1457, 766, 1624, 603, 297, 345, 715, 907, 85, 1037, 1581, 1406, 387, 994, 135, 1110, 896, 683, 1269, 1226, 1413, 723, 1143, 287, 844, 889, 1185, 1284, 164, 148, 172, 1030, 928, 1327, 77, 137, 1206, 1317, 1132, 552, 1157, 1369, 663, 418, 817, 1170, 597, 608, 1058, 369, 612, 541, 1348, 1450, 43, 1426, 212, 702, 629, 1274, 865, 1518, 238, 680, 136, 190, 45, 500, 158, 761, 691, 941, 1382, 1306, 701, 633, 1393, 435, 89, 1299, 1259, 1603, 200, 122, 550, 899, 1331, 984, 196, 1362, 278, 340, 1424, 625, 123, 224, 78, 381]. data [383, 684, 549, 491, 827, 925, 1349, 1495, 1502, 1569, 1420, 1642, 918, 1297, 1617, 288, 1414, 1599, 1006, 1024, 409, 166, 416, 460, 543, 1231, 730, 1186, 1335, 785, 1215, 814, 666, 745, 1063, 1573, 1583, 1294, 124, 1444, 708, 1342, 732, 1540, 176, 218, 274, 242, 255, 413, 379, 884, 1415, 1597, 93, 244, 285, 1621, 128, 573, 736, 717, 596, 1262, 1255, 1089, 741, 373, 1181, 1635, 354, 540, 772, 686, 589]. data-driven [925]. data-fusion [1013]. data-sharing [1529]. dataset [1389]. datasets [672, 750, 925]. days [340]. Dealing [1456, 1124, 298]. death [846, 144, 1285, 1638, 565, 489, 1501]. death-truncated [565]. deaths [1153]. debate [1079]. decades [800]. decay [42]. decipher [851]. Decision [314, 1319, 1241, 594, 788, 1434, 724, 111, 721, 496, 204, 877, 1264, 952, 1137]. decision-analytic [721]. decision-making [1434]. Decision-Theoretic [314]. decisions [425]. decline [1244]. decomposition [266, 385, 1654, 860, 400]. decorrelated [1044]. Deep [922, 986, 159, 1294, 846, 821, 302, 1616, 1060]. deepAFT [1609]. defibrillators [86]. defined [445]. Defining [1283, 1199, 1010, 1339, 1307, 1456]. Degrees [529]. delayed [1582, 794, 1604, 1535, 61, 792]. delays [964]. deletion [1024]. delirium [1536]. DeLong [1635]. demand [884]. dementia [1460]. demonstrated [971]. demonstration [246]. dengue [147, 754, 1292]. denominators [874]. densities [1480]. density [192, 873, 713, 1185, 1642, 1494, 335]. density-based [192, 335]. dental [1140]. department [419, 1337]. dependence [429]. dependencies [1011]. dependency [1257]. dependent [1650, 228, 1318, 856, 1350, 341, 130, 1478, 1093, 1515, 542, 470, 1357, 1172, 168, 743, 1206, 574, 1149, 1317, 688, 608, 959, 1169, 1456, 45, 631, 435, 160, 1323, 1580, 1231, 1125, 1007, 989, 1355]. depends [9, 581]. depth [421, 831]. depth-based [421, 831]. Depthgram [872]. derived [157]. descent [1658]. Describing [1313]. Design [1032, 96, 347, 1173, 1577, 42, 576, 211, 393, 257, 1568, 313, 1193, 1328, 1405, 607, 592, 1133, 1503, 885, 890, 840, 1057, 678, 661, 738, 1582, 1605, 292, 824, 1346, 498, 1630, 293, 1385, 1263, 1416, 504, 964, 521, 787, 463, 582, 170, 144, 1162, 486, 794, 1175, 1592, 155, 879, 1005, 1500, 427, 1387, 1126, 221, 351, 768, 346, 446, 913, 601, 572, 1535, 1496, 848, 1251, 243, 490, 431, 1447, 1248, 992, 1322, 793, 162, 1594, 5, 115, 813, 570, 611, 1647, 1494, 61, 1539, 503, 1340, 792, 751, 1180, 657, 1473, 1445, 632, 1417, 1195, 343, 867, 1272]. design-based [840]. Designing [1237, 1404, 1025, 1427, 348, 1604, 835, 1517]. designs [506, 1091, 173, 1533, 1591, 515, 1138, 1338, 318, 1511, 1526, 1127, 105, 487, 1560, 146, 252, 426, 174, 617, 769, 1254, 1410, 1425, 949, 1165, 1172, 869, 1492, 1509, 1242, 520, 1104, 1233, 1300, 574, 1149, 1656, 858, 435, 1289, 332, 41, 1007, 53, 4, 175, 999, 1469]. destructive [1375]. detect [900, 850, 1455, 1590, 835, 107]. Detecting [1510, 1587, 436, 671, 1627, 112, 279, 1621]. Detection [1350, 76, 1626, 87, 147, 1028, 70, 290, 1207, 129, 374, 950, 1490, 1185, 1544, 1482, 888, 1221, 1349, 1444, 708, 1272, 373]. Determinants [1514]. Determination [834, 385, 1574, 1093, 392, 162, 1160, 378, 923, 1568]. Determining [1632, 725]. deterministic [1203]. Developing [223, 230, 1270, 1575, 540, 957, 768, 377]. development [1096, 1052, 801, 1592, 507]. deviations [760]. device [725]. devices [1033, 1583]. diabetes [439, 1462, 716, 1471, 89]. diabetic [1320]. diagnose [1501]. diagnosis [977, 967, 771, 1471, 1575, 204, 1362]. diagnostic [619, 853, 536, 958, 887, 620, 764, 1256, 1346, 9, 581, 568, 1384, 1376, 618, 621, 625, 1373, 396, 804, 923, 388, 375, 1314, 53, 1142]. Diagnostics [734, 414, 1479]. dialysis [101, 1081, 564]. dichotomous [104]. dietary [12]. difference [24, 394, 548, 270, 352, 705, 1143, 1304, 609, 781, 757, 1194, 1150, 178]. difference-in-differences [1304]. differences [667, 334, 1576, 1106, 1483, 205, 773, 1412, 1026, 1304]. different [1097, 584, 832, 1179, 1461, 1047, 996, 1553, 980, 297, 1109]. Differential [1618, 956, 1508, 847, 1350, 1246, 1587, 169, 722, 785, 1542, 742]. differentially [1621]. diffusion [86]. dilution [1029, 485, 1022]. dimension [997]. dimensional [872, 1035, 1051, 1563, 1065, 1044, 740, 1004, 978, 316, 538, 991, 33, 188, 610, 59, 1268, 1606, 1212, 702, 693, 1518, 987, 941, 278, 1036, 1349, 1134, 1502, 213, 543, 730, 1186, 176, 413, 379, 517, 1291, 882, 686]. dimensions [807]. Direct [312, 464, 1527, 1644, 230, 1315, 1117, 298, 993, 860, 1495]. direct-acting [298]. directed [310, 931]. direction [237]. directionality [1631]. Dirichlet [1396]. disaggregation [747]. disclosure [90]. discontinuity [1338, 1496]. discover [767]. discoveries [1035]. discovery [1224, 428, 237, 645, 1176, 410, 1411, 1024]. Discrete [455, 288, 656, 95, 640, 833, 822, 577, 1108, 441, 1599, 460, 1540]. Discrete-time [288]. discretely [973]. discretization [267]. discriminant [316]. discrimination [546, 1652, 966, 1478, 1286]. Discussion [559, 558, 475, 325, 1076, 561, 478]. Disease [560, 1292, 1212, 423, 1090, 747, 536, 607, 559, 558, 87, 1504, 846, 1431, 946, 28, 703, 367, 704, 1346, 131, 557, 561, 290, 444, 977, 1023, 767, 1356, 523, 144, 771, 902, 1462, 92, 961, 449, 1302, 345, 1396, 995, 387, 1575, 1269, 1313, 676, 525, 1376, 1460, 366, 510, 441, 64, 1501, 1362, 532, 383, 811, 396, 213, 375, 543, 777, 60, 514, 951]. disease-19 [1396]. diseases [1418, 1153, 931, 1290, 505, 1103, 253, 1073, 1607, 193, 1501]. Dismantling [246]. disorder [953]. disorders [1558]. disparities [367, 13]. dispersed [1083]. dispersion [1621]. Displaying [445]. disseminated [953]. dissimilarities [656]. distance [605, 1573]. distributed [90, 1529, 1210, 216]. Distribution [938, 1493, 1225, 95, 640, 1118, 472, 1516, 898, 341, 406, 927, 1404, 1446, 1274, 307, 651]. Distribution-free [938, 1493, 927]. Distributional [1550, 1196]. distributions [1202, 1258, 822, 1474, 419, 123]. District [455]. District-level [455]. divergence [1386]. diversity [502]. DL [1060]. Do [1519, 1557]. document [153, 154]. done [1557]. don’t [300]. Dose [1091, 471, 937, 342, 384, 641, 261, 1069, 34, 647, 926, 1592, 65, 821, 1092, 783, 815, 979, 448, 1228, 490, 1607, 1085, 1647, 1541, 1568, 1445, 1598, 1067, 1066, 1070, 867]. dose-escalation [261]. dose-finding [342, 1069, 34, 647, 1092, 979, 1228, 490, 1647, 1445, 1067, 1066, 1070]. dose-optimization [1598]. dose-ranging [1607]. dose-response [65, 821, 783]. dose-toxicity [448]. dosing [982, 465]. Double [1107, 736, 1266, 837, 1298]. double-Cox [1266]. Double-wavelet [736]. Doubly [1503, 1239, 133, 1143, 158, 1034, 791, 1436, 695]. downscaling [994]. dramatically [832]. driven [583, 677, 174, 925, 396]. driver [1018]. drop [208]. drop-out [208]. dropout [1208, 839, 369, 2, 1253, 332]. dropouts [610, 1244, 918]. drug [1096, 304, 1173, 1592, 949, 768, 410, 507, 1487, 815, 1152, 465, 629, 1590, 78, 107]. drug-drug [1152, 1590, 78]. drugs [1486]. Dual [1647, 572]. Dual-agent [1647]. dual-criterion [572]. Duchenne [582]. due [1626, 450, 167, 349, 489, 1501, 1444, 373]. durability [690]. duration [794, 469, 1374, 611, 711]. during [1378, 147, 707, 477, 556, 764]. Dynamic [1049, 598, 444, 1064, 1659, 1552, 1634, 994, 959, 64, 1100, 852, 127, 656, 635, 1439, 272, 692, 1158, 1213, 1285, 922, 1154, 1053, 1476, 597, 294, 680, 247, 1555, 631, 701, 1494, 1034, 1399, 93, 244, 741]. dynamics [186, 1655, 914, 1320, 1302, 526]. dystrophy [582]. e-cigarette [1583]. each [334]. Early [147, 342, 521, 1221, 451, 641, 290, 1528, 232, 967, 507, 1607, 888]. early-phase [232, 1607]. Ecological [28, 52, 859]. Editorial [48, 320]. education [798, 800]. EEG [1584]. Effect [844, 1328, 334, 1585, 997, 677, 440, 464, 1527, 658, 1260, 567, 1332, 404, 30, 689, 1652, 646, 615, 1529, 229, 114, 1483, 930, 1003, 1659, 486, 794, 1455, 1649, 364, 1370, 1371, 485, 1457, 395, 86, 1021, 167, 1101, 995, 913, 600, 148, 50, 1300, 1604, 1022, 206, 860, 29, 489, 757, 1114, 835, 1281, 1282, 1408, 61, 792, 1125, 279, 842, 379, 1602, 1542, 882, 1099, 273, 1453]. Effective [676, 1481]. effectiveness [1310, 981, 1055, 111, 721, 1369, 975]. Effects [193, 1372, 102, 782, 415, 1283, 642, 1086, 1484, 1475, 515, 1138, 21, 953, 95, 640, 1098, 595, 1439, 1658, 948, 1344, 398, 786, 1177, 120, 894, 1278, 201, 266, 1173, 341, 1448, 551, 117, 462, 406, 1356, 1646, 259, 108, 924, 1410, 1572, 542, 494, 947, 1477, 427, 1046, 1648, 753, 1644, 1474, 773, 1412, 1608, 1315, 52, 359, 1267, 1567, 861, 671, 1073, 495, 1230, 1184, 1117, 1446, 921, 1432, 418, 81, 243, 366, 758, 554, 1409, 993, 860, 1443, 277, 1194, 1054, 857, 532, 145, 775, 1495, 425, 116, 280, 1598, 1136, 121, 1010, 837, 1192, 516, 1015, 522]. effects [1339, 296]. efficacy [223, 1504, 1532, 932, 292, 722, 867]. efficiency [1526, 96, 417, 1007]. Efficient [845, 592, 1470, 221, 921, 370, 196, 1573, 471, 1044, 737, 862, 1084, 496, 100, 547, 283, 1186, 1479, 343]. efficiently [236]. EHR [26]. EHR-based [26]. Eigen [1573]. Elasticity [378]. elderly [598, 449]. electrocardiogram [883]. electroencephalography [970]. electronic [173, 139, 66, 1071, 157, 210, 163, 494, 439, 952, 1141, 107]. elicitation [1311, 200]. eligibility [298, 1496, 1249]. Ellenberg [476, 475, 477, 478, 482]. emanating [95]. EMA’s [1486]. Embedded [1422, 1158]. emerge [932]. emergency [217, 419, 1607, 1337]. emerging [66]. Emission [612]. emotion [1555]. emphasis [1205]. Empirical [205, 1132, 1372, 1612, 90, 189, 1537, 587, 1230, 888, 51, 809]. emptive [995]. emulation [995]. emulators [236]. encoding [573]. end [1462, 36]. end-organ [1462]. endemic [946]. endpoint [1463, 1591, 885, 985, 908, 824, 689, 850, 1577, 240, 469, 254, 1037, 1237, 729, 913, 77, 1025, 575, 962, 1345, 1473, 808, 909, 945]. endpoints [1090, 1405, 1096, 1503, 1166, 1057, 1354, 847, 54, 299, 894, 252, 1659, 73, 886, 1492, 282, 260, 1313, 727, 858, 1594, 675, 1513, 532, 804, 1159, 1417]. enhance [1631, 1286]. Enhanced [587, 1578]. enhancement [293]. enhancements [934]. Enhancing [1549]. enrichment [506, 368, 588, 858]. enrolling [257]. enrollment [1588, 1647]. ensemble [892, 1505, 1154, 1214, 1282, 1352]. ensembles [746]. ensure [410]. entropy [591]. Envelope [1016, 421, 831]. Envelope-based [1016]. environment [850, 67, 55, 1627, 15, 965, 868]. environmental [117, 627, 1261, 861, 164, 1108, 405]. epidemic [1523, 1510, 895, 1137]. epidemics [378]. epidemiologic [649]. epidemiological [655]. epidemiology [153, 1156, 154, 405]. epidermal [720]. epithelial [444]. eQTL [1268]. equal [1246, 1469]. equality [1386]. equation [1508, 27, 227]. equations [256, 652, 191, 1243, 731, 1292, 820, 1485, 88, 136, 196, 1427]. equivalence [1452, 985, 809]. era [244, 107]. Erratum [1626]. Error [388, 1193, 1428, 636, 805, 1190, 450, 1256, 1605, 498, 493, 1033, 293, 153, 270, 234, 1423, 1645, 1486, 365, 1021, 103, 214, 1232, 1132, 1078, 154, 1316, 247, 1341, 1594, 632, 245, 1629]. error-contaminated [1316]. Error-corrected [388]. Error-prone [247, 1428, 636, 1256, 493, 1645]. errors [1235, 755, 874, 1210, 1500, 94, 896, 344, 370, 1599, 1353]. errors-in-variables [1599]. escalation [261]. establishing [395]. estimable [395]. Estimand [669, 1174, 459]. Estimands [1002, 1077, 555, 1276, 3, 1199, 1307, 496, 1080, 1074, 1640, 1076, 1075, 1079, 1456, 940, 91]. estimate [76, 584, 592, 642, 1475, 819, 849, 1187, 673, 1256, 818, 1270, 896, 680, 136, 1408, 425, 273]. estimated [1480, 539]. estimates [1225, 1198, 117, 167, 682, 751]. Estimating [667, 1017, 1086, 635, 744, 1118, 677, 120, 1290, 996, 615, 1525, 861, 896, 1073, 306, 1432, 399, 45, 29, 1377, 619, 256, 334, 127, 1283, 652, 1512, 191, 1243, 731, 493, 505, 874, 1538, 27, 1279, 259, 1572, 1292, 227, 820, 1651, 1360, 1191, 1608, 683, 1485, 377, 296, 418, 758, 136, 1456, 489, 196, 1345, 1427, 1559, 258, 389, 777, 837]. Estimation [24, 1504, 454, 932, 404, 1515, 486, 535, 829, 1330, 1644, 605, 1401, 1078, 1325, 1656, 193, 625, 760, 116, 1342, 1136, 882, 1570, 1328, 118, 1008, 1388, 428, 1612, 972, 291, 177, 1440, 805, 1040, 126, 1470, 1508, 138, 1044, 235, 658, 915, 308, 946, 1436, 567, 548, 838, 1147, 1177, 250, 372, 1361, 1061, 312, 16, 1537, 538, 1293, 114, 163, 1083, 110, 587, 930, 462, 109, 854, 174, 1553, 6, 1547, 494, 602, 739, 501, 1430, 497, 713, 820, 1167, 1398, 1488, 1505, 1651, 862, 1046, 1648, 1397, 311, 810, 886, 1492, 221, 1053, 1379, 346, 359, 1391, 1639, 891, 1238, 1143, 172]. estimation [50, 1104, 1233, 137, 1321, 1117, 1446, 1206, 921, 552, 1369, 1109, 38, 43, 1214, 1426, 294, 858, 158, 1411, 993, 1443, 1107, 1306, 1027, 1054, 1375, 984, 455, 1282, 195, 699, 1134, 1495, 1494, 942, 452, 388, 919, 280, 1125, 60, 639, 1034, 1192, 1479, 1353, 1099, 1015, 522, 354, 589]. estimator [1578, 398, 839, 445, 1553, 283, 1506, 1453]. estimators [76, 1428, 1585, 515, 1138, 906, 840, 464, 1343, 1448, 878, 352, 364, 62, 1001, 753, 743, 287, 1456, 633, 1326]. ethical [487]. Ethiopian [800]. Euclidean [605, 1573]. evaluate [997, 834, 205, 675, 181]. evaluated [1124]. Evaluating [87, 643, 1652, 1155, 878, 1419, 1454, 975, 1453, 641, 1384, 1649, 1277, 446, 609, 701, 1363]. Evaluation [1650, 1162, 537, 352, 1463, 428, 1168, 1116, 953, 1095, 1354, 69, 939, 1449, 146, 1478, 624, 1151, 1037, 1487, 1325, 570, 751, 1573, 1141]. Event [78, 825, 1636, 1234, 339, 1216, 361, 712, 118, 584, 1405, 1032, 1503, 1118, 1571, 1504, 846, 1472, 985, 1614, 1045, 856, 404, 261, 292, 1207, 870, 1064, 59, 1105, 174, 1659, 1093, 1515, 240, 588, 1398, 571, 1001, 753, 859, 469, 392, 260, 365, 766, 1406, 1404, 337, 577, 287, 844, 828, 77, 1184, 608, 1607, 858, 158, 761, 1393, 56, 1594, 1420, 11, 651, 218, 1018, 1182, 555]. event-contingent [859]. event-dependent [856]. event-history [218]. event-related [856]. Event-specific [825]. events [101, 1536, 1050, 1550, 30, 566, 1549, 1285, 497, 571, 1644, 1315, 1406, 1638, 1446, 206, 1274, 940, 988, 1468, 825, 1583, 91, 668, 1192, 1271, 1182, 107]. evidence [948, 783, 1640, 725]. evolution [1153]. Exact [30, 1278, 467, 62, 206, 1653, 36, 629, 394, 1537, 1117, 81, 1274]. examine [148, 1030]. exceed [584]. Excess [177, 562, 682, 1156, 452]. exchangeability [1334, 1005, 1539, 1617]. exchangeability-nonexchangeability [1334]. exclude [1183]. exclude-then-impute [1183]. Existing [66, 537, 446]. expanded [242]. expansion [1639]. expectation [442, 1023]. expectation-maximization [1023]. expected [1253]. expectile [829, 663]. expensive [236]. experiment [276, 665]. Experimental [1272, 1596, 472, 591, 750]. experiments [1276, 3, 1173, 270, 949, 865, 178]. experts [1351]. Explained [333, 1110, 882]. explanations [1012, 1524]. explanatory [748, 944]. Explicit [1068]. Exploiting [1631, 226, 15]. Exploration [1229, 872]. Exploratory [341, 73, 1622]. Exploring [1320, 1315, 139, 1659]. exponential [1516, 1620, 487]. exposed [887]. exposure [1163, 263, 1190, 677, 464, 1551, 210, 818, 117, 765, 696, 344, 1124, 1342]. exposures [1163, 636, 856, 1356, 627, 1456, 868]. expression [1246, 956, 785, 666, 1399]. Extended [304, 1445, 1345]. Extending [887, 982, 141, 424, 771, 968, 52, 1635, 1209]. extension [773, 1647]. Extensions [445, 1328, 529, 486, 1433]. external [336, 512, 1400, 586, 828, 630, 1590, 1617, 503, 616, 717]. extraction [1178, 1294]. extrapolation [710, 868]. extreme [671]. eye [771].

f [1582]. f-aCFO [1582]. facilities [101]. facility [683]. facility-based [683]. factor [714, 22, 1279, 1606, 916, 1267, 1273, 381, 1291]. factorial [1002, 617, 1509, 784]. factorization [1403, 1395]. factors [1284, 50, 1544, 1460, 1482]. failure [1234, 584, 636, 830, 1361, 130, 1547, 602, 1398, 1520, 1397, 1609, 357, 1316, 807, 1627, 675, 745, 1063, 91, 1355, 596]. failure-time [91]. failures [477]. fair [88]. false [887, 1176, 1411]. familial [1258]. families [1324, 1516]. family [407, 1605, 990, 164, 343]. family-wise [1605]. Familywise [1594, 103]. fansy [887]. Farrington [708]. Fast [1177, 1246, 1564, 1036]. fatality [854, 6]. favorable [1345]. FDA [78]. feasibility [711, 616]. Feature [826, 1176, 519, 1212, 1616, 23, 1294, 573]. Feature-specific [1176]. features [1548, 265, 1394]. Federated [1335, 1538]. Fetal [1461]. fever [683]. few [1419]. fiber [720]. fibrosis [64]. fields [311, 1060]. file [1364]. fill [954]. filter [1618]. find [867]. Finding [1542, 1091, 471, 937, 342, 1069, 34, 647, 1092, 815, 979, 1228, 490, 1647, 1541, 1445, 1067, 1066, 1070]. findings [391]. Fine [183, 738, 19, 584, 183, 1036]. Fine-Gray [183]. fine-mapping [738]. Finite [1452, 906, 294, 836, 812, 284]. Finite-sample [906]. Finkelstein [945]. first [442, 890, 556, 555]. first-order [442]. fit [1585, 1532, 1119, 1642, 517]. Fitting [1378, 652, 1148, 1170]. Fixed [947, 1658, 1278, 924, 395]. Fixed-effects [947]. Fleming [463]. Flexible [1040, 1449, 1548, 432, 233, 143, 1575, 38, 1195, 686, 1405, 1166, 1434, 512, 1421, 1038, 679, 222, 1549, 990, 1348, 162, 49, 145, 893, 1087, 1192]. flexible-hazards [1087]. fluorosis [1140]. fly [1312]. fMRI [872, 1555]. focus [1225]. follow [1189, 1083, 500, 616, 11]. follow-up [1189, 1083, 500, 616, 11]. following [667, 690, 572, 1325, 116]. food [577]. Forecasting [1337, 1508, 1019, 1552, 746, 1323]. forecasts [1352]. forest [883]. forests [788, 1184]. Foreword [796]. form [1122, 346]. formula [908, 1562, 1404, 808, 909]. formulae [275]. formulas [270, 1122]. Formulating [319]. formulation [419]. forward [800, 671, 160]. found [205]. founder [603]. Four [1068, 1585, 900]. four-level [900]. fraction [250, 463]. Fractional [1582, 782]. fractions [1456, 1056]. fractions-Revisiting [1456]. Fragility [246]. Frailty [716, 1266, 1324, 1166, 1536, 209, 723, 1156, 1108, 1299, 122]. Framework [669, 256, 428, 139, 714, 1174, 1655, 1467, 1565, 1434, 27, 1081, 771, 1360, 1459, 348, 1643, 544, 921, 1382, 695, 804, 91, 1390, 1089, 273]. frameworks [1620]. Fréchet [656]. free [305, 938, 1582, 824, 929, 1493, 927, 100, 11, 1186, 1089]. frequency [1374, 813, 363, 1182]. Frequentist [1358, 1591, 1416, 1569]. frequently [432]. friends [803]. full [334, 1457]. Fully [1263, 98, 785]. function [1388, 1431, 1516, 637, 347, 1630, 1343, 996, 1033, 873, 259, 1384, 1656, 274, 1629]. function-based [1033]. Functional [366, 369, 80, 872, 1224, 127, 970, 1153, 673, 1350, 444, 1239, 622, 644, 1356, 564, 947, 1477, 338, 883, 346, 1185, 612, 959, 500, 1491, 983, 736, 741, 1181]. functional-model-adjusted [80]. functions [1499, 1209, 189, 341, 421, 831, 395, 113, 45]. Further [529]. Fusing [1450]. Fusion [515, 1138, 1013, 82, 765, 437, 1295, 1610, 563]. futility [1437]. future [1345].

g [552]. g-estimation [552]. gain [541]. gait [1121]. Gamma [1404, 304, 209, 893, 1323]. gamma-frailty [209]. Ganju [908]. gap [1586, 866]. gaps [1167]. gated [1351]. Gaussian [1311, 786, 930, 1488, 311, 1048, 1214, 465, 1115, 777, 1099, 589]. Geertruida [175]. Gene [728, 920, 850, 67, 1360, 55, 1627, 15, 965, 1188, 666, 1399]. gene-based [920]. gene-environment [850, 67, 55, 1627, 15, 965]. Gene-gene [728, 920]. gene-level [1360]. general [617, 765, 762, 1509, 219, 200, 122]. generalizable [540]. Generalized [457, 466, 1112, 284, 65, 1000, 136, 1491, 1603, 1428, 256, 652, 1166, 1028, 191, 1187, 1436, 360, 1147, 353, 1243, 978, 731, 1207, 1628, 733, 706, 227, 820, 1172, 392, 1525, 1413, 1419, 913, 1485, 214, 1184, 552, 987, 277, 89, 196, 893, 1427, 1454, 1031, 708, 1291, 1415]. Generalizing [391]. genes [1020, 363, 573]. Genetic [245, 1546, 96, 898, 374, 164, 776, 1294, 563]. genetics [653]. Genome [302, 1399, 139, 125, 1014, 738, 1245, 850, 470, 1619]. Genome-wide [302, 1399, 1014, 1245, 850, 1619]. genomic [415, 1202, 554, 1382]. geo [387]. geo-statistical [387]. geocoding [868]. geographic [241, 754].

Geographically [708]. geometric [772]. Germany [1514]. gestational [1279]. Ghajari [434]. Gill [988]. given [548]. gland [447]. Glasgow [1345]. glaucoma [1471]. global [800, 892, 421, 831]. glycemic [687]. go [1560]. goals [725]. gold [619, 620, 1126, 618, 621, 625, 1373, 388, 1314, 1566]. gold-standard [620, 1126]. Goodness [1119, 1642, 517]. Goodness-of-fit [1119, 1642, 517]. gradient [1658, 341, 274]. graph [1620, 931, 1548, 666, 1390]. graph-theoretical [1548]. Graphical [187, 634, 214, 1098, 22, 991, 1634, 164, 1048, 1214, 1362, 1433, 777, 1380, 589]. graphs [1307]. Gray [183, 19, 584, 1036]. greater [1557, 997]. greedy [1633, 1502]. grid [1479]. Group [890, 1139, 165, 252, 560, 1425, 1304, 654, 1072, 1180, 124, 415, 559, 558, 1116, 1526, 524, 182, 932, 557, 561, 146, 787, 463, 739, 794, 467, 1126, 485, 349, 1535, 1022, 609, 1594, 1282, 1427, 1414, 1192]. group-sequential [787, 463, 1535]. group-specific [1414]. grouped [1658, 1239]. groups [445, 421, 831, 718]. growth [578, 1586, 1461, 773, 1567]. Guidance [339, 613, 153, 1394, 1233, 154]. guide [119, 1338, 194]. guided [1382, 1558, 274, 563]. guideline [1486]. Gumbel [789]. Guogen [207]. GUSTO [692]. gut [780, 1413]. GWAS [63].

H [175]. Haber [559]. Haenszel [1358, 264]. Hamiltonian [315]. handle [887, 791]. Handling [1354, 239, 930, 1501, 1287, 1189, 613, 691, 1006]. haplotype [15, 436]. haplotype-based [15]. harms [1082, 599]. Harrington [463]. hazard [19, 584, 126, 1499, 189, 182, 372, 183, 723, 1143, 1156, 399, 1150, 1559, 1602]. hazards [782, 1266, 1032, 1499, 726, 710, 1295, 950, 333, 386, 1088, 452, 1087, 354]. HCOMBS [504]. Health [791, 1133, 894, 489, 173, 139, 66, 1071, 1637, 157, 953, 210, 1560, 881, 163, 259, 494, 439, 1261, 1657, 1308, 861, 337, 1080, 50, 1077, 919, 975, 952, 1141, 107]. Healthcare [1295, 974, 179, 1537, 758, 215]. hearing [1058]. Heart [1421, 367, 1575, 366, 675]. heat [10, 580, 579]. heavy [542, 968]. heavy-tailed [542]. Heckman [1435]. hepatitis [1329]. hepatocellular [888]. heredity [1627]. heredity-constrained [1627]. Heritability [407, 256]. heterogeneity [1372, 1065, 658, 308, 1530, 519, 1478, 851, 1455, 754, 1649, 437, 439, 1371, 1581, 1230, 1300, 1012, 776, 630, 941, 277, 835, 1281, 122, 279, 563]. heterogeneous [1475, 1418, 1098, 595, 997, 1614, 567, 615, 371, 1515, 1622, 494, 1101, 137, 1281, 1282, 751, 730, 1335, 1136, 668, 1543]. heteroscedasticity [617]. heteroskedastic [1624]. Heuvel [175]. Hidden [536, 466, 510, 238, 128, 1543]. Hierarchical [931, 504, 1595, 1178, 179, 186, 658, 1530, 812, 1279, 542, 1351, 1437, 1552, 85, 896, 693, 763, 112, 784, 611, 1373, 63, 884]. hierarchically [1299, 1255]. High [316, 33, 188, 872, 1035, 1051, 1065, 223, 1044, 740, 1004, 856, 978, 538, 991, 610, 59, 1094, 1268, 1606, 81, 702, 865, 693, 1518, 807, 987, 941, 278, 1349, 1134, 1502, 213, 543, 730, 1186, 176, 413, 379, 517, 1291, 882, 686]. High-dimensional [316, 33, 188, 872, 1065, 1044, 1004, 978, 538, 991, 59, 1268, 1606, 702, 693, 1518, 987, 941, 278, 1134, 1502, 213, 543, 730, 176, 413, 379, 517, 1291, 882, 686]. high-efficacy [223]. high-throughput [1094, 865]. highly [95, 1486]. Hill [330, 322, 323, 324, 325, 326, 328, 329, 331]. Hispanic [1133]. histopathological [812]. Historical [1274, 1097, 531, 672, 1385, 623, 77, 1494]. history [536, 553, 1265, 1418, 846, 444, 253, 1329, 218]. hitting [890]. HIV [1570, 76, 1578, 583, 488, 838, 818, 719, 1496, 145]. HIV-1 [76]. HIV/AIDS [719]. Hochberg [654]. Holm [654]. home [1019]. homogeneity [1585, 123]. homogeneous [1204]. Hongtao [1069]. horizontal [817]. hormonal [1107]. horseshoe [1451]. hospital [86, 1131]. hospitalization [1378, 1081, 564]. hospitals [778]. House [1421]. Household [707]. HPTN [818]. Hua [207]. Hui [771]. Human [92, 449, 536, 210, 199, 221, 89, 1349]. hurdle [406]. Hybrid [446, 1405, 970, 1461, 644, 1252, 1539, 730]. Hygiene [799]. hyperparameter [1465]. hyperrectangular [1493]. hypersphere [385]. hypertension [1204]. hypervolume [1145, 1454]. hypotheses [249]. Hypothesis [870, 1322, 1202, 487, 892, 850, 1423, 156, 1227, 616, 809].

I-II [144]. I/II [1090, 607, 824, 351, 768, 490, 867]. IC [473]. ICI [187]. ICU [1536]. identifiability [1200]. Identification [559, 37, 560, 754, 1648, 993, 722, 1570, 451, 558, 998, 557, 561, 73, 1622, 535, 284, 660, 1359, 1382, 1088, 1347, 274]. identified [57]. identify [1163, 210]. Identifying [192, 10, 580, 1055, 579, 1384, 1284, 1188, 1020, 98, 1307, 1248, 1399]. IDNetwork [846]. ignorability [276, 1608]. ignorable [1298]. II [1090, 607, 471, 1534, 885, 1118, 937, 1434, 292, 824, 293, 1105, 144, 647, 1592, 1005, 351, 768, 601, 1233, 490, 611, 1568, 1417, 867]. II/III [1105, 1592]. III [1105, 1592, 1404, 576, 790]. IIIA [961]. illness [846, 1638]. illness-death [846, 1638]. image [1563, 381]. images [1497, 400]. imaging [1224, 127, 673, 812, 774, 1011, 653, 612, 400, 736]. imaging-based [812]. imbalance [170, 1469]. Imbalanced [152, 967]. immediate [1638]. immune [292, 1106, 1305]. immuno [257, 1367]. immuno-oncology [257, 1367]. immunodeficiency [210, 221, 89]. immunotherapy [292, 824, 1161, 315, 794, 351, 577, 1228, 61, 792]. Impact [799, 1187, 267, 1163, 464, 833, 898, 117, 934, 1588, 957, 253, 1496, 198]. impacting [1284]. impacts [1610]. impairment [1100]. imperfect [1198, 570]. implantable [86]. implement [552]. implementation [1092, 1469]. implications [25, 349]. importance [1170, 1141]. important [1188]. imprecise [1082]. improper [428]. improve [1368, 1503, 594, 665, 1486, 1042]. Improved [891, 1607, 984, 863, 249]. improvements [453]. improves [924]. Improving [779, 1069, 72, 919, 1066, 150, 173, 398, 999, 1067, 1070]. imputation [1428, 1208, 1612, 698, 1095, 194, 438, 1472, 1550, 843, 737, 248, 1222, 622, 822, 264, 613, 1389, 1435, 696, 473, 1039, 1441, 1157, 1325, 758, 88, 691, 2, 200, 1006, 1024, 814]. imputation-based [248, 1222]. Impute [1183]. Impute-then-exclude [1183]. imputing [667, 1183, 744]. inactivity [430]. incidence [1578, 177, 556, 1551, 838, 163, 221, 995, 683, 526, 45, 396, 79]. incidences [147]. incident [709]. include [530]. Inclusion [924, 1575]. income [1332, 420]. incomplete [1095, 1554, 426, 822, 1422, 1435, 500, 200, 1006, 1024, 416, 657]. Inconsistencies [270]. Inconsistency [1359, 10, 580, 579, 1363]. incorporate [1096]. incorporated [1530, 1392]. incorporates [1578]. Incorporating [34, 265, 709, 1261, 485, 727, 544, 611, 60, 244, 717, 1083, 623, 728, 1251, 1373, 1617, 452]. incorporation [1400]. increase [1642]. increases [263]. incubation [1225, 587]. independence [15, 925]. independent [1183, 513, 272, 691]. index [187, 428, 1086, 546, 1217, 648, 188, 966, 733, 417, 100, 1259, 1603, 1454, 918, 388, 1314, 749, 246]. indications [768]. indicator [401]. indicators [102, 556, 1112, 499]. indices [502, 1170]. indifferent [5, 115]. indirect [361, 464, 1315, 317, 496, 495, 844, 1074, 1076, 1117, 921, 1075, 993, 1443, 778, 1495]. indirectly [461]. Individual [512, 148, 1156, 845, 423, 90, 1418, 1584, 903, 643, 1548, 109, 1646, 13, 766, 167, 172, 1030, 296, 418, 1426, 1481, 122, 340, 1495, 1137, 884]. individual- [1495]. individual-level [423]. individual-specific [1548]. individualization [594]. Individualized [83, 1537, 1519, 849, 689, 1652, 926, 1325, 1262, 1147]. individually [528, 1427]. infant [780]. infants [771]. infarction [526]. infected [887]. infection [1225, 935, 1254, 1305, 345]. infections [488, 1377, 122]. infectious [423, 87, 1504, 1153, 946, 28, 1290, 345, 193]. Inference [1318, 1344, 276, 874, 1370, 1514, 612, 1463, 973, 1071, 669, 640, 935, 125, 126, 1571, 1203, 1044, 28, 191, 904, 30, 267, 1346, 1278, 1038, 1293, 1529, 1483, 679, 977, 1538, 27, 1218, 315, 133, 1564, 769, 31, 205, 197, 829, 947, 1422, 1438, 1046, 1556, 1084, 1176, 762, 106, 1615, 1226, 600, 1441, 1600, 204, 688, 206, 43, 1152, 1656, 770, 405, 158, 987, 860, 370, 1115, 456, 1077, 1362, 649, 684, 1236, 1420, 722, 919, 1231, 1335, 1215, 842, 825, 1294, 749, 422, 1296, 382, 999, 1271, 285, 864, 1089, 741]. inferences [90, 141, 1209, 673, 15]. inferiority [1041, 1532, 645, 1126, 1160, 1159]. Inferring [881, 519]. infimum [1205]. Inflated [1468, 95, 640, 938, 304, 1211, 1140, 1522, 571, 337, 1284, 1545, 1414, 307, 382, 685, 880]. inflation [912, 608]. influence [1631]. influential [1615]. influenza [707, 746, 994, 1352, 1137]. inform [377]. informant [1163]. Information [1530, 426, 1463, 1372, 119, 779, 1334, 1554, 1148, 114, 1400, 818, 587, 232, 672, 1385, 463, 769, 1422, 260, 728, 676, 1230, 1251, 630, 541, 1590, 1501, 1253, 611, 1494, 1445]. Information-incorporated [1530]. information-theoretic [1463]. Informative [509, 1216, 401, 943, 468, 1398, 604, 608, 369, 218]. informed [1103]. Informing [700, 426]. inhibition [1586]. initiated [859]. initiation [1404, 425]. INLA [1625]. inpatient [449]. insensitivity [971]. Insights [197, 1137, 599]. institutional [215]. Instrumental [669, 1332, 1288, 289, 1326, 1629]. instrumented [1304]. instruments [592, 1641, 659, 211]. integer [74]. integer-valued [74]. integers [337]. Integrated [606, 187, 1459, 785]. Integrating [22, 1290, 587, 1382]. integration [845, 1319, 793]. Integrative [458, 1616, 316, 851]. intensities [974]. intensity [339, 263, 1128, 432]. intensive [1378, 176, 772]. intention [1532]. intention-to-treat [1532]. inter [1584]. inter-trial [1584]. Interaction [730, 121, 900, 16, 920, 266, 20, 728, 1590, 196, 78, 1186]. interactions [658, 850, 67, 810, 55, 148, 1030, 525, 1152, 1627, 965, 107]. intercept [1557]. intercept-slope [1557]. intercurrent [940]. interference [1562, 1015]. interim [1437, 1072, 41]. intermediate [993]. intermittent [130, 966, 171]. interpretability [353, 624, 1141]. interpretable [705, 1390]. interpretation [655, 839, 1286]. interrupted [1168]. interruption [145]. interval [411, 457, 1512, 1546, 189, 1223, 548, 1119, 372, 1361, 16, 163, 955, 174, 521, 1122, 1398, 1520, 862, 62, 1379, 1567, 345, 135, 1653, 490, 1306, 1375, 195, 827, 1347, 942, 1144, 166, 1231, 745, 1573, 1583, 1342, 1540, 218, 255, 514, 596, 522]. interval-bounded [1347]. interval-censored [1546, 1223, 1119, 372, 1361, 163, 955, 1398, 1520, 862, 345, 135, 1306, 827, 1231, 745, 1540, 218, 255, 514, 596]. interval-censoring [189]. intervals [1276, 1198, 1440, 1534, 251, 3, 394, 286, 1200, 822, 1492, 605, 674, 297, 1341, 231, 633, 757, 1113, 1301, 178, 150]. intervention [900, 816, 12, 1410, 878, 753, 1307, 1374, 29, 975]. interventions [1532, 462, 895, 841, 503]. intra [1584]. intra-individual [1584]. intraclass [1331]. intracluster [1281]. Introduction [770, 359, 1060]. invalid [1326]. invasive [1303]. Inverse [1436, 497, 695, 700, 450, 1197, 1553, 262, 1327, 1169, 215, 760, 1125, 1296]. inverse-variance [760]. investigate [13]. Investigating [1496, 1567]. Investigation [518, 415, 405]. involved [1379]. involving [428, 974, 966, 1615, 691]. IPD [90]. irregular [1226]. Ising [1020]. Issues [1253, 875, 434, 433]. Italian [556]. Italians [1049]. Italy [13]. item [102, 692, 304, 613, 295]. item-response [692]. iterative [1490].

J [47, 908]. Japan [367]. Jason [183]. Jeffrey [476]. Jiang [207]. Jixian [1069]. Johnson [1129, 1130]. joinpoint [374, 934]. Joint [35, 1466, 1128, 1381, 1213, 1394, 719, 402, 608, 1374, 1444, 749, 732, 451, 585, 1586, 1166, 1366, 1658, 1536, 1260, 432, 767, 315, 1081, 1093, 748, 1462, 1201, 1315, 337, 1313, 866, 1039, 1476, 597, 1393, 1299, 1115, 1454, 1367, 382, 218, 1255, 741, 273, 951, 1626]. Jointly [917, 1636, 292]. Journal [894]. judgment [331]. jump [126]. Just [473]. justification [934].

Kaplan [76, 445, 303, 633]. Kappa [137, 688]. Kernel [713, 1407, 335, 802, 398, 462, 1381, 1349, 1494]. kidney [1572, 995, 532]. KNN [1308]. KNN-based [1308]. knockoff [1618]. knockoffs [527, 660]. know [300]. knowledge [1261]. known [713]. KORA [526]. Korean [854, 6]. kriging [1017, 1086].

l [59]. L1 [834]. labor [83]. laboratory [545, 1493]. lag [1210, 1404, 1602]. lagged [1072]. landmark [1213, 366]. Landmarking [1549, 866, 866]. landscape [66]. Laplacian [904]. Laplacian-P-splines [904]. large [527, 1564, 470, 1423, 1389, 683, 1441, 1460, 919, 1182, 540, 933]. large-scale [1423, 683, 1441, 1460, 919]. Lasso [658, 1416, 1282, 97, 814, 626, 1308, 1036]. Lasso-type [1036]. late [1091, 1205, 507, 867]. late-emphasis [1205]. late-onset [1091, 867]. late-phase [507]. latency [253]. Latent [1623, 499, 1636, 619, 102, 598, 620, 726, 26, 519, 27, 1356, 1275, 203, 439, 886, 1229, 1412, 402, 1313, 1413, 618, 621, 763, 1482, 362, 381, 670, 778, 375, 1146, 1291, 373]. lateral [510]. Latinos [1133]. Latouche [183]. lead [1519]. lean [1324]. learned [800]. Learning [439, 164, 628, 1558, 1519, 310, 1439, 1574, 236, 408, 210, 1061, 892, 991, 615, 1402, 926, 1175, 922, 986, 1268, 1154, 821, 1395, 1391, 1639, 1487, 1476, 492, 680, 159, 302, 1555, 701, 1616, 1580, 409, 952, 1141, 1060, 1262, 686]. learning-type [1574]. least [759, 978, 458, 1016, 1544]. least-squares [759]. leave [1458]. leave-one-out [1458]. lecture [328, 321, 323, 330]. Lee [854]. left [1421, 1106, 163, 662, 345, 337, 928, 1108, 45, 960]. left- [163]. left-censored [1106, 960]. left-truncated [662, 345, 337, 928]. left-truncation [45]. Leishmania [1305]. length [428, 1223, 449, 419]. length-biased [1223]. lengths [1083]. Leonard [1129]. leprosy [772]. Lessons [1183, 800]. letter [207]. leukemia [1471]. level [423, 900, 420, 114, 82, 227, 1201, 1360, 1037, 148, 238, 455, 1464, 1373, 1006]. Leveling [1486]. Leveling-off [1486]. levels [384]. Leveraging [725, 77]. life [1235, 733, 1112, 123]. lifestyle [1374]. Likelihood [1480, 714, 636, 805, 1062, 1050, 946, 367, 372, 1361, 1346, 1293, 587, 739, 878, 1384, 902, 1397, 1277, 810, 113, 1082, 921, 1496, 817, 1348, 1426, 1054, 984, 1589, 684, 942, 1144, 460, 751, 1231, 1215, 335, 363, 1142, 1015]. likelihood-based [810, 363, 1142]. likelihoods [1023]. limit [76, 1531, 373]. Limitations [819, 350]. limiting [1029]. limits [1626, 1028, 467, 1486, 1231, 1444]. linear [1234, 142, 442, 291, 126, 1028, 1258, 237, 1436, 1421, 786, 353, 466, 1207, 71, 955, 733, 426, 1489, 1645, 862, 1046, 1244, 1172, 392, 762, 1000, 461, 1021, 1419, 682, 1407, 429, 552, 987, 1491, 208, 1253, 362, 1362, 1464, 171, 965, 1569, 1506, 918, 1297, 97, 213, 543, 335, 517, 1415, 1479]. Link [416, 859, 99]. linkage [1364, 1597]. linkage-data [1597]. linked [66, 1424]. literature [551]. load [145]. Local [1411, 192, 407, 87, 1202, 1240, 1005, 1176, 1496, 732]. location [1177, 859, 1412, 1010]. location-scale [1177, 859]. log [126, 463, 780, 99, 1362]. log-contrast [780]. log-hazard [126]. log-linear [1362]. log-rank [463]. logistic [102, 299, 1069, 1245, 203, 813, 224, 649, 491, 1236, 1642, 1146, 1067, 75, 1066, 1070]. logit [136, 413]. logrank [1205, 905]. London [799]. long [182, 1161, 1549, 253, 1610, 576, 1109, 992]. long-term [182, 1161, 1549, 1610, 576, 1109, 992]. Longitudinal [225, 412, 1262, 1636, 1234, 585, 1208, 442, 21, 471, 95, 640, 1133, 1095, 938, 1366, 1466, 1467, 511, 740, 900, 1243, 267, 292, 894, 432, 466, 290, 996, 444, 789, 943, 610, 406, 767, 1140, 132, 1218, 706, 1275, 634, 1623, 1093, 385, 1489, 829, 947, 748, 20, 922, 820, 439, 752, 1394, 1201, 1606, 613, 1389, 810, 1365, 968, 719, 1000, 1229, 1273, 1305, 907, 1269, 1313, 1226, 889, 1601, 897, 1025, 1170, 597, 608, 366, 369, 702, 441, 1518, 492, 2, 1393, 435, 888, 1259, 1603, 362, 1115, 965, 1347, 1569, 918, 1297, 288, 1414, 1250, 460, 23, 814, 1367, 732, 176]. longitudinal [639, 1417, 244, 128, 1255]. longitudinal-survival [1367]. LOO [1458]. look [160]. loop [1363]. Lopez [831]. Lopez-Pintado [831]. loss [637, 1522, 1406]. lost [454]. Low [1291, 1017, 1086, 1045, 1332, 420, 1207, 505]. low-income [1332]. low-prevalence [505]. Low-rank [1291, 1017, 1086]. Lq [739, 1051]. Lq-based [1051]. Lq-likelihood [739]. LSHTM [797, 798]. lung [972, 75].

M&M [794]. M.Sc. [797]. Macaskill [115]. machine [1519, 802, 408, 892, 615, 1402, 926, 1154, 1391, 1639, 492, 1141]. magnetic [1224, 127, 673, 774, 736]. magnitude [1585, 1557, 1483]. main [779, 121]. maintain [1249]. Maintaining [191]. major [934, 298]. maker [328]. making [1241, 1434, 724, 496, 877]. malaria [1562, 1254]. Malinovsky [559]. maltreatment [169]. mammogram [1497]. mammography [1303]. management [1529]. managing [1528]. manifold [1145, 1454]. Manitoba [423]. Mann [1276, 3, 863, 424]. Mantel [264, 1358, 653]. many [519, 718]. mapping [747, 1311, 738, 931, 347, 1292, 1085]. MAR [208]. margin [985]. Marginal [1059, 468, 902, 1131, 778, 1585, 652, 640, 635, 677, 353, 1218, 600, 495, 38, 399, 1079, 1559, 918, 1271, 273]. marginalized [652]. margins [1078]. marked [1302, 1206]. marker [959]. markers [1449, 1055, 502, 687, 1094, 439, 1313, 278, 711, 166]. market [1556, 629]. Markov [442, 355, 1204, 126, 466, 111, 209, 311, 1302, 1401, 510, 238, 987, 491, 901, 57, 128, 1543]. Markov-modulated [1302]. Martingale [1502]. Martingale-residual-based [1502]. masked [1170]. mass [1221]. massive [59, 1063, 354]. matched [334, 177, 1312, 394, 1175, 1135, 813, 657, 952, 178]. matched-pair [657, 178]. Matching [361, 844, 718, 160, 118, 667, 334, 157, 678, 567, 408, 462, 143, 1157, 1559, 837, 1195]. Matching-adjusted [361, 844]. maternal [896]. mathematical [236]. MATLAB [1508]. matrices [1279, 907, 925, 1134]. matrix [1403, 1448, 1253, 1573, 1291]. matrix-factor [1291]. matrix-variate [1291]. matters [1229]. maturation [962]. Maxcombo [794]. Maximin [1300]. maximization [1023, 230]. maximizing [1052]. Maximum [372, 1361, 334, 440, 470, 739, 878, 794, 1397, 1082, 1426, 984, 1144, 1231, 1215, 1469, 1015]. MaxSPRT [1132]. Maxwell [406]. may [584]. MCAR [638]. McNemar [394, 936]. Mean [410, 1388, 24, 411, 1235, 7, 271, 454, 1219, 180, 1161, 733, 1155, 270, 634, 822, 1112, 1430, 424, 352, 94, 1267, 664, 969, 1237, 928, 1306, 781, 1114, 543, 255, 989]. mean-based [1155]. mean-variance [1430]. means [592, 839, 1576]. measles [1431]. Measure [639, 407, 1481, 378, 1349]. measured [300, 1374, 917, 1621]. Measurement [755, 247, 592, 805, 1235, 1190, 450, 498, 996, 1033, 130, 153, 234, 1021, 154, 370, 245, 1629]. measurements [819, 719, 32, 1347, 1427]. Measures [1110, 1523, 1318, 881, 270, 437, 62, 94, 705, 1000, 1657, 674, 1073, 1376, 541, 40, 1150, 1010, 599]. Measuring [1164, 568, 550, 1339, 1392]. mechanisms [1315, 1390, 285]. mechanistic [266, 20, 603, 1445]. median [640, 1392]. Mediation [1190, 1554, 993, 1120, 379, 1563, 1571, 1531, 165, 802, 1260, 266, 870, 1211, 1659, 104, 566, 1545, 499, 693, 606, 565, 670, 1329, 983, 790]. mediator [1563, 1531, 464, 870, 1117, 499, 1654, 565, 993, 1120]. mediator-outcome [993, 1120]. mediators [1211, 104, 693, 860, 1443, 1120]. Medical [797, 322, 192, 798, 799, 1338, 1332, 734, 432, 1302, 927, 1395, 1607, 725, 1264]. Medicare [449, 1460]. medication [881]. medications [953]. Medicine [799, 327, 788, 321, 854, 324, 326, 329, 798, 735, 1493, 492, 247, 1123, 47, 559, 325, 115, 331, 322]. medicines [1082]. Meier [76, 445, 303, 633]. melanoma [1324, 222, 942]. mellitus [1471]. membership [1017, 1086]. memberships [1006]. Memorial [321, 323]. memories [1130]. Mendelian [1220, 637, 1641, 681, 71, 659, 1288, 817, 1447, 899]. merging [750]. MERIT [1423]. MERS [854, 6]. mesothelioma [253]. Meta [502, 297, 1372, 415, 24, 735, 1096, 1512, 355, 876, 903, 512, 1050, 10, 580, 1147, 643, 30, 1346, 131, 1278, 648, 1576, 531, 1448, 1280, 551, 9, 581, 1122, 579, 1254, 542, 352, 623, 954, 1490, 1615, 766, 604, 85, 1037, 783, 671, 148, 172, 1030, 979, 1230, 1643, 1446, 1359, 1012, 418, 206, 212, 781, 1194, 1363, 226, 198, 775, 1373, 1009, 1598, 651, 668, 884, 1192, 685, 1142]. meta- [1194]. meta-analyses [24, 30, 1254, 1490, 206, 198, 668]. Meta-analysis [502, 297, 1372, 415, 735, 1512, 355, 876, 903, 512, 1050, 10, 580, 1147, 643, 1346, 1278, 648, 1448, 1280, 551, 9, 581, 1122, 579, 352, 954, 1615, 766, 604, 1037, 783, 671, 148, 172, 1030, 979, 1230, 1643, 1446, 1359, 1012, 418, 212, 781, 1363, 226, 775, 1373, 1009, 1598, 651, 884, 1192, 685, 1142]. meta-analytic [623, 85]. meta-analytic-predictive [531]. meta-analytical [1096]. meta-analyze [1576]. meta-regression [131, 542]. metabolic [1275]. metabolite [957]. metabolites [1626, 1444]. metabolomics [511]. method [1636, 305, 791, 453, 1334, 189, 1219, 724, 286, 10, 580, 1633, 856, 408, 863, 234, 579, 765, 67, 380, 929, 1522, 1520, 1360, 810, 72, 1092, 815, 702, 1590, 793, 56, 200, 1282, 32, 1345, 785, 1540, 1296, 176, 274, 331, 685, 573, 982]. Methodological [434, 433, 957, 1104]. methodologies [290]. methodology [471, 1062]. Methods [314, 365, 783, 89, 1495, 1283, 1168, 1519, 669, 583, 1475, 397, 1116, 1096, 1189, 1095, 963, 1334, 1179, 939, 237, 384, 843, 1256, 681, 267, 892, 304, 493, 1402, 232, 68, 153, 502, 170, 259, 205, 1572, 82, 926, 1394, 262, 762, 1486, 1229, 1395, 1657, 674, 1615, 85, 729, 317, 496, 495, 172, 750, 525, 1604, 418, 154, 1058, 38, 88, 858, 1072, 1124, 649, 1569, 1352, 409, 79, 842, 1583, 422, 1517]. methylated [1621]. metric [1434, 1379]. metrics [1650, 730]. micro [1247]. micro-randomized [1247]. microarrays [1178]. microbial [538]. microbiome [1260, 502, 380, 780, 916, 1413, 1348, 379, 1415]. middle [420]. migraine [340]. Milan [1265]. minimal [863, 170]. Minimization [1383, 170]. Minimum [336, 586, 828, 263, 470]. Mining [1182, 78]. mis [1621]. mis-measured [1621]. misaligned [387]. misclassification [958, 1071, 1439, 1511, 153, 896, 154, 43, 416, 245]. misclassified [536, 1336, 710, 1287]. misdiagnosis [599]. misdiagnosis-related [599]. mismeasured [1125]. Missing [25, 667, 513, 1163, 791, 1309, 938, 1354, 194, 1472, 120, 1197, 839, 239, 248, 1222, 37, 622, 826, 610, 1333, 1547, 537, 1112, 1645, 497, 820, 1520, 1001, 613, 473, 1457, 896, 1485, 1327, 1157, 1369, 865, 691, 1501, 1027, 171, 224, 381, 1506, 614, 1566, 176, 285, 128, 1262, 1635]. missingness [698, 119, 1203, 1139, 235, 43, 1253, 1390, 218]. misspecification [1172, 1229, 857, 121, 555]. misspecified [346]. Misspecifying [208]. mistakes [121]. mitigates [450]. Mitigating [130, 975]. Mixed [915, 532, 845, 1193, 142, 263, 1032, 21, 95, 640, 1035, 1095, 997, 1220, 1258, 1344, 401, 786, 1177, 353, 466, 551, 406, 426, 534, 1489, 748, 1477, 1046, 859, 1365, 94, 968, 886, 1000, 773, 1412, 52, 1267, 402, 1419, 429, 612, 1426, 492, 208, 1253, 362, 171, 145, 1408, 97, 460, 245, 1010, 516, 1181]. mixed-effect [997]. Mixed-effects [532, 21, 640, 1177, 1046, 773, 1412, 52, 1010]. Mixed-type [915, 1095]. Mixture [726, 1273, 1611, 1319, 1086, 1484, 21, 268, 569, 1004, 1119, 904, 898, 812, 610, 98, 284, 713, 156, 1396, 1110, 429, 691, 78, 942, 1215, 307]. mixtures [1202, 930, 627, 259, 1351, 1261, 861]. mobility [117]. Mod [471]. Model [414, 1422, 1172, 1593, 287, 897, 1524, 857, 999, 1234, 80, 217, 1097, 451, 585, 336, 536, 562, 1319, 102, 19, 782, 1183, 1585, 1378, 735, 1266, 1178, 887, 553, 1032, 1086, 1484, 1324, 21, 471, 1512, 125, 595, 126, 268, 569, 938, 890, 1028, 997, 1235, 1366, 1470, 289, 830, 948, 1472, 912, 1449, 1344, 512, 1536, 1336, 182, 1260, 1421, 786, 533, 1119, 734, 518, 904, 299, 30, 372, 1361, 261, 432, 726, 304, 239, 1652, 371, 1106, 1510, 789, 1069, 111, 1448, 1033, 955, 12, 1020, 610, 98, 406, 767, 1140, 914, 706, 234, 203, 1279, 1623, 222, 108]. model [356, 980, 1622, 149, 209, 566, 537, 535, 602, 710, 1549, 1645, 1210, 1295, 1477, 1522, 1201, 1606, 950, 1001, 859, 1357, 1244, 1389, 961, 94, 886, 183, 762, 1229, 773, 1412, 1042, 473, 52, 1609, 766, 1624, 1267, 1305, 113, 715, 1238, 1021, 357, 1037, 1396, 783, 74, 135, 1110, 896, 1404, 337, 1575, 577, 1039, 333, 1638, 1640, 889, 586, 377, 828, 386, 1108, 1109, 1288, 206, 1653, 817, 597, 776, 1460, 1348, 510, 1214, 99, 441, 1316, 238, 136, 1341, 1491, 763, 277, 781, 988, 1393, 1299, 1027, 208, 1194, 122, 1375, 1323, 350, 145, 78, 549, 1036, 965, 1088, 893, 1347, 1468]. model [1408, 1502, 1642, 942, 918, 1297, 452, 97, 1414, 396, 213, 460, 539, 777, 1087, 1186, 1215, 814, 1137, 1541, 79, 1367, 1146, 1141, 1067, 218, 413, 884, 1066, 1070, 989, 555, 1255, 1272, 741, 1181, 354, 1286, 951, 589]. Model-agnostic [1524]. Model-assisted [287, 897]. model-averaged [735]. Model-based [1422, 1593, 999, 471, 1652, 955, 12, 1640, 1541]. model-free [1186]. modelers [483]. Modeling [411, 1484, 1418, 83, 1133, 1451, 1153, 1584, 1019, 1497, 687, 1011, 1462, 86, 340, 1231, 307, 57, 951, 1636, 1234, 1049, 423, 1626, 1637, 640, 1098, 220, 1655, 147, 35, 318, 1658, 454, 692, 1466, 1467, 1431, 281, 843, 641, 420, 704, 292, 1173, 1239, 111, 721, 315, 1356, 1275, 720, 1320, 234, 523, 1081, 1093, 1410, 564, 1213, 1437, 748, 1244, 1302, 716, 1315, 1457, 907, 85, 1269, 1039, 1631, 928, 1135, 608, 1374, 1259, 1603, 917, 1367, 1444, 732, 1540, 1329, 382, 1102, 1291, 790, 273]. Modelling [125, 1611, 1579, 1128, 1539]. Models [1242, 845, 619, 1586, 339, 142, 187, 584, 415, 513, 973, 1163, 127, 407, 291, 598, 652, 974, 1368, 1017, 1557, 1174, 635, 186, 1166, 840, 1508, 1499, 1153, 1574, 620, 1258, 1148, 1465, 833, 903, 1164, 236, 22, 1620, 14, 367, 1004, 1436, 931, 1200, 1177, 353, 518, 300, 978, 466, 737, 1207, 1059, 991, 188, 71, 943, 733, 774, 1064, 551, 202, 1538, 1333, 426, 1218, 534, 634, 957, 447, 468, 1489, 542, 1112, 1351, 1645, 227, 713, 156, 1520, 1462, 862, 1046, 1397, 956, 1172, 1365, 392, 434, 358, 968, 1634, 990, 1000, 1435, 461, 674, 395, 1567, 603]. models [219, 1270, 1561, 1026, 1313, 866, 1413, 1419, 1073, 723, 164, 172, 1476, 682, 1601, 1407, 1156, 433, 618, 621, 429, 552, 612, 1607, 399, 1048, 693, 492, 1131, 1196, 807, 987, 1627, 1524, 1306, 89, 1253, 51, 362, 1115, 455, 1589, 532, 350, 171, 827, 1373, 1569, 1580, 1642, 1144, 1617, 1009, 543, 335, 1063, 390, 382, 216, 242, 517, 1018, 1010, 1415, 343, 128, 516, 1543, 540, 772, 1339]. moderate [307]. moderation [1123]. Modern [1649]. modification [86, 17, 844, 842]. Modified [127, 1334, 856, 1020, 385, 1384, 72, 1404, 638]. modulated [1302]. Momentary [859, 52]. monitor [64]. Monitoring [578, 1523, 583, 1204, 1467, 432, 144, 260, 1025, 1274, 1072, 611, 962, 1043]. Monotone [1499, 235, 237, 113, 448]. monotonic [1401, 1642]. monotonicity [237, 1131]. Monte [126, 315, 1423]. Monte-Carlo [315, 1423]. morbidity [1473]. Morris [475, 477, 476, 478, 482]. mortality [367, 856, 529, 420, 1320, 1505, 896, 1607, 1473, 1141]. motion [1121]. motivate [197]. motivating [76]. mover [536, 757]. mover-stayer [536]. MRI [1181]. MRP [1458]. MSc [799, 798]. much [528]. mucopolysaccaridosis [961]. Multi [1280, 1088, 973, 1418, 1596, 846, 1472, 849, 1434, 1145, 487, 1605, 304, 774, 1263, 1588, 613, 1459, 1487, 848, 1251, 211, 1450, 1322, 1085, 1257, 1617, 1180, 736, 864]. multi-arm [1596, 1434, 487, 1605, 1263, 848, 1251, 1180, 864]. multi-attribute [211]. multi-center [1588]. multi-class [1145, 1322]. multi-dose [1085]. multi-item [304, 613]. multi-metric [1434]. multi-omics [1459]. multi-parametric [774]. multi-sample [1257]. multi-site [849]. multi-source [1617]. multi-span [1450]. multi-stage [1596, 1434, 1605, 1263, 848, 1251, 1180]. Multi-state [1280, 973, 1418, 846, 1472]. multi-subject [736]. Multi-threshold [1088]. multi-valued [1487]. multiarm [84]. multiarmed [46]. Multiblock [1544]. multicase [929]. multicase-free-response [929]. multicategory [1034]. multicenter [1529, 1538, 230, 1058, 1323]. multidimensional [1356]. multidrug [1426]. multigroup [123]. Multikernel [97]. multikink [829]. Multilevel [970, 564, 1299, 1255, 652, 791, 1458, 1081, 1422, 1435, 1267, 1039, 1321, 382]. Multinomial [224, 472, 491]. multiomics [316]. Multiparameter [516, 135]. multiphase [275]. Multiple [438, 1472, 1202, 822, 264, 1435, 696, 758, 691, 1257, 1006, 1024, 1159, 814, 1428, 1310, 102, 1163, 592, 1017, 1086, 1475, 157, 1096, 179, 310, 1028, 472, 1546, 194, 912, 1121, 1480, 1565, 384, 398, 1290, 261, 894, 737, 347, 1630, 248, 1222, 1038, 33, 622, 672, 504, 502, 523, 911, 104, 1462, 249, 1394, 65, 613, 810, 434, 358, 1191, 886, 282, 762, 768, 1509, 473, 395, 896, 1039, 433, 1157, 1325, 718, 36, 499, 88, 1654, 405, 1382, 841, 606, 860, 1443, 888, 1594, 1259, 1603, 122, 400, 226, 32, 778, 965, 1301, 722, 452, 1227, 804, 116, 1120, 11, 666]. multiple [1568, 422, 1296, 1136, 1542, 514]. multiple-dose [1568]. multiple-type [810]. multiplex [509]. multiplicity [1380]. Multiply [1485, 1443, 918, 280, 667, 399]. multiply-imputing [667]. multireader [929]. multiregional [751]. multisample [927]. multisource [1539]. multistage [1534, 84, 1357]. Multistate [1174, 14, 39, 553, 966, 1462, 1561, 441, 452, 1540, 218]. Multistep [1448]. multistudy [1505]. Multithreshold [535]. multivariable [1585, 681, 674, 659, 350]. Multivariate [1489, 780, 1269, 1185, 1259, 965, 1181, 192, 102, 655, 1095, 1366, 1127, 1220, 915, 931, 432, 1059, 1106, 1239, 1448, 12, 767, 634, 911, 385, 1011, 542, 443, 922, 284, 1493, 1522, 752, 1001, 311, 968, 1624, 1305, 219, 907, 74, 1313, 597, 212, 1196, 781, 1393, 888, 362, 1111, 226, 171, 412, 1373, 1619, 732, 242, 563, 1621, 741, 951]. multiwave [318]. muscular [582]. myocardial [526].

N [594, 384]. N-of-1 [594]. NanoString [785]. national [896]. natural [536, 553, 1265, 1418, 464, 1117, 1443, 1329]. nCounter [785]. near [1556]. nearly [737]. necessity [529]. need [1501]. needed [834, 944]. Negative [1557, 105, 916, 411, 887, 1403, 472, 891, 715, 1375, 140]. neighborhood [50, 987]. neonatal [609]. nerve [720, 1320]. nested [564, 346, 1306, 701, 1289, 570]. nesting [349]. Net [981, 360, 1147, 10, 580, 1187, 1562, 579]. Net-benefit [981]. Network [876, 1050, 131, 1387, 186, 355, 846, 903, 1145, 10, 580, 518, 1290, 1280, 1618, 826, 1381, 1122, 579, 542, 92, 449, 1609, 86, 766, 1103, 783, 671, 728, 1643, 1359, 510, 1555, 781, 56, 1363, 226, 198, 1009, 416, 1188, 539, 666, 983]. networks [310, 1466, 518, 1548, 493, 1587, 1634, 1631, 777, 666]. neural [1609, 1103, 539, 666]. neurobehavioral [780]. neurocysticercosis [218]. neurodegeneration [186]. neurodegenerative [1103]. neuroimaging [1620, 1634, 1042, 543]. neuropathy [1320]. neuroscientific [338]. next [933]. no [620, 618, 621, 685]. node [826]. nodes [493, 1587]. noise [111, 447, 193]. noisy [1378, 533, 43]. Non [1204, 1600, 1570, 782, 1266, 1241, 1041, 1557, 1403, 1202, 1532, 645, 1579, 789, 1538, 1515, 1126, 1401, 1624, 1160, 1569, 1420, 1159, 825]. non-adherence [1532]. Non-homogeneous [1204]. non-inferiority [1041, 1532, 645, 1126, 1160, 1159]. non-linear [1569]. non-local [1202]. non-monotonic [1401]. non-negative [1403]. non-normal [789]. Non-parametric [1600, 1241]. non-proportional [782, 1266]. non-randomised [1579]. non-randomized [1420]. non-rare [1515]. non-regular [1557]. non-response [1570]. non-shared [1538]. non-stationary [1624]. non-terminal [825]. nonblinded [742]. noncompliance [894, 98, 601, 349, 295]. noncomplier [1608]. nonexchangeability [1334]. nonignorable [1208, 120, 2, 332, 381, 176, 1390, 128]. nonindependent [44]. Noninferiority [1045, 348]. noninformative [20]. Nonlinear [640, 315, 259, 1365, 1000, 1609, 145, 516]. nonmonotone [820, 1327, 1027]. Nonparametric [1612, 138, 873, 156, 1446, 1206, 492, 899, 1479, 1512, 453, 692, 1527, 977, 851, 359, 815, 1321, 1114, 984, 1236, 1188, 539, 732, 274]. nonpharmaceutical [895]. nonprobability [649]. nonrandomly [513]. nonrare [104]. nonresponders [257]. nonstationary [786]. nontraditional [1384]. normal [1106, 789, 73, 1492, 429, 1393]. normal/independent [513]. normalized [650]. normally [216]. normative [412]. Norwegian [1324]. nosocomial [125]. not-at-random [698]. Note [58, 182, 94, 377, 1012]. Novel [68, 577, 1303, 1569, 607, 157, 126, 1658, 1334, 1158, 721, 1211, 1140, 470, 929, 1522, 427, 459, 1073, 51, 1134, 1621]. Nowcasting [556, 994]. nuisance [1651]. null [892, 1537]. number [334, 1585, 834, 1122, 470, 944]. numbers [1467, 393]. numerical [181]. nutrient [1262]. Nutrition [791, 878].

obesity [1466]. Obituary [1129]. objectively [917]. observation [139, 943, 966, 244]. observational [313, 339, 1585, 173, 1533, 310, 1040, 963, 981, 7, 271, 615, 229, 1013, 462, 153, 1218, 1562, 1649, 106, 1487, 971, 750, 1321, 1132, 154, 680, 701, 1124, 383, 1335, 1195]. observations [826, 202, 137, 1306, 625, 614, 657, 960]. observed [536, 973, 447, 461, 167, 729, 760, 1339]. observers [834]. occasion [797]. occupancy [1378]. occurring [1270]. odds [530, 1534, 833, 1582, 863, 549, 1113, 1141, 1271]. off [854, 6, 348, 1486, 1082, 606]. older [1019, 1460]. omics [1459, 1048, 1272]. Omnibus [920]. on-the-fly [1312]. oncology [1205, 885, 1118, 641, 1069, 232, 1592, 979, 162, 257, 1445, 1367, 1067, 1066, 1070, 790]. One [172, 845, 1557, 1358, 1364, 1458, 633, 516]. one-sample [1358]. one-sided [516]. One-stage [172]. one-way [845]. ongoing [378]. Online [854, 6, 1046, 1556, 1232, 1154, 378]. only [90]. onset [1091, 87, 345, 867]. ontology [1182]. open [132]. operating [428, 228, 281, 1145, 312, 1052, 679, 110, 929, 168, 743, 258, 389, 1142, 1635]. operational [599]. operationalizing [1149]. operator [978]. opioid [953]. Opportunities [350, 66, 801]. Optimal [998, 318, 840, 661, 764, 487, 1645, 1505, 821, 869, 520, 574, 1113, 389, 1433, 1063, 4, 175, 958, 635, 272, 744, 1057, 1461, 235, 724, 1061, 37, 462, 374, 521, 926, 427, 1492, 1300, 1149, 294, 680, 1107, 1558, 216, 1034, 867]. Optimising [5, 115]. optimism [674]. Optimization [1126, 314, 1592, 1568, 1598, 1380]. Optimized [151]. Optimizing [506, 1476, 41, 1439]. options [172]. oral [577, 298]. Order [31, 1543, 442, 1173, 1263, 1228, 848, 772]. ordered [54, 33, 225, 104, 136, 1443, 1513]. ordering [1440, 548, 1293, 110, 1092, 1376, 1251]. Ordinal [99, 1208, 1637, 677, 138, 833, 1158, 408, 300, 131, 110, 1140, 568, 1269, 1601, 897, 2, 362, 684, 1159, 1043, 413]. ordinary [1508]. organ [1462]. organizations [997]. original [90]. orphan [507]. Orthogonal [949]. orthogonality [94]. osteoporosis [836]. other [1576, 1289]. out-of-hospital [1131]. outbreak [87, 147, 556, 854, 6, 708]. outbreaks [946]. Outcome [1345, 1463, 1097, 336, 263, 791, 1309, 677, 194, 438, 464, 1436, 1332, 641, 1158, 210, 1197, 648, 870, 1239, 519, 1623, 104, 82, 265, 765, 748, 902, 1167, 1172, 886, 348, 1308, 586, 377, 1030, 828, 1228, 1117, 574, 1149, 344, 691, 993, 370, 435, 1289, 1072, 1115, 224, 383, 1427, 1120, 416, 422, 1296, 273]. outcome-dependent [1172, 574, 1149, 435]. outcomes [1208, 1310, 852, 712, 118, 667, 1008, 1585, 1032, 636, 179, 744, 1511, 1546, 1247, 138, 1127, 1220, 1527, 833, 843, 1579, 191, 1187, 360, 404, 131, 731, 347, 1217, 1038, 229, 1140, 1333, 109, 911, 1279, 1093, 34, 1515, 566, 771, 780, 1371, 434, 358, 254, 1442, 461, 402, 1270, 507, 1485, 172, 50, 1601, 576, 1184, 433, 547, 897, 243, 43, 702, 294, 1545, 99, 441, 159, 1123, 992, 2, 1160, 226, 123, 198, 722, 1250, 657, 1287, 1136, 60, 1034, 1621, 1542, 685, 514, 790, 880, 1255]. outlier [950, 1490]. outliers [192, 872, 562, 331]. outlying [1615]. ovarian [444, 1221]. over-dispersed [1083]. Overall [1094, 1586, 273]. overdetection [57]. overdispersed [562]. overdose [1069, 1067, 1066, 1070]. overlap [977]. overpowered [1500]. overview [409].

P [183, 115, 904, 1365]. P-Splines [1365]. Package [286]. packages [1394, 1253]. pain [1273]. pair [657, 178]. Paired [1029, 1118, 1527, 691, 1322, 32, 749, 742]. pairs [488, 394, 1497]. pairwise [1187, 360, 1147, 1628, 1339]. pancreatic [958]. pandemic [1378, 477, 764, 1528, 975]. Pandemics [476, 479, 474, 480, 478]. panel [291, 401, 1398, 136]. Paper [324]. papillomavirus [536]. parallel [1437, 879, 913, 393, 864]. Parameter [1508, 338, 242, 997, 1516, 721, 859, 1238, 565, 757]. Parameter-expanded [242]. parameter-space [997]. parameterization [1093]. parameters [1651, 392, 94, 1406, 93]. Parametric [453, 359, 1321, 1241, 512, 1119, 774, 930, 202, 1365, 907, 889, 1600, 1460, 88, 888, 1063, 422, 1629, 150]. parametrization [1548]. part [300, 128, 153, 154]. partial [1570, 853, 1241, 458, 1092, 1016, 349, 1228, 544, 1544, 1409, 965, 1297]. Partially [1234, 237, 733, 536, 225, 1547, 1645, 1244, 729, 1407, 1170, 1107, 1289, 543]. participant [845, 512, 148, 172, 1030, 1426, 884]. participant-level [148]. participants [5, 115]. partition [541]. partly [1223, 1119, 1021, 827, 942, 1144, 213, 596]. partly-interval [1144]. passive [1019, 683]. Path [1654, 714, 1659, 1643]. Path-specific [1654]. pathogen [810]. pathogenicity [363]. pathogens [125]. paths [445, 1575]. Pathway [511, 1382, 63]. pathway-based [63]. pathway-guided [1382]. pathways [974]. patient [1418, 223, 903, 643, 518, 445, 1588, 34, 766, 1407, 418]. patient-reported [34]. patient-sharing [518]. patient-specific [418]. patients [239, 1081, 1285, 439, 1131, 1087]. pattern [1208, 119, 12, 610, 1588, 510, 2, 1390]. pattern-mixture [610]. patterns [186, 720, 564, 820, 684]. PD [834]. PD-L1 [834]. peak [1185]. Pearson [1522]. pediatric [624, 351]. pediatrics [465]. peer [86, 1631]. peer-influence [1631]. penalised [1397, 1144]. penalization [230, 775]. Penalized [1200, 950, 662, 766, 723, 702, 1518, 942, 1215, 589, 1050, 862, 1176, 653, 1575, 630, 984, 413]. penalty [437, 1295, 1488]. per-protocol [1532]. perfection [605]. Performance [259, 1573, 334, 1585, 305, 1164, 146, 1069, 1402, 1155, 537, 1471, 317, 496, 495, 725, 1067, 181, 639, 1066, 1070, 1339]. performed [1528]. period [367, 1200, 587, 426, 395, 253]. periodic [411, 144]. periodontal [149, 684]. periods [708]. Permutation [1224, 1301, 1219, 180, 762, 277, 668]. Permutation-based [1224, 1301, 668]. person [90, 773, 1412, 1274]. person-time [1274]. Personalised [1162]. personalize [1164, 1339]. Personalized [1154, 911, 986, 877, 1632, 777, 1141, 1479]. perspective [1652, 1084, 1077, 390]. PERT [1404]. Peter [183, 797]. pharmacodynamic [1445]. pharmacokinetic [261, 1593, 1445]. pharmacometric [471]. pharmacovigilance [1062]. Phase [1434, 451, 1428, 1090, 607, 471, 1534, 318, 885, 1118, 937, 342, 738, 1582, 641, 261, 292, 824, 293, 232, 1105, 521, 144, 647, 269, 1592, 1005, 1165, 351, 768, 1, 507, 1404, 601, 419, 576, 979, 490, 1607, 370, 435, 162, 611, 389, 1007, 1568, 1445, 57, 1417, 343, 790, 867, 982, 1092, 1647, 1517]. phase-III [1404]. phase-type [419, 57]. phenome [139]. phenome-wide [139]. phenotype [1071, 97]. phenotypes [49, 436]. phenotyping [26]. physical [823, 199, 1374, 307]. piecewise [126, 1421, 734]. pilot [313, 503, 616]. Pintado [831]. pipeline [405]. Pitfalls [1217]. placebo [384, 690, 182, 932, 879]. plane [535, 1297]. planned [803]. Planning [1455, 520, 148, 1030]. plasma [643]. plasmode [450, 1507]. plateau [451]. platelet [1254]. platform [1596, 1605, 351, 768, 1232, 1249, 1539]. pleiotropic [415, 637, 659]. Pleiotropy [681, 817]. plot [10, 580, 579]. plots [1561]. pneumococcal [810]. POCRM [448]. Point [174, 447, 1492, 1104, 1233, 451, 1470, 108, 1491, 89, 1124, 1349]. points [1068, 605, 36, 378]. Poisson [217, 472, 304, 406, 1302, 766, 1404, 1413, 629, 1274, 1323, 295, 307, 382]. Poisson-Gamma [1404]. policy [998, 1528, 1558]. political [1523]. pollution [405]. pollution-multiple [405]. polygenicity [1360]. polynomials [782]. polypharmacy [107]. polytomous [546, 966]. pool [527, 1022]. pooled [583, 1528, 221, 386, 540]. Pooling [883]. popular [832, 1253]. Population [1545, 391, 1570, 553, 598, 592, 141, 1209, 594, 250, 114, 874, 117, 133, 564, 1175, 1026, 317, 496, 495, 1074, 1076, 1075, 296, 630, 526, 238, 1248, 15, 649, 57, 1286]. population-adjusted [1074, 1076, 1075]. Population-average [1545]. population-based [1570, 250, 133, 15, 57]. population-level [114, 238]. populations [836, 367, 1515, 38]. poset [718]. poset-based [718]. positive [656, 124, 742]. positives [811]. positivity [1099]. Positron [612]. possibility [1535]. Post [1376, 738, 1572, 1556, 1460, 629]. post-acute [1460]. post-genome-wide [738]. post-kidney [1572]. post-market [1556, 629]. Post-test [1376]. Posteriors [1474]. postmarket [304]. postmarketing [69]. poststratification [1458, 1422]. potential [667, 66, 744, 427, 1284, 1544]. potentially [610]. Potentials [1507]. Power [1127, 731, 441, 841, 1097, 700, 650, 25, 1516, 900, 404, 241, 132, 803, 265, 1330, 392, 410, 1030, 81, 1642, 945, 1272, 933]. powerful [440, 517]. PRACTical [1632, 1092, 1233, 1149, 1160, 837, 119, 194, 985, 1394, 359, 632]. practice [220, 405, 899]. practices [997, 694]. pragmatic [265]. Prais [1168]. praise [1168]. pre [210, 818, 995]. pre-emptive [995]. pre-exposure [210, 818]. precancer [536]. Precision [647, 788, 398, 924, 492, 247, 1123, 1134]. preclinical [225, 410]. PreCoG [1633]. predict [655, 902, 434, 358, 433]. predictability [1588]. predicted [1646, 1600, 1339]. Predicting [1522, 469, 790, 1610, 1583]. Prediction [583, 142, 336, 1368, 83, 595, 1574, 1465, 833, 1164, 1548, 210, 444, 239, 689, 966, 1155, 1213, 1549, 1285, 922, 1477, 1505, 1438, 1459, 434, 358, 282, 674, 1270, 1561, 1026, 586, 377, 828, 433, 597, 959, 702, 302, 631, 1524, 340, 962, 97, 1141, 75, 1291, 244, 741, 540, 1286, 1339, 686]. prediction-based [282]. Prediction-driven [583]. predictions [712, 1652, 980, 1476, 1303, 1481, 416]. Predictive [1052, 706, 166, 1511, 939, 531, 1402, 1155, 957, 73, 537, 623, 1407, 660, 64, 44, 1031, 1367, 124, 181, 639, 933]. predictiveness [281]. predictor [239, 262, 1491]. predictors [1626, 464, 33, 527, 780, 1444]. preference [1425, 243]. preferences [5, 115]. Prefiltered [1633]. pregnancies [83, 1279]. pregnancy [83, 1108]. premature [771]. preplanned [1605]. Prescott [47]. presence [1235, 1190, 454, 847, 360, 498, 16, 37, 201, 1003, 537, 1330, 879, 1167, 1648, 1191, 525, 137, 1108, 1206, 1544, 817, 1170, 456, 614, 1031, 936]. presentation [219]. Preserving [849]. pressure [656]. preterm [780]. prevalence [1198, 1346, 505, 212, 60, 1353]. prevalent [709]. prevention [1504, 276, 818, 441]. previous [83]. primary [642, 1057, 1332, 673, 886, 1237, 804]. Prime [560, 559, 558, 557, 561, 1438]. prime-boost [1438]. prime-time [559, 558, 557, 561]. primer [103]. Principal [644, 1409, 970, 444, 930, 976, 1498, 1648, 338, 1608, 861, 544, 366, 369, 959]. Principled [575, 319]. Prior [1311, 1097, 1578, 650, 531, 672, 623, 56, 1616, 1494, 63]. prioritized [1187, 360]. priors [129, 1474, 198, 1598, 1099]. privacy [849]. Probabilistic [375, 1310, 1395, 56]. probabilities [1225, 414, 64, 258, 150]. Probability [721, 584, 700, 450, 1436, 1197, 1553, 1572, 497, 1438, 262, 1391, 1639, 1327, 490, 99, 303, 695, 215, 1085, 1345, 1125, 1296, 216]. probability-of-treatment [1553]. probit [1336, 1489, 242]. problem [1200, 1293]. problems [1264]. procedure [1199, 133, 1622, 665, 249, 954, 927, 1082, 1322, 1362, 809]. procedures [1465, 70, 1358, 705, 758, 1411, 654]. process [139, 1096, 1451, 846, 594, 1311, 432, 447, 1401, 1396, 1491, 400, 901, 57, 1099]. processes [1174, 1523, 677, 401, 444, 966, 1659, 1302, 465, 951]. processing [1555]. producing [189]. product [793]. Professional [801]. Professor [326, 323, 325]. Profile [817, 1595, 805, 946, 1573]. Profile-likelihood [817]. profiles [1365, 366]. Profiling [101, 102, 919]. prognosis [648]. Prognostic [1572, 1318, 512, 647, 106, 219]. prognostication [846]. program [1610]. progress [1596]. Progression [1073, 1090, 536, 607, 83, 824, 1280, 523, 961, 1212, 1269, 1313, 64, 1089]. progression-free [824, 1089]. progressive [1073]. project [800]. promising [869]. Promotion [539]. prone [1428, 636, 1256, 493, 1645, 247]. proof [301, 1417]. proof-of-concept [301, 1417]. Propensity [119, 397, 963, 450, 1240, 623, 750, 1157, 78, 1559, 1617, 590, 376, 1286, 118, 667, 334, 700, 1008, 157, 1040, 408, 1553, 155, 143, 1525, 1321, 399, 649, 425, 422, 1613]. Propensity-based [1286]. propensity-score [118]. Propensity-score-based [623]. proper [1089]. Properties [1457, 235, 1290, 170, 417, 1469, 697]. prophylaxis [210, 818]. proportion [1284, 882]. Proportional [460, 782, 1266, 833, 372, 726, 1295, 950, 386, 123, 549, 1088, 1056]. proportionality [99]. proportions [1224, 832, 394, 713, 1500, 231]. proposal [1334, 697]. Prospective [643, 707]. prostate [974, 774]. protection [1277, 632]. protective [1504]. protein [1178, 1221, 278]. protocol [1532, 716]. provided [1253]. provider [102, 953]. provider-based [953]. providers [758, 919]. proximal [1247]. pseudo [1216, 1346, 202, 902, 1277, 1306, 244]. pseudo-likelihood [1346, 1277]. pseudo-observation [244]. pseudo-observations [202, 1306]. pseudo-score [1277]. pseudo-value-based [1216]. psoriatic [527]. public [1560, 975]. publication [954, 1142]. publicly [1554]. published [587, 854, 106]. Publisher [58]. pulmonary [297]. pure [464]. purpose [1532]. puzzle [832]. Python [770].

Q [1439]. Q-learning [1439]. Qian [831]. QT [440]. QT/QTc [440]. QTc [440]. Qualifying [465]. qualitative [16, 1483]. quality [1537, 545, 1657, 123]. QuantDiffForecast [1508]. quantification [655]. quantified [1508]. quantify [714, 1150]. Quantifying [9, 167, 711, 721, 705, 1315, 100, 350, 581]. Quantile [1244, 1044, 14, 430, 297, 1308, 1518, 1409, 1347, 1188, 280]. Quantiles [911]. quantitative [843, 400]. quasi [636, 276, 1348]. quasi-experiment [276]. quasi-likelihood [636, 1348]. quaternion [1121]. questions [319, 405, 940]. Quicker [70].

R [908, 175, 853, 286, 1000, 770]. RABR [632]. radiomics [308]. Rakai [583]. Raking [365, 1428, 737]. Random [1567, 190, 857, 1372, 102, 698, 1557, 1470, 1658, 948, 876, 788, 1620, 403, 30, 1197, 839, 248, 1222, 341, 1448, 610, 542, 566, 794, 311, 883, 1457, 896, 1230, 1643, 1184, 1446, 206, 1027, 1253, 775, 792, 1566, 1598, 1192, 1298]. random-effect [30, 1457, 206]. random-effects [1372, 341, 1448, 1230, 775, 1598, 1192]. randomisation [1521]. randomised [1579, 5, 115, 1287]. Randomization [600, 456, 195, 46, 152, 1312, 1220, 637, 1641, 681, 276, 71, 863, 1588, 240, 665, 1084, 659, 1288, 817, 1447, 899, 657, 1473, 1043, 632, 1469, 864, 522]. Randomization-based [600, 456, 195]. Randomized [1162, 391, 1276, 1283, 1032, 1591, 1475, 1534, 1309, 890, 981, 678, 1383, 1247, 25, 661, 141, 1209, 985, 1219, 105, 384, 3, 690, 191, 900, 816, 292, 1560, 731, 1197, 347, 1630, 905, 528, 689, 180, 1333, 132, 68, 109, 1003, 1577, 787, 73, 205, 924, 1410, 42, 265, 414, 501, 1252, 1455, 227, 665, 753, 1371, 309, 593, 459, 1242, 729, 577, 1419, 600, 1485, 1030, 349, 1300, 895, 243, 940, 1248, 691, 835, 1281, 1054, 1072, 1632, 575, 195, 46, 1427, 1495, 1408, 1420, 393, 1301, 722, 1617, 1340, 279, 1568, 376, 1613, 275, 555, 685, 864, 880, 742]. range [306, 760]. ranging [1607]. Rank [1403, 1014, 1331, 1234, 1017, 1086, 1207, 955, 463, 1254, 424, 1544, 718, 491, 1291]. Rank-based [1014, 1254]. rank-type [955]. Ranked [836, 1611, 1242]. Ranking [1184, 355]. ranks [424, 522]. rapid [64, 1647, 708]. rare [1418, 1050, 30, 210, 1515, 67, 1446, 206, 436, 651, 668, 1192]. rate [1193, 887, 186, 1045, 734, 1510, 854, 222, 6, 1423, 156, 221, 715, 103, 1232, 1446, 1411, 56, 893, 61, 792, 539, 599]. rater [568]. raters [457, 1191, 32]. rates [935, 505, 874, 564, 1001, 753, 1176, 1401, 214, 988, 193, 1377, 616]. ratings [568, 688]. Ratio [753, 1109, 1150, 530, 1534, 1062, 1118, 1480, 182, 367, 908, 1293, 863, 133, 1384, 902, 1446, 399, 1274, 757, 56, 1085, 1113, 1559, 1642, 460, 335, 808, 909, 1141, 1271]. ratio-based [1085]. ratios [874, 825]. raw [199, 785]. RCRdiff [785]. RCTs [1432, 697]. Re [830, 1083, 501, 699]. re-estimation [1083, 501, 699]. Ready [560, 559, 558, 557, 561]. Real [962, 1519, 1055, 205, 623, 1556, 725, 1089]. real-time [1556]. real-world [1055, 205, 623, 725, 1089]. realistic [878]. reasoning [246]. Reassessment [982, 305, 72, 1092]. rebound [145]. recalculation [146]. recalculations [803]. recalibration [1368, 512]. recapture [1256]. receiver [428, 228, 281, 1145, 312, 1052, 679, 110, 929, 168, 743, 258, 389, 1142, 1635]. recency [838]. RECeUS [1109]. reciprocal [626]. Reclaiming [925]. recommendation [986]. recommendations [148, 1507, 837]. record [1071, 157, 1364, 494]. record-derived [157]. records [173, 139, 66, 210, 163, 439, 952, 1141, 107]. records-based [1141]. Recovery [90, 1460]. recruitment [173, 964, 1432, 1323, 616]. recurrence [1023]. recurrent [1571, 1504, 101, 1536, 1550, 444, 1064, 1285, 497, 571, 1001, 260, 1315, 1406, 158, 761, 988, 1468, 11, 1018, 1271, 555]. recurrent-event-free [11]. REDOMA [1598]. reduce [1651, 1517].

Reduced [491]. reduction [76, 1071, 714, 1515]. reestimation [140]. reference [1041, 1512, 843, 687, 1493, 604]. reference-based [843]. referenced [970]. Refined [1123]. Reflecting [327]. Reflections [324]. regarding [594, 1012]. regime [680]. regimen [37]. regimens [641, 465]. regimes [852, 1439, 272, 235, 1158, 1438, 1053, 294, 247, 701, 1034, 93]. region [970, 1454]. region-referenced [970]. regional [1308]. regions [687, 1493, 1621]. register [1203]. registers [995]. registration [896]. registry [1256, 133, 205, 1389, 190]. registry-based [1256]. Regression [401, 202, 1084, 262, 928, 344, 549, 11, 1146, 1216, 747, 562, 102, 19, 782, 1585, 1595, 805, 1338, 1451, 318, 906, 981, 840, 759, 1139, 802, 1044, 915, 1050, 1336, 14, 1223, 1128, 1177, 734, 299, 755, 131, 1207, 1222, 33, 789, 1069, 812, 943, 430, 873, 1245, 59, 1381, 203, 1458, 108, 356, 1547, 542, 414, 710, 829, 1422, 780, 929, 1268, 1520, 233, 143, 392, 662, 1176, 916, 485, 183, 365, 1267, 113, 891, 135, 1413, 1407, 1117, 663, 1022, 1496, 630, 612, 702, 1656, 1169, 1518, 492, 1196, 987, 1027, 1194, 1111, 944, 1424, 813, 1464, 224, 381, 491, 1056, 1236, 1347, 1468]. regression [1642, 375, 1188, 335, 745, 390, 1296, 1067, 756, 140, 75, 1066, 1070, 989, 1291, 717, 514, 596, 1629]. regression-based [1117]. regression-causal [802]. regressions [284, 44]. regular [1557]. regularization [524, 786, 1042]. regulations [563]. regulatory [98]. reinforcement [986, 821, 701]. Rejoinder [484, 561, 621, 1070]. relapses [405]. relapsing [226]. relapsing-remitting [226]. related [217, 856, 502, 1011, 716, 1460, 1060, 599]. relatedness [1258]. relation [1430]. relationship [1370, 143, 254, 1631, 405]. relationships [85, 226]. Relative [186, 1264, 1527, 1343, 133, 1167, 682, 695, 1113]. relatively [1622]. reliability [1657]. reliable [10, 580, 579]. Rematching [1312]. Remiro [1076]. Remiro-Azocar [1076]. remission [378]. remitting [226]. removed [887]. Renewable [1614]. reparameterization [990]. repeat [577]. repeated [83, 151, 270, 437, 94, 1000, 1073, 1377, 200, 625, 32, 1427]. repeatedly [300]. replacement [118]. Replicability [572]. replicated [443, 684]. replicates [589]. replication [933]. reply [580, 434]. reported [34, 1442, 377]. Reporting [78, 896, 1590, 960, 1182]. reports [859]. representation [1337, 1362]. reproducibility [1042, 410, 865]. reproducible [770]. repulsive [149]. required [895]. requirements [900, 1371, 279, 933]. resampling [350]. rescaling [1228]. rescue [459]. research [1585, 66, 1055, 1369, 53, 218]. residential [253]. residual [1235, 733, 1112, 1502]. resistance [1544]. resistant [1426]. Resolution [665]. resolve [1200]. resonance [1224, 127, 673, 774, 736]. resource [635]. resources [66]. respondents [1646]. responders [628, 257]. responding [483]. Response [816, 207, 496, 18, 761, 1570, 1193, 184, 102, 35, 692, 1584, 915, 292, 1197, 769, 1645, 879, 929, 65, 821, 1387, 783, 611, 1473, 1007, 1043, 632, 413, 1355, 790]. Response-adaptive [761, 769, 1387, 1043]. response-dependent [1007, 1355]. responses [836, 938, 188, 1239, 968, 1305, 1535, 1259, 1603, 362, 1111, 171, 1541, 1043, 245, 1355, 516]. responsible [1587]. responsivity [859]. restart [1535]. resting [127, 736]. Restricted [680, 255, 989, 782, 1388, 1499, 7, 271, 454, 1219, 180, 1161, 1263, 31, 928, 848, 781, 1114, 811]. restriction [1388, 1469]. results [411, 1578, 529, 591, 502, 576, 1007, 1528]. retention [1496]. retesting [60]. Retirement [489]. retrospective [1551]. reveal [186]. reveals [1137]. reversible [126]. Review [1095, 1620, 309, 296, 19, 1116, 963, 183, 1395, 1471, 1104, 160, 1124]. reviews [1490]. Revisiting [1456]. rhythmicity [1272]. rhythms [1275]. Ridge [653]. Ridge-penalized [653]. right [1090, 228, 1591, 1235, 1421, 873, 20, 262, 662, 345, 1406, 928, 1108, 663, 960, 1296]. right-censored [1090, 1591, 1235, 873, 262, 662, 345, 1406, 928, 663, 960]. right-censoring [1296]. Risk [788, 833, 19, 667, 334, 584, 263, 1368, 1017, 1086, 223, 823, 1148, 1614, 70, 210, 755, 1343, 163, 1356, 980, 1167, 434, 358, 348, 183, 1270, 1561, 1026, 682, 433, 1446, 298, 1460, 38, 366, 1303, 1194, 1113, 166, 244, 951]. risk-adjusted [70]. risk-benefit [348]. risks [974, 268, 569, 906, 759, 454, 438, 1059, 229, 980, 209, 1167, 1201, 719, 1053, 715, 723, 1143, 1638, 1600, 1170, 45, 40, 1036, 1227, 1599, 1329]. RMST [1509]. RMST-based [1509]. RNA [76, 1467, 956, 573]. RNA-seq [573, 956]. Robin [47]. Robust [291, 1571, 1065, 724, 740, 403, 538, 679, 1254, 494, 602, 739, 851, 1430, 1615, 346, 50, 1152, 941, 860, 1589, 1506, 382, 256, 1225, 428, 95, 791, 1051, 948, 1436, 681, 1061, 737, 1246, 462, 964, 133, 950, 1084, 1172, 705, 743, 1485, 1143, 399, 158, 695, 993, 1443, 1107, 1134, 918, 280, 730, 1034]. Robustness [871, 620, 878, 618, 621, 81, 606, 1007]. robustness-specificity [606]. ROC [1293, 605, 609, 1454, 178]. ROC-based [609]. Role [476, 479, 474, 182, 480, 478, 483, 13, 482]. room [217]. routine [309, 593]. routinely [953, 179]. routinely-collected [953]. Rule [625, 1282, 1061, 72]. rules [1519, 849, 545, 100, 1325, 611, 952, 1479, 1262]. run [854, 545, 6]. run-off [854, 6].

S [476, 831]. Sabour [434]. Saddlepoint [1245, 189]. SAEM [1365]. safe [1036]. Safety [38, 1482, 1429, 856, 304, 1556, 72, 1487, 1132, 629, 112]. Saharan [799]. Sample [1247, 1526, 1574, 985, 939, 908, 1197, 905, 163, 1402, 132, 109, 624, 1093, 240, 501, 227, 1371, 886, 1025, 988, 1289, 784, 1281, 699, 1566, 657, 279, 808, 909, 1380, 880, 336, 852, 700, 1405, 1452, 906, 819, 25, 1219, 738, 191, 900, 1243, 347, 1358, 1246, 146, 180, 1293, 1083, 59, 1646, 803, 1254, 1500, 392, 1404, 17, 586, 377, 828, 306, 659, 211, 817, 294, 441, 231, 633, 162, 1481, 899, 1160, 1632, 1257, 923, 1568, 140, 1102, 216, 933]. Sample-weighted [163]. samples [1611, 652, 1628, 49, 39]. Sampling [354, 836, 139, 1133, 1189, 318, 840, 946, 874, 1172, 716, 1242, 574, 1149, 757, 435, 15, 389, 1355, 1353, 1298]. Sampling-based [354]. sarcoma [1087]. sarcopenia [225]. SARS [1225, 1655, 1467, 1514]. SARS-CoV-2 [1225, 1467, 1514]. SARS-CoV2 [1655]. SAS [1569]. satisfied [94]. Scalable [807, 978]. Scalar [1656, 1629]. Scalar-on-function [1656, 1629]. Scale [429, 1097, 1516, 1177, 863, 1423, 859, 1412, 683, 1441, 1460, 554, 1194, 919, 1010, 1345]. Scale-Extended [1345]. Scaled [1486, 1494]. scales [523, 980, 613, 452]. scan [80, 1633, 850]. schedule [1092]. schedules [847, 261, 877]. schemes [1355]. Schoenfeld [945]. Schofield [620]. School [799]. science [425]. sclerosis [1121, 510, 405, 226]. scoping [1124]. Score [554, 231, 118, 667, 334, 700, 1008, 157, 397, 186, 1040, 963, 450, 823, 1044, 408, 146, 646, 493, 1239, 1245, 1553, 1572, 1240, 155, 623, 264, 1277, 143, 113, 750, 1321, 1157, 399, 1316, 78, 1559, 1617, 1599, 1031, 590, 422, 376, 1613, 837]. score-adjusted [78]. score-based [1572, 422]. score-matching [667]. scores [119, 223, 470, 1525, 638, 106, 1322, 425, 166, 181]. scoring [488]. screen [57, 742]. screen-identified [57]. screen-positive [742]. screened [553]. screening [411, 1202, 739, 967, 1175, 446, 1212, 238, 1303, 609, 888, 811, 1036, 936, 1566, 730, 1186, 960, 57, 742]. SD [115]. Seamless [607, 162, 1592, 1488, 992]. seamless-penalty [1488]. search [998, 671, 1359, 1399]. seasonal [87, 1352]. seasons [707]. Secondary [197, 49, 1355]. sectional [1578, 838, 731, 221, 383, 39]. segmentation [1461]. segmented [234]. SEIR} [887, 887]. SEIR}-fansy [887]. select [38, 573]. selected [1094]. Selecting [473]. Selection [545, 350, 1345, 343, 506, 1071, 1116, 1403, 151, 938, 1658, 289, 985, 1014, 237, 738, 740, 1004, 1260, 637, 1641, 518, 978, 1400, 943, 826, 1105, 174, 627, 911, 259, 385, 602, 926, 380, 1488, 364, 950, 1357, 638, 1435, 1042, 1593, 1575, 301, 723, 55, 1108, 1359, 243, 693, 858, 1411, 489, 456, 575, 196, 1616, 213, 1180, 23, 63, 1517, 216, 379, 1399, 1543]. Selective [173, 716, 1420]. Self [952, 960, 703, 856, 1515, 859, 1442, 243, 340, 1262]. self-controlled [703, 856, 1515]. self-initiated [859]. self-learning [1262]. Self-matched [952]. self-prediction [340]. self-reported [1442]. Self-reporting [960]. self-selection [243]. Semi [1121, 1236, 524, 466, 1292, 1365, 1401, 696, 907, 1638, 889, 1460, 1445, 1629]. semi-competing [1638]. semi-continuous [524, 1292, 696, 889]. semi-Markov [466, 1401]. semi-mechanistic [1445]. semi-parametric [1365, 907, 889, 1460, 1629]. Semi-supervised [1121, 1236]. semicompeting [209, 1329]. semicontinuous [12, 128]. Semiparametric [1388, 268, 569, 1028, 1366, 912, 1336, 1223, 250, 201, 1398, 571, 1277, 461, 55, 1393, 745, 555, 514, 596, 823, 1059, 991, 71, 114, 531, 163, 789, 943, 1064, 551, 1381, 468, 748, 1268, 1520, 862, 1397, 968, 1316, 807, 370, 675, 123, 1543]. senior [337]. sense [138]. sensitive [1347]. sensitivities [548, 742]. Sensitivity [194, 1209, 839, 229, 1490, 1608, 1326, 1208, 25, 548, 248, 1222, 610, 896, 2, 200, 1390, 1142]. sensor [1121, 1019]. Separable [1644]. seq [956, 573]. sequence [125]. sequences [1047]. sequencing [1621]. Sequential [1312, 843, 1218, 527, 1241, 890, 1526, 384, 347, 1630, 146, 252, 787, 463, 170, 794, 879, 1425, 467, 665, 1126, 913, 1304, 1132, 1535, 36, 629, 654, 1594, 1072, 116, 1180, 124, 1043]. sequentially [104]. Serial [362, 1029, 429]. series [1378, 1168, 1121, 703, 856, 1515, 443, 1624, 74, 925, 614]. seroconversion [193, 1377]. Set [756, 1611, 852, 836, 972, 1242, 399, 1464]. Set-regression [756]. set-valued [972]. sets [1538, 875, 1464, 540]. setting [737, 1493]. settings [1368, 1440, 398, 466, 1106, 895, 91]. several [1258]. severe [1345]. sham [533]. Shan [207]. Shape [1499]. shapes [400]. Shared [877, 1538, 209, 859, 93]. shared-parameter [859]. sharing [518, 1529]. shift [543]. shock [489]. short [182, 996, 1549, 1639, 576, 992, 708]. short- [992]. short-term [182, 996, 1549, 576]. shortening [711]. Shortest [1643, 605]. should [1557, 678, 1521, 258, 933]. Shrinkage [1343, 978, 129, 697, 1453]. shrinker [304]. sibling [1447]. sickness [1644]. sided [516]. sieve [842]. sign [394]. signal [1178, 129, 1482]. signals [1256, 883, 112]. signature [1052, 1042]. signatures [1188]. Significance [1314, 47, 8, 18, 363]. significant [1284]. SIM [795]. similar [1519]. Similarity [974]. Simple [1333, 398, 153, 1622, 121, 517]. simplified [395, 715]. Simulated [1171, 1082]. Simulating [1191, 526]. Simulation [396, 747, 652, 305, 1418, 450, 1465, 1579, 360, 1528, 1402, 259, 710, 317, 496, 1419, 495, 1124, 868, 1015]. simulation-extrapolation [710]. simulation-free [305]. simulations [127, 878, 1507]. simulations-Potentials [1507]. Simultaneous [1400, 523, 1488, 1227, 1658, 1279, 1114]. simultaneously [102, 584, 1086, 43, 399]. since [934]. Singapore [147]. Single [1041, 1307, 1450, 472, 22, 188, 733, 467, 956, 1259, 1603, 1124, 918, 573]. single-cell [956, 573]. single-index [188, 733, 918]. single-point [1124]. Single-stage [1041]. Single-world [1307]. Sir [324, 325, 326]. SIRV [1137]. site [849, 1404]. Six [266]. Six-way [266]. size [1216, 336, 852, 700, 1405, 819, 1247, 25, 661, 1526, 1574, 985, 939, 908, 900, 1197, 905, 146, 1083, 1402, 132, 59, 109, 803, 468, 624, 1093, 240, 501, 227, 1370, 1371, 392, 886, 1404, 17, 586, 377, 828, 306, 1025, 211, 895, 608, 441, 988, 1289, 162, 1481, 784, 1281, 1160, 1632, 699, 923, 1566, 657, 279, 1568, 808, 909, 140, 1380, 880, 933]. sizes [426, 231, 784, 835, 216]. skew [513, 429]. skew-normal [513, 429]. skew-normal/independent [513]. skewed [95, 1366, 1111]. skewed-longitudinal [1366]. slab [978, 1598]. Slamming [533]. sleep [656, 917]. slope [1557, 532]. slope-based [532]. slopes [1093]. Small [110, 652, 1219, 384, 191, 1243, 347, 180, 1628, 1122, 17, 682, 393]. small-sample [191, 1243]. SMART [1158]. SMARTs [852]. smoking [912]. smooth [189]. Smoothed [228, 1044, 1316, 1518, 277]. smoothing [1480, 1200, 732]. Snell [377]. social [1631]. socioeconomic [13]. soft [1486, 1087]. software [1394]. solution [1500]. solutions [121]. Some [1080, 633, 899, 1007, 121, 1328, 486, 62, 1486, 5, 115]. source [642, 1617]. sources [642, 1554, 1290]. South [1496, 800]. Space [420, 1484, 997, 505, 914, 1552, 1337]. Space-time [420]. span [1450]. Sparse [524, 1350, 1243, 1530, 59, 1279, 851, 458, 380, 1606, 1593, 1215, 75, 563]. sparsity [1239, 1264]. Spatial [423, 513, 720, 400, 80, 747, 1637, 1017, 1086, 186, 1311, 931, 1343, 1020, 774, 1357, 253, 510, 277, 455, 684, 925]. spatial-clustering [925]. Spatially [1103, 1484, 1020, 564, 1292, 719]. Spatio [387, 356, 1357, 1396]. Spatio-temporal [387, 356, 1396]. spatio-temporally [1357]. Spatiotemporal [129, 564]. speaking [797]. specific [937, 1548, 163, 679, 430, 1547, 1151, 1644, 168, 1176, 723, 418, 1654, 1414, 825]. specification [1174, 197]. specificity [548, 896, 606]. specify [1500]. specimens [1139]. spectra [1221]. spectral [443]. spherical [199]. SpiderLearner [1214]. Spike [978, 1598]. Spike-and-slab [978, 1598]. spine [462]. Spline [357, 1480, 830]. Spline-based [357, 830]. splines [782, 1200, 904, 1365]. Split [868]. splitting [151]. spontaneous [1590]. spreading [510]. squared [1010]. squares [759, 458, 1016, 1544]. stability [276, 350]. stability-controlled [276]. stable [364]. stacking [698, 746]. stage [1636, 1586, 506, 972, 1591, 1041, 1596, 595, 885, 1511, 1220, 1434, 1605, 293, 1263, 174, 170, 155, 1425, 1126, 1492, 520, 172, 1535, 848, 1251, 243, 1426, 992, 5, 115, 1180]. stages [1528, 996]. stagewise [272, 196]. Stan [704]. standard [619, 620, 270, 94, 1126, 618, 621, 625, 760, 1373, 388, 1314, 1566]. Standardization [1576, 38, 695, 778, 1286]. standardized [24, 874, 270, 352]. STAR [309, 593]. Start [17]. Stata [770]. State [425, 973, 127, 1418, 846, 1472, 111, 1280, 914, 1552, 1337, 1469, 736]. state-space [914]. state-transition [111]. states [897]. stationary [1624]. statistic [80, 863, 377, 971, 137, 945]. Statistical [477, 838, 900, 290, 1293, 1483, 1320, 170, 801, 1471, 1441, 148, 204, 1507, 992, 1420, 1541, 1583, 285, 599, 655, 66, 1116, 1189, 69, 800, 319, 1538, 387, 328, 246, 1631, 417, 525, 694, 987, 1253, 1345, 91, 331, 1621, 1469, 864, 1089, 373, 697]. Statistician [327, 321, 324, 325, 326, 329, 331]. Statisticians [476, 479, 474, 480, 478, 483, 482]. Statistics [854, 798, 1205, 799, 1258, 322, 1245, 82, 264, 896, 63, 47, 797, 559, 799, 115, 322]. status [1336, 468, 144, 710, 1165, 122, 811, 1502]. stay [449, 419]. stayer [536]. Steady [1469]. Steady-state [1469]. Step [701, 982, 850, 309, 593, 1362]. Step-adjusted [701]. step-up [982]. stepped [1127, 191, 816, 731, 426, 964, 68, 1003, 924, 42, 1455, 600, 841, 857, 1589, 275]. stepped-wedge [816, 731, 924]. stereotype [136]. still [1605]. STIR [1169]. Stochastic [20, 314, 1378, 1440, 1510, 110, 715, 1359]. straightforward [979]. strata [1648]. strategies [1041, 635, 1040, 438, 1472, 738, 28, 364, 613, 1149, 1107, 975]. strategy [764, 181]. stratification [157, 397, 408, 16, 976, 1498, 861, 544, 1409, 1287]. Stratified [1056, 1511, 498, 109, 1562, 309, 593, 574, 1149, 1521, 231, 277, 757]. STRATOS [153, 154]. streaming [1614]. streams [1154]. strength [86]. stress [1401]. strong [103]. structural [635, 1218, 1268, 1131, 1306]. structure [415, 42, 1229, 1267, 606, 208, 1188]. structure-aided [1188]. Structured [1169, 1580, 75, 1239, 1370, 728, 1196, 941, 1255]. structures [362]. Studentized [1219]. Student’s [968]. Student’s- [968]. studies [1650, 313, 339, 256, 853, 428, 1091, 173, 139, 1071, 177, 471, 1040, 963, 981, 7, 1209, 440, 1202, 1047, 1434, 738, 1579, 703, 286, 96, 1243, 1346, 834, 648, 1529, 229, 818, 1069, 955, 551, 1245, 1562, 502, 1122, 197, 20, 1384, 929, 1268, 65, 613, 709, 705, 230, 927, 106, 1615, 297, 1593, 1487, 682, 971, 572, 979, 1321, 547, 921, 1135, 1109, 211, 1058, 169, 500, 761, 370, 1323, 1124, 15, 383, 725, 804, 923, 245, 1018, 1066, 1070, 343, 1355, 514, 540, 933, 271, 1067]. Study [1133, 489, 1193, 1372, 747, 836, 1091, 1533, 652, 592, 1265, 1519, 1418, 157, 1354, 692, 912, 1465, 1223, 1579, 360, 894, 1346, 1528, 898, 1402, 1448, 117, 462, 624, 259, 470, 469, 1360, 821, 221, 1305, 1037, 1308, 995, 253, 671, 317, 496, 1419, 301, 495, 1230, 1131, 302, 1555, 340, 1373, 1195, 1015, 933, 1421]. study-based [302]. study-level [1373]. Study/Study [1133]. Studying [1108]. sub [780, 680, 799]. sub-compositional [780]. Sub-Saharan [799]. sub-tree [680]. subclasses [1379]. subdistribution [19, 584, 183]. Subgroup [937, 645, 371, 1351, 1297, 506, 998, 1622, 535, 1370, 1054, 1088, 918, 590, 745, 756, 274, 1542]. Subgroup-specific [937]. subgroups [223, 299, 504, 647, 851, 1370, 1382, 763, 670, 1347, 722]. subject [334, 1028, 677, 1139, 248, 1478, 773, 52, 1404, 1226, 429, 44, 245, 218, 736]. subjective [834]. subjects [1109]. subpopulation [174]. subsampling [1063]. subset [1442]. Substantive [1039]. substudies [29]. subtraction [819]. subtypes [224]. subtyping [1155]. success [1078, 216]. successes [477]. such [1557]. sufficiency [529]. sufficient [863, 170]. sum [70, 424, 409]. summaries [90]. Summarizing [1230, 1372]. Summary [648, 82, 705, 100, 817, 899, 63, 1142]. summary-data [899]. summary-level [82]. summed [656]. Sungim [854]. super [210, 1476, 107]. super-combo-drug [107]. superiority [1237]. superspreading [1514]. Supervised [1268, 1121, 519, 883, 959, 1236]. supplemental [642, 787]. supplementation [1262]. supported [190]. supremum [1205]. sure [566]. SuRF [380]. surface [605]. surgery [974, 744, 1551]. surgical [462, 409]. surrogacy [714, 254, 1581]. surrogate [1463, 1096, 1354, 1449, 1055, 59, 254, 85, 1037, 1101, 727, 711, 882]. Surv [1090]. Surv-CRM-12 [1090]. Surveillance [946, 87, 1429, 69, 304, 754, 1556, 994, 74, 683, 482, 676, 1132, 38, 629, 1274, 89, 877]. survey [391, 1133, 1189, 27, 424, 1525, 1656, 791]. surveys [1570, 1198, 1637, 287]. Survival [593, 314, 386, 565, 79, 508, 585, 1586, 1625, 1090, 187, 118, 1388, 1266, 972, 228, 553, 1368, 177, 1166, 35, 1366, 1511, 1203, 7, 271, 1179, 189, 1219, 1527, 724, 1579, 250, 824, 1217, 648, 180, 1161, 615, 1038, 229, 114, 1402, 1280, 445, 519, 767, 850, 1218, 133, 1081, 205, 1572, 414, 1112, 1549, 922, 1488, 1438, 709, 662, 168, 995, 1270, 664, 969, 1237, 135, 1212, 1110, 576, 928, 1407, 1108, 547, 1206, 1317, 552, 1170, 597, 959, 702, 294, 159, 302, 1627, 631, 1524, 695, 633, 781, 793, 1299, 1114, 1115, 984, 1150, 575, 944, 44, 549, 1506, 722]. survival [288, 166, 1392, 1063, 1367, 732, 1386, 274, 255, 1034, 989, 1089, 354, 273, 309]. survival-associated [1212]. survival-incorporated [1392]. survivor [1408]. survivors [1109, 1392]. Susan [476]. susceptibility [363]. susceptible [887]. susceptible-exposed-infected-removed [887]. swapped [993]. sweat [447]. switch [4, 175]. switching [201]. Symposium [797]. symptom [887, 1634, 777]. symptom-based [887]. synthesis [948, 1640]. Synthesizing [272, 1047, 630, 114]. System [78, 1590, 1182]. systematic [1116, 1490, 1132]. systems [896].

table [1577]. tables [47, 1278, 8, 18, 971, 340, 1182]. Tackling [1285]. tail [1225]. tailed [47, 542, 8, 18]. tailored [647]. Tailoring [1256, 1630, 247]. tails [789, 968]. taking [1344]. Tale [476, 474, 480, 478, 1458]. tales [899]. Tao [207]. Target [496, 1074, 1076, 1075, 391, 141, 1209, 1080, 1248, 1286]. Targeted [1487, 1015, 878, 1426]. targeting [364, 38]. TBI [1345]. techniques [755, 1270, 295]. technology [1019, 1080, 1077]. temperature [1451]. template [1195]. temporal [1368, 367, 356, 754, 1396, 387, 277]. temporally [1357]. term [182, 996, 1161, 1549, 1610, 576, 1109, 992]. terminal [1536, 1064, 571, 1315, 1638, 608, 761, 825]. terms [948]. ternary [1440]. test [1578, 1041, 1062, 440, 1258, 764, 105, 367, 838, 1119, 299, 834, 905, 1246, 870, 955, 1245, 9, 581, 197, 470, 794, 424, 1384, 421, 282, 831, 1101, 446, 653, 913, 1376, 1653, 81, 629, 169, 1341, 277, 609, 701, 625, 123, 811, 1642, 616, 1024, 1031, 936, 1386, 1102, 517, 945, 1142, 107, 373]. test-and-treat [701]. test-negative [105]. test-wise [1024]. Testing [892, 560, 108, 780, 1413, 659, 1298, 559, 558, 509, 1452, 583, 223, 289, 1139, 620, 165, 511, 487, 1045, 557, 561, 1528, 920, 850, 1423, 1398, 249, 1371, 392, 309, 593, 485, 762, 927, 1304, 618, 621, 1132, 1022, 243, 1322, 112, 1257, 1301, 722, 1227, 804, 901, 1192, 809, 1353]. tests [47, 1276, 536, 887, 1205, 1546, 832, 3, 394, 403, 1537, 180, 568, 463, 588, 947, 771, 1500, 156, 264, 65, 1509, 969, 8, 18, 1419, 554, 992, 877, 46, 811, 1373, 375, 936, 1314, 1566, 1433, 335, 516, 742]. their [1440, 934, 1273, 877, 5, 115, 1433]. Theoretic [314, 1463]. theoretical [1548]. Theory [883, 944, 102, 153, 535]. therapeutic [284]. therapies [1091, 768]. therapy [1496, 1107, 1150]. there [1368, 446, 1602]. thin [942]. thorough [440]. those [205]. thoughtfully [313]. Thoughts [476, 330]. Three [1428, 625, 1041, 1563, 227, 1126, 1379, 78, 393, 1006]. three-arm [1041, 1126, 393]. three-component [78]. three-dimensional [1563]. three-level [227, 1006]. Three-phase [1428]. threshold [673, 9, 581, 100, 1088, 1347, 918]. threshold-free [100]. thresholds [1082]. thrombosis [268, 569]. throughput [1094, 865]. tick [1305]. tick-borne [1305]. ties [530]. Time [560, 1555, 23, 1650, 1234, 339, 1216, 361, 1225, 712, 19, 118, 782, 1378, 1388, 411, 973, 228, 1405, 559, 558, 1032, 636, 1368, 656, 1168, 1017, 1086, 1503, 1204, 890, 830, 454, 985, 1318, 1614, 1121, 900, 420, 404, 1361, 261, 292, 557, 561, 180, 1161, 870, 881, 505, 430, 130, 1478, 59, 914, 1003, 1105, 174, 1659, 1081, 1093, 980, 1410, 240, 443, 602, 1520, 1556, 1397, 469, 392, 1634, 168, 183, 260, 1609, 743, 365, 766, 1624, 1021, 357, 664, 969, 1237, 74, 1404, 337, 577, 287, 844, 828, 50, 928, 77, 1184, 1108, 547, 298, 552, 1369, 959]. time [1607, 1274, 1316, 1169, 1456, 858, 807, 1627, 631, 1393, 1114, 160, 1594, 1337, 1513, 1072, 1323, 1249, 278, 857, 962, 1420, 1580, 1144, 452, 614, 288, 1599, 1006, 1231, 11, 539, 1125, 745, 1063, 1342, 91, 255, 639, 989, 1602, 555, 1355, 596]. time-course [278]. time-dependent [1650, 228, 1318, 130, 1478, 1093, 168, 743, 959, 1169, 1456, 631, 160, 1323, 1580, 1231, 1125, 989]. time-lag [1602]. time-lagged [1072]. time-to-event [1234, 339, 361, 712, 118, 1405, 1032, 1503, 985, 1614, 404, 261, 292, 870, 59, 1659, 1093, 240, 469, 392, 365, 766, 1404, 577, 287, 844, 828, 77, 1184, 1607, 858, 1393, 1594, 1420]. time-to-first-event [555]. time-to-pregnancy [1108]. Time-varying [1555, 23, 19, 782, 881, 914, 1003, 1081, 1410, 1397, 1634, 183, 1021, 337, 50, 298, 552, 1369, 1144, 1006, 639]. timeline [267]. times [1118, 1504, 7, 271, 1472, 1219, 943, 144, 337, 1226, 781]. timing [744, 1128, 241, 41]. tissue [1087]. TITE [867]. TITE-BOIN12 [867]. titer [550]. TMS [1311]. tobacco [98]. Tobit [1347]. tolerance [142, 251, 286, 1493, 150]. tolerate [286, 528]. tolerated [1469]. Tomography [612]. Tony [1130]. toolbox [1508]. top [488]. topics [154]. total [584]. toxicants [1108]. toxicities [1091]. toxicity [1090, 1582, 1068, 641, 551, 144, 417, 448, 490, 867]. Tracking [914]. tractable [585, 176]. trade [348, 1082, 606]. trade-off [348, 1082, 606]. Training [799, 798, 236]. trait [1441]. traits [965]. trajectories [656, 1344, 1462]. Trajectory [534, 1504, 996]. transcriptome [1360]. transcriptome-wide [1360]. transcriptomic [1014]. transcriptomics [1020]. Transfer [991]. transform [736]. transformation [1059, 71, 1064, 199, 468, 1520, 862, 461, 1238, 1393, 51, 827, 893, 1573]. transformed [1097]. transition [1431, 111, 1023, 1389, 1401, 1601]. Translating [940]. transmission [423, 1655, 707, 704, 1510, 914, 1514]. transplant [1572]. transplantation [222, 995]. Transportability [1640, 1013]. transportation [524]. Transporting [591]. treat [1532, 701, 944]. treated [334]. treatment [1328, 852, 700, 1519, 583, 1475, 515, 1138, 1596, 635, 1439, 272, 450, 677, 849, 440, 1047, 594, 1527, 235, 658, 1164, 724, 1436, 567, 398, 1158, 120, 894, 1061, 16, 689, 1652, 615, 1529, 37, 229, 201, 341, 462, 1003, 1105, 1659, 1553, 1646, 911, 486, 494, 794, 1455, 1649, 879, 986, 364, 1371, 1474, 705, 449, 459, 1053, 85, 167, 1101, 1487, 913, 600, 1073, 495, 844, 148, 172, 1030, 349, 1300, 1184, 100, 1604, 1317, 298, 1432, 1535, 296, 418, 1251, 1325, 243, 1450, 294, 680, 247, 761, 1107, 628, 701, 1194, 835, 1281, 857, 5, 115]. treatment [1282, 145, 1347, 1427, 425, 61, 722, 116, 792, 1180, 280, 952, 1125, 279, 1558, 1136, 1034, 1602, 93, 1542, 882, 1262, 522, 1339, 273, 1453]. treatment-covariate [1030]. treatment-dependent [341]. treatment-sensitive [1347]. treatments [157, 397, 355, 595, 472, 131, 1038, 911, 1487, 1369, 718, 758, 160, 116, 422, 1043, 1296, 1034]. tree [1440, 998, 548, 1370, 1376, 1135, 88, 680, 701, 274]. tree-based [1135, 88, 701]. tree-structured [1370]. trees [1319, 788, 492, 1111, 901, 717]. trend [374, 1384]. trends [1368, 367, 1000, 260, 79]. Trial [995, 391, 819, 223, 141, 1209, 1584, 384, 1630, 528, 98, 1263, 527, 787, 1588, 463, 1162, 1410, 794, 1252, 467, 1092, 351, 1404, 601, 1104, 1233, 77, 1604, 1025, 1535, 848, 1251, 243, 1450, 1374, 1248, 793, 162, 1632, 5, 115, 711, 1647, 1569, 1494, 61, 1539, 116, 792, 1180, 1367, 53, 275, 596, 867, 1262, 1089, 1162]. trial-An [1647]. trials [451, 712, 1283, 1032, 714, 1205, 1591, 1041, 1475, 982, 1174, 368, 1596, 1534, 1503, 152, 1309, 35, 272, 885, 890, 1118, 1504, 1057, 472, 678, 1383, 1247, 25, 661, 194, 937, 1334, 342, 985, 1047, 594, 1219, 1532, 1565, 105, 843, 1582, 645, 690, 191, 487, 641, 84, 900, 1147, 816, 404, 299, 261, 894, 824, 498, 1199, 1560, 731, 1197, 347, 905, 1358, 839, 689, 180, 252, 531, 293, 232, 1002, 1416, 1333, 132, 68, 109, 1003, 1218, 1105, 521, 1659, 1577, 73, 803, 582, 170, 205, 34, 924, 269, 42, 265, 240, 501, 878, 1437, 1455, 1210, 976, 1498, 1649]. trials [227, 1005, 1500, 427, 1370, 753, 1277, 469, 1371, 1474, 593, 1126, 459, 348, 351, 260, 1, 1307, 1237, 729, 17, 577, 1419, 913, 600, 1485, 815, 1030, 349, 1232, 576, 1228, 296, 1078, 895, 490, 1607, 431, 629, 441, 492, 940, 691, 841, 763, 628, 112, 784, 675, 835, 1281, 1054, 1513, 1072, 1160, 575, 1249, 857, 1589, 532, 123, 195, 46, 611, 699, 962, 1345, 1427, 1495, 1408, 393, 1301, 722, 1617, 1227, 1250, 503, 616, 1340, 751, 257, 1159, 279, 1568, 1445, 1287, 1043, 376, 1613, 1517, 1380, 1417, 1469, 555, 685, 717, 864, 790, 880, 522, 1453, 742, 309]. triangle [854, 6]. tribute [1130]. trilemmas [958]. trim [954]. trimmed [839]. trimming [450]. Tropical [799]. truck [1018]. True [258, 1463, 1224, 737, 254]. truncated [1421, 662, 345, 337, 928, 565]. truncation [1108, 45, 1298]. Tuberculosis [423, 276, 949, 297, 1317, 1426]. Tumor [308, 1234, 1586, 790]. tumour [1567]. tuning [1465]. Turner [115]. tutorial [853, 1508, 286, 689, 1553, 1462, 298, 770, 1569]. Tweedie [956]. Twenty [934]. twin [256, 898]. Two [47, 1591, 1511, 832, 800, 738, 476, 300, 850, 934, 1165, 8, 18, 243, 1426, 435, 1102, 128, 1636, 1586, 506, 428, 595, 318, 885, 819, 1220, 1219, 474, 548, 299, 480, 834, 1246, 1293, 478, 293, 174, 1577, 1254, 1213, 1437, 1549, 155, 1425, 1500, 421, 1201, 1046, 1126, 1492, 831, 1270, 520, 659, 1535, 817, 992, 763, 633, 370, 1501, 51, 899, 1362, 340, 5, 115, 893, 1340, 389, 936, 1566, 1007, 436, 178, 1602, 1469, 343, 945, 742]. two-arm [1340, 1469]. two-by-two [1577]. two-level [1201]. Two-part [300, 128]. Two-phase [738, 1165, 318, 370, 435, 389, 1007, 343]. Two-sample [1102, 1246, 1293, 1254, 659, 817, 899]. Two-stage [1591, 1511, 243, 1426, 1636, 1586, 506, 595, 885, 1220, 293, 174, 155, 1425, 1126, 1492, 520, 1535, 992, 5, 115]. Two-step [850, 1362]. Two-tailed [47, 8, 18]. Two-wave [435]. two-way [1046, 340, 893]. type [1193, 1095, 1574, 915, 403, 955, 463, 1500, 439, 1001, 810, 1486, 716, 419, 871, 1036, 57, 632]. types [1193, 584, 232, 1553, 748, 402, 1382, 436]. typhoid [683].

U [1499]. U-shape [1499]. U.S. [1352]. Uganda [583]. ultrahigh [1051, 1212, 1036]. ultrahigh-dimensional [1212]. ultrasound [1461]. umbrella [506, 1440]. unanchored [1009]. unbalanced [718]. unblinded [803]. uncertainties [1344]. Uncertainty [972, 155, 655, 471, 1508, 948, 721, 1481, 1541]. unconditional [1577, 813]. Uncovering [1275]. uncured [1109]. underdose [1068]. underlying [951]. underpowered [1500]. Understanding [1588, 301, 627]. undirected [1634]. unequal [661, 1200, 426, 784, 835]. unexposed [924]. unidirectional [4, 175]. Unified [1027, 1329, 256, 1565, 1199, 114, 67, 820, 1500, 1520, 392, 921]. uniformity [997]. Unifying [266]. unimodal [1499]. Unit [672, 218, 772]. unknown [1231, 363]. unmeasured [229, 971, 81, 993, 1120]. unscreened [553]. unstable [1041]. unstratified [1588]. Unsupervised [1395]. untestable [120]. updating [512]. upper [1225, 1231]. uptake [742]. urinary [957]. Use [714, 251, 839, 1037, 1513, 751, 1097, 339, 19, 953, 1062, 1562, 205, 183, 473, 409, 1583]. used [1183, 1163, 1168, 678, 62, 258]. Using [313, 782, 1368, 997, 594, 1641, 210, 1458, 976, 1101, 552, 660, 1642, 425, 1636, 391, 1570, 1234, 619, 256, 853, 118, 700, 1008, 415, 173, 139, 553, 119, 1086, 953, 1133, 635, 125, 1451, 890, 223, 1504, 472, 289, 1508, 677, 194, 937, 1209, 849, 1202, 1554, 1461, 1148, 1121, 1050, 946, 191, 360, 1147, 1019, 518, 290, 892, 996, 1161, 615, 71, 163, 341, 129, 519, 1023, 774, 930, 202, 462, 1416, 1140, 1218, 315, 854, 534, 957, 468, 1093, 6, 385, 82, 542, 494, 851, 926, 1351, 1210, 227, 879, 439, 954, 956, 1365, 262, 821, 886, 309, 593]. using [1176, 638, 314, 221, 348, 1000, 1435, 1315, 1242, 85, 1581, 387, 861, 1471, 1404, 1313, 1413, 676, 1185, 1476, 1327, 921, 1132, 959, 1607, 399, 702, 465, 770, 680, 1590, 992, 1107, 1306, 701, 888, 877, 1337, 400, 340, 1589, 649, 711, 778, 1427, 1569, 1340, 375, 1188, 785, 307, 814, 63, 1445, 1573, 1598, 1342, 242, 75, 1380, 1099, 1629, 1272, 951, 1498]. utilities [647]. Utility [926, 1101, 211, 490]. utility-based [490]. Utilizing [1496, 347, 787]. uTPI [490].

vaccinated [932]. vaccination [28, 1438, 455, 1495, 1137]. Vaccine [1560, 588, 1463, 690, 932, 856, 1207, 1556, 1277, 482, 629, 842, 1182]. vaccine-adverse [1207]. vaccines [1085]. vague [604]. validate [727]. validated [1442]. validating [239, 1384, 1026]. Validation [1646, 336, 512, 1458, 1053, 586, 828, 1341, 1286]. validity [191]. valuation [211]. value [1216, 124, 274, 809]. valued [972, 74, 1487]. values [1531, 239, 470, 1492, 1301, 1031]. variability [1584, 1192]. Variable [669, 943, 638, 1108, 213, 1141, 1636, 102, 1183, 598, 1116, 151, 289, 438, 1014, 237, 740, 1004, 1332, 637, 1020, 627, 259, 385, 602, 739, 380, 1488, 950, 1357, 886, 1486, 1412, 723, 55, 1359, 1288, 1411, 375, 63, 1399]. variables [1193, 1095, 726, 519, 1364, 153, 203, 1011, 1645, 748, 696, 402, 154, 1654, 944, 1326, 350, 1599, 917, 1146, 1287, 1629]. Variance [118, 1546, 137, 275, 1008, 1585, 1557, 906, 464, 617, 1553, 1430, 1651, 94, 773, 52, 1267, 306, 294, 760, 283]. Variance-components [1546]. variances [542, 1281, 751]. variant [67]. variants [932, 363]. variate [1291]. variation [1484, 1110, 333]. Variational [543, 1115]. various [352]. vary [85]. Varying [1308, 19, 782, 1484, 881, 914, 1003, 706, 1081, 356, 1410, 1549, 1397, 1634, 183, 1021, 337, 50, 298, 552, 1369, 1555, 400, 965, 1144, 1297, 1006, 23, 639]. varying-coefficient [706]. Varying-coefficients [1308]. vascular [268, 569].

Vector [398, 644]. Vector-based [398]. verification [853, 312, 811, 258, 1031, 936]. Versatile [969, 927]. versus [1183, 1579, 1577]. via [972, 998, 151, 186, 126, 1574, 1480, 567, 731, 1197, 646, 873, 199, 523, 414, 878, 1488, 1046, 230, 72, 1308, 728, 1359, 38, 56, 196, 1349, 730, 274, 181, 563, 1271, 1399, 1479, 717, 150]. videos [1461]. vigorous [307]. violation [1608]. violations [1099]. viral [145]. virology [342]. virus [210, 221, 603, 89]. visit [1189, 677, 847, 1332]. visits [217, 1501]. visualization [1136]. Visualizing [872, 340]. vital [896]. vol [854]. volatile [75]. volunteer [649]. volunteer-based [649]. voxel [774, 1619]. voxel-wise [774, 1619]. vs [1008, 220, 472, 205, 103, 298, 1450, 1274, 455, 1392, 555].

waiting [1337]. Wald [403]. Wald-type [403]. walk [896]. walks [876, 1643]. Walter [771, 115]. Wang [1069]. warfarin [298]. warping [656]. wastewater [1467]. Watson [321, 323, 330, 328]. wave [435]. wavelet [1221, 736]. way [845, 800, 266, 1046, 340, 893]. ways [1486]. weak [103, 659]. wealth [489]. wearable [1033, 1583]. wedge [1127, 191, 816, 731, 426, 964, 68, 1003, 924, 42, 1455, 600, 841, 857, 1589, 275]. weekly [307]. Weibull [95, 640, 1004]. weight [1150]. Weighted [76, 759, 820, 663, 1031, 936, 1636, 450, 1620, 1579, 1061, 163, 463, 497, 950, 1525, 1042, 913, 36, 1518, 1054, 1506, 1617, 1433, 1125, 708, 1296]. Weighting [1547, 700, 1008, 220, 235, 1436, 398, 1197, 1553, 197, 1001, 143, 262, 1327, 695, 649, 590, 376, 1613]. weighting-based [197]. weights [1533, 737, 1553, 1643, 215, 760, 925]. whether [997]. while [1605, 38]. Whitney [1276, 424, 3, 863]. whole [125]. whole-genome-sequence [125]. wide [139, 1014, 738, 1245, 850, 470, 1360, 302, 1619, 1399]. Wilcoxon [1276, 3, 863, 424]. Win [530, 1406, 908, 1513, 1056, 825, 808, 909]. win-fractions [1056]. Win-loss [1406]. Window [664, 969, 1237]. windows [1163, 11]. Winsten [1168]. wise [847, 1605, 774, 1024, 1619]. within [1585, 428, 953, 223, 1354, 707, 299, 545, 773, 52, 1037, 429, 218, 273]. within-cluster [1585]. within-run [545]. within-study [1354, 1037]. within-subject [773, 52, 429]. within-unit [218]. without [619, 276, 1474, 625, 1373]. Worcester [1421]. workflow [704, 1458]. world [1055, 205, 623, 1307, 725, 1089]. wrist [917].

X [908].

yearly [1431]. years [798, 322, 934, 896]. yield [832, 100]. Youden [428, 1454, 749]. young [79]. Yu [908].

zero [95, 640, 938, 304, 1211, 1522, 571, 337, 1284, 608, 1545, 1414, 307, 390, 382, 685, 880]. zero-augmented [390]. zero-inflated [640, 938, 304, 1211, 1522, 571, 337, 1284, 1545, 1414, 307, 382, 685, 880]. zero-inflation [608]. zeros [562]. Zhang [207, 1069]. zone [869].

References

Neuenschwander:2008:CAB

[1] Beat Neuenschwander, Michael Branson, and Thomas Gsponer. Critical aspects of the Bayesian approach to phase I cancer trials. Statistics in Medicine, 27(13):2420–2439, June 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See comments [269].

Tang:2018:CPI

[2] Yongqiang Tang. Controlled pattern imputation for sensitivity analysis of longitudinal binary and ordinal outcomes with nonignorable dropout. Statistics in Medicine, 37(9):1467–1481, April 30, 2018. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See correction [1208].

Fay:2018:CEC

[3] Michael P. Fay, Erica H. Brittain, Joanna H. Shih, Dean A. Follmann, and Erin E. Gabriel. Causal estimands and confidence intervals associated with Wilcoxon–Mann–Whitney tests in randomized experiments. Statistics in Medicine, 37(20):2923–2937, September 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See correction [1276].

Zhan:2018:OUS

[4] Zhuozhao Zhan, Geertruida H. de Bock, and Edwin R. van den Heuvel. Optimal unidirectional switch designs. Statistics in Medicine, 37(25): 3573–3588, November 10, 2018. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [175].

Walter:2019:OTS

[5] Stephen D. Walter, Robin M. Turner, and Petra Macaskill. Optimising the two-stage randomised trial design when some participants are indifferent in their treatment preferences. Statistics in Medicine, 38(13): 2317–2331, June 15, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [5].

Lee:2019:OEC

[6] Sungim Lee and Johan Lim. Online estimation of the case fatality rate using a run-off triangle data approach: an application to the Korean MERS outbreak in 2015. Statistics in Medicine, 38(14):2664–2679, June 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [854].

Conner:2019:ARM

[7] Sarah C. Conner, Lisa M. Sullivan, Emelia J. Benjamin, Michael P. LaValley, Sandro Galea, and Ludovic Trinquart. Adjusted restricted mean survival times in observational studies. Statistics in Medicine, 38 (20):3832–3860, September 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See correction [271].

Prescott:2019:TTS

[8] Robin J. Prescott. Two-tailed significance tests for 2 × 2 contingency tables: What is the alternative? Statistics in Medicine, 38(22):4264– 4269, September 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [47, 18, 8].

Jones:2019:QHD

[9] Hayley E. Jones, Constantine A. Gatsonsis, Thomas A. Trikalinos, Nicky J. Welton, and A. E. Ades. Quantifying how diagnostic test accuracy depends on threshold in a meta-analysis. Statistics in Medicine, 38 (24):4789–4803, October 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See correction [581].

Freeman:2019:IIN

[10] Suzanne C. Freeman, David Fisher, Ian R. White, Anne Auperin, and James R. Carpenter. Identifying inconsistency in network meta-analysis: Is the net heat plot a reliable method? Statistics in Medicine, 38(29): 5547–5564, December 20, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See comments [579] and reply [580].

Xia:2020:RAR

[11] Meng Xia, Susan Murray, and Nabihah Tayob. Regression analysis of recurrent-event-free time from multiple follow-up windows. Statistics in Medicine, 39(1):1–15, January 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jiang:2020:MBA

[12] Tao Jiang, Yahui Lu, Huimin Duan, Wei Zhang, and Aiyi Liu. A modelbased approach for clustering of multivariate semicontinuous data with application to dietary pattern analysis and intervention. Statistics in Medicine, 39(1):16–25, January 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mezzetti:2020:CIA

[13] Maura Mezzetti, Domenico Palli, and Francesca Dominici. Combining individual and aggregated data to investigate the role of socioeconomic disparities on cancer burden in Italy. Statistics in Medicine, 39(1):26–44, January 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Farcomeni:2020:MQR

[14] Alessio Farcomeni and Marco Geraci. Multistate quantile regression models. Statistics in Medicine, 39(1):45–56, January 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2020:EGE

[15] Lingxiao Wang, Daoying Lin, and Yan Li. Exploiting gene-environment independence in haplotype-based inferences for population-based casecontrol studies with complex sampling. Statistics in Medicine, 39(1): 57–69, January 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Haller:2020:CIE

[16] Bernhard Haller, Ulrich Mansmann, Dennis Dobler, Kurt Ulm, and Alexander Hapfelmeier. Confidence interval estimation for the changepoint of treatment stratification in the presence of a qualitative covariatetreatment interaction. Statistics in Medicine, 39(1):70–96, January 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pilz:2020:CAS

[17] Maximilian Pilz, Meinhard Kieser, and Kevin Kunzmann. Comments on “Adaptive sample size modification in clinical trials: Start small then ask for more?”. Statistics in Medicine, 39(1):97–98, January 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Prescott:2020:RCT

[18] Robin J. Prescott. Response to comments on “Two-tailed significance tests for 2 × 2 contingency tables: What is the alternative?”. Statistics in Medicine, 39(1):99–101, January 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [8, 47].

Austin:2020:RUT

[19] Peter C. Austin, Aurélien Latouche, and Jason P. Fine. A review of the use of time-varying covariates in the Fine–Gray subdistribution hazard competing risk regression model. Statistics in Medicine, 39(2):103–113, January 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See comment [183].

Lin:2020:SAM

[20] Jui-Hsiang Lin, Kuan-I Lin, Wen-Chung Lee, and Sheng-Hsuan Lin. Stochastic approach for mechanistic interaction under longitudinal studies with noninformative right censoring. Statistics in Medicine, 39(2): 114–128, January 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brisco:2020:NME

[21] Agnese Maria Di Brisco and Sonia Migliorati. A new mixed-effects mixture model for constrained longitudinal data. Statistics in Medicine, 39 (2):129–145, January 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fan:2020:IAS

[22] Xinyan Fan, Kuangnan Fang, Shuangge Ma, and Qingzhao Zhang. Integrating approximate single factor graphical models. Statistics in Medicine, 39(2):146–155, January 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xue:2020:TVF

[23] Lan Xue, Xinxin Shu, Peibei Shi, Colin O. Wu, and Annie Qu. Timevarying feature selection for longitudinal analysis. Statistics in Medicine, 39(2):156–170, January 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bakbergenuly:2020:EMA

[24] Ilyas Bakbergenuly, David C. Hoaglin, and Elena Kulinskaya. Estimation in meta-analyses of mean difference and standardized mean difference. Statistics in Medicine, 39(2):171–191, January 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cook:2020:MDS

[25] Thomas Cook and Ryan Zea. Missing data and sensitivity analysis for binary data with implications for sample size and power of randomized clinical trials. Statistics in Medicine, 39(2):192–204, January 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hubbard:2020:CBL

[26] Rebecca A. Hubbard. Correction to “A Bayesian latent class approach for EHR-based phenotyping”. Statistics in Medicine, 39(2):205, January 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kang:2020:CIL

[27] Joseph Kang, Yulei He, Jaeyoung Hong, Precious Esie, and Kyle T. Bernstein. Causal inference of latent classes in complex survey data with the estimating equation framework. Statistics in Medicine, 39(3):207– 219, February 10, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fisher:2020:EII

[28] Leigh H. Fisher and Jon Wakefield. Ecological inference for infectious disease data, with application to vaccination strategies. Statistics in Medicine, 39(3):220–238, February 10, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Talitman:2020:EIE

[29] Michal Talitman, Malka Gorfine, and David M. Steinberg. Estimating the intervention effect in calibration substudies. Statistics in Medicine, 39(3):239–251, February 10, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gronsbell:2020:EIR

[30] Jessica Gronsbell, Chuan Hong, Lei Nie, Ying Lu, and Lu Tian. Exact inference for the random-effect model for meta-analyses with rare events. Statistics in Medicine, 39(3):252–264, February 10, 2020. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See comment [206].

Larriba:2020:ORI

[31] Yolanda Larriba, Cristina Rueda, Miguel A. Fernández, and Shyamal D. Peddada. Order restricted inference in chronobiology. Statistics in Medicine, 39(3):265–278, February 10, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2020:AMA

[32] Wei Wang, Nan Lin, Jordan D. Oberhaus, and Michael S. Avidan. Assessing method agreement for paired repeated binary measurements administered by multiple raters. Statistics in Medicine, 39(3):279–293, February 10, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Huang:2020:HDR

[33] Lei Huang, Weiqiang Hang, and Yue Chao. High-dimensional regression with ordered multiple categorical predictors. Statistics in Medicine, 39 (3):294–309, February 10, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2020:IPR

[34] Shing M. Lee, Xiaoqi Lu, and Bin Cheng. Incorporating patient-reported outcomes in dose-finding clinical trials. Statistics in Medicine, 39(3):310– 325, February 10, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2020:JMB

[35] Bingshu E. Chen and Jia Wang. Joint modeling of binary response and survival for clustered data in clinical trials. Statistics in Medicine, 39 (3):326–339, February 10, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Silva:2020:ESA

[36] Ivair R. Silva, Joshua J. Gagne, Mehdi Najafzadeh, and Martin Kulldorff. Exact sequential analysis for multiple weighted binomial end points. Statistics in Medicine, 39(3):340–351, February 10, 2020. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Huang:2020:IOT

[37] Ying Huang and Xiao-Hua Zhou. Identification of the optimal treatment regimen in the presence of missing covariates. Statistics in Medicine, 39 (4):353–368, February 20, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shi:2020:SSE

[38] Xu Shi, Robert Wellman, Patrick J. Heagerty, Jennifer C. Nelson, and Andrea J. Cook. Safety surveillance and the estimation of risk in select populations: Flexible methods to control for confounding while targeting marginal comparisons via standardization. Statistics in Medicine, 39(4): 369–386, February 20, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zeng:2020:MAC

[39] Leilei Zeng, Richard J. Cook, and Jooyoung Lee. Multistate analysis from cross-sectional and auxiliary samples. Statistics in Medicine, 39 (4):387–408, February 20, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Su:2020:AMC

[40] Chien-Lin Su and Lajmi Lakhal-Chaieb. Association measures for clustered competing risks. Statistics in Medicine, 39(4):409–423, February 20, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wu:2020:OIA

[41] Xiao Wu, Yi Xu, and Bradley P. Carlin. Optimizing interim analysis timing for Bayesian adaptive commensurate designs. Statistics in Medicine, 39(4):424–437, February 20, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2020:DAC

[42] Fan Li. Design and analysis considerations for cohort stepped wedge cluster randomized trials with a decay correlation structure. Statistics in Medicine, 39(4):438–455, February 20, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shu:2020:CIN

[43] Di Shu and Grace Y. Yi. Causal inference with noisy data: Bias analysis and estimation approaches to simultaneously addressing missingness and misclassification in binary outcomes. Statistics in Medicine, 39(4):456– 468, February 20, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2020:APA

[44] Ming Wang, Qi Long, Chixiang Chen, and Lijun Zhang. Assessing predictive accuracy of survival regressions subject to nonindependent censoring. Statistics in Medicine, 39(4):469–480, February 20, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stegherr:2020:ECI

[45] Regina Stegherr, Arthur Allignol, Reinhard Meister, Christof Schaefer, and Jan Beyersmann. Estimating cumulative incidence functions in competing risks data with dependent left-truncation. Statistics in Medicine, 39(4):481–493, February 20, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2020:RTM

[46] Yanying Wang, William F. Rosenberger, and Diane Uschner. Randomization tests for multiarmed randomized clinical trials. Statistics in Medicine, 39(4):494–509, February 20, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Andres:2020:CTT

[47] A. Martín Andrés, I. Herranz Tejedor, and F. Gayá Moreno. Comments on “Two-tailed significance tests for 2 × 2 contingency tables: What is the alternative?” by Robin J. Prescott, Statistics in Medicine 2019; 38:4264–4269. Statistics in Medicine, 39(4):510–513, February 20, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [8, 18].

Day:2020:E

[48] Simon Day, Els Goetghebeur, and Joel Greenhouse. Editorial. Statistics in Medicine, 39(5):515–516, February 28, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tounkara:2020:FCB

[49] Fodé Tounkara, Geneviève Lefebvre, Celia Greenwood, and Karim Oualkacha. A flexible copula-based approach for the analysis of secondary phenotypes in ascertained samples. Statistics in Medicine, 39 (5):517–543, February 28, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Robbins:2020:REC

[50] Michael W. Robbins, Beth Ann Griffin, Regina A. Shih, and Mary Ellen Slaughter. Robust estimation of the causal effect of time-varying neighborhood factors on health outcomes. Statistics in Medicine, 39(5):544– 561, February 28, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tian:2020:ECT

[51] Yuqi Tian, Torsten Hothorn, Chun Li, Frank E. Harrell, Jr., and Bryan E. Shepherd. An empirical comparison of two novel transformation models. Statistics in Medicine, 39(5):562–576, February 28, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nordgren:2020:EME

[52] Rachel Nordgren, Donald Hedeker, Genevieve Dunton, and Chih-Hsiang Yang. Extending the mixed-effects model to consider within-subject variance for ecological momentary assessment data. Statistics in Medicine, 39(5):577–590, February 28, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zapf:2020:ATD

[53] Antonia Zapf, Maria Stark, Oke Gerke, Christoph Ehret, Norbert Benda, Patrick Bossuyt, Jon Deeks, Johannes Reitsma, Todd Alonzo, and Tim Friede. Adaptive trial designs in diagnostic accuracy research. Statistics in Medicine, 39(5):591–601, February 28, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Follmann:2020:AOC

[54] Dean Follmann, Michael P. Fay, Toshimitsu Hamasaki, and Scott Evans. Analysis of ordered composite endpoints. Statistics in Medicine, 39(5): 602–616, February 28, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ren:2020:SBV

[55] Jie Ren, Fei Zhou, Xiaoxi Li, Qi Chen, Hongmei Zhang, Shuangge Ma, Yu Jiang, and Cen Wu. Semiparametric Bayesian variable selection for gene-environment interactions. Statistics in Medicine, 39(5):617–638, February 28, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thommes:2020:APE

[56] Edward W. Thommes, Salaheddin M. Mahmud, Yinong Young-Xu, Julia Thornton Snider, Robertus van Aalst, Jason K. H. Lee, Yuliya Halchenko, Ellyn Russo, and Ayman Chit. Assessing the prior event rate ratio method via probabilistic bias analysis on a Bayesian network. Statistics in Medicine, 39(5):639–659, February 28, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yen:2020:MOS

[57] Amy Ming-Fang Yen and Hsiu-Hsi Chen. Modeling the overdetection of screen-identified cancers in population-based cancer screening with the Coxian phase-type Markov process. Statistics in Medicine, 39(5):660– 673, February 28, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2020:PN

[58] Anonymous. Publisher’s note. Statistics in Medicine, 39(5):674, February 28, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kawaguchi:2020:SSC

[59] Eric S. Kawaguchi, Marc A. Suchard, Zhenqiu Liu, and Gang Li. A surrogate l_{0} sparse Cox’s regression with applications to sparse high-dimensional massive sample size time-to-event data. Statistics in Medicine, 39(6):675–686, March 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2020:IRO

[60] Wei Zhang, Aiyi Liu, Qizhai Li, and Paul S. Albert. Incorporating retesting outcomes for estimation of disease prevalence. Statistics in Medicine, 39(6):687–697, March 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wei:2020:CIT

[61] Jing Wei and Jianrong Wu. Cancer immunotherapy trial design with cure rate and delayed treatment effect. Statistics in Medicine, 39(6):698–708, March 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lui:2020:EIE

[62] Kung-Jong Lui. Exact interval estimators for some commonly used measures of binary agreement. Statistics in Medicine, 39(6):709–723, March 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yang:2020:BHV

[63] Yi Yang, Saonli Basu, and Lin Zhang. A Bayesian hierarchical variable selection prior for pathway-based GWAS using summary statistics. Statistics in Medicine, 39(6):724–739, March 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Szczesniak:2020:DPP

[64] Rhonda D. Szczesniak, Weiji Su, Cole Brokamp, Ruth H. Keogh, John P. Pestian, Michael Seid, Peter J. Diggle, and John P. Clancy. Dynamic predictive probabilities to monitor rapid cystic fibrosis disease progression. Statistics in Medicine, 39(6):740–756, March 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ma:2020:GMC

[65] Shiyang Ma and Michael P. McDermott. Generalized multiple contrast tests in dose-response studies. Statistics in Medicine, 39(6):757–772, March 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Beesley:2020:ELH

[66] Lauren J. Beesley, Maxwell Salvatore, Lars G. Fritsche, Anita Pandit, Arvind Rao, Chad Brummett, Cristen J. Willer, Lynda D. Lisabeth, and Bhramar Mukherjee. The emerging landscape of health research based on biobanks linked to electronic health records: Existing resources, statistical challenges, and potential opportunities. Statistics in Medicine, 39(6):773–800, March 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lim:2020:UMR

[67] Elise Lim, Han Chen, Josée Dupuis, and Ching-Ti Liu. A unified method for rare variant analysis of gene-environment interactions. Statistics in Medicine, 39(6):801–813, March 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kennedy-Shaffer:2020:NMA

[68] Lee Kennedy-Shaffer, Victor de Gruttola, and Marc Lipsitch. Novel methods for the analysis of stepped wedge cluster randomized trials. Statistics in Medicine, 39(7):815–844, March 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ding:2020:ESA

[69] Yuxin Ding, Marianthi Markatou, and Robert Ball. An evaluation of statistical approaches to postmarketing surveillance. Statistics in Medicine, 39(7):845–874, March 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gan:2020:QDR

[70] Fah Fatt Gan, Jing Sheng Yuen, and Sven Knoth. Quicker detection riskadjusted cumulative sum charting procedures. Statistics in Medicine, 39(7):875–889, March 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Huang:2020:MRU

[71] Yen-Tsung Huang. Mendelian randomization using semiparametric linear transformation models. Statistics in Medicine, 39(7):890–905, March 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mozgunov:2020:ISC

[72] Pavel Mozgunov and Thomas Jaki. Improving safety of the continual reassessment method via a modified allocation rule. Statistics in Medicine, 39(7):906–922, March 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Krzykalla:2020:EIP

[73] Julia Krzykalla, Axel Benner, and Annette Kopp-Schneider. Exploratory identification of predictive biomarkers in randomized trials with normal endpoints. Statistics in Medicine, 39(7):923–939, March 30, 2020. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pedeli:2020:IVT

[74] Xanthi Pedeli and Dimitris Karlis. An integer-valued time series model for multivariate surveillance. Statistics in Medicine, 39(7):940–954, March 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2020:SSL

[75] Xiaochen Zhang, Qingzhao Zhang, Xiaofeng Wang, Shuangge Ma, and Kuangnan Fang. Structured sparse logistic regression with application to lung cancer prediction using breath volatile biomarkers. Statistics in Medicine, 39(7):955–967, March 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ahmed:2020:WKM

[76] Ismaïl Ahmed and Philippe Flandre. Weighted Kaplan–Meier estimators motivating to estimate HIV-1 RNA reduction censored by a limit of detection. Statistics in Medicine, 39(7):968–983, March 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Roychoudhury:2020:BLH

[77] Satrajit Roychoudhury and Beat Neuenschwander. Bayesian leveraging of historical control data for a clinical trial with time-to-event endpoint. Statistics in Medicine, 39(7):984–995, March 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2020:PSA

[78] Xueying Wang, Lang Li, Lei Wang, Weixing Feng, and Pengyue Zhang. Propensity score-adjusted three-component mixture model for drug-drug interaction data mining in FDA adverse event reporting system. Statistics in Medicine, 39(7):996–1010, March 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yang:2020:SMM

[79] Wan Yang, Rebecca D. Kehm, and Mary Beth Terry. Survival model methods for analyses of cancer incidence trends in young adults. Statistics in Medicine, 39(7):1011–1024, March 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ahmed:2020:FMA

[80] Mohamed-Salem Ahmed and Michaël Genin. A functional-model-adjusted spatial scan statistic. Statistics in Medicine, 39(8):1025–1040, April 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shauly-Aharonov:2020:ETH

[81] Michal Shauly-Aharonov. An exact test with high power and robustness to unmeasured confounding effects. Statistics in Medicine, 39(8):1041– 1053, April 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2020:CDF

[82] Hongkai Li, Wang Miao, Zheng Cai, Xinhui Liu, Tao Zhang, Fuzhong Xue, and Zhi Geng. Causal data fusion methods using summary-level statistics for a continuous outcome. Statistics in Medicine, 39(8):1054– 1067, April 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Buhule:2020:MRL

[83] Olive D. Buhule, Hyoyoung Choo-Wosoba, and Paul S. Albert. Modeling repeated labor curves in consecutive pregnancies: Individualized prediction of labor progression from previous pregnancy data. Statistics in Medicine, 39(8):1068–1083, April 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ghosh:2020:AMM

[84] Pranab Ghosh, Lingyun Liu, and Cyrus Mehta. Adaptive multiarm multistage clinical trials. Statistics in Medicine, 39(8):1084–1102, April 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Papanikos:2020:BHM

[85] Tasos Papanikos, John R. Thompson, Keith R. Abrams, Nicolas Städler, Oriana Ciani, Rod Taylor, and Sylwia Bujkiewicz. Bayesian hierarchical meta-analytic methods for modeling surrogate relationships that vary across treatment classes using aggregate data. Statistics in Medicine, 39(8):1103–1124, April 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

OMalley:2020:MPE

[86] A. James O’Malley, Erika L. Moen, Julie P. W. Bynum, Andrea M. Austin, and Jonathan S. Skinner. Modeling peer effect modification by network strength: The diffusion of implantable cardioverter defibrillators in the US hospital network. Statistics in Medicine, 39(8):1125–1144, April 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brown:2020:EAA

[87] Alexandria C. Brown, Stephen A. Lauer, Christine C. Robinson, Ann-Christine Nyquist, Suchitra Rao, and Nicholas G. Reich. Evaluating the ALERT algorithm for local outbreak onset detection in seasonal infectious disease surveillance data. Statistics in Medicine, 39(8):1145–1155, April 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Slade:2020:FCT

[88] Emily Slade and Melissa G. Naylor. A fair comparison of tree-based and parametric methods in multiple imputation by chained equations. Statistics in Medicine, 39(8):1156–1166, April 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tapsoba:2020:MGC

[89] Jean de Dieu Tapsoba, Ching-Yun Wang, Sahar Zangeneh, and Ying Qing Chen. Methods for generalized change-point models: with applications to human immunodeficiency virus surveillance and diabetes data. Statistics in Medicine, 39(8):1167–1182, April 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bonofiglio:2020:ROI

[90] Federico Bonofiglio, Martin Schumacher, and Harald Binder. Recovery of original individual person data (IPD) inferences from empirical IPD summaries only: Applications to distributed computing under disclosure constraints. Statistics in Medicine, 39(8):1183–1198, April 15, 2020. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Young:2020:CFC

[91] Jessica G. Young, Mats J. Stensrud, Eric J. Tchetgen Tchetgen, and Miguel A. Hernán. A causal framework for classical statistical estimands in failure-time settings with competing events. Statistics in Medicine, 39(8):1199–1236, April 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ma:2020:HDC

[92] Chenjin Ma, Yang Li, BenChang Shia, and Shuangge Ma. Human disease cost network analysis. Statistics in Medicine, 39(9):1237–1249, April 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhao:2020:CDT

[93] Ying-Qi Zhao, Ruoqing Zhu, Guanhua Chen, and Yingye Zheng. Constructing dynamic treatment regimes with shared parameters for censored data. Statistics in Medicine, 39(9):1250–1263, April 30, 2020. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Maruo:2020:NBS

[94] Kazushi Maruo, Ryota Ishii, Yusuke Yamaguchi, Masaaki Doi, and Masahiko Gosho. A note on the bias of standard errors when orthogonality of mean and variance parameters is not satisfied in the mixed model for repeated measures analysis. Statistics in Medicine, 39(9):1264–1274, April 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Burger:2020:RBM

[95] Divan Aristo Burger, Robert Schall, Johannes Theodorus Ferreira, and Ding-Geng Chen. A robust Bayesian mixed effects approach for zero inflated and highly skewed longitudinal count data emanating from the zero inflated discrete Weibull distribution. Statistics in Medicine, 39(9): 1275–1291, April 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gjerdevik:2020:DEG

[96] Miriam Gjerdevik, Håkon K. Gjessing, Julia Romanowska, Øystein A. Haaland, Astanand Jugessur, Nikolai O. Czajkowski, and Rolv T. Lie. Design efficiency in genetic association studies. Statistics in Medicine, 39(9):1292–1310, April 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wen:2020:MLM

[97] Yalu Wen and Qing Lu. Multikernel linear mixed model with adaptive lasso for complex phenotype prediction. Statistics in Medicine, 39(9): 1311–1327, April 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kaizer:2020:FBM

[98] Alexander M. Kaizer and Joseph S. Koopmeiners. A fully Bayesian mixture model approach for identifying noncompliance in a regulatory tobacco clinical trial. Statistics in Medicine, 39(9):1328–1342, April 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Singh:2020:OOC

[99] Gurbakhshash Singh and Gordon Hilton Fick. Ordinal outcomes: a cumulative probability model with the log link and an assumption of proportionality. Statistics in Medicine, 39(9):1343–1361, April 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sadatsafavi:2020:TFS

[100] Mohsen Sadatsafavi, Mohammad Ali Mansournia, and Paul Gustafson. A threshold-free summary index for quantifying the capacity of covariates to yield efficient treatment rules. Statistics in Medicine, 39(9):1362–1373, April 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Estes:2020:PDF

[101] Jason P. Estes, Yanjun Chen, Damla Sentürk, Connie M. Rhee, Esra Kürüm, Amy S. You, Elani Streja, Kamyar Kalantar-Zadeh, and Danh V. Nguyen. Profiling dialysis facilities for adverse recurrent events. Statistics in Medicine, 39(9):1374–1389, April 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Austin:2020:CMR

[102] Peter C. Austin, Douglas S. Lee, and George Leckie. Comparing a multivariate response Bayesian random effects logistic regression model with a latent variable item response theory model for provider profiling on multiple binary indicators simultaneously. Statistics in Medicine, 39(9): 1390–1406, April 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Proschan:2020:PSV

[103] Michael A. Proschan and Erica H. Brittain. A primer on strong vs weak control of familywise error rate. Statistics in Medicine, 39(9):1407–1413, April 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lai:2020:MAN

[104] En-Yu Lai, Stephannie Shih, Yen-Tsung Huang, and Shunping Wang. A mediation analysis for a nonrare dichotomous outcome with sequentially ordered multiple mediators. Statistics in Medicine, 39(10):1415–1428, May 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dufault:2020:ACC

[105] Suzanne M. Dufault and Nicholas P. Jewell. Analysis of counts for cluster randomized trials: Negative controls and test-negative designs. Statistics in Medicine, 39(10):1429–1439, May 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nguyen:2020:APP

[106] Tri-Long Nguyen, Gary S. Collins, Fabio Pellegrini, Karel G. M. Moons, and Thomas P. A. Debray. On the aggregation of published prognostic scores for causal inference in observational studies. Statistics in Medicine, 39(10):1440–1457, May 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhu:2020:SCD

[107] Anqi Zhu, Donglin Zeng, Li Shen, Xia Ning, Lang Li, and Pengyue Zhang. A super-combo-drug test to detect adverse drug events and drug interactions from electronic health records in the era of polypharmacy. Statistics in Medicine, 39(10):1458–1472, May 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2020:TCP

[108] Chun Yin Lee, Xuerong Chen, and Kwok Fai Lam. Testing for changepoint in the covariate effects based on the Cox regression model. Statistics in Medicine, 39(10):1473–1488, May 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kennedy-Shaffer:2020:SSE

[109] Lee Kennedy-Shaffer and Michael D. Hughes. Sample size estimation for stratified individual and cluster randomized trials with binary outcomes. Statistics in Medicine, 39(10):1489–1513, May 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jang:2020:SAE

[110] Eun Jin Jang, Balgobin Nandram, Yousun Ko, and Dal Ho Kim. Small area estimation of receiver operating characteristic curves for ordinal data under stochastic ordering. Statistics in Medicine, 39(10):1514–1528, May 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Iskandar:2020:ANM

[111] Rowan Iskandar. Adding noise to Markov cohort state-transition model in decision modeling and cost-effectiveness analysis. Statistics in Medicine, 39(10):1529–1540, May 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tan:2020:HTA

[112] Xianming Tan, Bingshu E. Chen, Jianping Sun, Tejendra Patel, and Joseph G. Ibrahim. A hierarchical testing approach for detecting safety signals in clinical trials. Statistics in Medicine, 39(10):1541–1557, May 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pagui:2020:ASF

[113] Euloge C. Kenne Pagui and Enrico A. Colosimo. Adjusted score functions for monotone likelihood in the Cox regression model. Statistics in Medicine, 39(10):1558–1572, May 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Huang:2020:UAS

[114] Chiung-Yu Huang and Jing Qin. A unified approach for synthesizing population-level covariate effect information in semiparametric estimation with survival data. Statistics in Medicine, 39(10):1573–1590, May 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Walter:2020:CWS

[115] Stephen D. Walter. Correction to Walter SD, Turner RM, Macaskill P. Optimising the two-stage randomised trial design when some participants are indifferent in their treatment preferences (2019). Statistics in Medicine 38, 2317–2331. Statistics in Medicine, 39(10):1591, May 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Whitehead:2020:ETE

[116] John Whitehead, Yasin Desai, and Thomas Jaki. Estimation of treatment effects following a sequential trial of multiple treatments. Statistics in Medicine, 39(11):1593–1609, May 20, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Joseph:2020:IPM

[117] Anny-Claude Joseph, Montserrat Fuentes, and David C. Wheeler. The impact of population mobility on estimates of environmental exposure effects in a case-control study. Statistics in Medicine, 39(11):1610–1622, May 20, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Austin:2020:VEW

[118] Peter C. Austin and Guy Cafri. Variance estimation when using propensity-score matching with replacement with survival or time-to-event outcomes. Statistics in Medicine, 39(11):1623–1640, May 20, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Blake:2020:PSU

[119] Helen A. Blake, Clémence Leyrat, Kathryn E. Mansfield, Shaun Seaman, Laurie A. Tomlinson, James Carpenter, and Elizabeth J. Williamson. Propensity scores using missingness pattern information: a practical guide. Statistics in Medicine, 39(11):1641–1657, May 20, 2020. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gomes:2020:ETE

[120] Manuel Gomes, Michael G. Kenward, Richard Grieve, and James Carpenter. Estimating treatment effects under untestable assumptions with nonignorable missing data. Statistics in Medicine, 39(11):1658–1674, May 20, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2020:IAU

[121] Min Zhang, Youfei Yu, Shikun Wang, Maxwell Salvatore, Lars G. Fritsche, Zihuai He, and Bhramar Mukherjee. Interaction analysis under misspecification of main effects: Some common mistakes and simple solutions. Statistics in Medicine, 39(11):1675–1694, May 20, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tran:2020:GFM

[122] Thao M. P. Tran, Steven Abrams, and Roel Braekers. A general frailty model to accommodate individual heterogeneity in the acquisition of multiple infections: an application to bivariate current status data. Statistics in Medicine, 39(12):1695–1714, May 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2020:BSH

[123] Chunlin Wang and Dongsheng Tu. A bootstrap semiparametric homogeneity test for the distributions of multigroup proportional data, with applications to analysis of quality of life outcomes in clinical trials. Statistics in Medicine, 39(12):1715–1731, May 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ye:2020:GSC

[124] Xuan Ye, Larry L. Tang, and Xiaochen Zhu. Group sequential comparison of positive predictive value curves for correlated biomarker data. Statistics in Medicine, 39(12):1732–1745, May 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cassidy:2020:MBI

[125] Rosanna Cassidy, Theodore Kypraios, and Philip D. O’Neill. Modelling, Bayesian inference, and model assessment for nosocomial pathogens using whole-genome-sequence data. Statistics in Medicine, 39(12):1746– 1765, May 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chapple:2020:NBC

[126] Andrew G. Chapple, Taylor Peak, and Ashok Hemal. A novel Bayesian continuous piecewise linear log-hazard model, with estimation and inference via reversible jump Markov chain Monte Carlo. Statistics in Medicine, 39(12):1766–1780, May 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Behboudi:2020:MMS

[127] Maryam Behboudi and Rahman Farnoosh. Modified models and simulations for estimating dynamic functional connectivity in resting state functional magnetic resonance imaging. Statistics in Medicine, 39(12): 1781–1800, May 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhou:2020:TPH

[128] Xiaoxiao Zhou, Kai Kang, and Xinyuan Song. Two-part hidden Markov models for semicontinuous longitudinal data with nonignorable missing covariates. Statistics in Medicine, 39(13):1801–1816, June 15, 2020. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jhuang:2020:SSD

[129] An-Ting Jhuang, Montserrat Fuentes, Dipankar Bandyopadhyay, and Brian J. Reich. Spatiotemporal signal detection using continuous shrinkage priors. Statistics in Medicine, 39(13):1817–1832, June 15, 2020. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jiang:2020:MBI

[130] Shu Jiang, Richard J. Cook, and Leilei Zeng. Mitigating bias from intermittent measurement of time-dependent covariates in failure time analysis. Statistics in Medicine, 39(13):1833–1845, June 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gwon:2020:NMR

[131] Yeongjin Gwon, May Mo, Ming-Hui Chen, Zhiyi Chi, Juan Li, Amy H. Xia, and Joseph G. Ibrahim. Network meta-regression for ordinal outcomes: Applications in comparing Crohn’s disease treatments. Statistics in Medicine, 39(13):1846–1870, June 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kasza:2020:SSP

[132] Jessica Kasza, Richard Hooper, Andrew Copas, and Andrew B. Forbes. Sample size and power calculations for open cohort longitudinal cluster randomized trials. Statistics in Medicine, 39(13):1871–1883, June 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Komukai:2020:DRI

[133] Sho Komukai and Satoshi Hattori. Doubly robust inference procedure for relative survival ratio in population-based cancer registry data. Statistics in Medicine, 39(13):1884–1900, June 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wynants:2020:C

[134] Laure Wynants. Correction. Statistics in Medicine, 39(13):1901–1902, June 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Peng:2020:MRM

[135] Defen Peng, Gilbert MacKenzie, and Kevin Burke. A multiparameter regression model for interval-censored survival data. Statistics in Medicine, 39(14):1903–1918, June 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Spiess:2020:GEE

[136] Martin Spiess, Daniel Fernández, Thuong Nguyen, and Ivy Liu. Generalized estimating equations to estimate the ordered stereotype logit model for panel data. Statistics in Medicine, 39(14):1919–1940, June 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ryan:2020:VEK

[137] Mary M. Ryan, William D. Spotnitz, and Daniel L. Gillen. Variance estimation for the kappa statistic in the presence of clustered data and heterogeneous observations. Statistics in Medicine, 39(14):1941–1951, June 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dai:2020:NEB

[138] Tian Dai, Ying Guo, Limin Peng, and Amita Manatunga. Nonparametric estimation of broad sense agreement between ordinal and censored continuous outcomes. Statistics in Medicine, 39(14):1952–1964, June 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Beesley:2020:AFE

[139] Lauren J. Beesley, Lars G. Fritsche, and Bhramar Mukherjee. An analytic framework for exploring sampling and observation process biases in genome and phenome-wide association studies using electronic health records. Statistics in Medicine, 39(14):1965–1979, June 30, 2020. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zapf:2020:BSS

[140] Antonia Zapf, Thomas Asendorf, Christoph Anten, Tobias Mütze, and Tim Friede. Blinded sample size reestimation for negative binomial regression with baseline adjustment. Statistics in Medicine, 39(14):1980– 1998, June 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dahabreh:2020:EIR

[141] Issa J. Dahabreh, Sarah E. Robertson, Jon A. Steingrimsson, Elizabeth A. Stuart, and Miguel A. Hernán. Extending inferences from a randomized trial to a new target population. Statistics in Medicine, 39 (14):1999–2014, June 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2020:CPT

[142] Anonymous. Confidence, prediction, and tolerance in linear mixed models. Statistics in Medicine, 39(14):2015, June 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mao:2020:FRA

[143] Huzhang Mao and Liang Li. Flexible regression approach to propensity score analysis and its relationship with matching and weighting. Statistics in Medicine, 39(15):2017–2034, July 10, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2020:PID

[144] Juhee Lee, Peter F. Thall, and Pavlos Msaouel. A phase I-II design based on periodic and continuous monitoring of disease status and the times to toxicity and death. Statistics in Medicine, 39(15):2035–2050, July 10, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2020:FNM

[145] Rui Wang, Ante Bing, Cathy Wang, Yuchen Hu, Ronald J. Bosch, and Victor DeGruttola. A flexible nonlinear mixed effects model for HIV viral load rebound after treatment interruption. Statistics in Medicine, 39(15):2051–2066, July 10, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Herrmann:2020:NCP

[146] Carolin Herrmann, Maximilian Pilz, Meinhard Kieser, and Geraldine Rauch. A new conditional performance score for the evaluation of adaptive group sequential designs with sample size recalculation. Statistics in Medicine, 39(15):2067–2100, July 10, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2020:EDO

[147] Piao Chen, Xiuju Fu, Stefan Ma, Hai-Yan Xu, Wanbing Zhang, Gaoxi Xiao, Rick Siow Mong Goh, George Xu, and Lee Ching Ng. Early dengue outbreak detection modeling based on dengue incidences in Singapore during 2012 to 2017. Statistics in Medicine, 39(15):2101–2114, July 10, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Riley:2020:IPD

[148] Richard D. Riley, Thomas P. A. Debray, David Fisher, Miriam Hattle, Nadine Marlin, Jeroen Hoogland, Francois Gueyffier, Jan A. Staessen, Jiguang Wang, Karel G. M. Moons, Johannes B. Reitsma, and Joie Ensor. Individual participant data meta-analysis to examine interactions between treatment effect and participant-level covariates: Statistical recommendations for conduct and planning. Statistics in Medicine, 39(15): 2115–2137, July 10, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2020:BBR

[149] Yuliang Li, Dipankar Bandyopadhyay, Fangzheng Xie, and Yanxun Xu. BAREB: a Bayesian repulsive biclustering model for periodontal data. Statistics in Medicine, 39(16):2139–2151, July 20, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zou:2020:ICP

[150] Yixuan Zou and Derek S. Young. Improving coverage probabilities for parametric tolerance intervals via bootstrap calibration. Statistics in Medicine, 39(16):2152–2166, July 20, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Capanu:2020:OVS

[151] Marinela Capanu, Mihai Giurcanu, Colin B. Begg, and Mithat Gönen. Optimized variable selection via repeated data splitting. Statistics in Medicine, 39(16):2167–2184, July 20, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chandereng:2020:IRC

[152] Thevaa Chandereng, Xiaodan Wei, and Rick Chappell. Imbalanced randomization in clinical trials. Statistics in Medicine, 39(16):2185–2196, July 20, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Keogh:2020:SGD

[153] Ruth H. Keogh, Pamela A. Shaw, Paul Gustafson, Raymond J. Carroll, Veronika Deffner, Kevin W. Dodd, Helmut Küchenhoff, Janet A.

Tooze, Michael P. Wallace, Victor Kipnis, and Laurence S. Freedman. STRATOS guidance document on measurement error and misclassification of variables in observational epidemiology: Part 1 — basic theory and simple methods of adjustment. Statistics in Medicine, 39(16): 2197–2231, July 20, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shaw:2020:SGD

[154] Pamela A. Shaw, Paul Gustafson, Raymond J. Carroll, Veronika Deffner, Kevin W. Dodd, Ruth H. Keogh, Victor Kipnis, Janet A. Tooze, Michael P. Wallace, Helmut Küchenhoff, and Laurence S. Freedman. STRATOS guidance document on measurement error and misclassification of variables in observational epidemiology: Part 2 — more complex methods of adjustment and advanced topics. Statistics in Medicine, 39 (16):2232–2263, July 20, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liao:2020:UDS

[155] Shirley X. Liao and Corwin M. Zigler. Uncertainty in the design stage of two-stage Bayesian propensity score analysis. Statistics in Medicine, 39(17):2265–2290, July 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lopez-Cheda:2020:NCH

[156] Ana López-Cheda, Maria Amalia Jácome, Ingrid Van Keilegom, and Ricardo Cao. Nonparametric covariate hypothesis tests for the cure rate in mixture cure models. Statistics in Medicine, 39(17):2291–2307, July 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brown:2020:NAP

[157] Derek W. Brown, Stacia M. DeSantis, Thomas J. Greene, Vahed Maroufy, Ashraf Yaseen, Hulin Wu, George Williams, and Michael D. Swartz. A novel approach for propensity score matching and stratification for multiple treatments: Application to an electronic health recordderived study. Statistics in Medicine, 39(17):2308–2323, July 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Su:2020:DRE

[158] Chien-Lin Su, Russell Steele, and Ian Shrier. Doubly robust estimation and causal inference for recurrent event data. Statistics in Medicine, 39(17):2324–2338, July 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Steingrimsson:2020:DLS

[159] Jon Arni Steingrimsson and Samantha Morrison. Deep learning for survival outcomes. Statistics in Medicine, 39(17):2339–2349, July 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thomas:2020:MTD

[160] Laine E. Thomas, Siyun Yang, Daniel Wojdyla, and Douglas E. Schaubel. Matching with time-dependent treatments: a review and look forward. Statistics in Medicine, 39(17):2350–2370, July 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhu:2020:C

[161] Rui Zhu. Correction. Statistics in Medicine, 39(17):2371–2372, July 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Teng:2020:SPO

[162] Zhaoyang Teng, Yuan Tian, Yi Liu, and Guohui Liu. Seamless phase 2/3 oncology trial design with flexible sample size determination. Statistics in Medicine, 39(18):2373–2386, August 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hyun:2020:SWS

[163] Noorie Hyun, Hormuzd A. Katki, and Barry I. Graubard. Sampleweighted semiparametric estimation of cause-specific cumulative risk and incidence using left- or interval-censored data from electronic health records. Statistics in Medicine, 39(18):2387–2402, August 15, 2020. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ribeiro:2020:LGE

[164] Adèle H. Ribeiro and Júlia Maria Pavan Soler. Learning genetic and environmental graphical models from family data. Statistics in Medicine, 39(18):2403–2422, August 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Derkach:2020:GTM

[165] Andriy Derkach, Steven C. Moore, Simina M. Boca, and Joshua N. Sampson. Group testing in mediation analysis. Statistics in Medicine, 39 (18):2423–2436, August 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wu:2020:PAM

[166] Yuan Wu, Xiaofei Wang, Jiaxing Lin, Beilin Jia, and Kouros Owzar. Predictive accuracy of markers or risk scores for interval censored survival data. Statistics in Medicine, 39(18):2437–2446, August 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Parast:2020:QBD

[167] Layla Parast and Beth Ann Griffin. Quantifying the bias due to observed individual confounders in causal treatment effect estimates. Statistics in Medicine, 39(18):2447–2476, August 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Meddis:2020:CST

[168] Alessandra Meddis, Paul Blanche, François C. Bidard, and Aurélien Latouche. A covariate-specific time-dependent receiver operating characteristic curve for correlated survival data. Statistics in Medicine, 39 (19):2477–2489, August 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sordello:2020:TDA

[169] Matteo Sordello and Dylan S. Small. A test for differential ascertainment in case-control studies with application to child maltreatment. Statistics in Medicine, 39(19):2490–2505, August 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lauzon:2020:SPM

[170] Steven D. Lauzon, Viswanathan Ramakrishnan, Paul J. Nietert, Jody D. Ciolino, Michael D. Hill, and Wenle Zhao. Statistical properties of minimal sufficient balance and minimization as methods for controlling baseline covariate imbalance at the design stage of sequential clinical trials. Statistics in Medicine, 39(19):2506–2517, August 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2020:BAM

[171] Wan-Lun Wang. Bayesian analysis of multivariate linear mixed models with censored and intermittent missing responses. Statistics in Medicine, 39(19):2518–2535, August 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Riley:2020:OSI

[172] Richard D. Riley, Amardeep Legha, Dan Jackson, Tim P. Morris, Joie Ensor, Kym I. E. Snell, Ian R. White, and Danielle L. Burke. One-stage individual participant data meta-analysis models for continuous and binary outcomes: Comparison of treatment coding options and estimation methods. Statistics in Medicine, 39(19):2536–2555, August 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Barrett:2020:SRD

[173] James E. Barrett, Aylin Cakiroglu, Catey Bunce, Anoop Shah, and Spiros Denaxas. Selective recruitment designs for improving observational studies using electronic health records. Statistics in Medicine, 39 (19):2556–2567, August 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kimani:2020:PIE

[174] Peter K. Kimani, Susan Todd, Lindsay A. Renfro, Ekkehard Glimm, Josephine N. Khan, John A. Kairalla, and Nigel Stallard. Point and interval estimation in two-stage adaptive designs with time to event data and biomarker-driven subpopulation selection. Statistics in Medicine, 39 (19):2568–2586, August 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhan:2020:COU

[175] Zhuozhao Zhan and Edwin R. van den Heuvel. Correction for “Optimal unidirectional switch designs” by Zhuozhao Zhan, Geertruida H. de Bock, and Edwin R. van den Heuvel. Statistics in Medicine, 39(19):2587, August 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [4].

Yuan:2020:TMA

[176] Chengbo Yuan, Donald Hedeker, Robin Mermelstein, and Hui Xie. A tractable method to account for high-dimensional nonignorable missing data in intensive longitudinal data. Statistics in Medicine, 39(20):2589– 2605, September 10, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Boschini:2020:ECI

[177] Cristina Boschini, Klaus K. Andersen, Hélène Jacqmin-Gadda, Pierre Joly, and Thomas H. Scheike. Excess cumulative incidence estimation for matched cohort survival studies. Statistics in Medicine, 39(20):2606– 2620, September 10, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2020:CID

[178] Yunqi Zhang and Niansheng Tang. Confidence intervals of the difference between areas under two ROC curves in matched-pair experiments. Statistics in Medicine, 39(20):2621–2638, September 10, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Carragher:2020:BHA

[179] Raymond Carragher, Tanja Mueller, Marion Bennie, and Chris Robertson. A Bayesian hierarchical approach for multiple outcomes in routinely collected healthcare data. Statistics in Medicine, 39(20):2639– 2654, September 10, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Horiguchi:2020:PTC

[180] Miki Horiguchi and Hajime Uno. On permutation tests for comparing restricted mean survival time with small sample from randomized trials. Statistics in Medicine, 39(20):2655–2670, September 10, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2020:NSE

[181] Yilong Zhang and Yongzhao Shao. A numerical strategy to evaluate performance of predictive scores via a copula-based approach. Statistics in Medicine, 39(20):2671–2684, September 10, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Flandre:2020:NRP

[182] Philippe Flandre and John O’Quigley. Note on the role of the placebo group in the short-term and long-term hazard ratio model. Statistics in Medicine, 39(20):2685–2688, September 10, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mori:2020:CRU

[183] Shunsuke Mori. Comment on: A review of the use of time-varying covariates in the Fine-Gray subdistribution hazard competing risk regression model by Peter C. Austin, Aurélien Latouche, and Jason P. Fine. Statistics in Medicine, 39(20):2689–2691, September 10, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [19].

Austin:2020:AR

[184] Peter C. Austin and Aurélien Latouche. Authors’ response. Statistics in Medicine, 39(20):2692, September 10, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Baiocchi:2020:C

[185] Mike Baiocchi, Jing Cheng, and Dylan Small. Correction. Statistics in Medicine, 39(20):2693, September 10, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cespedes:2020:RRC

[186] Marcela I. Cespedes, James M. McGree, Christopher C. Drovandi, Kerrie Mengersen, Jurgen Fripp, and James D. Doecke. Relative rate of change in cognitive score network dynamics via Bayesian hierarchical models reveal spatial patterns of neurodegeneration. Statistics in Medicine, 39 (21):2695–2713, September 20, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Austin:2020:GCC

[187] Peter C. Austin, Frank E. Harrell, Jr., and David van Klaveren. Graphical calibration curves and the integrated calibration index (ICI) for survival models. Statistics in Medicine, 39(21):2714–2742, September 20, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Huang:2020:HDS

[188] Hailin Huang, Jizi Shangguan, Xinmin Li, and Hua Liang. Highdimensional single-index models with censored responses. Statistics in Medicine, 39(21):2743–2754, September 20, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dissanayake:2020:ESA

[189] Manjari Dissanayake and A. Alexandre Trindade. An empirical saddlepoint approximation method for producing smooth survival and hazard functions under interval-censoring. Statistics in Medicine, 39(21):2755– 2766, September 20, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stare:2020:RCS

[190] Janez Stare, Robin Henderson, and Nina Ruzić Gorenjec. Random cancers as supported by registry data. Statistics in Medicine, 39(21):2767– 2778, September 20, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ford:2020:MVI

[191] Whitney P. Ford and Philip M. Westgate. Maintaining the validity of inference in small-sample stepped wedge cluster randomized trials with binary outcomes when using generalized estimating equations. Statistics in Medicine, 39(21):2779–2792, September 20, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Abuzaid:2020:IDB

[192] Ali H. Abuzaid. Identifying density-based local outliers in medical multivariate circular data. Statistics in Medicine, 39(21):2793–2798, September 20, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Teunis:2020:ESR

[193] P. F. M. Teunis and J. C. H. van Eijkeren. Estimation of seroconversion rates for infectious diseases: Effects of age and noise. Statistics in Medicine, 39(21):2799–2814, September 20, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cro:2020:SAC

[194] Suzie Cro, Tim P. Morris, Michael G. Kenward, and James R. Carpenter. Sensitivity analysis for clinical trials with missing continuous outcome data using controlled multiple imputation: a practical guide. Statistics in Medicine, 39(21):2815–2842, September 20, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2020:RBI

[195] Yanying Wang and William F. Rosenberger. Randomization-based interval estimation in randomized clinical trials. Statistics in Medicine, 39 (21):2843–2854, September 20, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vaughan:2020:EIS

[196] Gregory Vaughan, Robert Aseltine, Kun Chen, and Jun Yan. Efficient interaction selection for clustered data via stagewise generalized estimating equations. Statistics in Medicine, 39(22):2855–2868, September 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2020:SAC

[197] Fan Li and Andrew S. Allen. Secondary analysis of case-control association studies: Insights on weighting-based inference motivate a new specification test. Statistics in Medicine, 39(22):2869–2882, September 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2020:ICP

[198] Zhenxun Wang, Lifeng Lin, James S. Hodges, and Haitao Chu. The impact of covariance priors on arm-based Bayesian network meta-analyses with binary outcomes. Statistics in Medicine, 39(22):2883–2900, September 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kos:2020:CHP

[199] Michał Kos, Małgorzata Bogdan, Nancy W. Glynn, and Jaroslaw Harezlak. Classification of human physical activity based on raw accelerometry data via spherical coordinate transformation. Statistics in Medicine, 39(22):2901–2920, September 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tompsett:2020:GME

[200] Daniel Tompsett, Stephen Sutton, Shaun R. Seaman, and Ian R. White. A general method for elicitation, imputation, and sensitivity analysis for incomplete repeated binary data. Statistics in Medicine, 39(22):2921– 2935, September 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Huang:2020:SCB

[201] Chia-Hui Huang, Yi-Hau Chen, Jinn-Li Wang, and Mey Wang. Semiparametric copula-based analysis for treatment effects in the presence of treatment switching. Statistics in Medicine, 39(22):2936–2948, September 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Johansen:2020:RMU

[202] Martin Nygård Johansen, Søren Lundbye-Christensen, and Erik Thorlund Parner. Regression models using parametric pseudo-observations. Statistics in Medicine, 39(22):2949–2961, September 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kristensen:2020:BLR

[203] Simon Bang Kristensen and Bo Martin Bibby. A bivariate logistic regression model based on latent variables. Statistics in Medicine, 39(22): 2962–2979, September 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sande:2020:SID

[204] Sumaiya Z. Sande, Jialiang Li, Ralph D’Agostino, Tien Yin Wong, and Ching-Yu Cheng. Statistical inference for decision curve analysis, with applications to cataract diagnosis. Statistics in Medicine, 39(22):2980– 3002, September 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2020:EUC

[205] Hui-Jie Lee, John B. Wong, Beilin Jia, Xinyue Qi, and Elizabeth R. DeLong. Empirical use of causal inference methods to evaluate survival differences in a real-world registry vs those found in randomized clinical trials. Statistics in Medicine, 39(22):3003–3021, September 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shan:2020:CEI

[206] Guogen Shan, Hua Zhang, and Tao Jiang. Comments on “Exact inference for the random-effect model for meta-analyses with rare events”. Statistics in Medicine, 39(22):3022–3023, September 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [30].

Gronsbell:2020:RLG

[207] Jessica Gronsbell and Lu Tian. Response to the letter by Guogen Shan, Hua Zhang, and Tao Jiang. Statistics in Medicine, 39(22):3024–3025, September 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thomadakis:2020:MCS

[208] Christos Thomadakis, Loukia Meligkotsidou, Nikos Pantazis, and Giota Touloumi. Misspecifying the covariance structure in a linear mixed model under MAR drop-out. Statistics in Medicine, 39(23):3027–3041, October 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2020:SGF

[209] Jing Li, Ying Zhang, Giorgos Bakoyannis, and Sujuan Gao. On shared gamma-frailty conditional Markov model for semicompeting risks data. Statistics in Medicine, 39(23):3042–3058, October 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gruber:2020:UEH

[210] Susan Gruber, Douglas Krakower, John T. Menchaca, Katherine Hsu, Rebecca Hawrusik, Judith C. Maro, Noelle M. Cocoros, Benjamin A.

Kruskal, Ira B. Wilson, Kenneth H. Mayer, and Michael Klompas. Using electronic health records to identify candidates for human immunodeficiency virus pre-exposure prophylaxis: an application of super learning to risk prediction when the outcome is rare. Statistics in Medicine, 39 (23):3059–3073, October 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shams:2020:DSS

[211] Shahriar Shams and Eleanor Pullenayegum. Design and sample size considerations for valuation studies of multi-attribute utility instruments. Statistics in Medicine, 39(23):3074–3104, October 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Siegel:2020:BMM

[212] Lianne Siegel, Kyle Rudser, Siobhan Sutcliffe, Alayne Markland, Linda Brubaker, Sheila Gahagan, Ann E. Stapleton, and Haitao Chu. A Bayesian multivariate meta-analysis of prevalence data. Statistics in Medicine, 39(23):3105–3119, October 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wu:2020:VSH

[213] Qiwei Wu, Hui Zhao, Liang Zhu, and Jianguo Sun. Variable selection for high-dimensional partly linear additive Cox model with application to Alzheimer’s disease. Statistics in Medicine, 39(23):3120–3134, October 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Robertson:2020:GAC

[214] David S. Robertson, James M. S. Wason, and Frank Bretz. Graphical approaches for the control of generalized error rates. Statistics in Medicine, 39(23):3135–3155, October 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tang:2020:CIP

[215] Thai-Son Tang, Peter C. Austin, Keith A. Lawson, Antonio Finelli, and Olli Saarela. Constructing inverse probability weights for institutional comparisons in healthcare. Statistics in Medicine, 39(23):3156–3172, October 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2020:CAS

[216] Ying-Ying Zhang, Teng-Zhong Rong, and Man-Man Li. The contemplated average success probability for normally distributed models with an application to optimal sample sizes selection. Statistics in Medicine, 39(23):3173–3183, October 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Al-Wahsh:2020:BAP

[217] Huda Al-Wahsh and Abdulkadir Hussein. A bivariate autoregressive Poisson model and its application to asthma-related emergency room visits. Statistics in Medicine, 39(23):3184–3194, October 15, 2020. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2020:MJM

[218] Hongbin Zhang, Elizabeth A. Kelvin, Arturo Carpio, and W. Allen Hauser. A multistate joint model for interval-censored event-history data subject to within-unit clustering and informative missingness, with application to neurocysticercosis research. Statistics in Medicine, 39(23):3195– 3206, October 15, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Palazon-Bru:2020:GPH

[219] Antonio Palazón-Bru, Francisco Martín-Pérez, Emma Mares-García, Concepción Beneyto-Ripoll, Vicente Francisco Gil-Guillén, Ángel Pérez-Sempere, and María Ángeles Carbonell-Torregrosa. A general presentation on how to carry out a CHARMS analysis for prognostic multivariate models. Statistics in Medicine, 39(23):3207–3225, October 15, 2020. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chattopadhyay:2020:BVM

[220] Ambarish Chattopadhyay, Christopher H. Hase, and José R. Zubizarreta. Balancing vs modeling approaches to weighting in practice. Statistics in Medicine, 39(24):3227–3254, October 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Molebatsi:2020:EEH

[221] Kesaobaka Molebatsi, Lesego Gabaitiri, Lucky Mokgatlhe, Sikhulile Moyo, Simani Gaseitsiwe, Kathleen E. Wirth, Victor DeGruttola, and Eric Tchetgen Tchetgen. Efficient estimation of human immunodeficiency virus incidence rate using a pooled cross-sectional cohort study design. Statistics in Medicine, 39(24):3255–3271, October 30, 2020. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Leao:2020:NCR

[222] Jeremias Leão, Marcelo Bourguignon, Diego I. Gallardo, Ricardo Rocha, and Vera Tomazella. A new cure rate model with flexible competing causes with applications to melanoma and transplantation data. Statistics in Medicine, 39(24):3272–3284, October 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cherlin:2020:DTH

[223] Svetlana Cherlin and James M. S. Wason. Developing and testing high-efficacy patient subgroups within a clinical trial using risk scores. Statistics in Medicine, 39(24):3285–3298, October 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2020:MLR

[224] Ching-Yun Wang and Li Hsu. Multinomial logistic regression with missing outcome data: an application to cancer subtypes. Statistics in Medicine, 39(24):3299–3312, October 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ip:2020:LPO

[225] Edward H. Ip, Shyh-Huei Chen, Karen Bandeen-Roche, Jaime L. Speiser, Li Cai, and Denise K. Houston. Longitudinal partially ordered data analysis for preclinical sarcopenia. Statistics in Medicine, 39(24):3313– 3328, October 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Waddingham:2020:ERB

[226] Ed Waddingham, Paul M. Matthews, and Deborah Ashby. Exploiting relationships between outcomes in Bayesian multivariate network metaanalysis with an application to relapsing-remitting multiple sclerosis. Statistics in Medicine, 39(24):3329–3346, October 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2020:SSC

[227] Jingxia Liu and Graham A. Colditz. Sample size calculation in three-level cluster randomized trials using generalized estimating equation models. Statistics in Medicine, 39(24):3347–3372, October 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Beyene:2020:STD

[228] Kassu Mehari Beyene and Anouar El Ghouch. Smoothed time-dependent receiver operating characteristic curve for right censored survival data. Statistics in Medicine, 39(24):3373–3396, October 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Huang:2020:SAT

[229] Rong Huang, Ronghui Xu, and Parambir S. Dulai. Sensitivity analysis of treatment effect to unmeasured confounding in observational studies with survival and competing risks outcomes. Statistics in Medicine, 39 (24):3397–3411, October 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Meisner:2020:DBC

[230] Allison Meisner, Chirag R. Parikh, and Kathleen F. Kerr. Developing biomarker combinations in multicenter studies via direct maximization and penalization. Statistics in Medicine, 39(24):3412–3426, October 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tang:2020:SCI

[231] Yongqiang Tang. Score confidence intervals and sample sizes for stratified comparisons of binomial proportions. Statistics in Medicine, 39 (24):3427–3457, October 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jin:2020:BMA

[232] Jin Jin, Marie-Karelle Riviere, Xiaodong Luo, and Yingwen Dong. Bayesian methods for the analysis of early-phase oncology basket trials with information borrowing across cancer types. Statistics in Medicine, 39(25):3459–3475, November 10, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ma:2020:FBC

[233] Zichen Ma, Timothy E. Hanson, and Yen-Yi Ho. Flexible bivariate correlated count data regression. Statistics in Medicine, 39(25):3476–3490, November 10, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kotinkaduwa:2020:SME

[234] Lak N. Kotinkaduwa and Pankaj K. Choudhary. A segmented measurement error model for modeling and analysis of method comparison data. Statistics in Medicine, 39(25):3491–3502, November 10, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dong:2020:APW

[235] Lin Dong, Eric Laber, Yair Goldberg, Rui Song, and Shu Yang. Ascertaining properties of weighting in the estimation of optimal treatment regimes under monotone missingness. Statistics in Medicine, 39(25): 3503–3520, November 10, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ellis:2020:ALE

[236] Alexandra G. Ellis, Rowan Iskandar, Christopher H. Schmid, John B. Wong, and Thomas A. Trikalinos. Active learning for efficiently training emulators of computationally expensive mathematical models. Statistics in Medicine, 39(25):3521–3548, November 10, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Engebretsen:2020:PLM

[237] Solveig Engebretsen and Ingrid K. Glad. Partially linear monotone methods with automatic variable selection and monotonicity direction discovery. Statistics in Medicine, 39(25):3549–3568, November 10, 2020. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Soper:2020:HMM

[238] Braden C. Soper, Mari Nygård, Ghaleb Abdulla, Rui Meng, and Jan F. Nygård. A hidden Markov model for population-level cervical cancer screening data. Statistics in Medicine, 39(25):3569–3590, November 10, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hoogland:2020:HMP

[239] Jeroen Hoogland, Marit van Barreveld, Thomas P. A. Debray, Johannes B. Reitsma, Tom E. Verstraelen, Marcel G. W. Dijkgraaf, and Aeilko H. Zwinderman. Handling missing predictor values when validating and applying a prediction model to new patients. Statistics in Medicine, 39(25):3591–3607, November 10, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2020:SSC

[240] Jianghao Li and Sin-Ho Jung. Sample size calculation for cluster randomization trials with a time-to-event endpoint. Statistics in Medicine, 39(25):3608–3623, November 10, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Joseph:2020:TGP

[241] Anny-Claude Joseph, Catherine A. Calder, and David C. Wheeler. The timing of geographic power. Statistics in Medicine, 39(25):3624–3636, November 10, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2020:PED

[242] Xiao Zhang. Parameter-expanded data augmentation for analyzing correlated binary data using multivariate probit models. Statistics in Medicine, 39(25):3637–3652, November 10, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shi:2020:TSR

[243] Yu Shi, Briana Cameron, Xian Gu, Michael Kane, Peter Peduzzi, and Denise A. Esserman. Two-stage randomized trial design for testing treatment, preference, and self-selection effects for count outcomes. Statistics in Medicine, 39(25):3653–3683, November 10, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhao:2020:ILB

[244] Lili Zhao, Susan Murray, Laura H. Mariani, and Wenjun Ju. Incorporating longitudinal biomarkers for dynamic risk prediction in the era of big data: a pseudo-observation approach. Statistics in Medicine, 39(26): 3685–3699, November 20, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2020:GAS

[245] Qihuang Zhang and Grace Y. Yi. Genetic association studies with bivariate mixed responses subject to measurement error and misclassification. Statistics in Medicine, 39(26):3700–3719, November 20, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Potter:2020:DFI

[246] Gail E. Potter. Dismantling the Fragility Index: a demonstration of statistical reasoning. Statistics in Medicine, 39(26):3720–3731, November 20, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Spicker:2020:MEP

[247] Dylan Spicker and Michael P. Wallace. Measurement error and precision medicine: Error-prone tailoring covariates in dynamic treatment regimes. Statistics in Medicine, 39(26):3732–3755, November 20, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hsu:2020:MIB

[248] Chiu-Hsieh Hsu, Yulei He, Chengcheng Hu, and Wei Zhou. A multiple imputation-based sensitivity analysis approach for data subject to missing not at random. Statistics in Medicine, 39(26):3756–3771, November 20, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lu:2020:ICP

[249] Zeng-Hua Lu. An improved closed procedure for testing multiple hypotheses. Statistics in Medicine, 39(26):3772–3786, November 20, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gu:2020:SEC

[250] Ennan Gu, Jiajia Zhang, Wenbin Lu, Lianming Wang, and Federico Felizzi. Semiparametric estimation of the cure fraction in populationbased cancer survival analysis. Statistics in Medicine, 39(26):3787–3805, November 20, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2020:UTI

[251] Chian Chen and Chin-Fu Hsiao. Use of tolerance intervals for assessing biosimilarity. Statistics in Medicine, 39(26):3806–3822, November 20, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hu:2020:GSD

[252] Junxiao Hu, Patrick J. Blatchford, Neil A. Goldenberg, and John M. Kittelson. Group sequential designs for clinical trials with bivariate endpoints. Statistics in Medicine, 39(26):3823–3839, November 20, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Petrof:2020:IRH

[253] Oana Petrof, Thomas Neyens, Valerie Nuyts, Kristiaan Nackaerts, Benoit Nemery, and Christel Faes. On the impact of residential history in the spatial analysis of diseases with a long latency period: a study of mesothelioma in Belgium. Statistics in Medicine, 39(26):3840–3866, November 20, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Meyvisch:2020:RBA

[254] Paul Meyvisch, Ariel Alonso, Wim Van der Elst, and Geert Molenberghs. On the relationship between association and surrogacy when both the surrogate and true endpoint are binary outcomes. Statistics in Medicine, 39(26):3867–3878, November 20, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2020:RMS

[255] Chenyang Zhang, Yuanshan Wu, and Guosheng Yin. Restricted mean survival time for interval-censored data. Statistics in Medicine, 39(26): 3879–3895, November 20, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Arbet:2020:RUF

[256] Jaron Arbet, Matt McGue, and Saonli Basu. A robust and unified framework for estimating heritability in twin studies using generalized estimating equations. Statistics in Medicine, 39(27):3897–3913, November 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xu:2020:DIO

[257] Zhenzhen Xu, Bin Zhu, and Yongsoek Park. Design for immuno-oncology clinical trials enrolling both responders and nonresponders. Statistics in Medicine, 39(27):3914–3936, November 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wu:2020:TVP

[258] Yougui Wu. True verification probabilities should not be used in estimating the area under receiver operating characteristic curve. Statistics in Medicine, 39(27):3937–3946, November 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lazarevic:2020:PVF

[259] Nina Lazarevic, Luke D. Knibbs, Peter D. Sly, and Adrian G. Barnett. Performance of variable and function selection methods for estimating the nonlinear health effects of correlated chemical mixtures: a simulation study. Statistics in Medicine, 39(27):3947–3967, November 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mutze:2020:BCI

[260] Tobias Mütze, Susanna Salem, Norbert Benda, Heinz Schmidli, and Tim Friede. Blinded continuous information monitoring of recurrent event endpoints with time trends in clinical trials. Statistics in Medicine, 39 (27):3968–3985, November 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gunhan:2020:BTE

[261] Burak Kürsad Günhan, Sebastian Weber, and Tim Friede. A Bayesian time-to-event pharmacokinetic model for phase I dose-escalation trials with multiple schedules. Statistics in Medicine, 39(27):3986–4000, November 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Matsouaka:2020:RRC

[262] Roland A. Matsouaka and Folefac D. Atem. Regression with a rightcensored predictor using inverse probability weighting methods. Statistics in Medicine, 39(27):4001–4015, November 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bebu:2020:MIM

[263] Ionut Bebu, Barbara H. Braffett, and John M. Lachin. The minimum intensity of a mixed exposure that increases the risk of an outcome. Statistics in Medicine, 39(27):4016–4024, November 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lu:2020:MIS

[264] Kaifeng Lu. Multiple imputation score tests and an application to Cochran–Mantel–Haenszel statistics. Statistics in Medicine, 39(27): 4025–4036, November 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2020:IPF

[265] Dateng Li, Song Zhang, and Jing Cao. Incorporating pragmatic features into power analysis for cluster randomized trials with a count outcome. Statistics in Medicine, 39(27):4037–4050, November 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Huang:2020:SWD

[266] Yen-Tsung Huang, An-Shun Tai, Meng-Ying Chou, Geng-Xian Lin, and Sheng-Hsuan Lin. Six-way decomposition of causal effects: Unifying mediation and mechanistic interaction. Statistics in Medicine, 39(27): 4051–4068, November 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Guerra:2020:IDT

[267] Steve Ferreira Guerra, Mireille E. Schnitzer, Amélie Forget, and Lucie Blais. Impact of discretization of the timeline for longitudinal causal inference methods. Statistics in Medicine, 39(27):4069–4085, November 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2020:SMC

[268] Chyong-Mei Chen, Pao sheng Shen, Chih-Ching Lin, and Chih-Cheng Wu. Semiparametric mixture cure model analysis with competing risks data: Application to vascular access thrombosis data. Statistics in Medicine, 39(27):4086–4099, November 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2020:CCA

[269] Xiang Li, Kevin Liu, and Hong Tian. Comments on “Critical aspects of the Bayesian approach to phase I cancer trials”. Statistics in Medicine, 39(27):4100, November 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kitchenham:2020:IFS

[270] Barbara Kitchenham and Lech Madeyski. Inconsistencies with formulas for the standard error of the standardized mean difference of repeated measures experiments. Statistics in Medicine, 39(27):4101–4104, November 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Conner:2020:CAR

[271] Sarah C. Conner and Ludovic Trinquart. Correction: Adjusted restricted mean survival times in observational studies. Statistics in Medicine, 39 (27):4105, November 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [7].

Chen:2020:SIS

[272] Yuan Chen, Yuanjia Wang, and Donglin Zeng. Synthesizing independent stagewise trials for optimal dynamic treatment regimes. Statistics in Medicine, 39(28):4107–4119, December 10, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

vanOudenhoven:2020:MEO

[273] Floor M. van Oudenhoven, Sophie H. N. Swinkels, Joseph G. Ibrahim, and Dimitris Rizopoulos. A marginal estimate for the overall treatment effect on a survival outcome within the joint modeling framework. Statistics in Medicine, 39(28):4120–4132, December 10, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2020:NMV

[274] Pingye Zhang, Junshui Ma, Xinqun Chen, and Yue Shentu. A nonparametric method for value function guided subgroup identification via gradient tree boosting for censored survival data. Statistics in Medicine, 39(28):4133–4146, December 10, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2020:VFM

[275] Pengyue Zhang, Abigail Shoben, Rebecca Jackson, and Soledad Fernandez. Variance formulae for multiphase stepped wedge cluster randomized trial. Statistics in Medicine, 39(28):4147–4168, December 10, 2020. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hazlett:2020:IRI

[276] Chad Hazlett, Werner Maokola, and David Ami Wulf. Inference without randomization or ignorability: a stability-controlled quasi-experiment on the prevention of tuberculosis. Statistics in Medicine, 39(28):4169–4186, December 10, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tang:2020:SGA

[277] Yannan Tang, Veronica M. Vieira, Scott M. Bartell, and Daniel L. Gillen. A stratified generalized additive model and permutation test for temporal heterogeneity of smoothed bivariate spatial effects. Statistics in Medicine, 39(28):4187–4200, December 10, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vishwakarma:2020:CAH

[278] Gajendra K. Vishwakarma, Atanu Bhattacharjee, Souvik Banerjee, and Benoit Liquet. Classification algorithm for high-dimensional protein markers in time-course data. Statistics in Medicine, 39(28):4201–4217, December 10, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yang:2020:SSR

[279] Siyun Yang, Fan Li, Monique A. Starks, Adrian F. Hernandez, Robert J. Mentz, and Kingshuk R. Choudhury. Sample size requirements for detecting treatment effect heterogeneity in cluster randomized trials. Statistics in Medicine, 39(28):4218–4237, December 10, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xie:2020:MRE

[280] Yuying Xie, Cecilia Cotton, and Yeying Zhu. Multiply robust estimation of causal quantile treatment effects. Statistics in Medicine, 39(28):4238– 4251, December 10, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Escarela:2020:CMR

[281] Gabriel Escarela, Carlos Erwin Rodríguez, and Gabriel Núñez-Antonio. Copula modeling of receiver operating characteristic and predictiveness curves. Statistics in Medicine, 39(28):4252–4266, December 10, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Montgomery:2020:PBT

[282] Robert N. Montgomery and Jonathan D. Mahnken. A prediction-based test for multiple endpoints. Statistics in Medicine, 39(28):4267–4280, December 10, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2020:EVE

[283] Qing Wang and Alexandria Guo. An efficient variance estimator of AUC and its applications to binary classification. Statistics in Medicine, 39 (28):4281–4300, December 10, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2020:GFM

[284] Hongmei Liu and J. Sunil Rao. Generalized finite mixture of multivariate regressions with applications to therapeutic biomarker identification. Statistics in Medicine, 39(28):4301–4324, December 10, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhao:2020:SIM

[285] Yang Zhao. Statistical inference for missing data mechanisms. Statistics in Medicine, 39(28):4325–4333, December 10, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Francq:2020:TAT

[286] Bernard G. Francq, Marion Berger, and Charles Boachie. To tolerate or to agree: a tutorial on tolerance intervals in method comparison studies with BivRegBLS R package. Statistics in Medicine, 39(28):4334–4349, December 10, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Reist:2020:MAE

[287] Benjamin M. Reist and Richard Valliant. Model-assisted estimators for time-to-event data from complex surveys. Statistics in Medicine, 39(29): 4351–4371, December 20, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wen:2020:DTS

[288] Chi-Chung Wen and Yi-Hau Chen. Discrete-time survival data with longitudinal covariates. Statistics in Medicine, 39(29):4372–4385, December 20, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Choi:2020:BTB

[289] Byeong Yeob Choi, Jason P. Fine, and M. Alan Brookhart. Bias testing, bias correction, and confounder selection using an instrumental variable model. Statistics in Medicine, 39(29):4386–4404, December 20, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Han:2020:SAU

[290] Yongli Han, Paul S. Albert, Christine D. Berg, Nicolas Wentzensen, Hormuzd A. Katki, and Danping Liu. Statistical approaches using longitudinal biomarkers for disease early detection: a comparison of methodologies. Statistics in Medicine, 39(29):4405–4420, December 20, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Beyaztas:2020:REL

[291] Beste Hamiye Beyaztas and Soutir Bandyopadhyay. Robust estimation for linear panel data models. Statistics in Medicine, 39(29):4421–4438, December 20, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Guo:2020:BBR

[292] Beibei Guo and Yong Zang. BILITE: a Bayesian randomized phase II design for immunotherapy by jointly modeling the longitudinal immune response and time-to-event efficacy. Statistics in Medicine, 39(29):4439– 4451, December 20, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jin:2020:BET

[293] Huaqing Jin and Guosheng Yin. Bayesian enhancement two-stage design with error control for phase II clinical trials. Statistics in Medicine, 39 (29):4452–4465, December 20, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Simoneau:2020:FSV

[294] Gabrielle Simoneau, Erica E. M. Moodie, Jagtar S. Nijjar, and Robert W. Platt. Finite sample variance estimation for optimal dynamic treatment regimes of survival outcomes. Statistics in Medicine, 39(29):4466–4479, December 20, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wu:2020:PIC

[295] Qin Wu, Man-Lai Tang, Derrick Wing-Hong Fung, and Guo-Liang Tian. Poisson item count techniques with noncompliance. Statistics in Medicine, 39(29):4480–4498, December 20, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Senarathne:2020:BRB

[296] Siththara Gedara J. Senarathne, Antony M. Overstall, and James M. McGree. Book review: Bayesian adaptive N-of-1 trials for estimating population and individual treatment effects. Statistics in Medicine, 39 (29):4499–4518, December 20, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ozturk:2020:MAQ

[297] Omer Ozturk and Narayanaswamy Balakrishnan. Meta-analysis of quantile intervals from different studies with an application to a pulmonary tuberculosis data. Statistics in Medicine, 39(29):4519–4537, December 20, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schnitzer:2020:TDT

[298] Mireille E. Schnitzer, Robert W. Platt, and Madeleine Durand. A tutorial on dealing with time-varying eligibility for treatment: Comparing the risk of major bleeding with direct-acting oral anticoagulants vs warfarin. Statistics in Medicine, 39(29):4538–4550, December 20, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Grill:2020:ACC

[299] Susann Grill, Arne Ring, Werner Brannath, and Martin Scharpenberg. Assessing consistency in clinical trials with two subgroups and binary endpoints: a new test within the logistic regression model. Statistics in Medicine, 39(30):4551–4573, December 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gueorguieva:2020:TPM

[300] Ralitza Gueorguieva, Eugenia Buta, Meghan Morean, and Suchitra Krishnan-Sarin. Two-part models for repeatedly measured ordinal data with “don’t know” category. Statistics in Medicine, 39(30):4574–4592, December 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Qu:2020:UAS

[301] Yongming Qu, Yu Du, Ying Zhang, and Lei Shen. Understanding and adjusting for the selection bias from a proof-of-concept study to a more confirmatory study. Statistics in Medicine, 39(30):4593–4604, December 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sun:2020:GWA

[302] Tao Sun, Yue Wei, Wei Chen, and Ying Ding. Genome-wide association study-based deep learning for survival prediction. Statistics in Medicine, 39(30):4605–4620, December 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stevens:2020:CKM

[303] Nathaniel T. Stevens and Lu Lu. Comparing Kaplan–Meier curves with the probability of agreement. Statistics in Medicine, 39(30):4621–4635, December 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Heo:2020:EMI

[304] Seok-Jae Heo and Inkyung Jung. Extended multi-item gamma Poisson shrinker methods based on the zero-inflated Poisson model for postmarket drug safety surveillance. Statistics in Medicine, 39(30):4636–4650, December 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Braun:2020:SFA

[305] Thomas M. Braun. A simulation-free approach to assessing the performance of the continual reassessment method. Statistics in Medicine, 39 (30):4651–4666, December 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rychtar:2020:ESV

[306] Jan Rychtár and Dewey T. Taylor. Estimating the sample variance from the sample size and range. Statistics in Medicine, 39(30):4667–4686, December 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xue:2020:MDW

[307] Xiaonan Xue, Qibin Qi, Daniela Sotres-Alvarez, Scott C. Roesch, Maria M. Llabre, Sierra A. Bainter, Yasmin Mossavar-Rahmani, Robert Kaplan, and Tao Wang. Modeling daily and weekly moderate and vigorous physical activity using zero-inflated mixture Poisson distribution. Statistics in Medicine, 39(30):4687–4703, December 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Eloyan:2020:THE

[308] Ani Eloyan, Mun Sang Yue, and Davit Khachatryan. Tumor heterogeneity estimation for radiomics in cancer. Statistics in Medicine, 39(30): 4704–4723, December 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mehrotra:2020:BRS

[309] Devan V. Mehrotra and Rachel Marceau West. Book review: Survival analysis using a 5-step stratified testing and amalgamation routine (5-STAR) in randomized clinical trials. Statistics in Medicine, 39(30):4724– 4744, December 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Castelletti:2020:BLM

[310] Federico Castelletti, Luca La Rocca, Stefano Peluso, Francesco C. Stingo, and Guido Consonni. Bayesian learning of multiple directed networks from observational data. Statistics in Medicine, 39(30):4745–4766, December 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

MacNab:2020:BEM

[311] Ying C. MacNab. Bayesian estimation of multivariate Gaussian Markov random fields with constraint. Statistics in Medicine, 39(30):4767–4788, December 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hai:2020:DEA

[312] Yan Hai and Gengsheng Qin. Direct estimation of the area under the receiver operating characteristic curve with verification biased data. Statistics in Medicine, 39(30):4789–4820, December 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Aikens:2020:PDO

[313] Rachael C. Aikens, Dylan Greaves, and Michael Baiocchi. A pilot design for observational studies: Using abundant data thoughtfully. Statistics in Medicine, 39(30):4821–4840, December 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mitra:2020:BDT

[314] Riten Mitra, Peter Müller, and Arinjita Bhattacharyya. Bayesian decision-theoretic methods for survival data using stochastic optimization. Statistics in Medicine, 39(30):4841–4852, December 30, 2020. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kerioui:2020:BIU

[315] Marion Kerioui, Francois Mercier, Julie Bertrand, Coralie Tardivon, René Bruno, Jérémie Guedj, and Solène Desmée. Bayesian inference using Hamiltonian Monte-Carlo algorithm for nonlinear joint modeling in the context of cancer immunotherapy. Statistics in Medicine, 39(30): 4853–4868, December 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

He:2020:HDI

[316] Yong He, Hao Chen, Hao Sun, Jiadong Ji, Yufeng Shi, Xinsheng Zhang, and Lei Liu. High-dimensional integrative copula discriminant analysis for multiomics data. Statistics in Medicine, 39(30):4869–4884, December 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Phillippo:2020:APP

[317] David M. Phillippo, Sofia Dias, A. E. Ades, and Nicky J. Welton. Assessing the performance of population adjustment methods for anchored indirect comparisons: a simulation study. Statistics in Medicine, 39(30): 4885–4911, December 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See comments [495, 496].

Chen:2020:OMS

[318] Tong Chen and Thomas Lumley. Optimal multiwave sampling for regression modeling in two-phase designs. Statistics in Medicine, 39(30): 4912–4921, December 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Goetghebeur:2020:FCQ

[319] Els Goetghebeur, Saskia le Cessie, Bianca De Stavola, Erica Em Moodie, Ingeborg Waernbaum, and “On behalf of” the topic group Causal Inference (Tg7) of the Stratos initiative”. Formulating causal questions and principled statistical answers. Statistics in Medicine, 39(30):4922–4948, December 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Day:2021:E

[320] Simon Day, Els Goetghebeur, Joel Greenhouse, and Robert Platt. Editorial. Statistics in Medicine, 40(1):1–2, January 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hill:2021:AWM

[321] Austin Bradford Hill. Alfred Watson Memorial Lecture: The statistician in medicine. Statistics in Medicine, 40(1):3–16, January 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See commentary [323, 324, 322, 323, 324, 325, 326, 327, 328, 329, 330, 331].

Farewell:2021:MSA

[322] Vern Farewell and Tony Johnson. Medical statistics, Austin Bradford Hill, and a celebration of 40 years of Statistics in Medicine. Statistics in Medicine, 40(1):17–28, January 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hubbard:2021:CPA

[323] Rebecca A. Hubbard. Commentary on Professor Austin Bradford Hill’s Alfred Watson Memorial Lecture. Statistics in Medicine, 40(1):29–31, January 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [321].

Lin:2021:RPS

[324] Xihong Lin. Reflections on the 1962 paper “The Statistician in Medicine” by Sir Austin Bradford Hill. Statistics in Medicine, 40(1):32–34, January 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [321].

Louis:2021:DSM

[325] Thomas A. Louis. Discussion of The Statistician in Medicine, by Professor Sir Austin Bradford Hill. Statistics in Medicine, 40(1):35–36, January 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [321].

Moodie:2021:CSM

[326] Erica E. M. Moodie and David A. Stephens. Commentary on “The Statistician in Medicine” by Professor Sir Austin Bradford Hill. Statistics in Medicine, 40(1):37–41, January 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [321].

Dempsey:2021:RSM

[327] Walter Dempsey and Bhramar Mukherjee. Reflecting on “A Statistician in Medicine” in 2020. Statistics in Medicine, 40(1):42–48, January 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [321].

Phillips:2021:BHW

[328] Christopher J. Phillips. A. B. Hill’s 1962 Watson lecture: The statistical consultant as consensus maker. Statistics in Medicine, 40(1):49–51, January 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [321].

Ryan:2021:CSM

[329] Louise M. Ryan. Comments on “The Statistician in Medicine” by Austin Bradford Hill. Statistics in Medicine, 40(1):52–54, January 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [321].

Wittes:2021:TBH

[330] Janet T. Wittes. Thoughts on A. B. Hill’s Watson Lecture. Statistics in Medicine, 40(1):55–57, January 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [321].

Zhang:2021:MAS

[331] Jeffrey Zhang, Bo Zhang, and Dylan S. Small. A method to aid statistical judgment on outliers: Comment on Hill’s The Statistician in Medicine. Statistics in Medicine, 40(1):58–63, January 15, 2021. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [321].

Wang:2021:ACD

[332] Xi Wang and Vernon M. Chinchilli. Analysis of crossover designs with nonignorable dropout. Statistics in Medicine, 40(1):64–84, January 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rava:2021:EVU

[333] Denise Rava and Ronghui Xu. Explained variation under the additive hazards model. Statistics in Medicine, 40(1):85–100, January 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Austin:2021:ECM

[334] Peter C. Austin and Elizabeth A. Stuart. The effect of a constraint on the maximum number of controls matched to each treated subject on the performance of full matching on the propensity score when estimating risk differences. Statistics in Medicine, 40(1):101–118, January 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yan:2021:KDB

[335] Feifei Yan, Qing-Song Xu, Man-Lai Tang, and Ziqi Chen. Kernel densitybased likelihood ratio tests for linear regression models. Statistics in Medicine, 40(1):119–132, January 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Archer:2021:MSS

[336] Lucinda Archer, Kym I. E. Snell, Joie Ensor, Mohammed T. Hudda, Gary S. Collins, and Richard D. Riley. Minimum sample size for external validation of a clinical prediction model with a continuous outcome. Statistics in Medicine, 40(1):133–146, January 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Piulachs:2021:BJM

[337] Xavier Piulachs, Eleni-Rosalina Andrinopoulou, Montserrat Guillén, and Dimitris Rizopoulos. A Bayesian joint model for zero-inflated integers and left-truncated event times with a time-varying association: Applications to senior health care. Statistics in Medicine, 40(1):147–166, January 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Margaritella:2021:PCB

[338] Nicolò Margaritella, Vanda Inácio, and Ruth King. Parameter clustering in Bayesian functional principal component analysis of neuroscientific data. Statistics in Medicine, 40(1):167–184, January 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Andersen:2021:ATE

[339] Per Kragh Andersen, Maja Pohar Perme, Hans C. van Houwelingen, Richard J. Cook, Pierre Joly, Torben Martinussen, Jeremy M. G. Taylor, Michal Abrahamowicz, and Terry M. Therneau. Analysis of time-to-event for observational studies: Guidance to the use of intensity models. Statistics in Medicine, 40(1):185–211, January 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vives-Mestres:2021:MVT

[340] Marina Vives-Mestres and Amparo Casanova. Modeling and visualizing two-way contingency tables using compositional data analysis: a case-study on individual self-prediction of migraine days. Statistics in Medicine, 40(2):213–225, January 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Imai:2021:EAT

[341] Takumi Imai, Shiro Tanaka, and Koji Kawakami. Exploratory assessment of treatment-dependent random-effects distribution using gradient functions. Statistics in Medicine, 40(2):226–239, January 30, 2021. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dehbi:2021:EPD

[342] Hakim-Moulay Dehbi, David M. Lowe, and John O’Quigley. Early phase dose-finding trials in virology. Statistics in Medicine, 40(2):240–253, January 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhong:2021:SME

[343] Yujie Zhong and Richard J. Cook. Selection models for efficient twophase design of family studies. Statistics in Medicine, 40(2):254–270, January 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shaw:2021:RCC

[344] Pamela A. Shaw, Jiwei He, and Bryan E. Shepherd. Regression calibration to correct correlated errors in outcome and exposure. Statistics in Medicine, 40(2):271–286, January 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pak:2021:ALT

[345] Daewoo Pak, Jun Liu, Jing Ning, Guadalupe Gómez Melis, and Yu Shen. Analyzing left-truncated and right-censored infectious disease cohort data with interval-censored infection onset. Statistics in Medicine, 40 (2):287–298, January 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nuno:2021:REN

[346] Michelle M. Nuño and Daniel L. Gillen. Robust estimation in the nested case-control design under a misspecified covariate functional form. Statistics in Medicine, 40(2):299–311, January 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hartman:2021:DAC

[347] Holly Hartman, Roy N. Tamura, Matthew J. Schipper, and Kelley M. Kidwell. Design and analysis considerations for utilizing a mapping function in a small sample, sequential, multiple assignment, randomized trials with continuous outcomes. Statistics in Medicine, 40(2):312–326, January 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Montepiedra:2021:FCR

[348] Grace Montepiedra, Ritesh Ramchandani, Sachiko Miyahara, and Soyeon Kim. A framework for considering the risk-benefit trade-off in designing noninferiority trials using composite outcome approaches. Statistics in Medicine, 40(2):327–348, January 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Roberts:2021:INR

[349] Chris Roberts. The implications of noncompliance for randomized trials with partial nesting due to group treatment. Statistics in Medicine, 40 (2):349–368, January 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wallisch:2021:SVM

[350] Christine Wallisch, Daniela Dunkler, Geraldine Rauch, Riccardo de Bin, and Georg Heinze. Selection of variables for multivariable models: Opportunities and limitations in quantifying model stability by resampling. Statistics in Medicine, 40(2):369–381, January 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mu:2021:BAP

[351] Rongji Mu, Haitao Pan, and Guoying Xu. A Bayesian adaptive phase I/II platform trial design for pediatric immunotherapy trials. Statistics in Medicine, 40(2):382–402, January 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lin:2021:EVE

[352] Lifeng Lin and Ariel M. Aloe. Evaluation of various estimators for standardized mean difference in meta-analysis. Statistics in Medicine, 40(2): 403–426, January 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gory:2021:CGL

[353] Jeffrey J. Gory, Peter F. Craigmile, and Steven N. MacEachern. A class of generalized linear mixed models adjusted for marginal interpretability. Statistics in Medicine, 40(2):427–440, January 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zuo:2021:SBE

[354] Lulu Zuo, Haixiang Zhang, HaiYing Wang, and Lei Liu. Sampling-based estimation for massive survival data with additive hazards model. Statistics in Medicine, 40(2):441–450, January 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chaimani:2021:MCA

[355] Anna Chaimani, Raphaël Porcher, Émilie Sbidian, and Dimitris Mavridis. A Markov chain approach for ranking treatments in network meta-analysis. Statistics in Medicine, 40(2):451–464, January 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2021:CST

[356] Junho Lee, Maria E. Kamenetsky, Ronald E. Gangnon, and Jun Zhu. Clustered spatio-temporal varying coefficient regression model. Statistics in Medicine, 40(2):465–480, January 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pang:2021:SBA

[357] Menglan Pang, Robert W. Platt, Tibor Schuster, and Michal Abrahamowicz. Spline-based accelerated failure time model. Statistics in Medicine, 40(2):481–497, January 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See comment [830].

Martin:2021:CPM

[358] Glen P. Martin, Matthew Sperrin, Kym I. E. Snell, Iain Buchan, and Richard D. Riley. Clinical prediction models to predict the risk of multiple binary outcomes: a comparison of approaches. Statistics in Medicine, 40(2):498–517, January 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Oganisian:2021:PIB

[359] Arman Oganisian and Jason A. Roy. A practical introduction to Bayesian estimation of causal effects: Parametric and nonparametric approaches. Statistics in Medicine, 40(2):518–551, January 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Giai:2021:NBP

[360] Joris Giai, Delphine Maucort-Boulch, Brice Ozenne, Jean-Christophe Chiêm, Marc Buyse, and Julien Péron. Net benefit in the presence of correlated prioritized outcomes using generalized pairwise comparisons: a simulation study. Statistics in Medicine, 40(3):553–565, February 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Aouni:2021:MAI

[361] Jihane Aouni, Nadia Gaudel-Dedieu, and Bernard Sebastien. Matchingadjusted indirect comparisons: Application to time-to-event data. Statistics in Medicine, 40(3):566–577, February 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See comments [844].

Tran:2021:SCS

[362] Trung Dung Tran, Emmanuel Lesaffre, Geert Verbeke, and Geert Molenberghs. Serial correlation structures in latent linear mixed models for analysis of multivariate longitudinal ordinal responses. Statistics in Medicine, 40(3):578–592, February 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yang:2021:LBA

[363] Yunqi Yang, Christine Hong, Jane W. Liang, Stephen Gruber, Giovanni Parmigiani, Gregory Idos, and Danielle Braun. A likelihood-based approach to assessing frequency of pathogenicity among variants of unknown significance in susceptibility genes. Statistics in Medicine, 40(3): 593–606, February 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Loh:2021:CSS

[364] Wen Wei Loh and Stijn Vansteelandt. Confounder selection strategies targeting stable treatment effect estimators. Statistics in Medicine, 40 (3):607–630, February 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Oh:2021:RRC

[365] Eric J. Oh, Bryan E. Shepherd, Thomas Lumley, and Pamela A. Shaw. Raking and regression calibration: Methods to address bias from correlated covariate and time-to-event error. Statistics in Medicine, 40(3): 631–649, February 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shi:2021:FPCa

[366] Bin Shi, Peng Wei, and Xuelin Huang. Functional principal component based landmark analysis for the effects of longitudinal cholesterol profiles on the risk of coronary heart disease. Statistics in Medicine, 40(3):650– 667, February 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fu:2021:LRT

[367] Wenjiang Fu, Junyu Ding, Kuikui Gao, Shuangge Ma, and Lu Tian. A likelihood ratio test on temporal trends in age-period-cohort models with applications to the disparities of heart disease mortality among US populations and comparison with Japan. Statistics in Medicine, 40(3): 668–689, February 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Burnett:2021:AET

[368] Thomas Burnett and Christopher Jennison. Adaptive enrichment trials: What are the benefits? Statistics in Medicine, 40(3):690–711, February 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shi:2021:FPCb

[369] Haolun Shi, Jianghu Dong, Liangliang Wang, and Jiguo Cao. Functional principal component analysis for longitudinal data with informative dropout. Statistics in Medicine, 40(3):712–724, February 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tao:2021:ESI

[370] Ran Tao, Sarah C. Lotspeich, Gustavo Amorim, Pamela A. Shaw, and Bryan E. Shepherd. Efficient semiparametric inference for two-phase studies with outcome and covariate measurement errors. Statistics in Medicine, 40(3):725–738, February 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hu:2021:SAH

[371] Xiangbin Hu, Jian Huang, Li Liu, Defeng Sun, and Xingqiu Zhao. Subgroup analysis in the heterogeneous Cox model. Statistics in Medicine, 40(3):739–757, February 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Guan:2021:MAB

[372] Zhong Guan. Maximum approximate Bernstein likelihood estimation in proportional hazard model for interval-censored data. Statistics in Medicine, 40(3):758–778, February 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zou:2021:NST

[373] Yuhan Zou, Zuoxiang Peng, Jerry Cornell, Peng Ye, and Hua He. A new statistical test for latent class in censored data due to detection limit. Statistics in Medicine, 40(3):779–798, February 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kim:2021:BGA

[374] Seongyoon Kim, Sanghee Lee, Jung-Il Choi, and Hyunsoon Cho. Binary genetic algorithm for optimal joinpoint detection: Application to cancer trend analysis. Statistics in Medicine, 40(3):799–822, February 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wu:2021:PCD

[375] Zhenke Wu and Irena Chen. Probabilistic cause-of-disease assignment using case-control diagnostic tests: a latent variable regression approach. Statistics in Medicine, 40(4):823–841, February 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zeng:2021:PSW

[376] Shuxi Zeng, Fan Li, Rui Wang, and Fan Li. Propensity score weighting for covariate adjustment in randomized clinical trials. Statistics in Medicine, 40(4):842–858, February 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See correction [1613].

Riley:2021:NEC

[377] Richard D. Riley, Ben Van Calster, and Gary S. Collins. A note on estimating the Cox–Snell R^{2} from a reported C statistic (AUROC) to inform sample size calculations for developing a prediction model with a binary outcome. Statistics in Medicine, 40(4):859–864, February 20, 2021. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Veres-Ferrer:2021:EMO

[378] Ernesto J. Veres-Ferrer and Jose M. Pavía. Elasticity as a measure for online determination of remission points in ongoing epidemics. Statistics in Medicine, 40(4):865–884, February 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2021:MES

[379] Haixiang Zhang, Jun Chen, Yang Feng, Chan Wang, Huilin Li, and Lei Liu. Mediation effect selection in high-dimensional and compositional microbiome data. Statistics in Medicine, 40(4):885–896, February 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2021:SNM

[380] Lihui Liu, Hong Gu, Johan Van Limbergen, and Toby Kenney. SuRF: a new method for sparse variable selection, with application in microbiome data analysis. Statistics in Medicine, 40(4):897–919, February 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2021:BLF

[381] Xiaoqing Wang, Xinyuan Song, and Hongtu Zhu. Bayesian latent factor on image regression with nonignorable missing data. Statistics in Medicine, 40(4):920–932, February 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zandkarimi:2021:RIJ

[382] Eghbal Zandkarimi, Abbas Moghimbeigi, and Hossein Mahjub. Robust inference in the joint modeling of multilevel zero-inflated Poisson and Cox models. Statistics in Medicine, 40(4):933–949, February 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2021:CBC

[383] Mei-Cheng Wang and Yuchen Yang. Complexity and bias in crosssectional data with binary disease outcome in observational studies. Statistics in Medicine, 40(4):950–962, February 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fang:2021:BMC

[384] Fang Fang, Kimberly A. Hochstedler, Roy N. Tamura, Thomas M. Braun, and Kelley M. Kidwell. Bayesian methods to compare dose levels with placebo in a small n, sequential, multiple assignment, randomized trial. Statistics in Medicine, 40(4):963–977, February 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2021:DCM

[385] Kuo-Jung Lee, Ray-Bing Chen, Min-Sun Kwak, and Keunbaik Lee. Determination of correlations in multivariate longitudinal data with modified Cholesky and hypersphere decomposition using Bayesian variable selection approach. Statistics in Medicine, 40(4):978–997, February 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Saha-Chaudhuri:2021:SAU

[386] Paramita Saha-Chaudhuri and Lamin Juwara. Survival analysis under the Cox proportional hazards model with pooled covariates. Statistics in Medicine, 40(4):998–1020, February 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Parsaeian:2021:STA

[387] Mahboubeh Parsaeian, Majid Jafari Khaledi, Farshad Farzadfar, Mahdi Mahdavi, Hojjat Zeraati, Mahmood Mahmoudi, Ardeshir Khosravi, and Kazem Mohammad. Spatio-temporal analysis of misaligned burden of disease data using a geo-statistical approach. Statistics in Medicine, 40 (4):1021–1033, February 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wu:2021:ECE

[388] Mixia Wu, Xiaoyu Zhang, Wei Zhang, Xu Zhang, and Aiyi Liu. Errorcorrected estimation of a diagnostic accuracy index of a biomarker against a continuous gold standard. Statistics in Medicine, 40(4):1034– 1058, February 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wu:2021:OTP

[389] Yougui Wu. Optimal two-phase sampling for estimating the area under the receiver operating characteristic curve. Statistics in Medicine, 40 (4):1059–1071, February 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ye:2021:CZA

[390] Tairan Ye, Victor H. Lachos, Xiaojing Wang, and Dipak K. Dey. Comparisons of zero-augmented continuous regression models from a Bayesian perspective. Statistics in Medicine, 40(5):1073–1100, February 28, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ackerman:2021:GRT

[391] Benjamin Ackerman, Catherine R. Lesko, Juned Siddique, Ryoko Susukida, and Elizabeth A. Stuart. Generalizing randomized trial findings to a target population using complex survey population data. Statistics in Medicine, 40(5):1101–1120, February 28, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Martens:2021:UAS

[392] Michael J. Martens and Brent R. Logan. A unified approach to sample size and power determination for testing parameters in generalized linear and time-to-event regression models. Statistics in Medicine, 40(5):1121– 1132, February 28, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Watson:2021:DAT

[393] Samuel I. Watson, Alan Girling, and Karla Hemming. Design and analysis of three-arm parallel cluster randomized trials with small numbers of clusters. Statistics in Medicine, 40(5):1133–1146, February 28, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fay:2021:CID

[394] Michael P. Fay and Keith Lumbard. Confidence intervals for difference in proportions for matched pairs compatible with exact McNemar’s or sign tests. Statistics in Medicine, 40(5):1147–1159, February 28, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

OBrien:2021:SAE

[395] Robert M. O’Brien. A simplified approach for establishing estimable functions in fixed effect age-period-cohort multiple classification models. Statistics in Medicine, 40(5):1160–1171, February 28, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Westerberg:2021:SMD

[396] Marcus Westerberg, Rolf Larsson, Lars Holmberg, Pär Stattin, and Hans Garmo. Simulation model of disease incidence driven by diagnostic activity. Statistics in Medicine, 40(5):1172–1188, February 28, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brown:2021:PSS

[397] Derek W. Brown, Thomas J. Greene, Michael D. Swartz, Anna V. Wilkinson, and Stacia M. DeSantis. Propensity score stratification methods for continuous treatments. Statistics in Medicine, 40(5):1189–1203, February 28, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Garrido:2021:VBK

[398] Melissa M. Garrido, Jessica Lum, and Steven D. Pizer. Vector-based kernel weighting: a simple estimator for improving precision and bias of average treatment effects in multiple treatment settings. Statistics in Medicine, 40(5):1204–1223, February 28, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shu:2021:EMH

[399] Di Shu, Peisong Han, Rui Wang, and Sengwee Toh. Estimating the marginal hazard ratio by simultaneously using a set of propensity score models: a multiply robust approach. Statistics in Medicine, 40(5):1224– 1242, February 28, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tzeng:2021:SPD

[400] ShengLi Tzeng, Jun Zhu, Amy J. Weisman, Tyler J. Bradshaw, and Robert Jeraj. Spatial process decomposition for quantitative imaging biomarkers using multiple images of varying shapes. Statistics in Medicine, 40(5):1243–1261, February 28, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ge:2021:RAM

[401] Lei Ge, Liang Zhu, and Jianguo Sun. Regression analysis of mixed panel count data with informative indicator processes. Statistics in Medicine, 40(5):1262–1271, February 28, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pan:2021:JAM

[402] Deng Pan, Yingying Wei, and Xinyuan Song. Joint analysis of mixed types of outcomes with latent variables. Statistics in Medicine, 40(5): 1272–1284, February 28, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ghosh:2021:RWT

[403] Abhik Ghosh, Ayanendranath Basu, and Leandro Pardo. Robust Waldtype tests under random censoring. Statistics in Medicine, 40(5):1285– 1305, February 28, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Greene:2021:EAB

[404] Erich J. Greene, Peter Peduzzi, James Dziura, Can Meng, Michael E. Miller, Thomas G. Travison, and Denise Esserman. Estimation of ascertainment bias and its effect on power in clinical trials with time-to-event outcomes. Statistics in Medicine, 40(5):1306–1320, February 28, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sommer:2021:AEE

[405] Alice J. Sommer, Emmanuelle Leray, Young Lee, and Marie-Abèle C. Bind. Assessing environmental epidemiology questions in practice with a causal inference pipeline: an investigation of the air pollution-multiple sclerosis relapses relationship. Statistics in Medicine, 40(6):1321–1335, March 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kang:2021:LBM

[406] Tong Kang, Jeremy Gaskins, Steven Levy, and Somnath Datta. A longitudinal Bayesian mixed effects model with hurdle Conway–Maxwell– Poisson distribution. Statistics in Medicine, 40(6):1336–1356, March 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Berentsen:2021:HCL

[407] Geir D. Berentsen, Francesca Azzolini, Hans J. Skaug, Rolv T. Lie, and Håkon K. Gjessing. Heritability curves: a local measure of heritability in family models. Statistics in Medicine, 40(6):1357–1382, March 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Greene:2021:MLC

[408] Thomas J. Greene, Stacia M. DeSantis, Derek W. Brown, Anna V. Wilkinson, and Michael D. Swartz. A machine learning compatible method for ordinal propensity score stratification and matching. Statistics in Medicine, 40(6):1383–1399, March 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Woodall:2021:OCU

[409] William H. Woodall, George Rakovich, and Stefan H. Steiner. An overview and critique of the use of cumulative sum methods with surgical learning curve data. Statistics in Medicine, 40(6):1400–1413, March 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Novick:2021:MCP

[410] Steven Novick and Tianhui Zhang. Mean comparisons and power calculations to ensure reproducibility in preclinical drug discovery. Statistics in Medicine, 40(6):1414–1428, March 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Baker:2021:MMT

[411] Stuart G. Baker. Modeling the mean time to interval cancer after negative results of periodic cancer screening. Statistics in Medicine, 40(6): 1429–1439, March 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2021:LMN

[412] Zheng Wang, Yu Cheng, Eric C. Seaberg, Leah H. Rubin, Andrew J. Levine, James T. Becker, and Neuropsychology Working Group of the Macs. Longitudinal multivariate normative comparisons. Statistics in Medicine, 40(6):1440–1452, March 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2021:BPC

[413] Yiran Zhang and Kellie J. Archer. Bayesian penalized cumulative logit model for high-dimensional data with an ordinal response. Statistics in Medicine, 40(6):1453–1481, March 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2021:MDC

[414] Longhai Li, Tingxuan Wu, and Cindy Feng. Model diagnostics for censored regression via randomized survival probabilities. Statistics in Medicine, 40(6):1482–1497, March 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Baghfalaki:2021:BMA

[415] Taban Baghfalaki, Pierre-Emmanuel Sugier, Therese Truong, Anthony N. Pettitt, Kerrie Mengersen, and Benoit Liquet. Bayesian metaanalysis models for cross cancer genomic investigation of pleiotropic effects using group structure. Statistics in Medicine, 40(6):1498–1518, March 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wu:2021:LPI

[416] Qiong Wu, Zhen Zhang, Tianzhou Ma, James Waltz, Donald Milton, and Shuo Chen. Link predictions for incomplete network data with outcome misclassification. Statistics in Medicine, 40(6):1519–1534, March 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Razaee:2021:PTI

[417] Zahra S. Razaee, Arash A. Amini, Márcio A. Diniz, Mourad Tighiouart, Greg Yothers, and André Rogatko. On the properties of the toxicity index and its statistical efficiency. Statistics in Medicine, 40(6):1535– 1552, March 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Seo:2021:CME

[418] Michael Seo, Ian R. White, Toshi A. Furukawa, Hissei Imai, Marco Valgimigli, Matthias Egger, Marcel Zwahlen, and Orestis Efthimiou. Comparing methods for estimating patient-specific treatment effects in individual patient data meta-analysis. Statistics in Medicine, 40(6):1553– 1573, March 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rizk:2021:AFC

[419] Jean Rizk, Cathal Walsh, and Kevin Burke. An alternative formulation of Coxian phase-type distributions with covariates: Application to emergency department length of stay. Statistics in Medicine, 40(6):1574–1592, March 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Godwin:2021:STM

[420] Jessica Godwin and Jon Wakefield. Space-time modeling of child mortality at the Admin-2 level in a low and middle income countries context. Statistics in Medicine, 40(7):1593–1638, March 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lopez-Pintado:2021:DBG

[421] Sara Lopez-Pintado and Kun Qian. A depth-based global envelope test for comparing two groups of functions with applications to biomedical data. Statistics in Medicine, 40(7):1639–1652, March 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See comment [831].

Yu:2021:CPP

[422] Youfei Yu, Min Zhang, Xu Shi, Megan E. V. Caram, Roderick J. A. Little, and Bhramar Mukherjee. A comparison of parametric propensity score-based methods for causal inference with multiple treatments and a binary outcome. Statistics in Medicine, 40(7):1653–1677, March 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Amiri:2021:SMI

[423] Leila Amiri, Mahmoud Torabi, Rob Deardon, and Michael Pickles. Spatial modeling of individual-level infectious disease transmission: Tuberculosis data in Manitoba, Canada. Statistics in Medicine, 40(7):1678–1704, March 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lin:2021:EMW

[424] Tuo Lin, Tian Chen, Jinyuan Liu, and Xin M. Tu. Extending the Mann– Whitney–Wilcoxon rank sum test to survey data for comparing mean ranks. Statistics in Medicine, 40(7):1705–1717, March 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Webster-Clark:2021:UPS

[425] Michael Webster-Clark, Til Stürmer, Tiansheng Wang, Kenneth Man, Danica Marinac-Dabic, Kenneth J. Rothman, Alan R. Ellis, Mugdha Gokhale, Mark Lunt, Cynthia Girman, and Robert J. Glynn. Using propensity scores to estimate effects of treatment initiation decisions: State of the science. Statistics in Medicine, 40(7):1718–1735, March 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kasza:2021:ICS

[426] Jessica Kasza, Rhys Bowden, and Andrew B. Forbes. Information content of stepped wedge designs with unequal cluster-period sizes in linear mixed models: Informing incomplete designs. Statistics in Medicine, 40(7): 1736–1751, March 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lu:2021:ODC

[427] Yitao Lu, Julie Zhou, Li Xing, and Xuekui Zhang. The optimal design of clinical trials with potential biomarker effects: a novel computational approach. Statistics in Medicine, 40(7):1752–1766, March 30, 2021. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bantis:2021:LRO

[428] Leonidas E. Bantis, John V. Tsimikas, Gregory R. Chambers, Michela Capello, Samir Hanash, and Ziding Feng. The length of the receiver operating characteristic curve and the two cutoff Youden index within a robust framework for discovery, evaluation, and cutoff estimation in biomarker studies involving improper receiver operating characteristic curves. Statistics in Medicine, 40(7):1767–1789, March 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schumacher:2021:SMS

[429] Fernanda L. Schumacher, Victor H. Lachos, and Larissa A. Matos. Scale mixture of skew-normal linear mixed models with within-subject serial dependence. Statistics in Medicine, 40(7):1790–1810, March 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jia:2021:CSQ

[430] Yichen Jia and Jong-Hyeon Jeong. Cause-specific quantile regression on inactivity time. Statistics in Medicine, 40(7):1811–1824, March 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shrestha:2021:FMJ

[431] Sama Shrestha and Sonia Jain. A Bayesian-bandit adaptive design for N-of-1 clinical trials. Statistics in Medicine, 40(7):1825–1844, March 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gupta:2021:FMJ

[432] Resmi Gupta, Jane C. Khoury, Mekibib Altaye, Roman Jandarov, and Rhonda D. Szczesniak. Flexible multivariate joint model of longitudinal intensity and binary process for medical monitoring of frequently collected data. Statistics in Medicine, 40(7):1845–1858, March 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sabour:2021:CPM

[433] Siamak Sabour and Hadis Ghajari. Clinical prediction models to predict the risk of multiple binary outcomes: Methodological issues. Statistics in Medicine, 40(7):1859–1860, March 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See reply [434].

Martin:2021:ARS

[434] Glen Philip Martin, Matthew Sperrin, Kym I. E. Snell, Iain Buchan, and Richard D. Riley. Authors’ reply to Sabour and Ghajari “Clinical prediction models to predict the risk of multiple binary outcomes: Methodological issues”. Statistics in Medicine, 40(7):1861–1862, March 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [433].

Tao:2021:TWT

[435] Ran Tao, Nathaniel D. Mercaldo, Sebastien Haneuse, Jacob M. Maronge, Paul J. Rathouz, Patrick J. Heagerty, and Jonathan S. Schildcrout. Twowave two-phase outcome-dependent sampling designs, with applications to longitudinal binary data. Statistics in Medicine, 40(8):1863–1876, April 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yuan:2021:DRH

[436] Xiaochen Yuan and Swati Biswas. Detecting rare haplotype association with two correlated phenotypes of binary and continuous types. Statistics in Medicine, 40(8):1877–1900, April 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2021:CHR

[437] Lili Liu, Mae Gordon, J. Philip Miller, Michael Kass, Lu Lin, Shujie Ma, and Lei Liu. Capturing heterogeneity in repeated measures data by fusion penalty. Statistics in Medicine, 40(8):1901–1916, April 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Curnow:2021:MIS

[438] Elinor Curnow, Rachael A. Hughes, Kate Birnie, Michael J. Crowther, Margaret T. May, and Kate Tilling. Multiple imputation strategies for a bounded outcome variable in a competing risks analysis. Statistics in Medicine, 40(8):1917–1929, April 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lou:2021:LLH

[439] Jitong Lou, Yuanjia Wang, Lang Li, and Donglin Zeng. Learning latent heterogeneity for type 2 diabetes patients using longitudinal health markers in electronic health records. Statistics in Medicine, 40(8):1930–1946, April 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Deng:2021:PTM

[440] Yuhao Deng, Fangyi Chen, Yang Li, Kaihuan Qian, Rui Wang, and Xiao-Hua Zhou. A powerful test for the maximum treatment effect in thorough QT/QTc studies. Statistics in Medicine, 40(8):1947–1959, April 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Smith:2021:PSS

[441] Isabelle L. Smith, Jane E. Nixon, and Linda Sharples. Power and sample size for multistate model analysis of longitudinal discrete outcomes in disease prevention trials. Statistics in Medicine, 40(8):1960–1971, April 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bender:2021:FOM

[442] Shaun Bender, Victoria Gamerman, Peter P. Reese, Daniel Lloyd Gray, Yimei Li, and Justine Shults. The first-order Markov conditional linear expectation approach for analysis of longitudinal data. Statistics in Medicine, 40(8):1972–1988, April 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2021:CAB

[443] Zeda Li, Scott A. Bruce, Clinton J. Wutzke, and Yang Long. Conditional adaptive Bayesian spectral analysis of replicated multivariate time series. Statistics in Medicine, 40(8):1989–2005, April 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hong:2021:DPD

[444] Yizhou Hong, Liwen Su, Siyi Song, and Fangrong Yan. Dynamic prediction of disease processes based on recurrent history and functional principal component analysis of longitudinal biomarkers: Application for ovarian epithelial cancer. Statistics in Medicine, 40(8):2006–2023, April 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jay:2021:DSP

[445] Melissa Jay and Rebecca A. Betensky. Displaying survival of patient groups defined by covariate paths: Extensions of the Kaplan–Meier estimator. Statistics in Medicine, 40(8):2024–2036, April 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Park:2021:HDE

[446] Yeonhee Park, Ying Yuan, Jing Ning, Suyu Liu, and Ziding Feng. Hybrid design evaluating new biomarkers when there is an existing screening test. Statistics in Medicine, 40(8):2037–2054, April 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kuronen:2021:PPM

[447] Mikko Kuronen, Mari Myllymäki, Adam Loavenbruck, and Aila Särkkä. Point process models for sweat gland activation observed with noise. Statistics in Medicine, 40(8):2055–2072, April 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Saha:2021:CCW

[448] Pooja T. Saha, Jason P. Fine, and Anastasia Ivanova. Consistency of the CRM when the dose-toxicity curve is not monotone and its application to the POCRM. Statistics in Medicine, 40(8):2073–2082, April 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mei:2021:HDC

[449] Hao Mei, Ruofan Jia, Guanzhong Qiao, Zhenqiu Lin, and Shuangge Ma. Human disease clinical treatment network for the elderly: The analysis of Medicare inpatient length of stay data. Statistics in Medicine, 40(8): 2083–2099, April 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Conover:2021:PST

[450] Mitchell M. Conover, Kenneth J. Rothman, Til Stürmer, Alan R. Ellis, Charles Poole, and Michele Jonsson Funk. Propensity score trimming mitigates bias due to covariate measurement error in inverse probability of treatment weighted analyses: a plasmode simulation. Statistics in Medicine, 40(9):2101–2112, April 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Altzerinakou:2021:CPJ

[451] Maria-Athina Altzerinakou and Xavier Paoletti. Change-point joint model for identification of plateau of activity in early phase trials. Statistics in Medicine, 40(9):2113–2138, April 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Weibull:2021:MMI

[452] Caroline E. Weibull, Paul C. Lambert, Sandra Eloranta, Therese M. L. Andersson, Paul W. Dickman, and Michael J. Crowther. A multistate model incorporating estimation of excess hazards and multiple time scales. Statistics in Medicine, 40(9):2139–2154, April 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2021:PNI

[453] Lin-An Chen and Chu-Lan Kao. Parametric and nonparametric improvements in Bland and Altman’s assessment of agreement method. Statistics in Medicine, 40(9):2155–2176, April 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Conner:2021:EMR

[454] Sarah C. Conner and Ludovic Trinquart. Estimation and modeling of the restricted mean time lost in the presence of competing risks. Statistics in Medicine, 40(9):2177–2196, April 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Utazi:2021:DLE

[455] C. Edson Utazi, Kristine Nilsen, Oliver Pannell, Winfred Dotse-Gborgbortsi, and Andrew J. Tatem. District-level estimation of vaccination coverage: Discrete vs continuous spatial models. Statistics in Medicine, 40(9):2197–2211, April 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Uschner:2021:RBI

[456] Diane Uschner. Randomization-based inference in the presence of selection bias. Statistics in Medicine, 40(9):2212–2229, April 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bhaumik:2021:GCI

[457] Dulal K. Bhaumik, Hairong Shi, Domenic J. Reda, and Bikas K. Sinha. Generalized confidence interval for an agreement between raters. Statistics in Medicine, 40(9):2230–2238, April 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liang:2021:ISP

[458] Weijuan Liang, Shuangge Ma, Qingzhao Zhang, and Tingyu Zhu. Integrative sparse partial least squares. Statistics in Medicine, 40(9):2239– 2256, April 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Michiels:2021:NEA

[459] Hege Michiels, Cristina Sotto, An Vandebosch, and Stijn Vansteelandt. A novel estimand to adjust for rescue treatment in randomized clinical trials. Statistics in Medicine, 40(9):2257–2271, April 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wu:2021:PLR

[460] Hongqian Wu and Michael P. Jones. Proportional likelihood ratio mixed model for discrete longitudinal data. Statistics in Medicine, 40(9):2272– 2285, April 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Neve:2021:SLT

[461] Jan De Neve and Heidelinde Dehaene. Semiparametric linear transformation models for indirectly observed outcomes. Statistics in Medicine, 40(9):2286–2303, April 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kallus:2021:MRE

[462] Nathan Kallus, Brenton Pennicooke, and Michele Santacatterina. More robust estimation of average treatment effects using kernel optimal matching in an observational study of spine surgical interventions. Statistics in Medicine, 40(10):2305–2320, May 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kundu:2021:IFF

[463] Madan G. Kundu and Jyotirmoy Sarkar. On information fraction for Fleming–Harrington type weighted log-rank tests in a group-sequential clinical trial design. Statistics in Medicine, 40(10):2321–2338, May 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Diop:2021:IAP

[464] Awa Diop, Geneviève Lefebvre, Caroline S. Duchaine, Danielle Laurin, and Denis Talbot. The impact of adjusting for pure predictors of exposure, mediator, and outcome on the variance of natural direct and indirect effect estimators. Statistics in Medicine, 40(10):2339–2354, May 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Siivola:2021:QDD

[465] Eero Siivola, Sebastian Weber, and Aki Vehtari. Qualifying drug dosing regimens in pediatrics using Gaussian processes. Statistics in Medicine, 40(10):2355–2372, May 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Haji-Maghsoudi:2021:GLM

[466] Saiedeh Haji-Maghsoudi, Jan Bulla, Majid Sadeghifar, Ghodratollah Roshanaei, and Hossein Mahjub. Generalized linear mixed hidden semi-Markov models in longitudinal settings: a Bayesian approach. Statistics in Medicine, 40(10):2373–2388, May 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lloyd:2021:ECL

[467] Chris J. Lloyd. Exact confidence limits after a group sequential single arm binary trial. Statistics in Medicine, 40(10):2389–2399, May 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lam:2021:MAC

[468] Kwok Fai Lam, Chun Yin Lee, Kin Yau Wong, and Dipankar Bandyopadhyay. Marginal analysis of current status data with informative cluster size using a class of semiparametric transformation cure models. Statistics in Medicine, 40(10):2400–2412, May 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Machida:2021:PSD

[469] Ryunosuke Machida, Yosuke Fujii, and Takashi Sozu. Predicting study duration in clinical trials with a time-to-event endpoint. Statistics in Medicine, 40(10):2413–2421, May 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2021:NTC

[470] Zhengbang Li, Sanan Qin, and Qizhai Li. A novel test by combining the maximum and minimum values among a large number of dependent Z-scores with application to genome wide association study. Statistics in Medicine, 40(10):2422–2434, May 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Buatois:2021:CME

[471] Simon Buatois, Sebastian Ueckert, Nicolas Frey, Sylvie Retout, and France Mentré. cLRT-Mod: an efficient methodology for pharmacometric model-based analysis of longitudinal phase II dose finding studies under model uncertainty. Statistics in Medicine, 40(10):2435–2451, May 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chiarappa:2021:CPC

[472] Joseph A. Chiarappa and Donald R. Hoover. Comparative Poisson clinical trials of multiple experimental treatments vs a single control using the negative multinomial distribution. Statistics in Medicine, 40(10): 2452–2466, May 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Noghrehchi:2021:SMM

[473] Firouzeh Noghrehchi, Jakub Stoklosa, Spiridon Penev, and David I. Warton. Selecting the model for multiple imputation of missing data: Just use an IC! Statistics in Medicine, 40(10):2467–2497, May 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ellenberg:2021:ACT

[474] Susan S. Ellenberg and Jeffrey S. Morris. AIDS and COVID: a tale of two pandemics and the role of statisticians. Statistics in Medicine, 40(11): 2499–2510, May 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See discussion [475, 476, 477, 478, 479, 480, 481, 482, 483] and rejoinder [484].

DeGruttola:2021:DAE

[475] Victor De Gruttola, Ravi Goyal, and Natasha K. Martin. Discussion of article by Ellenberg and Morris. Statistics in Medicine, 40(11):2511–2512, May 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [474].

Gail:2021:TAC

[476] Mitchell H. Gail. Thoughts on “AIDS and COVID-19: a Tale of Two Pandemics and the Role of Statisticians” by Susan S. Ellenberg and Jeffrey S. Morris. Statistics in Medicine, 40(11):2513–2514, May 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [474].

Dean:2021:SSF

[477] Natalie Dean. Statistical successes and failures during the COVID-19 pandemic: Comments on Ellenberg and Morris. Statistics in Medicine, 40(11):2515–2517, May 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [474].

Isham:2021:CDA

[478] Valerie Isham. Contribution to the discussion of AIDS and Covid-19: a tale of two pandemics and the role of statisticians by Ellenberg and Morris. Statistics in Medicine, 40(11):2518–2520, May 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [474].

Brookmeyer:2021:CRS

[479] Ron Brookmeyer. Commentary on the role of statisticians in pandemics. Statistics in Medicine, 40(11):2521–2523, May 20, 2021. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [474].

Halloran:2021:CAC

[480] M. Elizabeth Halloran. Comment on AIDS and COVID-19: a tale of two pandemics and the role of statisticians. Statistics in Medicine, 40(11): 2524–2525, May 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [474].

Follmann:2021:C

[481] Dean Follmann. Comment. Statistics in Medicine, 40(11):2526–2527, May 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [474].

Platt:2021:CEM

[482] Robert W. Platt. Comment on Ellenberg and Morris: The role of statisticians in vaccine surveillance. Statistics in Medicine, 40(11):2528–2529, May 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [474].

Jewell:2021:RSM

[483] Britta L. Jewell and Nicholas P. Jewell. On the role of statisticians and modelers in responding to AIDS and COVID-19. Statistics in Medicine, 40(11):2530–2535, May 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [474].

Ellenberg:2021:R

[484] Susan S. Ellenberg and Jeffrey S. Morris. Rejoinder. Statistics in Medicine, 40(11):2536–2539, May 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [474].

Mokalled:2021:IDE

[485] Stefani C. Mokalled, Christopher S. McMahan, Joshua M. Tebbs, Derek Andrew Brown, and Christopher R. Bilder. Incorporating the dilution effect in group testing regression. Statistics in Medicine, 40(11): 2540–2555, May 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2021:ETE

[486] Wen Li, Xiaofei Bai, Qiqi Deng, Fang Liu, and Cong Chen. Estimation of treatment effect in 2-in-1 adaptive design and some of its extensions. Statistics in Medicine, 40(11):2556–2577, May 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Frieri:2021:OED

[487] Rosamarie Frieri and Maroussa Zagoraiou. Optimal and ethical designs for hypothesis testing in multi-arm exponential trials. Statistics in Medicine, 40(11):2578–2603, May 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2021:TSP

[488] Athena Chen, Oliver Laeyendecker, Susan H. Eshleman, Daniel R. Monaco, Kai Kammers, Harry Benjamin Larman, and Ingo Ruczinski. A top scoring pairs classifier for recent HIV infections. Statistics in Medicine, 40(11):2604–2612, May 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tan:2021:ASB

[489] Yaoyuan Vincent Tan, Carol A. C. Flannagan, Lindsay R. Pool, and Michael R. Elliott. Accounting for selection bias due to death in estimating the effect of wealth shock on cognition for the health and retirement study. Statistics in Medicine, 40(11):2613–2625, May 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shi:2021:UUB

[490] Haolun Shi, Jiguo Cao, Ying Yuan, and Ruitao Lin. uTPI: a utilitybased toxicity probability interval design for phase I/II dose-finding trials. Statistics in Medicine, 40(11):2626–2649, May 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2021:RRM

[491] Pei Wang, Erin L. Abner, David W. Fardo, Frederick A. Schmitt, Gregory A. Jicha, Linda J. Van Eldik, and Richard J. Kryscio. Reduced rank multinomial logistic regression in Markov chains with application to cognitive data. Statistics in Medicine, 40(11):2650–2664, May 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Spanbauer:2021:NML

[492] Charles Spanbauer and Rodney Sparapani. Nonparametric machine learning for precision medicine with longitudinal clinical trials and Bayesian additive regression trees with mixed models. Statistics in Medicine, 40(11):2665–2691, May 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Huang:2021:CSM

[493] Xianzheng Huang and Hongmei Zhang. Corrected score methods for estimating Bayesian networks with error-prone nodes. Statistics in Medicine, 40(11):2692–2712, May 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2021:REH

[494] Ruohong Li, Honglang Wang, and Wanzhu Tu. Robust estimation of heterogeneous treatment effects using electronic health record data. Statistics in Medicine, 40(11):2713–2752, May 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Remiro-Azocar:2021:CMC

[495] Antonio Remiro-Azócar, Anna Heath, and Gianluca Baio. Conflating marginal and conditional treatment effects: Comments on “Assessing the performance of population adjustment methods for anchored indirect comparisons: a simulation study”. Statistics in Medicine, 40(11):2753– 2758, May 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [317].

Phillippo:2021:TEE

[496] David M. Phillippo, Sofia Dias, Anthony E. Ades, and Nicky J. Welton. Target estimands for efficient decision making: Response to comments on “Assessing the performance of population adjustment methods for anchored indirect comparisons: a simulation study”. Statistics in Medicine, 40(11):2759–2763, May 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [317].

Lin:2021:IPW

[497] Feng-Chang Lin, Jianwen Cai, Yu Deng, and Charles R. Esther, Jr. Inverse probability weighted estimation for recurrent events data with missing category. Statistics in Medicine, 40(12):2765–2782, May 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Halabi:2021:DAS

[498] Susan Halabi, Chen-Yen Lin, and Aiyi Liu. On the design and the analysis of stratified biomarker trials in the presence of measurement error. Statistics in Medicine, 40(12):2783–2799, May 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sint:2021:LCM

[499] Kyaw Sint, Robert Rosenheck, and Haiqun Lin. Latent class mediator for multiple indicators of mediation. Statistics in Medicine, 40(12):2800– 2820, May 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Strzalkowska-Kominiak:2021:CFD

[500] Ewa Strzalkowska-Kominiak and Juan Romo. Censored functional data for incomplete follow-up studies. Statistics in Medicine, 40(12):2821– 2838, May 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2021:SSR

[501] Xin Li, Wei Ma, and Feifang Hu. Sample size re-estimation for covariateadaptive randomized clinical trials. Statistics in Medicine, 40(12):2839– 2858, May 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Koh:2021:MAM

[502] Hyunwook Koh, Susan Tuddenham, Cynthia L. Sears, and Ni Zhao. Meta-analysis methods for multiple related markers: Applications to microbiome studies with the results on multiple α-diversity indices. Statistics in Medicine, 40(12):2859–2876, May 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wilson:2021:BDA

[503] Duncan T. Wilson, James M. S. Wason, Julia Brown, Amanda J. Farrin, and Rebecca E. A. Walwyn. Bayesian design and analysis of external pilot trials for complex interventions. Statistics in Medicine, 40(12): 2877–2892, May 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kang:2021:HBC

[504] Daniel Kang, Christopher S. Coffey, Brian J. Smith, Ying Yuan, Qian Shi, and Jun Yin. Hierarchical Bayesian clustering design of multiple biomarker subgroups (HCOMBS). Statistics in Medicine, 40(12):2893– 2921, May 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jay:2021:BAE

[505] Melissa Jay, Jacob Oleson, Mary Charlton, and Ali Arab. A Bayesian approach for estimating age-adjusted rates for low-prevalence diseases over space and time. Statistics in Medicine, 40(12):2922–2938, May 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ballarini:2021:OSS

[506] Nicolás M. Ballarini, Thomas Burnett, Thomas Jaki, Christoper Jennison, Franz König, and Martin Posch. Optimizing subgroup selection in two-stage adaptive enrichment and umbrella designs. Statistics in Medicine, 40(12):2939–2956, May 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pateras:2021:CAE

[507] Konstantinos Pateras, Stavros Nikolakopoulos, and Kit C. B. Roes. Combined assessment of early and late-phase outcomes in orphan drug development. Statistics in Medicine, 40(12):2957–2974, May 30, 2021. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Alvares:2021:BSA

[508] Danilo Alvares, Elena Lázaro, Virgilio Gómez-Rubio, and Carmen Armero. Bayesian survival analysis with BUGS. Statistics in Medicine, 40(12):2975–3020, May 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bilder:2021:IAT

[509] Christopher R. Bilder, Joshua M. Tebbs, and Christopher S. McMahan. Informative array testing with multiplex assays. Statistics in Medicine, 40(13):3021–3034, June 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shin:2021:BHM

[510] Yei Eun Shin, Dawei Liu, Huiyan Sang, Toby A. Ferguson, and Peter X. K. Song. A binary hidden Markov model on spatial network for amyotrophic lateral sclerosis disease spreading pattern analysis. Statistics in Medicine, 40(13):3035–3052, June 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ebrahimpoor:2021:PTL

[511] Mitra Ebrahimpoor, Pietro Spitali, Jelle J. Goeman, and Roula Tsonaka. Pathway testing for longitudinal metabolomics. Statistics in Medicine, 40(13):3053–3065, June 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ensor:2021:IPD

[512] Joie Ensor, Kym I. E. Snell, Thomas P. A. Debray, Paul C. Lambert, Maxime P. Look, Mamas A. Mamas, Karel G. M. Moons, and Richard D. Riley. Individual participant data meta-analysis for external validation, recalibration, and updating of a flexible parametric prognostic model. Statistics in Medicine, 40(13):3066–3084, June 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bandyopadhyay:2021:SSN

[513] Dipankar Bandyopadhyay, Marcos O. Prates, Xiaoyue Zhao, and Victor H. Lachos. Spatial skew-normal/independent models for nonrandomly missing clustered data. Statistics in Medicine, 40(13):3085–3105, June 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhou:2021:SRAa

[514] Qingning Zhou, Jianwen Cai, and Haibo Zhou. Semiparametric regression analysis of case-cohort studies with multiple interval-censored disease outcomes. Statistics in Medicine, 40(13):3106–3123, June 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Breskin:2021:FDE

[515] Alexander Breskin, Stephen R. Cole, Jessie K. Edwards, Ron Brookmeyer, Joseph J. Eron, and Adimora A. Adimora. Fusion designs and estimators for treatment effects. Statistics in Medicine, 40(13):3124–3137, June 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See corrigendum [1138].

Zhou:2021:MOS

[516] Guohai Zhou and Lang Wu. Multiparameter one-sided tests for nonlinear mixed effects models with censored responses. Statistics in Medicine, 40 (13):3138–3152, June 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2021:SYP

[517] Qi Zhang, Feifei Chen, Shunyao Wu, and Hua Liang. A simple yet powerful test for assessing goodness-of-fit of high-dimensional linear models. Statistics in Medicine, 40(13):3153–3166, June 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Goyal:2021:IPS

[518] Ravi Goyal and Victor De Gruttola. Investigation of patient-sharing networks using a Bayesian network model selection approach for congruence class models. Statistics in Medicine, 40(13):3167–3180, June 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jia:2021:ILH

[519] Beilin Jia, Donglin Zeng, Jason J. Z. Liao, Guanghan F. Liu, Xianming Tan, Guoqing Diao, and Joseph G. Ibrahim. Inferring latent heterogeneity using many feature variables supervised by survival outcome. Statistics in Medicine, 40(13):3181–3195, June 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pilz:2021:OPA

[520] Maximilian Pilz, Kevin Kunzmann, Carolin Herrmann, Geraldine Rauch, and Meinhard Kieser. Optimal planning of adaptive two-stage designs. Statistics in Medicine, 40(13):3196–3213, June 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kojima:2021:ECP

[521] Masahiro Kojima. Early completion of phase I cancer clinical trials with Bayesian optimal interval design. Statistics in Medicine, 40(14):3215– 3226, June 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zou:2021:CIE

[522] Guangyong Zou. Confidence interval estimation for treatment effects in cluster randomization trials based on ranks. Statistics in Medicine, 40 (14):3227–3250, June 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kuhnel:2021:SMA

[523] Line Kühnel, Anna-Karin Berger, Bo Markussen, and Lars L. Raket. Simultaneous modeling of Alzheimer’s disease progression via multiple cognitive scales. Statistics in Medicine, 40(14):3251–3266, June 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Feng:2021:SGR

[524] Tianshu Feng and Linda Ng Boyle. Sparse group regularization for semicontinuous transportation data. Statistics in Medicine, 40(14):3267– 3285, June 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rossel:2021:CSM

[525] Jean-Benoît Rossel, Valentin Rousson, and Yves Eggli. A comparison of statistical methods for allocating disease costs in the presence of interactions. Statistics in Medicine, 40(14):3286–3298, June 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Simonetto:2021:SDA

[526] Cristoforo Simonetto, Susanne Rospleszcz, Margit Heier, Christa Meisinger, Annette Peters, and Jan Christian Kaiser. Simulating the dynamics of atherosclerosis to the incidence of myocardial infarction, applied to the KORA population. Statistics in Medicine, 40(14):3299–3312, June 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kormaksson:2021:SKC

[527] Matthias Kormaksson, Luke J. Kelly, Xuan Zhu, Sibylle Haemmerle, Luminita Pricop, and David Ohlssen. Sequential knockoffs for continuous and categorical predictors: With application to a large psoriatic arthritis clinical trial pool. Statistics in Medicine, 40(14):3313–3328, June 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hemming:2021:CHM

[528] Karla Hemming, Monica Taljaard, Mirjam Moerbeek, and Andrew Forbes. Contamination: How much can an individually randomized trial tolerate? Statistics in Medicine, 40(14):3329–3351, June 30, 2021. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gleiss:2021:DNS

[529] Andreas Gleiss, Robin Henderson, and Michael Schemper. Degrees of necessity and of sufficiency: Further results and extensions, with an application to COVID-19 mortality in Austria. Statistics in Medicine, 40 (14):3352–3366, June 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brunner:2021:WOA

[530] Edgar Brunner, Marc Vandemeulebroecke, and Tobias Mütze. Win odds: an adaptation of the win ratio to include ties. Statistics in Medicine, 40 (14):3367–3384, June 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hupf:2021:BSM

[531] Bradley Hupf, Veronica Bunn, Jianchang Lin, and Cheng Dong. Bayesian semiparametric meta-analytic-predictive prior for historical control borrowing in clinical trials. Statistics in Medicine, 40(14):3385–3399, June 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vonesh:2021:CME

[532] Edward F. Vonesh and Tom Greene. Correction: Mixed-effects models for slope-based endpoints in clinical trials of chronic kidney disease. Statistics in Medicine, 40(14):3400–3401, June 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gelman:2021:SSB

[533] Andrew Gelman and Matthijs Vákár. Slamming the sham: a Bayesian model for adaptive adjustment with noisy control data. Statistics in Medicine, 40(15):3403–3424, July 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Klich:2021:TCU

[534] Amna Klich, René Ecochard, and Fabien Subtil. Trajectory clustering using mixed classification models. Statistics in Medicine, 40(15):3425– 3439, July 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2021:MCP

[535] Jialiang Li, Yaguang Li, Baisuo Jin, and Michael R. Kosorok. Multithreshold change plane model: Estimation theory and applications in subgroup identification. Statistics in Medicine, 40(15):3440–3459, July 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Aron:2021:HMS

[536] Jordan Aron, Paul S. Albert, Nicolas Wentzensen, and Li C. Cheung. Hidden mover-stayer model for disease progression accounting for misclassified and partially observed diagnostic tests: Application to the natural history of human papillomavirus and cervical precancer. Statistics in Medicine, 40(15):3460–3476, July 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2021:EPM

[537] Pin Li, Jeremy M. G. Taylor, Daniel E. Spratt, R. Jeffery Karnes, and Matthew J. Schipper. Evaluation of predictive model performance of an existing model in the presence of missing data. Statistics in Medicine, 40(15):3477–3498, July 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

He:2021:RCE

[538] Yong He, Pengfei Liu, Xinsheng Zhang, and Wang Zhou. Robust covariance estimation for high-dimensional compositional data with application to microbial communities analysis. Statistics in Medicine, 40(15): 3499–3515, July 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xie:2021:PTC

[539] Yujing Xie and Zhangsheng Yu. Promotion time cure rate model with a neural network estimated nonparametric component. Statistics in Medicine, 40(15):3516–3532, July 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

deJong:2021:DMG

[540] Valentijn M. T. de Jong, Karel G. M. Moons, Marinus J. C. Eijkemans, Richard D. Riley, and Thomas P. A. Debray. Developing more generalizable prediction models from pooled studies and large clustered data sets. Statistics in Medicine, 40(15):3533–3559, July 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shi:2021:NPB

[541] Daoyuan Shi, Ming-Hui Chen, Lynn Kuo, and Paul O. Lewis. New partition based measures for data compatibility and information gain. Statistics in Medicine, 40(15):3560–3581, July 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2021:BNM

[542] Hao Li, Daeyoung Lim, Ming-Hui Chen, Joseph G. Ibrahim, Sungduk Kim, Arvind K. Shah, and Jianxin Lin. Bayesian network metaregression hierarchical models using heavy-tailed multivariate random effects with covariate-dependent variances. Statistics in Medicine, 40 (15):3582–3603, July 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wu:2021:VBP

[543] Ying Wu and Niansheng Tang. Variational Bayesian partially linear mean shift models for high-dimensional Alzheimer’s disease neuroimaging data. Statistics in Medicine, 40(15):3604–3624, July 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sanders:2021:IPA

[544] Eric Sanders, Paul Gustafson, and Mohammad Ehsanul Karim. Incorporating partial adherence into the principal stratification analysis framework. Statistics in Medicine, 40(15):3625–3644, July 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Laubender:2021:SWR

[545] Ruediger Paul Laubender and Andrea Geistanger. Selection of within-run quality control rules for laboratory biomarkers. Statistics in Medicine, 40(16):3645–3666, July 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dover:2021:CPD

[546] Douglas C. Dover, Sunjidatul Islam, Cynthia M. Westerhout, Linn E. Moore, Padma Kaul, and Anamaria Savu. Computing the polytomous discrimination index. Statistics in Medicine, 40(16):3667–3681, July 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sandoval:2021:CEC

[547] Grecio J. Sandoval, Ionut Bebu, and John M. Lachin. Cost-efficient clinical studies with continuous time survival outcomes. Statistics in Medicine, 40(16):3682–3694, July 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gao:2021:CIE

[548] Yi Gao and Lili Tian. Confidence interval estimation for sensitivity and difference between two sensitivities at a given specificity under tree ordering. Statistics in Medicine, 40(16):3695–3723, July 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2021:RAA

[549] Lu Wang and Lianming Wang. Regression analysis of arbitrarily censored survival data under the proportional odds model. Statistics in Medicine, 40(16):3724–3739, July 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tran:2021:MAA

[550] Thao M. P. Tran, Steven Abrams, Marc Aerts, Kirsten Maertens, and Niel Hens. Measuring association among censored antibody titer data. Statistics in Medicine, 40(16):3740–3761, July 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jo:2021:BSM

[551] Seongil Jo, Beomjo Park, Yeonseung Chung, Jeongseon Kim, Eunji Lee, Jangwon Lee, and Taeryon Choi. Bayesian semiparametric mixed effects models for meta-analysis of the literature data: an application to cadmium toxicity studies. Statistics in Medicine, 40(16):3762–3778, July 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Seaman:2021:UGL

[552] Shaun R. Seaman, Ruth H. Keogh, Oliver Dukes, and Stijn Vansteelandt. Using generalized linear models to implement g-estimation for survival data with time-varying confounding. Statistics in Medicine, 40(16):3779– 3790, July 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bhatt:2021:MSM

[553] Rikesh Bhatt, Ardo van den Hout, and Nora Pashayan. A multistate survival model of the natural history of cancer using data from screened and unscreened population. Statistics in Medicine, 40(16):3791–3807, July 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Soave:2021:STS

[554] David Soave, Jerald F. Lawless, and Philip Awadalla. Score tests for scale effects, with application to genomic analysis. Statistics in Medicine, 40 (16):3808–3822, July 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhong:2021:SRE

[555] Yujie Zhong and Richard J. Cook. Semiparametric recurrent event vs time-to-first-event analyses in randomized trials: Estimands and model misspecification. Statistics in Medicine, 40(16):3823–3842, July 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

DiLoro:2021:NCI

[556] Pierfrancesco Alaimo Di Loro, Fabio Divino, Alessio Farcomeni, Giovanna Jona Lasinio, Gianfranco Lovison, Antonello Maruotti, and Marco Mingione. Nowcasting COVID-19 incidence indicators during the Italian first outbreak. Statistics in Medicine, 40(16):3843–3864, July 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Haber:2021:GTR

[557] Gregory Haber, Yaakov Malinovsky, and Paul S. Albert. Is group testing ready for prime-time in disease identification? Statistics in Medicine, 40(17):3865–3880, July 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See discussion [558, 559, 560] and rejoinder [561].

Bilder:2021:DGT

[558] Christopher R. Bilder, Joshua M. Tebbs, and Christopher S. McMahan. Discussion on “Is group testing ready for prime-time in disease identification”. Statistics in Medicine, 40(17):3881–3886, July 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [557].

Biggerstaff:2021:DGT

[559] Brad J. Biggerstaff. Discussion of “Is group testing ready for prime-time in disease Identification?” by Haber, Malinovsky, and Albert, Statistics in Medicine, 2021. Statistics in Medicine, 40(17):3887–3888, July 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [557].

Johnson:2021:CGT

[560] Wesley O. Johnson. Comment on “Is Group Testing Ready for Prime Time in Disease Identification?”. Statistics in Medicine, 40(17):3889– 3891, July 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [557].

Haber:2021:RDG

[561] Gregory Haber, Yaakov Malinovsky, and Paul S. Albert. Rejoinder to discussion on Is group testing ready for prime-time in disease identification? Statistics in Medicine, 40(17):3892–3894, July 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [557].

Ascari:2021:NRM

[562] Roberto Ascari and Sonia Migliorati. A new regression model for overdispersed binomial data accounting for outliers and an excess of zeros. Statistics in Medicine, 40(17):3895–3914, July 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2021:BGH

[563] Sanguo Zhang, Xiaonan Hu, Ziye Luo, Yu Jiang, Yifan Sun, and Shuangge Ma. Biomarker-guided heterogeneity analysis of genetic regulations via multivariate sparse fusion. Statistics in Medicine, 40(17): 3915–3936, July 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2021:MMS

[564] Yihao Li, Danh V. Nguyen, Sudipto Banerjee, Connie M. Rhee, Kamyar Kalantar-Zadeh, Esra Kürüm, and Damla Sentürk. Multilevel modeling of spatially nested functional data: Spatiotemporal patterns of hospitalization rates in the US dialysis population. Statistics in Medicine, 40(17): 3937–3952, July 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tai:2021:SMA

[565] An-Shun Tai, Chun-An Tsai, and Sheng-Hsuan Lin. Survival mediation analysis with the death-truncated mediator: The completeness of the survival mediation parameter. Statistics in Medicine, 40(17):3953–3974, July 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2021:CMA

[566] Wei Li, Zhi Geng, and Xiao-Hua Zhou. Causal mediation analysis with sure outcomes of random events model. Statistics in Medicine, 40(17): 3975–3989, July 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gao:2021:AHT

[567] Zijun Gao, Trevor Hastie, and Robert Tibshirani. Assessment of heterogeneous treatment effect estimation accuracy via matching. Statistics in Medicine, 40(17):3990–4013, July 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kim:2021:MRB

[568] Chanmin Kim, Xiaoyan Lin, and Kerrie P. Nelson. Measuring rater bias in diagnostic tests with ordinal ratings. Statistics in Medicine, 40(17): 4014–4033, July 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2021:SMC

[569] Chyong-Mei Chen, Pao sheng Shen, Chih-Ching Lin, and Chih-Cheng Wu. Semiparametric mixture cure model analysis with competing risks data: Application to vascular access thrombosis data. Statistics in Medicine, 40(17):4034, July 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2021:BEU

[570] Xuan Wang, Yingye Zheng, Majken Karoline Jensen, Zeling He, and Tianxi Cai. Biomarker evaluation under imperfect nested case-control design. Statistics in Medicine, 40(18):4035–4052, August 15, 2021. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ma:2021:SAZ

[571] Chenchen Ma, Tao Hu, and Zhantao Lin. Semiparametric analysis of zero-inflated recurrent events with a terminal event. Statistics in Medicine, 40(18):4053–4067, August 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rosenkranz:2021:RSF

[572] Gerd K. Rosenkranz. Replicability of studies following a dual-criterion design. Statistics in Medicine, 40(18):4068–4076, August 15, 2021. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhou:2021:CEM

[573] Yan Zhou, Li Zhang, Jinfeng Xu, Jun Zhang, and Xiaodong Yan. Category encoding method to select feature genes for the classification of bulk and single-cell RNA-seq data. Statistics in Medicine, 40(18):4077–4089, August 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sauer:2021:OAS

[574] Sara Sauer, Bethany Hedt-Gauthier, and Sebastien Haneuse. Optimal allocation in stratified cluster-based outcome-dependent sampling designs. Statistics in Medicine, 40(18):4090–4107, August 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

VanLancker:2021:PSB

[575] Kelly Van Lancker, Oliver Dukes, and Stijn Vansteelandt. Principled selection of baseline covariates to account for censoring in randomized trials with a survival endpoint. Statistics in Medicine, 40(18):4108–4121, August 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Roig:2021:DPI

[576] Marta Bofill Roig, Yu Shen, and Guadalupe Gómez Melis. Design of phase III trials with long-term survival outcomes based on short-term binary results. Statistics in Medicine, 40(18):4122–4135, August 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Purington:2021:NAD

[577] Natasha Purington, Sandra Andorf, Bryan Bunning, Sharon Chinthrajah, Kari Nadeau, and Manisha Desai. Novel application of a discrete time-to-event model for randomized oral immunotherapy clinical trials with repeat food challenges. Statistics in Medicine, 40(18):4136–4149, August 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Agosto:2021:MCC

[578] Arianna Agosto, Alexandra Campmas, Paolo Giudici, and Andrea Renda. Monitoring COVID-19 contagion growth. Statistics in Medicine, 40(18):4150–4160, August 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Krahn:2021:CII

[579] Ulrike Krahn, Harald Binder, Gerta Rücker, and Jochem König. Comments on “Identifying inconsistency in network meta-analysis: Is the net heat plot a reliable method?”. Statistics in Medicine, 40(18):4161–4163, August 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [10, 580].

Freeman:2021:ARC

[580] Suzanne C. Freeman, David Fisher, Ian R. White, and James R. Carpenter. Authors’ reply to “Comments on Identifying inconsistency in network meta-analysis: Is the net heat plot a reliable method?”. Statistics in Medicine, 40(18):4164–4165, August 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [10, 579].

Jones:2021:CQH

[581] Hayley E. Jones. Correction: Quantifying how diagnostic test accuracy depends on threshold in a meta-analysis. Statistics in Medicine, 40(18): 4166, August 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [9].

Lake:2021:BAD

[582] Stephen L. Lake, Melanie A. Quintana, Kristine Broglio, Jennifer Panagoulias, Scott M. Berry, and Michael A. Panzara. Bayesian adaptive design for clinical trials in Duchenne muscular dystrophy. Statistics in Medicine, 40(19):4167–4184, August 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brand:2021:PDP

[583] Adam Brand, Susanne May, James P. Hughes, Gertrude Nakigozi, Steven J. Reynolds, and Erin E. Gabriel. Prediction-driven pooled testing methods: Application to HIV treatment monitoring in Rakai, Uganda. Statistics in Medicine, 40(19):4185–4199, August 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Austin:2021:FGS

[584] Peter C. Austin, Ewout W. Steyerberg, and Hein Putter. Fine–Gray subdistribution hazard models to simultaneously estimate the absolute risk of different event types: Cumulative total failure probability may exceed 1. Statistics in Medicine, 40(19):4200–4212, August 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Alvares:2021:TBJ

[585] Danilo Alvares and Francisco J. Rubio. A tractable Bayesian joint model for longitudinal and survival data. Statistics in Medicine, 40(19):4213– 4229, August 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Riley:2021:MSS

[586] Richard D. Riley, Thomas P. A. Debray, Gary S. Collins, Lucinda Archer, Joie Ensor, Maarten van Smeden, and Kym I. E. Snell. Minimum sample size for external validation of a clinical prediction model with a binary outcome. Statistics in Medicine, 40(19):4230–4251, August 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jiang:2021:EEL

[587] Zhongfeng Jiang, Baoying Yang, Jing Qin, and Yong Zhou. Enhanced empirical likelihood estimation of incubation period of COVID-19 by integrating published information. Statistics in Medicine, 40(19):4252– 4268, August 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2021:VAE

[588] Shuoran Li and Lili Zhao. Vaccine adverse event enrichment tests. Statistics in Medicine, 40(19):4269–4278, August 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

vanWieringen:2021:PEG

[589] Wessel N. van Wieringen and Yao Chen. Penalized estimation of the Gaussian graphical model from data with replicates. Statistics in Medicine, 40(19):4279–4293, August 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yang:2021:PSW

[590] Siyun Yang, Elizabeth Lorenzi, Georgia Papadogeorgou, Daniel M. Wojdyla, Fan Li, and Laine E. Thomas. Propensity score weighting for causal subgroup analysis. Statistics in Medicine, 40(19):4294–4309, August 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Josey:2021:TER

[591] Kevin P. Josey, Seth A. Berkowitz, Debashis Ghosh, and Sridharan Raghavan. Transporting experimental results with entropy balancing. Statistics in Medicine, 40(19):4310–4326, August 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bitan:2021:ESD

[592] Michal Bitan, Malka Gorfine, Laura Rosen, and David M. Steinberg. Efficient study design to estimate population means with multiple measurement instruments. Statistics in Medicine, 40(19):4327–4340, August 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mehrotra:2021:SAU

[593] Devan V. Mehrotra and Rachel Marceau West. Survival analysis using a 5-step stratified testing and amalgamation routine (5-STAR) in randomized clinical trials. Statistics in Medicine, 40(19):4341–4343, August 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Diaz:2021:UPC

[594] Francisco J. Diaz. Using population crossover trials to improve the decision process regarding treatment individualization in n-of-1 trials.

Statistics in Medicine, 40(20):4345–4361, September 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chalkou:2021:TSP

[595] Konstantina Chalkou, Ewout Steyerberg, Matthias Egger, Andrea Manca, Fabio Pellegrini, and Georgia Salanti. A two-stage prediction model for heterogeneous effects of treatments. Statistics in Medicine, 40 (20):4362–4375, September 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhou:2021:SRAb

[596] Qingning Zhou, Yanqing Sun, and Peter B. Gilbert. Semiparametric regression analysis of partly interval-censored failure time data with application to an AIDS clinical trial. Statistics in Medicine, 40(20):4376– 4394, September 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shen:2021:BJM

[597] Fan Shen and Liang Li. Backward joint model and dynamic prediction of survival with multivariate longitudinal data. Statistics in Medicine, 40(20):4395–4409, September 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bianconcini:2021:DLV

[598] Silvia Bianconcini and Silvia Cagnone. Dynamic latent variable models for the analysis of cognitive abilities in the elderly population. Statistics in Medicine, 40(20):4410–4429, September 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhu:2021:SIC

[599] Yuxin Zhu, Zheyu Wang, Ava L. Liberman, Tzu-Pu Chang, and David Newman-Toker. Statistical insights for crude-rate-based operational measures of misdiagnosis-related harms. Statistics in Medicine, 40(20):4430– 4441, September 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rabideau:2021:RBI

[600] Dustin J. Rabideau and Rui Wang. Randomization-based inference for a marginal treatment effect in stepped wedge cluster randomized trials. Statistics in Medicine, 40(20):4442–4456, September 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ren:2021:BPI

[601] Tingyang Ren, Weining Shen, Liwen Zhang, and Haibing Zhao. Bayesian phase II clinical trial design with noncompliance. Statistics in Medicine, 40(20):4457–4472, September 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2021:REV

[602] Yi Li, Muxuan Liang, Lu Mao, and Sijian Wang. Robust estimation and variable selection for the accelerated failure time model. Statistics in Medicine, 40(20):4473–4491, September 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ortega-Villa:2021:MAF

[603] Ana M. Ortega-Villa, Martha C. Nason, and Dean Follmann. The mechanistic analysis of founder virus data in challenge models. Statistics in Medicine, 40(20):4492–4504, September 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ott:2021:HVV

[604] Manuela Ott, Martyn Plummer, and Małgorzata Roos. How vague is vague? How informative is informative? Reference analysis for Bayesian meta-analysis. Statistics in Medicine, 40(20):4505–4521, September 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mosier:2021:ECC

[605] Brian R. Mosier and Leonidas E. Bantis. Estimation and construction of confidence intervals for biomarker cutoff-points under the shortest Euclidean distance from the ROC surface to the perfection corner. Statistics in Medicine, 40(20):4522–4539, September 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tai:2021:IMM

[606] An-Shun Tai and Sheng-Hsuan Lin. Integrated multiple mediation analysis: a robustness-specificity trade-off in causal structure. Statistics in Medicine, 40(21):4541–4567, September 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Biard:2021:SPI

[607] Lucie Biard, Shing M. Lee, and Bin Cheng. Seamless phase I/II design for novel anticancer agents with competing disease progression. Statistics in Medicine, 40(21):4568–4581, September 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shen:2021:JML

[608] Biyi Shen, Chixiang Chen, Danping Liu, Somnath Datta, Nasrollah Ghahramani, Vernon M. Chinchilli, and Ming Wang. Joint modeling of longitudinal data with informative cluster size adjusted for zero-inflation and a dependent terminal event. Statistics in Medicine, 40(21):4582– 4596, September 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tang:2021:RBT

[609] Larry L. Tang, Zhen Meng, and Qizhai Li. A ROC-based test for evaluating the group difference with an application to neonatal audiology screening. Statistics in Medicine, 40(21):4597–4608, September 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kaciroti:2021:BSA

[610] Niko A. Kaciroti and Roderick J. A. Little. Bayesian sensitivity analyses for longitudinal data with dropouts that are potentially missing not at random: a high dimensional pattern-mixture model. Statistics in Medicine, 40(21):4609–4628, September 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2021:BHM

[611] Jian Wang, Junsheng Ma, Chunyan Cai, Naval Daver, and Jing Ning. A Bayesian hierarchical monitoring design for phase II cancer clinical trials: Incorporating information on response duration into monitoring rules. Statistics in Medicine, 40(21):4629–4639, September 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shi:2021:IFM

[612] Baoyi Shi and R. Todd Ogden. Inference in functional mixed regression models with applications to Positron Emission Tomography imaging data. Statistics in Medicine, 40(21):4640–4659, September 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mainzer:2021:CMI

[613] Rheanna Mainzer, Jemishabye Apajee, Cattram D. Nguyen, John B. Carlin, and Katherine J. Lee. A comparison of multiple imputation strategies for handling missing data in multi-item scales: Guidance for longitudinal studies. Statistics in Medicine, 40(21):4660–4674, September 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Weiss:2021:ACT

[614] Christian H. Weiß. Analyzing categorical time series in the presence of missing observations. Statistics in Medicine, 40(21):4675–4690, September 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hu:2021:EHS

[615] Liangyuan Hu, Jiayi Ji, and Fan Li. Estimating heterogeneous survival treatment effect in observational data using machine learning. Statistics in Medicine, 40(21):4691–4713, September 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wilson:2021:HTF

[616] Duncan T. Wilson, Julia Brown, Amanda J. Farrin, and Rebecca E. A. Walwyn. A hypothesis test of feasibility for external pilot trials assessing recruitment, follow-up, and adherence rates. Statistics in Medicine, 40 (21):4714–4731, September 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Konietschke:2021:ACU

[617] Frank Konietschke, Cong Cao, Asanka Gunawardana, and Georg Zimmermann. Analysis of covariance under variance heteroscedasticity in general factorial designs. Statistics in Medicine, 40(21):4732–4749, September 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schofield:2021:RLC

[618] Matthew R. Schofield, Michael J. Maze, John A. Crump, Matthew P. Rubach, Renee Galloway, and Katrina J. Sharples. On the robustness of latent class models for diagnostic testing with no gold standard. Statistics in Medicine, 40(22):4751–4763, September 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See comments [620] and rejoinder [621]..

Albert:2021:CCU

[619] Paul S. Albert. Continued controversy in using latent class models for estimating diagnostic accuracy without a gold standard. Statistics in Medicine, 40(22):4764–4765, September 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dendukuri:2021:CRL

[620] Nandini Dendukuri. Commentary on “On the robustness of latent class models for diagnostic testing with no gold-standard” by Schofield et al. Statistics in Medicine, 40(22):4766–4769, September 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [618].

Schofield:2021:RRL

[621] Matthew R. Schofield, Michael J. Maze, John A. Crump, Matthew P. Rubach, Renee L. Galloway, and Katrina J. Sharples. Rejoinder to “On the robustness of latent class models for diagnostic testing with no gold standard”. Statistics in Medicine, 40(22):4770–4771, September 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [620, 618].

Jang:2021:BMI

[622] Jeong Hoon Jang, Amita K. Manatunga, Changgee Chang, and Qi Long. A Bayesian multiple imputation approach to bivariate functional data with missing components. Statistics in Medicine, 40(22):4772–4793, September 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2021:PSB

[623] Meizi Liu, Veronica Bunn, Bradley Hupf, Junjing Lin, and Jianchang Lin. Propensity-score-based meta-analytic predictive prior for incorporating real-world and historical data. Statistics in Medicine, 40(22):4794–4808, September 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lawrence:2021:SSA

[624] John Lawrence. Sample size adaptations and evaluation of pediatric study interpretability. Statistics in Medicine, 40(22):4809–4814, September 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Walter:2021:EDT

[625] Stephen D. Walter. Estimation of diagnostic test accuracy: a “Rule of Three” for data with repeated observations but without a gold standard. Statistics in Medicine, 40(22):4815–4829, September 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mallick:2021:RBL

[626] Himel Mallick, Rahim Alhamzawi, Erina Paul, and Vladimir Svetnik. The reciprocal Bayesian LASSO. Statistics in Medicine, 40(22):4830– 4849, September 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kowal:2021:BVS

[627] Daniel R. Kowal, Mercedes Bravo, Henry Leong, Alexander Bui, Robert J. Griffin, Katherine B. Ensor, and Marie Lynn Miranda. Bayesian variable selection for understanding mixtures in environmental exposures. Statistics in Medicine, 40(22):4850–4871, September 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Talisa:2021:LCC

[628] Victor B. Talisa and Chung-Chou H. Chang. Learning and confirming a class of treatment responders in clinical trials. Statistics in Medicine, 40(22):4872–4889, September 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Silva:2021:EST

[629] Ivair R. Silva, Judith Maro, and Martin Kulldorff. Exact sequential test for clinical trials and post-market drug and vaccine safety surveillance with Poisson and binary data. Statistics in Medicine, 40(22):4890–4913, September 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sheng:2021:SEA

[630] Ying Sheng, Yifei Sun, Chiung-Yu Huang, and Mi-Ok Kim. Synthesizing external aggregated information in the penalized Cox regression under population heterogeneity. Statistics in Medicine, 40(23):4915–4930, October 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Suresh:2021:CBA

[631] Krithika Suresh, Jeremy M. G. Taylor, and Alexander Tsodikov. A copula-based approach for dynamic prediction of survival with a binary time-dependent covariate. Statistics in Medicine, 40(23):4931–4946, October 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhan:2021:PRA

[632] Tianyu Zhan, Lu Cui, Ziqian Geng, Lanju Zhang, Yihua Gu, and Ivan S. F. Chan. A practical response adaptive block randomization (RABR) design with analytic type I error protection. Statistics in Medicine, 40 (23):4947–4960, October 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tang:2021:SNC

[633] Yongqiang Tang. Some new confidence intervals for Kaplan–Meier based estimators from one and two sample survival data. Statistics in Medicine, 40(23):4961–4976, October 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kohli:2021:GMM

[634] Priya Kohli, Xinyu Du, and Haoyang Shen. Graphical models for mean and covariance of multivariate longitudinal data. Statistics in Medicine, 40(23):4977–4995, October 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Caniglia:2021:EOD

[635] Ellen C. Caniglia, Eleanor J. Murray, Miguel A. Hernán, and Zach Shahn. Estimating optimal dynamic treatment strategies under resource constraints using dynamic marginal structural models. Statistics in Medicine, 40(23):4996–5005, October 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Boe:2021:AQL

[636] Lillian A. Boe, Lesley F. Tinker, and Pamela A. Shaw. An approximate quasi-likelihood approach for error-prone failure time outcomes and exposures. Statistics in Medicine, 40(23):5006–5024, October 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gkatzionis:2021:BVS

[637] Apostolos Gkatzionis, Stephen Burgess, David V. Conti, and Paul J. Newcombe. Bayesian variable selection with a pleiotropic loss function in Mendelian randomization. Statistics in Medicine, 40(23):5025–5045, October 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mimi:2021:VSC

[638] Afsana Mimi and Md Hasinur Rahaman Khan. Variable selection for censored data using Modified Correlation Adjusted coRrelation (MCAR) scores. Statistics in Medicine, 40(23):5046–5064, October 15, 2021. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2021:TVP

[639] Jing Zhang, Jing Ning, Xuelin Huang, and Ruosha Li. On the timevarying predictive performance of longitudinal biomarkers: Measure and estimation. Statistics in Medicine, 40(23):5065–5077, October 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Burger:2021:NME

[640] Divan A. Burger and Emmanuel Lesaffre. Nonlinear mixed-effects modeling of longitudinal count data: Bayesian inference about median counts based on the marginal zero-inflated discrete Weibull distribution. Statistics in Medicine, 40(23):5078–5095, October 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gerard:2021:BMB

[641] Emma Gerard, Sarah Zohar, Christelle Lorenzato, Moreno Ursino, and Marie-Karelle Riviere. Bayesian modeling of a bivariate toxicity outcome for early phase oncology trials evaluating dose regimens. Statistics in Medicine, 40(23):5096–5114, October 15, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Boatman:2021:BSS

[642] Jeffrey A. Boatman, David M. Vock, and Joseph S. Koopmeiners. Borrowing from supplemental sources to estimate causal effects from a primary data source. Statistics in Medicine, 40(24):5115–5130, October 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Goldfeld:2021:PIP

[643] Keith S. Goldfeld, Danni Wu, Thaddeus Tarpey, Mengling Liu, Yinxiang Wu, Andrea B. Troxel, and Eva Petkova. Prospective individual patient data meta-analysis: Evaluating convalescent plasma for COVID-19. Statistics in Medicine, 40(24):5131–5151, October 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jang:2021:PCA

[644] Jeong Hoon Jang. Principal component analysis of hybrid functional and vector data. Statistics in Medicine, 40(24):5152–5173, October 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fazzari:2021:SDN

[645] Melissa J. Fazzari and Mimi Y. Kim. Subgroup discovery in noninferiority trials. Statistics in Medicine, 40(24):5174–5187, October 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hu:2021:ACC

[646] Zonghui Hu. Assessing conditional causal effect via characteristic score. Statistics in Medicine, 40(24):5188–5198, October 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2021:PBP

[647] Juhee Lee, Peter F. Thall, and Pavlos Msaouel. Precision Bayesian phase I–II dose-finding based on utilities tailored to prognostic subgroups. Statistics in Medicine, 40(24):5199–5217, October 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hattori:2021:SCI

[648] Satoshi Hattori and Xiao-Hua Zhou. Summary concordance index for meta-analysis of prognosis studies with a survival outcome. Statistics in Medicine, 40(24):5218–5236, October 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2021:ALP

[649] Lingxiao Wang, Richard Valliant, and Yan Li. Adjusted logistic propensity weighting methods for population inference using nonprobability volunteer-based epidemiologic cohorts. Statistics in Medicine, 40(24): 5237–5250, October 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Carvalho:2021:NPP

[650] Luiz Max Carvalho and Joseph G. Ibrahim. On the normalized power prior. Statistics in Medicine, 40(24):5251–5275, October 30, 2021. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zabriskie:2021:CCD

[651] Brinley N. Zabriskie, Chris Corcoran, and Pralay Senchaudhuri. A comparison of confidence distribution approaches for rare event metaanalysis. Statistics in Medicine, 40(24):5276–5297, October 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bie:2021:FMM

[652] Ruofan Bie, Sebastien Haneuse, Nathan Huey, Jonathan Schildcrout, and Glen McGee. Fitting marginal models in small samples: a simulation study of marginalized multilevel models and generalized estimating equations. Statistics in Medicine, 40(24):5298–5312, October 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pluta:2021:RPA

[653] Dustin Pluta, Tong Shen, Gui Xue, Chuansheng Chen, Hernando Ombao, and Zhaoxia Yu. Ridge-penalized adaptive Mantel test and its application in imaging genetics. Statistics in Medicine, 40(24):5313–5332, October 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tamhane:2021:GSH

[654] Ajit C. Tamhane, Dong Xi, and Jiangtao Gou. Group sequential Holm and Hochberg procedures. Statistics in Medicine, 40(24):5333–5350, October 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bartolucci:2021:MSA

[655] Francesco Bartolucci, Fulvia Pennoni, and Antonietta Mira. A multivariate statistical approach to predict COVID-19 count data with epidemiological interpretation and uncertainty quantification. Statistics in Medicine, 40(24):5351–5372, October 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bottaz-Bosson:2021:CPA

[656] Guillaume Bottaz-Bosson, Agnès Hamon, Jean-Louis Pépin, Sébastien Bailly, and Adeline Samson. Continuous positive airway pressure adherence trajectories in sleep apnea: Clustering with summed discrete Fréchet and dynamic time warping dissimilarities. Statistics in Medicine, 40 (24):5373–5396, October 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xu:2021:SSC

[657] Xiaohan Xu, Hong Zhu, Anh Q. Hoang, and Chul Ahn. Sample size considerations for matched-pair cluster randomization design with incomplete observations of binary outcomes. Statistics in Medicine, 40 (24):5397–5416, October 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Du:2021:LEH

[658] Yu Du, Huan Chen, and Ravi Varadhan. Lasso estimation of hierarchical interactions for analyzing heterogeneity of treatment effect. Statistics in Medicine, 40(25):5417–5433, November 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sanderson:2021:TCW

[659] Eleanor Sanderson, Wes Spiller, and Jack Bowden. Testing and correcting for weak and pleiotropic instruments in two-sample multivariable Mendelian randomization. Statistics in Medicine, 40(25):5434–5452, November 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sechidis:2021:UKC

[660] Konstantinos Sechidis, Matthias Kormaksson, and David Ohlssen. Using knockoffs for controlled predictive biomarker identification. Statistics in Medicine, 40(25):5453–5473, November 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Copas:2021:ODC

[661] Andrew J. Copas and Richard Hooper. Optimal design of cluster randomized trials allowing unequal allocation of clusters and unequal cluster size between arms. Statistics in Medicine, 40(25):5474–5486, November 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McGough:2021:PRL

[662] Sarah F. McGough, Devin Incerti, Svetlana Lyalina, Ryan Copping, Balasubramanian Narasimhan, and Robert Tibshirani. Penalized regression for left-truncated and right-censored survival data. Statistics in Medicine, 40(25):5487–5500, November 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Seipp:2021:WER

[663] Alexander Seipp, Verena Uslar, Dirk Weyhe, Antje Timmer, and Fabian Otto-Sobotka. Weighted expectile regression for right-censored data. Statistics in Medicine, 40(25):5501–5520, November 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Paukner:2021:WMS

[664] Mitchell Paukner and Richard Chappell. Window mean survival time. Statistics in Medicine, 40(25):5521–5533, November 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Long:2021:RSR

[665] Mingya Long, Liuquan Sun, and Qizhai Li. k-resolution sequential randomization procedure to improve covariates balance in a randomized experiment. Statistics in Medicine, 40(25):5534–5546, November 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yang:2021:GCN

[666] Hu Yang, Zhong Zhuang, and Wei Pan. A graph convolutional neural network for gene expression data analysis with multiple gene networks. Statistics in Medicine, 40(25):5547–5564, November 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Austin:2021:EAR

[667] Peter C. Austin, Donald B. Rubin, and Neal Thomas. Estimating adjusted risk differences by multiply-imputing missing control binary potential outcomes following propensity score-matching. Statistics in Medicine, 40(25):5565–5586, November 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zabriskie:2021:PBA

[668] Brinley N. Zabriskie, Chris Corcoran, and Pralay Senchaudhuri. A permutation-based approach for heterogeneous meta-analyses of rare events. Statistics in Medicine, 40(25):5587–5604, November 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bowden:2021:CIV

[669] Jack Bowden, Björn Bornkamp, Ekkehard Glimm, and Frank Bretz. Connecting instrumental variable methods for causal inference to the estimand framework. Statistics in Medicine, 40(25):5605–5627, November 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2021:CMA

[670] WenWu Wang, Jinfeng Xu, Joel Schwartz, Andrea Baccarelli, and Zhonghua Liu. Causal mediation analysis with latent subgroups. Statistics in Medicine, 40(25):5628–5641, November 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Petropoulou:2021:FSA

[671] Maria Petropoulou, Georgia Salanti, Gerta Rücker, Guido Schwarzer, Irini Moustaki, and Dimitris Mavridis. A forward search algorithm for detecting extreme study effects in network meta-analysis. Statistics in Medicine, 40(25):5642–5656, November 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jin:2021:UIP

[672] Huaqing Jin and Guosheng Yin. Unit information prior for adaptive information borrowing from multiple historical datasets. Statistics in Medicine, 40(25):5657–5672, November 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ge:2021:BEP

[673] Yunjiang Ge, Stephanie Hare, Gang Chen, James A. Waltz, Peter Kochunov, L. Elliot Hong, and Shuo Chen. Bayes estimate of primary threshold in clusterwise functional magnetic resonance imaging inferences. Statistics in Medicine, 40(25):5673–5689, November 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Noma:2021:CIP

[674] Hisashi Noma, Tomohiro Shinozaki, Katsuhiro Iba, Satoshi Teramukai, and Toshi A. Furukawa. Confidence intervals of prediction accuracy measures for multivariable prediction models based on the bootstrap-based optimism correction methods. Statistics in Medicine, 40(26):5691–5701, November 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Toenges:2021:CSA

[675] Gerrit Toenges, Tobias Mütze, and Antje Jahn-Eimermacher. A comparison of semiparametric approaches to evaluate composite endpoints in heart failure trials. Statistics in Medicine, 40(26):5702–5724, November 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Qiu:2021:EDS

[676] Peihua Qiu and Kai Yang. Effective disease surveillance by using covariate information. Statistics in Medicine, 40(26):5725–5745, November 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Coulombe:2021:EME

[677] Janie Coulombe, Erica E. M. Moodie, and Robert W. Platt. Estimating the marginal effect of a continuous exposure on an ordinal outcome using data subject to covariate-driven treatment and visit processes. Statistics in Medicine, 40(26):5746–5764, November 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chondros:2021:WSM

[678] Patty Chondros, Obioha C. Ukoumunne, Jane M. Gunn, and John B. Carlin. When should matching be used in the design of cluster randomized trials? Statistics in Medicine, 40(26):5765–5778, November 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Inacio:2021:RFI

[679] Vanda Inácio, Vanda M. Lourenço, Miguel de Carvalho, Richard A. Parker, and Vincent Gnanapragasam. Robust and flexible inference for the covariate-specific receiver operating characteristic curve. Statistics in Medicine, 40(26):5779–5795, November 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Speth:2021:RST

[680] Kelly Speth and Lu Wang. Restricted sub-tree learning to estimate an optimal dynamic treatment regime using observational data. Statistics in Medicine, 40(26):5796–5812, November 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Grant:2021:PRM

[681] Andrew J. Grant and Stephen Burgess. Pleiotropy robust methods for multivariable Mendelian randomization. Statistics in Medicine, 40(26): 5813–5830, November 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Roberti:2021:BCE

[682] Sander Roberti, Flora E. van Leeuwen, Michael Hauptmann, and Ruth M. Pfeiffer. Bias correction for estimates from linear excess relative risk models in small case-control studies. Statistics in Medicine, 40(26):5831–5852, November 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Phillips:2021:BAE

[683] Maile T. Phillips, James E. Meiring, Merryn Voysey, Joshua L. Warren, Stephen Baker, Buddha Basnyat, John D. Clemens, Christiane Dolecek, Sarah J. Dunstan, Gordon Dougan, Melita A. Gordon, Deus Thindwa, Robert S. Heyderman, Kathryn E. Holt, Firdausi Qadri, Andrew J. Pollard, Virginia E. Pitzer, and the Strataa Study Group. A Bayesian approach for estimating typhoid fever incidence from large-scale facilitybased passive surveillance data. Statistics in Medicine, 40(26):5853–5870, November 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2021:CLI

[684] Pingping Wang, Ting Fung Ma, Dipankar Bandyopadhyay, Yincai Tang, and Jun Zhu. Composite likelihood inference for ordinal periodontal data with replicated spatial patterns. Statistics in Medicine, 40(26): 5871–5893, November 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhou:2021:BCM

[685] Zhengyang Zhou, Minge Xie, David Huh, and Eun-Young Mun. A bias correction method in meta-analysis of randomized clinical trials with no adjustments for zero-inflated outcomes. Statistics in Medicine, 40(26): 5894–5909, November 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

vanNee:2021:FCD

[686] Mirrelijn M. van Nee, Lodewyk F. A. Wessels, and Mark A. van de Wiel. Flexible co-data learning for high-dimensional prediction. Statistics in Medicine, 40(26):5910–5925, November 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lado-Baleato:2021:MCR

[687] Óscar Lado-Baleato, Javier Roca-Pardiñas, Carmen Cadarso-Suárez, and Francisco Gude. Modeling conditional reference regions: Application to glycemic markers. Statistics in Medicine, 40(26):5926–5946, November 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sen:2021:BID

[688] Ananda Sen, Pin Li, Wen Ye, and Alfred Franzblau. Bayesian inference of dependent kappa for binary ratings. Statistics in Medicine, 40(26): 5947–5960, November 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hoogland:2021:TIT

[689] Jeroen Hoogland, Joanna IntHout, Michail Belias, Maroeska M. Rovers, Richard D. Riley, Frank E. Harrell Jr, Karel G. M. Moons, Thomas P. A. Debray, and Johannes B. Reitsma. A tutorial on individualized treatment effect prediction from randomized trials with a binary endpoint. Statistics in Medicine, 40(26):5961–5981, November 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fintzi:2021:AVD

[690] Jonathan Fintzi and Dean Follmann. Assessing vaccine durability in randomized trials following placebo crossover. Statistics in Medicine, 40 (27):5983–6007, November 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sullivan:2021:MIH

[691] Thomas R. Sullivan, Lisa N. Yelland, Margarita Moreno-Betancur, and Katherine J. Lee. Multiple imputation for handling missing outcome data in randomized trials involving a mixture of independent and paired data. Statistics in Medicine, 40(27):6008–6020, November 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cremaschi:2021:BNA

[692] Andrea Cremaschi, Maria De Iorio, Yap Seng Chong, Birit Broekman, Michael J. Meaney, and Michelle Z. L. Kee. A Bayesian nonparametric approach to dynamic item-response modeling: an application to the GUSTO cohort study. Statistics in Medicine, 40(27):6021–6037, November 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Song:2021:BHM

[693] Yanyi Song, Xiang Zhou, Jian Kang, Max T. Aung, Min Zhang, Wei Zhao, Belinda L. Needham, Sharon L. R. Kardia, Yongmei Liu, John D. Meeker, Jennifer A. Smith, and Bhramar Mukherjee. Bayesian hierarchical models for high-dimensional mediation analysis with coordinated selection of correlated mediators. Statistics in Medicine, 40(27):6038– 6056, November 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sanchez:2021:BPS

[694] Ricardo Sanchez, Beth Ann Griffin, Joseph Pane, and Daniel F. McCaffrey. Best practices in statistical computing. Statistics in Medicine, 40 (27):6057–6068, November 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Syriopoulou:2021:IPW

[695] Elisavet Syriopoulou, Mark J. Rutherford, and Paul C. Lambert. Inverse probability weighting and doubly robust standardization in the relative survival framework. Statistics in Medicine, 40(27):6069–6092, November 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nguyen:2021:MIS

[696] Cattram D. Nguyen, Margarita Moreno-Betancur, Laura Rodwell, Helena Romaniuk, John B. Carlin, and Katherine J. Lee. Multiple imputation of semi-continuous exposure variables that are categorized for analysis. Statistics in Medicine, 40(27):6093–6106, November 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

vanZwet:2021:SPR

[697] Erik van Zwet, Simon Schwab, and Stephen Senn. The statistical properties of RCTs and a proposal for shrinkage. Statistics in Medicine, 40 (27):6107–6117, November 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Beesley:2021:ARM

[698] Lauren J. Beesley and Jeremy M. G. Taylor. Accounting for not-at-random missingness through imputation stacking. Statistics in Medicine, 40(27):6118–6132, November 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2021:SSR

[699] Peijin Wang and Shein-Chung Chow. Sample size re-estimation in clinical trials. Statistics in Medicine, 40(27):6133–6149, November 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Austin:2021:IPS

[700] Peter C. Austin. Informing power and sample size calculations when using inverse probability of treatment weighting using the propensity score. Statistics in Medicine, 40(27):6150–6163, November 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tang:2021:SAT

[701] Ming Tang, Lu Wang, Michael A. Gorin, and Jeremy M. G. Taylor. Stepadjusted tree-based reinforcement learning for evaluating nested dynamic treatment regimes using test-and-treat observational data. Statistics in Medicine, 40(27):6164–6177, November 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Signorelli:2021:PRC

[702] Mirko Signorelli, Pietro Spitali, Cristina Al-Khalili Szigyarto, The Mark-Md Consortium, and Roula Tsonaka. Penalized regression calibration: a method for the prediction of survival outcomes using complex longitudinal and high-dimensional data. Statistics in Medicine, 40(27):6178–6196, November 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fonseca-Rodriguez:2021:ABS

[703] Osvaldo Fonseca-Rodríguez, Anne-Marie Fors Connolly, Ioannis Katsoularis, Krister Lindmark, and Paddy Farrington. Avoiding bias in self-controlled case series studies of coronavirus disease 2019. Statistics in Medicine, 40(27):6197–6208, November 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Grinsztajn:2021:BWD

[704] Léo Grinsztajn, Elizaveta Semenova, Charles C. Margossian, and Julien Riou. Bayesian workflow for disease transmission modeling in Stan. Statistics in Medicine, 40(27):6209–6234, November 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McCaw:2021:CCI

[705] Zachary R. McCaw, Lu Tian, Jiawei Wei, Brian Lee Claggett, Frank Bretz, Garrett Fitzmaurice, and Lee-Jen Wei. Choosing clinically interpretable summary measures and robust analytic procedures for quantifying the treatment difference in comparative clinical studies. Statistics in Medicine, 40(28):6235–6242, December 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kim:2021:PGV

[706] Seonjin Kim, Hyunkeun Ryan Cho, and Mi-Ok Kim. Predictive generalized varying-coefficient longitudinal model. Statistics in Medicine, 40 (28):6243–6259, December 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dahlgren:2021:HTI

[707] F. Scott Dahlgren, Ivo M. Foppa, Melissa S. Stockwell, Celibell Y. Vargas, Philip LaRussa, and Carrie Reed. Household transmission of influenza A and B within a prospective cohort during the 2013–2014 and 2014–2015 seasons. Statistics in Medicine, 40(28):6260–6276, December 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yoneoka:2021:GWG

[708] Daisuke Yoneoka, Takayuki Kawashima, Koji Makiyama, Yuta Tanoue, Shuhei Nomura, and Akifumi Eguchi. Geographically weighted generalized farrington algorithm for rapid outbreak detection over short data accumulation periods. Statistics in Medicine, 40(28):6277–6294, December 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mandal:2021:ISD

[709] Soutrik Mandal, Jing Qin, and Ruth M. Pfeiffer. Incorporating survival data into case-control studies with incident and prevalent cases. Statistics in Medicine, 40(28):6295–6308, December 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2021:SEA

[710] Shuwei Li, Tian Tian, Tao Hu, and Jianguo Sun. A simulationextrapolation approach for regression analysis of misclassified current status data with the additive hazards model. Statistics in Medicine, 40 (28):6309–6320, December 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2021:QFS

[711] Xuan Wang, Tianxi Cai, Lu Tian, Florence Bourgeois, and Layla Parast. Quantifying the feasibility of shortening clinical trial duration using surrogate markers. Statistics in Medicine, 40(28):6321–6343, December 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Aubel:2021:BAE

[712] Paul Aubel, Marine Antigny, Ronan Fougeray, Frédéric Dubois, and Gaëlle Saint-Hilary. A Bayesian approach for event predictions in clinical trials with time-to-event outcomes. Statistics in Medicine, 40(28): 6344–6359, December 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2021:KDE

[713] Siyun Liu and Tao Yu. Kernel density estimation in mixture models with known mixture proportions. Statistics in Medicine, 40(28):6360– 6372, December 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bloore:2021:ULR

[714] Katherine Bloore, Yang Song, Howard Cabral, Joseph Massaro, and Michael LaValley. Use of the likelihood reduction factor in a path analysis framework to quantify surrogacy in clinical trials. Statistics in Medicine, 40(28):6373–6386, December 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pal:2021:SSE

[715] Suvra Pal. A simplified stochastic EM algorithm for cure rate model with negative binomial competing risks: an application to breast cancer data. Statistics in Medicine, 40(28):6387–6409, December 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nevalainen:2021:FMU

[716] Jaakko Nevalainen, Somnath Datta, Jorma Toppari, Jorma Ilonen, Heikki Hyöty, Riitta Veijola, Mikael Knip, and Suvi M. Virtanen. Frailty modeling under a selective sampling protocol: an application to type 1 diabetes related autoantibodies. Statistics in Medicine, 40(28):6410–6420, December 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhou:2021:IED

[717] Tianjian Zhou and Yuan Ji. Incorporating external data into the analysis of clinical trials via Bayesian additive regression trees. Statistics in Medicine, 40(28):6421–6442, December 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Silan:2021:MPB

[718] Margherita Silan, Giovanna Boccuzzo, and Bruno Arpino. Matching on poset-based average rank for multiple treatments to compare many unbalanced groups. Statistics in Medicine, 40(28):6443–6458, December 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Momenyan:2021:JAL

[719] Somayeh Momenyan. Joint analysis of longitudinal measurements and spatially clustered competing risks HIV/AIDS data. Statistics in Medicine, 40(28):6459–6477, December 10, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Konstantinou:2021:SME

[720] Konstantinos Konstantinou and Aila Särkkä. Spatial modeling of epidermal nerve fiber patterns. Statistics in Medicine, 40(29):6479–6500, December 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Iskandar:2021:PBA

[721] Rowan Iskandar. Probability bound analysis: a novel approach for quantifying parameter uncertainty in decision-analytic modeling and costeffectiveness analysis. Statistics in Medicine, 40(29):6501–6522, December 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wei:2021:IIS

[722] Yue Wei, Jason C. Hsu, Wei Chen, Emily Y. Chew, and Ying Ding. Identification and inference for subgroups with differential treatment efficacy from randomized controlled trials with survival outcomes through multiple testing. Statistics in Medicine, 40(29):6523–6540, December 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rakhmawati:2021:PVS

[723] Trias W. Rakhmawati, Il Do Ha, Hangbin Lee, and Youngjo Lee. Penalized variable selection for cause-specific hazard frailty models with clustered competing-risks data. Statistics in Medicine, 40(29):6541–6557, December 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fang:2021:RMO

[724] Yuexin Fang, Baqun Zhang, and Min Zhang. Robust method for optimal treatment decision making based on survival data. Statistics in Medicine, 40(29):6558–6576, December 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2021:LRW

[725] Chenguang Wang, Gary L. Rosner, Tingting Bao, Nelson Lu, Wei-Chen Chen, Heng Li, Ram Tiwari, Yunling Xu, and Lilly Q. Yue. Leveraging real-world evidence for determining performance goals for medical device studies. Statistics in Medicine, 40(29):6577–6589, December 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

He:2021:MPH

[726] Haijin He, Dongxiao Han, Xinyuan Song, and Liuquan Sun. Mixture proportional hazards cure model with latent variables. Statistics in Medicine, 40(29):6590–6604, December 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Roberts:2021:IBC

[727] Emily K. Roberts, Michael R. Elliott, and Jeremy M. G. Taylor. Incorporating baseline covariates to validate surrogate endpoints with a constant biomarker under control arm. Statistics in Medicine, 40(29): 6605–6618, December 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Qin:2021:GGI

[728] Xing Qin, Shuangge Ma, and Mengyun Wu. Gene-gene interaction analysis incorporating network information via a structured Bayesian approach. Statistics in Medicine, 40(29):6619–6633, December 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pham:2021:CMA

[729] Tra My Pham, Ian R. White, Brennan C. Kahan, Tim P. Morris, Simon J. Stanworth, and Gordon Forbes. A comparison of methods for analyzing a binary composite endpoint with partially observed components in randomized controlled trials. Statistics in Medicine, 40(29): 6634–6650, December 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xiong:2021:ISH

[730] Wei Xiong and Han Pan. Interaction screening for high-dimensional heterogeneous data via robust hybrid metrics. Statistics in Medicine, 40 (29):6651–6673, December 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Harrison:2021:PCA

[731] Linda J. Harrison and Rui Wang. Power calculation for analyses of crosssectional stepped-wedge cluster randomized trials with binary outcomes via generalized estimating equations. Statistics in Medicine, 40(29):6674– 6688, December 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

You:2021:JMM

[732] Lu You and Peihua Qiu. Joint modeling of multivariate nonparametric longitudinal data and survival data: a local smoothing approach. Statistics in Medicine, 40(29):6689–6706, December 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jin:2021:PLS

[733] Peng Jin and Mengling Liu. Partially linear single-index generalized mean residual life models. Statistics in Medicine, 40(29):6707–6722, December 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gomez:2021:NPR

[734] Yolanda M. Gómez, Diego I. Gallardo, Jeremias Leão, and Vinicius F. Calsavara. On a new piecewise regression model with cure rate: Diagnostics and application to medical data. Statistics in Medicine, 40(29): 6723–6742, December 20, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bartos:2021:BMA

[735] Frantisek Bartos, Quentin F. Gronau, Bram Timmers, Willem M. Otte, Alexander Ly, and Eric-Jan Wagenmakers. Bayesian model-averaged meta-analysis in medicine. Statistics in Medicine, 40(30):6743–6761, December 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhou:2021:DWT

[736] Minchun Zhou, Brian D. Boyd, Warren D. Taylor, and Hakmook Kang. Double-wavelet transform for multi-subject resting state functional magnetic resonance imaging data. Statistics in Medicine, 40(30):6762–6776, December 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Han:2021:CMI

[737] Kyunghee Han, Pamela A. Shaw, and Thomas Lumley. Combining multiple imputation with raking of weights: an efficient and robust approach in the setting of nearly true models. Statistics in Medicine, 40(30):6777– 6791, December 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Espin-Garcia:2021:TPS

[738] Osvaldo Espin-Garcia, Radu V. Craiu, and Shelley B. Bull. Two-phase sample selection strategies for design and analysis in post-genome-wide association fine-mapping studies. Statistics in Medicine, 40(30):6792– 6817, December 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2021:RGV

[739] Yang Li, Rong Li, Yichen Qin, Cunjie Lin, and Yuhong Yang. Robust group variable screening based on maximum lq-likelihood estimation. Statistics in Medicine, 40(30):6818–6834, December 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fu:2021:RAV

[740] Liya Fu, Jiaqi Li, and You-Gan Wang. Robust approach for variable selection with high dimensional longitudinal data analysis. Statistics in Medicine, 40(30):6835–6854, December 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zou:2021:BID

[741] Haotian Zou, Kan Li, Donglin Zeng, Sheng Luo, and Alzheimer’s Disease Neuroimaging Initiative. Bayesian inference and dynamic prediction of multivariate joint model with functional data: an application to Alzheimer’s disease. Statistics in Medicine, 40(30):6855–6872, December 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

vandeVen:2021:CST

[742] Peter M. van de Ven, Andrea Bassi, and Johannes Berkhof. Comparing the sensitivities of two screening tests in nonblinded randomized paired screen-positive trials with differential screening uptake. Statistics in Medicine, 40(30):6873–6884, December 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nuno:2021:CRT

[743] Michelle M. Nuño and Daniel L. Gillen. Censoring-robust timedependent receiver operating characteristic curve estimators. Statistics in Medicine, 40(30):6885–6899, December 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2021:EOT

[744] Xiaofei Chen, Daniel F. Heitjan, Gerald Greil, and Haekyung Jeon-Slaughter. Estimating the optimal timing of surgery by imputing potential outcomes. Statistics in Medicine, 40(30):6900–6917, December 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yang:2021:SRA

[745] Dian Yang, Mingyue Du, and Jianguo Sun. Semiparametric regression analysis of clustered interval-censored failure time data with a cured subgroup. Statistics in Medicine, 40(30):6918–6930, December 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McAndrew:2021:ASE

[746] Thomas McAndrew and Nicholas G. Reich. Adaptively stacking ensembles for influenza forecasting. Statistics in Medicine, 40(30):6931–6952, December 30, 2021. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Arambepola:2022:SSD

[747] Rohan Arambepola, Tim C. D. Lucas, Anita K. Nandi, Peter W. Gething, and Ewan Cameron. A simulation study of disaggregation regression for spatial disease mapping. Statistics in Medicine, 41(1):1–16, January 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lim:2022:BSJ

[748] Woobeen Lim, Michael L. Pennell, Michelle J. Naughton, and Electra D. Paskett. Bayesian semiparametric joint modeling of longitudinal explanatory variables of mixed types and a binary outcome. Statistics in Medicine, 41(1):17–36, January 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yin:2022:JIA

[749] Jingjing Yin, Hani Samawi, and Lili Tian. Joint inference about the AUC and Youden index for paired biomarkers. Statistics in Medicine, 41(1): 37–64, January 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rosenman:2022:PSM

[750] Evan T. R. Rosenman, Art B. Owen, Mike Baiocchi, and Hailey R. Banack. Propensity score methods for merging observational and experimental datasets. Statistics in Medicine, 41(1):65–86, January 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wu:2022:ULE

[751] Yuh-Jenn Wu, Yu-Chieh Cheng, Chieh Chiang, Li-Hsueh Cheng, Ching-Ti Liu, and Chin-Fu Hsiao. Use of likelihood estimates for variances for the design and evaluation of multiregional clinical trials with heterogeneous variances. Statistics in Medicine, 41(1):87–107, January 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lu:2022:BCC

[752] Zihang Lu and Wendy Lou. Bayesian consensus clustering for multivariate longitudinal data. Statistics in Medicine, 41(1):108–127, January 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ma:2022:REI

[753] Xiangmei Ma, Paul Milligan, Kwok Fai Lam, and Yin Bun Cheung. Ratio estimators of intervention effects on event rates in cluster randomized trials. Statistics in Medicine, 41(1):128–145, January 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lin:2022:IGC

[754] Pei-Sheng Lin. Identification of geographic clusters for temporal heterogeneity with application to dengue surveillance. Statistics in Medicine, 41(1):146–162, January 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Guolo:2022:MEC

[755] Annamaria Guolo. Measurement errors in control risk regression: a comparison of correction techniques. Statistics in Medicine, 41(1):163–179, January 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yuan:2022:SRA

[756] Ao Yuan, Lida Wang, and Ming T. Tan. Set-regression with applications to subgroup analysis. Statistics in Medicine, 41(1):180–193, January 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tang:2022:MCI

[757] Yongqiang Tang. MOVER confidence intervals for a difference or ratio effect parameter under stratified sampling. Statistics in Medicine, 41(1):

194–207, January 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Silva:2022:MIP

[758] Gabriella C. Silva and Roee Gutman. Multiple imputation procedures for estimating causal effects with multiple treatments with application to the comparison of healthcare providers. Statistics in Medicine, 41(1): 208–226, January 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Choi:2022:WLS

[759] Sangbum Choi, Taehwa Choi, Hyunsoon Cho, and Dipankar Bandyopadhyay. Weighted least-squares regression with competing risks data. Statistics in Medicine, 41(2):227–241, January 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Walter:2022:ESD

[760] Stephen D. Walter, Jan Rychtár, Dewey Taylor, and Narayanaswamy Balakrishnan. Estimation of standard deviations and inverse-variance weights from an observed range. Statistics in Medicine, 41(2):242–257, January 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Su:2022:RAT

[761] Pei-Fang Su. Response-adaptive treatment allocation for clinical studies with recurrent event and terminal event data. Statistics in Medicine, 41(2):258–275, January 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mrkvicka:2022:NMM

[762] Tomás Mrkvicka, Mari Myllymäki, Mikko Kuronen, and Naveen Naidu Narisetty. New methods for multiple testing in permutation inference for the general linear model. Statistics in Medicine, 41(2):276–297, January 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Takeda:2022:CHB

[763] Kentaro Takeda, Shufang Liu, and Alan Rong. Constrained hierarchical Bayesian model for latent subgroups in basket trials with two classifiers. Statistics in Medicine, 41(2):298–309, January 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Du:2022:ODT

[764] Jiacong Du, Lauren J. Beesley, Seunggeun Lee, Xiang Zhou, Walter Dempsey, and Bhramar Mukherjee. Optimal diagnostic test allocation strategy during the COVID-19 pandemic and beyond. Statistics in Medicine, 41(2):310–327, January 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2022:CDF

[765] Hongkai Li, Jinzhu Jia, Ran Yan, Fuzhong Xue, and Zhi Geng. A causal data fusion method for the general exposure and outcome. Statistics in Medicine, 41(2):328–339, January 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ollier:2022:PPM

[766] Edouard Ollier, Pierre Blanchard, Gwénaël Le Teuff, and Stefan Michiels. Penalized Poisson model for network meta-analysis of individual patient time-to-event data. Statistics in Medicine, 41(2):340–355, January 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kang:2022:JMM

[767] Kai Kang, Deng Pan, and Xinyuan Song. A joint model for multivariate longitudinal and survival data to discover the conversion to Alzheimer’s disease. Statistics in Medicine, 41(2):356–373, January 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mu:2022:BPI

[768] Rongji Mu, Jin Xu, Rui (Sammi) Tang, Scott Kopetz, and Ying Yuan. A Bayesian phase I/II platform design for co-developing drug combination therapies for multiple indications. Statistics in Medicine, 41(2):374–389, January 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lane:2022:CII

[769] Adam Lane. Conditional information and inference in response-adaptive allocation designs. Statistics in Medicine, 41(2):390–406, January 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Smith:2022:ICC

[770] Matthew J. Smith, Mohammad A. Mansournia, Camille Maringe, Paul N. Zivich, Stephen R. Cole, Clémence Leyrat, Aurélien Belot, Bernard Rachet, and Miguel A. Luque-Fernandez. Introduction to computational causal inference using reproducible Stata, R, and Python code: a tutorial. Statistics in Medicine, 41(2):407–432, January 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2022:EHW

[771] Yu-Lun Liu, Gui-Shuang Ying, Graham E. Quinn, Xiao-Hua Zhou, and Yong Chen. Extending Hui–Walter framework to correlated outcomes with application to diagnosis tests of an eye disease among premature infants. Statistics in Medicine, 41(3):433–448, February 10, 2022. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

deSouza:2022:CGM

[772] Roberta de Souza and Carlos Alberto Ribeiro Diniz. Correlated geometric models of order k and its application to intensive care unit and leprosy data. Statistics in Medicine, 41(3):449–470, February 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nestler:2022:EME

[773] Steffen Nestler. An extension of the mixed-effects growth model that considers between-person differences in the within-subject variance and the autocorrelation. Statistics in Medicine, 41(3):471–482, February 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jin:2022:BSM

[774] Jin Jin, Lin Zhang, Ethan Leng, Gregory J. Metzger, and Joseph S. Koopmeiners. Bayesian spatial models for voxel-wise prostate cancer classification using multi-parametric magnetic resonance imaging data. Statistics in Medicine, 41(3):483–499, February 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2022:PAR

[775] Yipeng Wang, Lifeng Lin, Christopher G. Thompson, and Haitao Chu. A penalization approach to random-effects meta-analysis. Statistics in Medicine, 41(3):500–516, February 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shen:2022:CAM

[776] Xiaoxi Shen, Yalu Wen, Yuehua Cui, and Qing Lu. A conditional autoregressive model for genetic association analysis accounting for genetic heterogeneity. Statistics in Medicine, 41(3):517–542, February 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xie:2022:CGG

[777] Shanghong Xie, Erin McDonnell, and Yuanjia Wang. Conditional Gaussian graphical model for estimating personalized disease symptom networks. Statistics in Medicine, 41(3):543–553, February 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2022:MIS

[778] Yifei Wang, Daniel J. Tancredi, and Diana L. Miglioretti. Marginal indirect standardization using latent clustering on multiple hospitals. Statistics in Medicine, 41(3):554–566, February 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2022:IMA

[779] Chixiang Chen, Peisong Han, and Fan He. Improving main analysis by borrowing information from auxiliary data. Statistics in Medicine, 41 (3):567–579, February 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2022:MLC

[780] Xiaokang Liu, Xiaomei Cong, Gen Li, Kendra Maas, and Kun Chen. Multivariate log-contrast regression with sub-compositional predictors: Testing the association between preterm infants’ gut microbiome and neurobehavioral outcomes. Statistics in Medicine, 41(3):580–594, February 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tang:2022:BMN

[781] Xiaoyu Tang and Ludovic Trinquart. Bayesian multivariate network meta-analysis model for the difference in restricted mean survival times. Statistics in Medicine, 41(3):595–611, February 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Austin:2022:UFP

[782] Peter C. Austin, Jiming Fang, and Douglas S. Lee. Using fractional polynomials and restricted cubic splines to model non-proportional hazards or time-varying covariate effects in the Cox regression model. Statistics in Medicine, 41(3):612–624, February 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pedder:2022:MAE

[783] Hugo Pedder, Sofia Dias, Martin Boucher, Meg Bennetts, David Mawdsley, and Nicky J. Welton. Methods to assess evidence consistency in doseresponse model based network meta-analysis. Statistics in Medicine, 41 (4):625–644, February 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tian:2022:SSC

[784] Zizhong Tian, Denise Esserman, Guangyu Tong, Ondrej Blaha, James Dziura, Peter Peduzzi, and Fan Li. Sample size calculation in hierarchical 2 × 2 factorial trials with unequal cluster sizes. Statistics in Medicine, 41(4):645–664, February 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xu:2022:RFI

[785] Can Xu, Xinlei Wang, Johan Lim, Guanghua Xiao, and Yang Xie. RCRdiff: a fully integrated Bayesian method for differential expression analysis using raw NanoString nCounter data. Statistics in Medicine, 41 (4):665–680, February 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gecili:2022:BRN

[786] Emrah Gecili, Siva Sivaganesan, Ozgur Asar, John P. Clancy, Assem Ziady, and Rhonda D. Szczesniak. Bayesian regularization for a nonstationary Gaussian linear mixed effects model. Statistics in Medicine, 41 (4):681–697, February 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kotalik:2022:GSR

[787] Ales Kotalik, David M. Vock, Brian P. Hobbs, and Joseph S. Koopmeiners. A group-sequential randomized trial design utilizing supplemental trial data. Statistics in Medicine, 41(4):698–718, February 20, 2022. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Doubleday:2022:RCD

[788] Kevin Doubleday, Jin Zhou, Hua Zhou, and Haoda Fu. Risk controlled decision trees and random forests for precision medicine. Statistics in Medicine, 41(4):719–735, February 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hyun:2022:SGR

[789] Noorie Hyun, David J. Couper, and Donglin Zeng. A semiparametric Gumbel regression model for analyzing longitudinal data with nonnormal tails. Statistics in Medicine, 41(4):736–750, February 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhou:2022:POP

[790] Jie Zhou, Xun Jiang, Hong Amy Xia, Peng Wei, and Brian P. Hobbs. Predicting outcomes of phase III oncology trials with Bayesian mediation modeling of tumor response. Statistics in Medicine, 41(4):751–768, February 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Butera:2022:DRM

[791] Nicole M. Butera, Donglin Zeng, Annie Green Howard, Penny Gordon-Larsen, and Jianwen Cai. A doubly robust method to handle missing multilevel outcome data with application to the China Health and Nutrition Survey. Statistics in Medicine, 41(4):769–785, February 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wu:2022:CIT

[792] Jianrong Wu and Jing Wei. Cancer immunotherapy trial design with random delayed treatment effect and cure rate. Statistics in Medicine, 41(4):786–797, February 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tang:2022:CST

[793] Yongqiang Tang. Complex survival trial design by the product integration method. Statistics in Medicine, 41(4):798–814, February 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2022:MMD

[794] Bosheng Li, Liwen Su, Yuqing Ye, and Fangrong Yan. M&M: a maximum duration design with the Maxcombo test for a group sequential trial of an immunotherapy with a random delayed treatment effect. Statistics in Medicine, 41(4):815–830, February 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2022:SA

[795] Anonymous. SIM 40th anniversary. Statistics in Medicine, 41(5):831, February 28, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2022:F

[796] Anonymous. Foreword. Statistics in Medicine, 41(5):832, February 28, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Armitage:2022:PAS

[797] Peter Armitage. Peter Armitage speaking on the occasion of the 50th anniversary of the M.Sc. in Medical Statistics, LSHTM (for the symposium on April 11–12, 2019). Statistics in Medicine, 41(5):833–834, February 28, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ashby:2022:CTE

[798] Deborah Ashby. Commentary on training and education in medical statistics, in celebration of 40 years of statistics in medicine and 50 years of the MSc medical statistics at LSHTM. Statistics in Medicine, 41(5): 835–837, February 28, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Carpenter:2022:TCB

[799] James R. Carpenter, Jim Todd, Kathy Baisley, John Bradley, Nazarius Mbona Tumwesigye, Patrick Musonda, and Tobias Chirwa. Training and capacity building in medical statistics in Sub-Saharan Africa: Impact of the London School of Hygiene and Tropical Medicine MSc in Medical Statistics, 1969 to 2021. Statistics in Medicine, 41(5): 838–844, February 28, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Duchateau:2022:TDS

[800] Luc Duchateau and Paul Janssen. Two decades of statistical education collaboration in the global South: Lessons learned from an Ethiopian project and the way forward. Statistics in Medicine, 41(5):845–846, February 28, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

LeBlanc:2022:SAP

[801] Marissa LeBlanc, Corina S. Rueegg, Nural Bekiroglu, Tonya M. Esterhuizen, Morten W. Fagerland, Ragnhild S. Falk, Kathrine F. Frøslie, Erika Graf, Georg Heinze, Ulrike Held, René Holst, Theis Lange, Madhu Mazumdar, Ida H. Myrberg, Martin Posch, Jamie C. Sergeant, Werner Vach, Eric A. Vance, Harald Weedon-Fekjær, and Manuela Zucknick. Statistical advising: Professional development opportunities for the biostatistician. Statistics in Medicine, 41(5):847–859, February 28, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Devick:2022:BKM

[802] Katrina L. Devick, Jennifer F. Bobb, Maitreyi Mazumdar, Birgit Claus Henn, David C. Bellinger, David C. Christiani, Robert O. Wright, Paige L. Williams, Brent A. Coull, and Linda Valeri. Bayesian kernel machine regression-causal mediation analysis. Statistics in Medicine, 41 (5):860–876, February 28, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kunzmann:2022:CPF

[803] Kevin Kunzmann, Michael J. Grayling, Kim May Lee, David S. Robertson, Kaspar Rufibach, and James M. S. Wason. Conditional power and friends: The why and how of (un)planned, unblinded sample size recalculations in confirmatory trials. Statistics in Medicine, 41(5):877–890, February 28, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Westphal:2022:MTF

[804] Max Westphal, Antonia Zapf, and Werner Brannath. A multiple testing framework for diagnostic accuracy studies with co-primary endpoints. Statistics in Medicine, 41(5):891–909, February 28, 2022. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cao:2022:APL

[805] Zhiqiang Cao and Man Yu Wong. Approximate profile likelihood estimation for Cox regression with covariate measurement error. Statistics in Medicine, 41(5):910–931, February 28, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tang:2022:C

[806] Man Lai Tang and Hon Keung Tony Ng. Correction. Statistics in Medicine, 41(5):932, February 28, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Suder:2022:SAS

[807] Piotr M. Suder and Aaron J. Molstad. Scalable algorithms for semiparametric accelerated failure time models in high dimensions. Statistics in Medicine, 41(6):933–949, March 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yu:2022:SSFa

[808] Ron Xiaolong Yu and Jitendra Ganju. Sample size formula for a win ratio endpoint. Statistics in Medicine, 41(6):950–963, March 15, 2022.

CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See comments [908, 909].

Zhao:2022:BVE

[809] Yi Zhao, Brian S. Caffo, and Joshua B. Ewen. B-value and empirical equivalence bound: a new procedure of hypothesis testing. Statistics in Medicine, 41(6):964–980, March 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Man:2022:ALB

[810] Irene Man, Johannes A. Bogaards, Kishan Makwana, Krzysztof Trzciński, and Kari Auranen. Approximate likelihood-based estimation method of multiple-type pathogen interactions: an application to longitudinal pneumococcal carriage data. Statistics in Medicine, 41(6): 981–993, March 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2022:CAS

[811] Lu Wang and Xiao-Hua Zhou. Comparing the accuracy of screening tests with verification of disease status restricted to test positives. Statistics in Medicine, 41(6):994–1008, March 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Im:2022:BHF

[812] Yunju Im, Yuan Huang, Jian Huang, and Shuangge Ma. Bayesian hierarchical finite mixture of regression for histopathological imaging-based cancer data analysis. Statistics in Medicine, 41(6):1009–1022, March 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wan:2022:CUL

[813] Fei Wan. Conditional or unconditional logistic regression for frequency matched case-control design? Statistics in Medicine, 41(6):1023–1041, March 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yamaguchi:2022:MIL

[814] Yusuke Yamaguchi, Satoshi Yoshida, Toshihiro Misumi, and Kazushi Maruo. Multiple imputation for longitudinal data using Bayesian lasso imputation model. Statistics in Medicine, 41(6):1042–1058, March 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Razaee:2022:NBM

[815] Zahra S. Razaee, Galen Cook-Wiens, and Mourad Tighiouart. A nonparametric Bayesian method for dose finding in drug combinations cancer trials. Statistics in Medicine, 41(6):1059–1080, March 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Grayling:2022:RAI

[816] Michael J. Grayling, James M. S. Wason, and Sofía S. Villar. Response adaptive intervention allocation in stepped-wedge cluster randomized trials. Statistics in Medicine, 41(6):1081–1099, March 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shapland:2022:PLB

[817] Chin Yang Shapland, Qingyuan Zhao, and Jack Bowden. Profilelikelihood Bayesian model averaging for two-sample summary data Mendelian randomization in the presence of horizontal pleiotropy. Statistics in Medicine, 41(6):1100–1119, March 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hughes:2022:CIE

[818] James P. Hughes, Brian D. Williamson, Chloe Krakauer, Gordon Chau, Brayan Ortiz, Jon Wakefield, Craig Hendrix, K. Rivet Amico, Timothy H. Holtz, Linda-Gail Bekker, and Robert Grant. Combining information to estimate adherence in studies of pre-exposure prophylaxis for HIV prevention: Application to HPTN 067. Statistics in Medicine, 41 (6):1120–1136, March 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2022:LCT

[819] Kewei Chen, Xiaojuan Guo, Rong Pan, Chengjie Xiong, Danielle J. Harvey, Yinghua Chen, Li Yao, Yi Su, Eric M. Reiman, and for the Alzheimer’s Disease Neuroimaging Initiative. Limitations of clinical trial sample size estimate by subtraction of two measurements. Statistics in Medicine, 41(7):1137–1147, March 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2022:WGE

[820] Meng Liu and Yang Zhao. Weighted generalized estimating equations and unified estimation for longitudinal data with nonmonotone missing data patterns. Statistics in Medicine, 41(7):1148–1156, March 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Matsuura:2022:OAA

[821] Kentaro Matsuura, Junya Honda, Imad El Hanafi, Takashi Sozu, and Kentaro Sakamaki. Optimal adaptive allocation using deep reinforcement learning in a dose-response study. Statistics in Medicine, 41(7):1157– 1171, March 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2022:MIC

[822] Chung-Han Lee and Hsiuying Wang. Multiple imputation confidence intervals for the mean of the discrete distributions for incomplete data. Statistics in Medicine, 41(7):1172–1190, March 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cui:2022:SRS

[823] Erjia Cui, E. Christi Thompson, Raymond J. Carroll, and David Ruppert. A semiparametric risk score for physical activity. Statistics in Medicine, 41(7):1191–1204, March 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Guo:2022:BBB

[824] Beibei Guo and Yong Zang. BIPSE: a biomarker-based phase I/II design for immunotherapy trials with progression-free survival endpoint. Statistics in Medicine, 41(7):1205–1224, March 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yang:2022:ESW

[825] Song Yang, James Troendle, Daewoo Pak, and Eric Leifer. Event-specific win ratios for inference with terminal and non-terminal events. Statistics in Medicine, 41(7):1225–1241, March 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jin:2022:FSC

[826] Zhuxuan Jin, Jian Kang, and Tianwei Yu. Feature selection and classification over the network with missing node observations. Statistics in Medicine, 41(7):1242–1262, March 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2022:BTM

[827] Chunjie Wang, Jingjing Jiang, and Xinyuan Song. Bayesian transformation models with partly interval-censored data. Statistics in Medicine, 41(7):1263–1279, March 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Riley:2022:MSS

[828] Richard D. Riley, Gary S. Collins, Joie Ensor, Lucinda Archer, Sarah Booth, Sarwar I. Mozumder, Mark J. Rutherford, Maarten van Smeden, Paul C. Lambert, and Kym I. E. Snell. Minimum sample size calculations for external validation of a clinical prediction model with a time-to-event outcome. Statistics in Medicine, 41(7):1280–1295, March 30, 2022. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2022:EIM

[829] Dongyu Li, Lei Wang, and Weihua Zhao. Estimation and inference for multikink expectile regression with longitudinal data. Statistics in Medicine, 41(7):1296–1313, March 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Clements:2022:RSB

[830] Mark Clements, Benjamin Christoffersen, Patrick Royston, and Michael Crowther. Re: Spline-based accelerated failure time model. Statistics in Medicine, 41(7):1314–1315, March 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [357].

Mrkvicka:2022:CDB

[831] Tomás Mrkvicka and Mari Myllymäki. Comment on “A depth-based global envelope test for comparing two groups of functions with applications to biomedical data” by s. lopez-pintado and k. qian. Statistics in Medicine, 41(7):1316–1317, March 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [421].

Dablander:2022:PPT

[832] Fabian Dablander, Karoline Huth, Quentin F. Gronau, Alexander Etz, and Eric-Jan Wagenmakers. A puzzle of proportions: Two popular Bayesian tests can yield dramatically different conclusions. Statistics in Medicine, 41(8):1319–1333, April 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Edlinger:2022:RPM

[833] Michael Edlinger, Maarten van Smeden, Hannes F. Alber, Maria Wanitschek, and Ben Van Calster. Risk prediction models for discrete ordinal outcomes: Calibration and the impact of the proportional odds assumption. Statistics in Medicine, 41(8):1334–1360, April 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Han:2022:DNO

[834] Gang Han, Michael J. Schell, Emily S. Reisenbichler, Bohong Guo, and David L. Rimm. Determination of the number of observers needed to evaluate a subjective test and its application in two PD-L1 studies. Statistics in Medicine, 41(8):1361–1375, April 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tong:2022:AUC

[835] Guangyu Tong, Denise Esserman, and Fan Li. Accounting for unequal cluster sizes in designing cluster randomized trials to detect treatment effect heterogeneity. Statistics in Medicine, 41(8):1376–1396, April 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Azimian:2022:RSS

[836] Masoud Azimian, Mohammad Moradi, Mohammad Jafari Jozani, and William D. Leslie. Ranked set sampling in finite populations with bivariate responses: an application to an osteoporosis study. Statistics in Medicine, 41(8):1397–1420, April 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2022:PRD

[837] Yunshu Zhang, Shu Yang, Wenyu Ye, Douglas E. Faries, Ilya Lipkovich, and Zbigniew Kadziola. Practical recommendations on double score matching for estimating causal effects. Statistics in Medicine, 41(8): 1421–1445, April 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gao:2022:SCC

[838] Fei Gao and Marlena Bannick. Statistical considerations for crosssectional HIV incidence estimation based on recency test. Statistics in Medicine, 41(8):1446–1461, April 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hazewinkel:2022:SMR

[839] Audinga-Dea Hazewinkel, Jack Bowden, Kaitlin H. Wade, Tom Palmer, Nicola J. Wiles, and Kate Tilling. Sensitivity to missing not at random dropout in clinical trials: Use and interpretation of the trimmed means estimator. Statistics in Medicine, 41(8):1462–1481, April 15, 2022. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2022:OSD

[840] Tong Chen and Thomas Lumley. Optimal sampling for design-based estimators of regression models. Statistics in Medicine, 41(8):1482–1497, April 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sundin:2022:PAS

[841] Phillip Sundin and Catherine M. Crespi. Power analysis for stepped wedge trials with multiple interventions. Statistics in Medicine, 41(8): 1498–1512, April 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yang:2022:CIM

[842] Guandong Yang, Laura B. Balzer, and David Benkeser. Causal inference methods for vaccine sieve analysis with effect modification. Statistics in Medicine, 41(8):1513–1524, April 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fang:2022:SMC

[843] Yixin Fang and Man Jin. Sequential modeling for a class of referencebased imputation methods in clinical trials with quantitative or binary outcomes. Statistics in Medicine, 41(8):1525–1540, April 15, 2022. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Remiro-Azocar:2022:EMA

[844] Antonio Remiro-Azócar, Anna Heath, and Gianluca Baio. Effect modification in anchored indirect treatment comparison: Comments on “Matching-adjusted indirect comparisons: Application to time-to-event data”. Statistics in Medicine, 41(8):1541–1553, April 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [361].

Agarwala:2022:EIA

[845] Neha Agarwala, Junyong Park, and Anindya Roy. Efficient integration of aggregate data and individual participant data in one-way mixed models. Statistics in Medicine, 41(9):1555–1572, April 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cottin:2022:IDI

[846] Aziliz Cottin, Nicolas Pecuchet, Marine Zulian, Agathe Guilloux, and Sandrine Katsahian. IDNetwork: a deep illness-death network based on multi-state event history process for disease prognostication. Statistics in Medicine, 41(9):1573–1598, April 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Eaton:2022:ACE

[847] Anne A. Eaton and Emily C. Zabor. Analysis of composite endpoints with component-wise censoring in the presence of differential visit schedules. Statistics in Medicine, 41(9):1599–1612, April 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Serra:2022:ORM

[848] Alessandra Serra, Pavel Mozgunov, and Thomas Jaki. An order restricted multi-arm multi-stage clinical trial design. Statistics in Medicine, 41(9): 1613–1626, April 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Danieli:2022:PDP

[849] Coraline Danieli and Erica E. M. Moodie. Preserving data privacy when using multi-site data to estimate individualized treatment rules. Statistics in Medicine, 41(9):1627–1643, April 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kawaguchi:2022:TSH

[850] Eric S. Kawaguchi, Gang Li, Juan Pablo Lewinger, and W. James Gauderman. Two-step hypothesis testing to detect gene-environment interactions in a genome-wide scan with a survival endpoint. Statistics in Medicine, 41(9):1644–1657, April 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2022:RNI

[851] Zihan Li, Ziye Luo, and Yifan Sun. Robust nonparametric integrative analysis to decipher heterogeneity and commonality across subgroups using sparse boosting. Statistics in Medicine, 41(9):1658–1687, April 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Artman:2022:BSB

[852] William J. Artman, Brent A. Johnson, Kevin G. Lynch, James R. McKay, and Ashkan Ertefaie. Bayesian set of best dynamic treatment regimes: Construction and sample size calculation for SMARTs with binary outcomes. Statistics in Medicine, 41(9):1688–1708, April 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Arifin:2022:CPV

[853] Wan Nor Arifin and Umi Kalsom Yusof. Correcting for partial verification bias in diagnostic accuracy studies: a tutorial using R. Statistics in Medicine, 41(9):1709–1727, April 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kim:2022:COE

[854] Byungwon Kim, Seonghong Kim, Sungkyu Jung, Woncheol Jang, and Johan Lim. Comments on “Online estimation of the case fatality rate using a run-off triangle data approach: an application to the Korean MERS outbreak in 2015” by Sungim Lee and Johan Lim published in Statistics in Medicine (vol. 38, 2644–2679, 2019). Statistics in Medicine, 41(9):1728–1732, April 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [6].

Lake:2022:C

[855] Stephen Lake. Correction. Statistics in Medicine, 41(9):1733–1734, April 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ghebremichael-Weldeselassie:2022:MSC

[856] Yonas Ghebremichael-Weldeselassie, Marie Joëlle Jabagi, Jérémie Botton, Marion Bertrand, Bérangère Baricault, Jérôme Drouin, Alain Weill, Mahmoud Zureik, Rosemary Dray-Spira, and Paddy Farrington. A modified self-controlled case series method for event-dependent exposures and high event-related mortality, with application to COVID-19 vaccine safety. Statistics in Medicine, 41(10):1735–1750, May 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Voldal:2022:MMS

[857] Emily C. Voldal, Fan Xia, Avi Kenny, Patrick J. Heagerty, and James P. Hughes. Model misspecification in stepped wedge trials: Random effects for time or treatment. Statistics in Medicine, 41(10):1751–1766, May 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stefano:2022:CEM

[858] Fulvio Di Stefano, Matthieu Pannaux, Anne Correges, Stephanie Galtier, Veronique Robert, and Gaelle Saint-Hilary. A comparison of estimation methods adjusting for selection bias in adaptive enrichment designs with time-to-event endpoints. Statistics in Medicine, 41(10):1767–1779, May 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ma:2022:SPL

[859] Qianheng Ma, Robin J. Mermelstein, and Donald Hedeker. A sharedparameter location-scale mixed model to link the responsivity in selfinitiated event reports and the event-contingent Ecological Momentary Assessments. Statistics in Medicine, 41(10):1780–1796, May 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tai:2022:RIE

[860] An-Shun Tai, Yi-Juan Du, and Sheng-Hsuan Lin. Robust inference on effects attributable to mediators: a controlled-direct-effect-based approach for causal effect decomposition with multiple mediators. Statistics in Medicine, 41(10):1797–1814, May 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Peng:2022:EHE

[861] Roger D. Peng, Jia C. Liu, Meredith C. McCormack, Loretta J. Mickley, and Michelle L. Bell. Estimating the health effects of environmental mixtures using principal stratification. Statistics in Medicine, 41(10): 1815–1828, May 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lu:2022:EPE

[862] Minggen Lu, Yan Liu, Chin-Shang Li, and Jianguo Sun. An efficient penalized estimation approach for semiparametric linear transformation models with interval-censored data. Statistics in Medicine, 41(10):1829– 1845, May 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Johns:2022:CSM

[863] Hannah Johns, Dominic Italiano, Bruce Campbell, and Leonid Churilov. Common scale minimal sufficient balance: an improved method for covariate-adaptive randomization based on the Wilcoxon–Mann– Whitney odds ratio statistic. Statistics in Medicine, 41(10):1846–1861, May 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhou:2022:CRS

[864] Yunji Zhou, Elizabeth L. Turner, Ryan A. Simmons, and Fan Li. Constrained randomization and statistical inference for multi-arm parallel cluster randomized controlled trials. Statistics in Medicine, 41(10):1862– 1883, May 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Singh:2022:ARH

[865] Roopali Singh, Feipeng Zhang, and Qunhua Li. Assessing reproducibility of high-throughput experiments in the case of missing data. Statistics in Medicine, 41(10):1884–1899, May 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Putter:2022:LBG

[866] Hein Putter and Hans C. van Houwelingen. Landmarking 2.0: Bridging the gap between joint models and landmarking. Statistics in Medicine, 41(11):1901–1917, May 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhou:2022:TBB

[867] Yanhong Zhou, Ruitao Lin, J. Jack Lee, Daniel Li, Li Wang, Ruobing Li, and Ying Yuan. TITE-BOIN12: a Bayesian phase I/II trial design to find the optimal biological dose with late-onset toxicity and efficacy. Statistics in Medicine, 41(11):1918–1931, May 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Won:2022:SCS

[868] Jung Y. Won, Emma V. Sanchez-Vaznaugh, Yuqi Zhai, and Brisa N. Sánchez. Split and combine simulation extrapolation algorithm to correct geocoding coarsening of built environment exposures. Statistics in Medicine, 41(11):1932–1949, May 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mehta:2022:OAP

[869] Cyrus Mehta, Apurva Bhingare, Lingyun Liu, and Pralay Senchaudhuri. Optimal adaptive promising zone designs. Statistics in Medicine, 41(11): 1950–1970, May 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Huang:2022:HTC

[870] Yen-Tsung Huang. Hypothesis test for causal mediation of time-to-event mediator and outcome. Statistics in Medicine, 41(11):1971–1985, May 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vanacore:2022:RTC

[871] Amalia Vanacore and Maria Sole Pellegrino. Robustness of κ-type coefficients for clinical agreement. Statistics in Medicine, 41(11):1986–2004, May 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Aleman-Gomez:2022:DVO

[872] Yasser Alemán-Gómez, Ana Arribas-Gil, Manuel Desco, Antonio Elías, and Juan Romo. Depthgram: Visualizing outliers in high-dimensional functional data with application to fMRI data exploration. Statistics in Medicine, 41(11):2005–2024, May 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jiang:2022:NRR

[873] Hui Jiang, Lei Huang, and Yingcun Xia. Nonparametric regression with right-censored covariate via conditional density function. Statistics in Medicine, 41(11):2025–2051, May 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jiang:2022:IAR

[874] Jiming Jiang, Yuanyuan Li, Thuan Nguyen, and Mandi Yu. Inference about ratios of age-standardized rates with sampling errors in the population denominators for estimating both rates. Statistics in Medicine, 41(11):2052–2068, May 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kao:2022:CIC

[875] Chu-Lan Kao, Lin-An Chen, and Chi-An Hu. Critical issues concerning the assessment sets in agreement assessment. Statistics in Medicine, 41 (11):2069–2089, May 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Davies:2022:NMA

[876] Annabel L. Davies, Theodoros Papakonstantinou, Adriani Nikolakopoulou, Gerta Rücker, and Tobias Galla. Network meta-analysis and random walks. Statistics in Medicine, 41(12):2091–2114, May 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tomer:2022:SDM

[877] Anirudh Tomer, Daan Nieboer, Monique J. Roobol, Ewout W. Steyerberg, and Dimitris Rizopoulos. Shared decision making of burdensome surveillance tests using personalized schedules and their burden and benefit. Statistics in Medicine, 41(12):2115–2131, May 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2022:ERT

[878] Haodong Li, Sonali Rosete, Jeremy Coyle, Rachael V. Phillips, Nima S. Hejazi, Ivana Malenica, Benjamin F. Arnold, Jade Benjamin-Chung, Andrew Mertens, John M. Colford Jr, Mark J. van der Laan, and Alan E. Hubbard. Evaluating the robustness of targeted maximum likelihood estimators via realistic simulations in nutrition intervention trials. Statistics in Medicine, 41(12):2132–2165, May 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2022:ATB

[879] Xiaoyan Liu, Chanmin Kim, Zifei Han, Pilar Lim, Satrajit Roychoudhury, Maurizio Fava, and Gheorghe Doros. Assessing treatment benefit in the presence of placebo response using the sequential parallel comparison design. Statistics in Medicine, 41(12):2166–2190, May 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhou:2022:SSC

[880] Zhengyang Zhou, Dateng Li, and Song Zhang. Sample size calculation for cluster randomized trials with zero-inflated count outcomes. Statistics in Medicine, 41(12):2191–2204, May 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hunter:2022:IMA

[881] Kristen B. Hunter, Mark E. Glickman, and Luis F. Campos. Inferring medication adherence from time-varying health measures. Statistics in Medicine, 41(12):2205–2226, May 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhou:2022:EPT

[882] Ruixuan Rachel Zhou, Sihai Dave Zhao, and Layla Parast. Estimation of the proportion of treatment effect explained by a high-dimensional surrogate. Statistics in Medicine, 41(12):2227–2246, May 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Maturo:2022:PRF

[883] Fabrizio Maturo and Rosanna Verde. Pooling random forest and functional data analysis for biomedical signals supervised classification: Theory and application to electrocardiogram data. Statistics in Medicine, 41(12):2247–2275, May 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2022:BHM

[884] Shengwei Zhang, Haitao Chu, Warren K. Bickel, Chap T. Le, Tracy T. Smith, Janet L. Thomas, Eric C. Donny, Dorothy K. Hatsukami, and Xianghua Luo. A Bayesian hierarchical model for individual participant data meta-analysis of demand curves. Statistics in Medicine, 41(12): 2276–2290, May 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2022:BTS

[885] Lichang Chen, Jianhong Pan, Yanpeng Wu, Jingxian Wang, Fangyao Chen, Jun Zhao, and Pingyan Chen. Bayesian two-stage design for phase II oncology trials with binary endpoint. Statistics in Medicine, 41(12): 2291–2301, May 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McMenamin:2022:SSE

[886] Martina E. McMenamin, Jessica K. Barrett, Anna Berglind, and James M. S. Wason. Sample size estimation using a latent variable model for mixed outcome co-primary, multiple primary and composite endpoints. Statistics in Medicine, 41(13):2303–2316, June 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bhaduri:2022:ESE

[887] Ritwik Bhaduri, Ritoban Kundu, Soumik Purkayastha, Michael Kleinsasser, Lauren J. Beesley, Bhramar Mukherjee, and Jyotishka Datta. Extending the susceptible-exposed-infected-removed (SEIR) model to handle the false negative rate and symptom-based administration of COVID-19 diagnostic tests: SEIR-fansy. Statistics in Medicine, 41(13):2317– 2337, June 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tayob:2022:MPE

[888] Nabihah Tayob, Anna S. F. Lok, and Ziding Feng. A multivariate parametric empirical Bayes screening approach for early detection of hepatocellular carcinoma using multiple longitudinal biomarkers. Statistics in Medicine, 41(13):2338–2353, June 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ren:2022:SPB

[889] Junting Ren, Susan Tapert, Chun Chieh Fan, and Wesley K. Thompson. A semi-parametric Bayesian model for semi-continuous longitudinal data. Statistics in Medicine, 41(13):2354–2374, June 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2022:GSD

[890] Yiming Chen, John Lawrence, and Mei-Ling Ting Lee. Group sequential design for randomized trials using “first hitting time” model. Statistics in Medicine, 41(13):2375–2402, June 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pagui:2022:IEN

[891] Euloge Clovis Kenne Pagui, Alessandra Salvan, and Nicola Sartori. Improved estimation in negative binomial regression. Statistics in Medicine, 41(13):2403–2416, June 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Han:2022:TGN

[892] Sunwoo Han, Youyi Fong, and Ying Huang. Testing a global null hypothesis using ensemble machine learning methods. Statistics in Medicine, 41(13):2417–2426, June 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2022:TWF

[893] Pei Wang and Suvra Pal. A two-way flexible generalized gamma transformation cure rate model. Statistics in Medicine, 41(13):2427–2447, June 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Guo:2022:ACA

[894] Lulu Guo, Yi Qian, and Hui Xie. Assessing complier average causal effects from longitudinal trials with multiple endpoints and treatment noncompliance: an application to a study of Arthritis Health Journal. Statistics in Medicine, 41(13):2448–2465, June 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sheen:2022:RSC

[895] Justin K. Sheen, Johannes Haushofer, C. Jessica E. Metcalf, and Lee Kennedy-Shaffer. The required size of cluster randomized trials of nonpharmaceutical interventions in epidemic settings. Statistics in Medicine, 41(13):2466–2482, June 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Peterson:2022:EME

[896] Emily Peterson, Doris Chou, Ann-Beth Moller, Alison Gemmill, Lale Say, and Leontine Alkema. Estimating misclassification errors in the reporting of maternal mortality in national civil registration vital statistics systems: a Bayesian hierarchical bivariate random walk model to estimate sensitivity and specificity for multiple countries and years with missing data. Statistics in Medicine, 41(14):2483–2496, June 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schildcrout:2022:MAA

[897] Jonathan S. Schildcrout, Frank E. Harrell, Jr., Patrick J. Heagerty, Sebastien Haneuse, Chiara Di Gravio, Shawn P. Garbett, Paul J. Rathouz, and Bryan E. Shepherd. Model-assisted analyses of longitudinal, ordinal outcomes with absorbing states. Statistics in Medicine, 41(14): 2497–2512, June 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hu:2022:MDA

[898] Zonghui Hu, Pengfei Li, Dean Follmann, and Jing Qin. A mixture distribution approach for assessing genetic impact from twin study. Statistics in Medicine, 41(14):2513–2522, June 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Trane:2022:NBT

[899] Ralph Møller Trane and Hyunseung Kang. Nonparametric bounds in two-sample summary-data Mendelian randomization: Some cautionary tales for practice. Statistics in Medicine, 41(14):2523–2541, June 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ghosh:2022:SPS

[900] Samiran Ghosh, Siuli Mukhopadhyay, Priyanka Majumder, and Bo Wang. Statistical power and sample size requirements to detect an intervention by time interaction in four-level longitudinal cluster randomized trials. Statistics in Medicine, 41(14):2542–2556, June 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wu:2022:ATB

[901] Xiaowei Wu and Hongxiao Zhu. Association testing for binary trees — a Markov branching process approach. Statistics in Medicine, 41(14):

2557–2573, June 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2022:MCP

[902] Danping Liu, Yongli Han, and Aiyi Liu. Marginal, conditional, and pseudo likelihood ratio approaches for biomarker combination to predict a binary disease outcome. Statistics in Medicine, 41(14):2574–2585, June 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Efthimiou:2022:BMA

[903] Orestis Efthimiou, Michael Seo, Eirini Karyotaki, Pim Cuijpers, Toshi A. Furukawa, Guido Schwarzer, Gerta Rücker, and Dimitris Mavridis. Bayesian models for aggregate and individual patient data component network meta-analysis. Statistics in Medicine, 41(14):2586–2601, June 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gressani:2022:LPS

[904] Oswaldo Gressani, Christel Faes, and Niel Hens. Laplacian-P-splines for Bayesian inference in the mixture cure model. Statistics in Medicine, 41(14):2602–2626, June 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hattori:2022:SSC

[905] Satoshi Hattori, Sho Komukai, and Tim Friede. Sample size calculation for the augmented logrank test in randomized clinical trials. Statistics in Medicine, 41(14):2627–2644, June 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2022:FSA

[906] Xinyuan Chen and Fan Li. Finite-sample adjustments in variance estimators for clustered competing risks regression. Statistics in Medicine, 41(14):2645–2664, June 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Papageorgiou:2022:BSP

[907] Georgios Papageorgiou. Bayesian semi-parametric modeling of covariance matrices for multivariate longitudinal data. Statistics in Medicine, 41(14):2665–2687, June 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gasparyan:2022:CSS

[908] Samvel B. Gasparyan, Elaine K. Kowalewski, and Gary G. Koch. Comments on “Sample size formula for a win ratio endpoint” by R. X. Yu and J. Ganju. Statistics in Medicine, 41(14):2688–2690, June 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [808, 909].

Yu:2022:SSFb

[909] Ron Xiaolong Yu and Jitendra Ganju. Sample size formula for a win ratio endpoint. Statistics in Medicine, 41(14):2691–2692, June 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [808, 908].

Anonymous:2022:C

[910] Anonymous. Correction. Statistics in Medicine, 41(14):2693, June 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kulasekera:2022:QBP

[911] Karunarathna B. Kulasekera and Chathura Siriwardhana. Quantiles based personalized treatment selection for multivariate outcomes and multiple treatments. Statistics in Medicine, 41(15):2695–2710, July 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Das:2022:SMI

[912] Ujjwal Das and Ranojoy Basu. Semiparametric multiple inflation count model with application to a smoking cessation study. Statistics in Medicine, 41(15):2711–2724, July 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Quan:2022:GWC

[913] Hui Quan, Xiaofei Chen, Junxiang Luo, and Xun Chen. A generalized weighted combination test of treatment effect for clinical trials with a sequential parallel comparison design and binary endpoint. Statistics in Medicine, 41(15):2725–2744, July 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Keller:2022:TTD

[914] Joshua P. Keller, Tianjian Zhou, Andee Kaplan, G. Brooke Anderson, and Wen Zhou. Tracking the transmission dynamics of COVID-19 with a time-varying coefficient state-space model. Statistics in Medicine, 41(15):

2745–2767, July 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ekvall:2022:MTM

[915] Karl Oskar Ekvall and Aaron J. Molstad. Mixed-type multivariate response regression with covariance estimation. Statistics in Medicine, 41 (15):2768–2785, July 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mishra:2022:NBF

[916] Aditya K. Mishra and Christian L. Müller. Negative binomial factor regression with application to microbiome data analysis. Statistics in Medicine, 41(15):2786–2803, July 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xue:2022:JMS

[917] Xiaonan Xue, Simin Hua, Kathleen Weber, Qibin Qi, Robert Kaplan, Deborah R. Gustafson, Anjali Sharma, Audrey French, and Helen J. Burgess. Jointly modeling of sleep variables that are objectively measured by wrist actigraphy. Statistics in Medicine, 41(15):2804–2821, July 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wei:2022:MRS

[918] Kecheng Wei, Huichen Zhu, Guoyou Qin, Zhongyi Zhu, and Dongsheng Tu. Multiply robust subgroup analysis based on a single-index threshold linear marginal model for longitudinal data with dropouts. Statistics in Medicine, 41(15):2822–2839, July 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wu:2022:ILS

[919] Wenbo Wu, Yuan Yang, Jian Kang, and Kevin He. Improving large-scale estimation and inference for profiling health care providers. Statistics in Medicine, 41(15):2840–2853, July 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hebert:2022:OTA

[920] Florian Hébert, David Causeur, and Mathieu Emily. Omnibus testing approach for gene-based gene-gene interaction. Statistics in Medicine, 41(15):2854–2878, July 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Satten:2022:EEI

[921] Glen A. Satten, Sarah W. Curtis, Claudia Solis-Lemus, Elizabeth J. Leslie, and Michael P. Epstein. Efficient estimation of indirect effects in case-control studies using a unified likelihood framework. Statistics in Medicine, 41(15):2879–2893, July 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lin:2022:DLD

[922] Jeffrey Lin and Sheng Luo. Deep learning for the dynamic prediction of multivariate longitudinal and survival data. Statistics in Medicine, 41(15):2894–2907, July 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wilson:2022:BSS

[923] Kevin J. Wilson, S. Faye Williamson, A. Joy Allen, Cameron J. Williams, Thomas P. Hellyer, and B. Clare Lendrem. Bayesian sample size determination for diagnostic accuracy studies. Statistics in Medicine, 41(15): 2908–2922, July 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2022:IUC

[924] Kenneth Menglin Lee, Xiangmei Ma, Grace Meijuan Yang, and Yin Bun Cheung. Inclusion of unexposed clusters improves the precision of fixed effects analysis of stepped-wedge cluster randomized trials. Statistics in Medicine, 41(15):2923–2938, July 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2022:RIS

[925] Wei Wang, Xiong Xiao, Jian Qian, Shiqi Chen, Fang Liao, Fei Yin, Tao Zhang, Xiaosong Li, and Yue Ma. Reclaiming independence in spatial-clustering datasets: a series of data-driven spatial weights matrices. Statistics in Medicine, 41(15):2939–2956, July 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2022:UBA

[926] Pin Li, Jeremy M. G. Taylor, Philip S. Boonstra, Theodore S. Lawrence, and Matthew J. Schipper. Utility based approach in individualized optimal dose selection using machine learning methods. Statistics in Medicine, 41(16):2957–2977, July 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mukherjee:2022:DFP

[927] Amitava Mukherjee, Wolfgang Kössler, and Marco Marozzi. A distribution-free procedure for testing versatile alternative in medical multisample comparison studies. Statistics in Medicine, 41(16):2978– 3002, July 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rong:2022:RMR

[928] Rong Rong, Jing Ning, and Hong Zhu. Regression modeling of restricted mean survival time for left-truncated right-censored data. Statistics in Medicine, 41(16):3003–3021, July 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2022:NRM

[929] Xueqing Liu, Jiarui Sun, and Xiao-Hua Zhou. A novel regression method for the analysis of multireader multicase-free-response receiver operating characteristics studies. Statistics in Medicine, 41(16):3022–3038, July 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jo:2022:HPA

[930] Booil Jo. Handling parametric assumptions in principal causal effect estimation using Gaussian mixtures. Statistics in Medicine, 41(16):3039– 3056, July 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gao:2022:HMD

[931] Leiwen Gao, Abhirup Datta, and Sudipto Banerjee. Hierarchical multivariate directed acyclic graph autoregressive models for spatial diseases mapping. Statistics in Medicine, 41(16):3057–3075, July 20, 2022. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Follmann:2022:EVE

[932] Dean Follmann, Michael Fay, and Craig Magaret. Estimation of vaccine efficacy for variants that emerge after the placebo group is vaccinated. Statistics in Medicine, 41(16):3076–3089, July 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

vanZwet:2022:HLS

[933] Erik W. van Zwet and Steven N. Goodman. How large should the next study be? Predictive power and sample size requirements for replication studies. Statistics in Medicine, 41(16):3090–3101, July 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kim:2022:TYS

[934] Hyune-Ju Kim, Huann-Sheng Chen, Jeffrey Byrne, Bill Wheeler, and Eric J. Feuer. Twenty years since Joinpoint 1.0: Two major enhancements, their justification, and impact. Statistics in Medicine, 41(16): 3102–3130, July 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cahoy:2022:BIA

[935] Dexter Cahoy and Joseph Sedransk. Bayesian inference for asymptomatic COVID-19 infection rates. Statistics in Medicine, 41(16):3131–3148, July 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wu:2022:WMT

[936] Yougui Wu. Weighted McNemar’s test for the comparison of two screening tests in the presence of verification bias. Statistics in Medicine, 41 (16):3149–3163, July 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Curtis:2022:SSD

[937] Alexandra Curtis, Brian Smith, and Andrew G. Chapple. Subgroupspecific dose finding for phase I–II trials using Bayesian clustering. Statistics in Medicine, 41(16):3164–3179, July 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2022:DFM

[938] Chun-Shu Chen and Chung-Wei Shen. Distribution-free model selection for longitudinal zero-inflated count data with missing responses and covariates. Statistics in Medicine, 41(16):3180–3198, July 20, 2022. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dobbin:2022:SSM

[939] Kevin K. Dobbin and Lisa M. McShane. Sample size methods for evaluation of predictive biomarkers. Statistics in Medicine, 41(16):3199–3210, July 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stensrud:2022:TQE

[940] Mats J. Stensrud and Oliver Dukes. Translating questions to estimands in randomized clinical trials with intercurrent events. Statistics in Medicine, 41(16):3211–3228, July 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sun:2022:RSH

[941] Yifan Sun, Ziye Luo, and Xinyan Fan. Robust structured heterogeneity analysis approach for high-dimensional data. Statistics in Medicine, 41 (17):3229–3259, July 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Webb:2022:PLE

[942] Annabel Webb, Jun Ma, and Serigne N. Lô. Penalized likelihood estimation of a mixture cure Cox model with partly interval censoring — an application to thin melanoma. Statistics in Medicine, 41(17):3260–3280, July 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jazi:2022:VSS

[943] Omidali Aghababaei Jazi and Eleanor Pullenayegum. Variable selection in semiparametric regression models for longitudinal data with informative observation times. Statistics in Medicine, 41(17):3281–3298, July 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vancak:2022:NNT

[944] Valentin Vancak, Yair Goldberg, and Stephen Z. Levine. The number needed to treat adjusted for explanatory variables in regression and survival analysis: Theory and application. Statistics in Medicine, 41(17): 3299–3320, July 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhou:2022:CPF

[945] Thomas J. Zhou, Michael P. LaValley, Kerrie P. Nelson, Howard J. Cabral, and Joseph M. Massaro. Calculating power for the Finkelstein and Schoenfeld test statistic for a composite endpoint with two components. Statistics in Medicine, 41(17):3321–3335, July 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fairley:2022:SEI

[946] Michael Fairley, Isabelle J. Rao, Margaret L. Brandeau, Gary L. Qian, and Gregg S. Gonsalves. Surveillance for endemic infectious disease outbreaks: Adaptive sampling using profile likelihood estimation. Statistics in Medicine, 41(17):3336–3348, July 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2022:FEI

[947] Ruonan Li, Luo Xiao, Ekaterina Smirnova, Erjia Cui, Andrew Leroux, and Ciprian M. Crainiceanu. Fixed-effects inference and tests of correlation for longitudinal functional data. Statistics in Medicine, 41(17): 3349–3364, July 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cruz:2022:RBB

[948] Ivette Raices Cruz, Matthias C. M. Troffaes, Johan Lindström, and Ullrika Sahlin. A robust Bayesian bias-adjusted random effects model for consideration of uncertainty about bias terms in evidence synthesis. Statistics in Medicine, 41(17):3365–3379, July 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Luna:2022:OAC

[949] Jose Luna, Jessica Jaynes, Hongquan Xu, and Weng Kee Wong. Orthogonal array composite designs for drug combination experiments with applications for tuberculosis. Statistics in Medicine, 41(17):3380–3397, July 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Luo:2022:PWP

[950] Bin Luo, Xiaoli Gao, and Susan Halabi. Penalized weighted proportional hazards model for robust variable selection and outlier detection. Statistics in Medicine, 41(17):3398–3420, July 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

vanOudenhoven:2022:MUB

[951] Floor M. van Oudenhoven, Sophie H. N. Swinkels, Tobias Hartmann, and Dimitris Rizopoulos. Modeling the underlying biological processes in Alzheimer’s disease using a multivariate competing risk joint model. Statistics in Medicine, 41(17):3421–3433, July 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xu:2022:SML

[952] Tianchen Xu, Yuan Chen, Donglin Zeng, and Yuanjia Wang. Selfmatched learning to construct treatment decision rules from electronic health records. Statistics in Medicine, 41(17):3434–3447, July 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Buchanan:2022:TED

[953] Ashley Buchanan, Tianyu Sun, Jing Wu, Hilary Aroke, Jeffrey Bratberg, Josiah Rich, Stephen Kogut, and Joseph Hogan. Toward evaluation of disseminated effects of medications for opioid use disorder within provider-based clusters using routinely-collected health data. Statistics in Medicine, 41(18):3449–3465, August 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Luo:2022:APB

[954] Chongliang Luo, Arielle Marks-Anglin, Rui Duan, Lifeng Lin, Chuan Hong, Haitao Chu, and Yong Chen. Accounting for publication bias using a bivariate trim and fill meta-analysis procedure. Statistics in Medicine, 41(18):3466–3478, August 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jantzen:2022:ACS

[955] Rodolphe Jantzen, Pascale Tubert-Bitter, and Philippe Broët. Analyzing cohort studies with interval-censored data: a new model-based linear rank-type test. Statistics in Medicine, 41(18):3479–3491, August 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mallick:2022:DES

[956] Himel Mallick, Suvo Chatterjee, Shrabanti Chowdhury, Saptarshi Chatterjee, Ali Rahnavard, and Stephanie C. Hicks. Differential expression of single-cell RNA-seq data using Tweedie models. Statistics in Medicine, 41(18):3492–3510, August 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Krstic:2022:IMC

[957] Nikolas Krstic, Kevin Multani, David S. Wishart, Tom Blydt-Hansen, and Gabriela V. Cohen Freue. The impact of methodological choices when developing predictive models using urinary metabolite data. Statistics in Medicine, 41(18):3511–3526, August 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bantis:2022:OBC

[958] Leonidas E. Bantis and John V. Tsimikas. On optimal biomarker cutoffs accounting for misclassification costs in diagnostic trilemmas with applications to pancreatic cancer. Statistics in Medicine, 41(18):3527–3546, August 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shi:2022:DPT

[959] Haolun Shi, Shu Jiang, and Jiguo Cao. Dynamic prediction with timedependent marker in survival analysis using supervised functional principal component analysis. Statistics in Medicine, 41(18):3547–3560, August 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yefenof:2022:SRS

[960] Jonathan Yefenof, Yair Goldberg, Jennifer Wiler, Avishai Mandelbaum, and Ya’acov Ritov. Self-reporting and screening: Data with right-censored, left-censored, and complete observations. Statistics in Medicine, 41(18):3561–3578, August 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Marion:2022:BMD

[961] Joe Marion, Juan Ruiz, and Benjamin R. Saville. Bayesian model of disease progression in mucopolysaccaridosis IIIA. Statistics in Medicine, 41(18):3579–3595, August 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2022:RTM

[962] Li Wang, Yang Liu, Xiaotian Chen, and Erik Pulkstenis. Real time monitoring and prediction of time to endpoint maturation in clinical trials. Statistics in Medicine, 41(18):3596–3611, August 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chang:2022:PSM

[963] Ting-Hsuan Chang and Elizabeth A. Stuart. Propensity score methods for observational studies with clustered data: a review. Statistics in Medicine, 41(18):3612–3626, August 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kasza:2022:BSW

[964] Jessica Kasza, Rhys Bowden, Richard Hooper, and Andrew B. Forbes. The batched stepped wedge design: a design robust to delays in cluster recruitment. Statistics in Medicine, 41(18):3627–3641, August 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2022:MPL

[965] Honglang Wang, Jingyi Zhang, Kelly L. Klump, Sybil Alexandra Burt, and Yuehua Cui. Multivariate partial linear varying coefficients model for gene-environment interactions with multiple longitudinal traits. Statistics in Medicine, 41(19):3643–3660, August 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jiang:2022:PDI

[966] Shu Jiang and Richard J. Cook. The polytomous discrimination index for prediction involving multistate processes under intermittent observation. Statistics in Medicine, 41(19):3661–3678, August 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2022:CCI

[967] Yiming Li, Wei-Wen Hsu, and for the Alzheimer’s Disease Neuroimaging Initiative. A classification for complex imbalanced data in disease screening and early diagnosis. Statistics in Medicine, 41(19):3679–3695, August 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mattos:2022:EMS

[968] Thalita B. Mattos, Victor H. Lachos, Luis M. Castro, and Larissa A. Matos. Extending multivariate Student’s-t semiparametric mixed models for longitudinal data with censored responses and heavy tails. Statistics in Medicine, 41(19):3696–3719, August 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Paukner:2022:VTW

[969] Mitchell Paukner and Richard Chappell. Versatile tests for window mean survival time. Statistics in Medicine, 41(19):3720–3736, August 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Campos:2022:MHP

[970] Emilie Campos, Aaron Wolfe Scheffler, Donatello Telesca, Catherine Sugar, Charlotte DiStefano, Shafali Jeste, April R. Levin, Adam Naples, Sara J. Webb, Frederick Shic, Geraldine Dawson, Susan Faja, James C. McPartland, Damla Sentürk, and Autism Biomarkers Consortium for Clinical Trials. Multilevel hybrid principal components analysis for region-referenced functional electroencephalography data. Statistics in Medicine, 41(19):3737–3757, August 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rosenbaum:2022:SDI

[971] Paul R. Rosenbaum. A statistic with demonstrated insensitivity to unmeasured bias for 2 × 2 × S tables in observational studies. Statistics in Medicine, 41(19):3758–3771, August 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bergquist:2022:ULC

[972] Savannah Bergquist, Gabriel A. Brooks, Mary Beth Landrum, Nancy L. Keating, and Sherri Rose. Uncertainty in lung cancer stage for survival estimation via set-valued classification. Statistics in Medicine, 41 (19):3772–3788, August 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Barone:2022:BID

[973] Rosario Barone and Andrea Tancredi. Bayesian inference for discretely observed continuous time multi-state models. Statistics in Medicine, 41 (19):3789–3803, August 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Binder:2022:SCR

[974] Nadine Binder, Kathrin Möllenhoff, August Sigle, and Holger Dette. Similarity of competing risks models with constant intensities in an application to clinical healthcare pathways involving prostate cancer surgery. Statistics in Medicine, 41(19):3804–3819, August 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xie:2022:EEP

[975] Shanghong Xie, Wenbo Wang, Qinxia Wang, Yuanjia Wang, and Donglin Zeng. Evaluating effectiveness of public health intervention strategies for mitigating COVID-19 pandemic. Statistics in Medicine, 41(19):3820– 3836, August 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lipkovich:2022:UPS

[976] Ilya Lipkovich, Bohdana Ratitch, Yongming Qu, Xiang Zhang, Mingyang Shan, and Craig Mallinckrodt. Using principal stratification in analysis of clinical trials. Statistics in Medicine, 41(19):3837–3877, August 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See correction [1498].

Inacio:2022:BNI

[977] Vanda Inácio and Javier E. Garrido Guillén. Bayesian nonparametric inference for the overlap coefficient: With an application to disease diagnosis. Statistics in Medicine, 41(20):3879–3898, September 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Guo:2022:SSL

[978] Boyi Guo, Byron C. Jaeger, A. K. M. Fazlur Rahman, D. Leann Long, and Nengjun Yi. Spike-and-slab least absolute shrinkage and selection operator generalized additive models and scalable algorithms for highdimensional data analysis. Statistics in Medicine, 41(20):3899–3914, September 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rover:2022:SMA

[979] Christian Röver, Moreno Ursino, Tim Friede, and Sarah Zohar. A straightforward meta-analysis approach for oncology phase I dose-finding studies. Statistics in Medicine, 41(20):3915–3940, September 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2022:CRP

[980] Minjung Lee and Jason P. Fine. Competing risks predictions with different time scales under the additive risk model. Statistics in Medicine, 41(20):3941–3957, September 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2022:NBR

[981] Shuai Chen and Jeffrey S. Hoch. Net-benefit regression with censored cost-effectiveness data from randomized or observational studies. Statistics in Medicine, 41(20):3958–3974, September 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Braun:2022:ECR

[982] Thomas M. Braun and Francois Mercier. Extending the Continual Reassessment Method to accommodate step-up dosing in Phase I trials. Statistics in Medicine, 41(20):3975–3990, September 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhao:2022:BNM

[983] Yize Zhao, Tianqi Chen, Jiachen Cai, Sarah Lichenstein, Marc N. Potenza, and Sarah W. Yip. Bayesian network mediation analysis with application to the brain functional connectome. Statistics in Medicine, 41(20):3991–4005, September 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tubbs:2022:INP

[984] Justin D. Tubbs, Lane G. Chen, Thuan-Quoc Thach, and Pak C. Sham. Improved nonparametric penalized maximum likelihood estimation for arbitrarily censored survival data. Statistics in Medicine, 41(20):4006– 4021, September 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dehbi:2022:SSC

[985] Hakim-Moulay Dehbi, Andrew Embleton-Thirsk, and Zachary Ryan Mc-Caw. Sample size calculation for randomized selection trials with a timeto-event endpoint and a margin of practical equivalence. Statistics in Medicine, 41(20):4022–4033, September 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2022:DRL

[986] Mingyang Liu, Xiaotong Shen, and Wei Pan. Deep reinforcement learning for personalized treatment recommendation. Statistics in Medicine, 41 (20):4034–4056, September 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sun:2022:MNR

[987] Lizhe Sun and Faming Liang. Markov neighborhood regression for statistical inference of high-dimensional generalized linear models. Statistics in Medicine, 41(20):4057–4078, September 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tang:2022:SSC

[988] Yongqiang Tang and Ronan Fitzpatrick. Sample size calculation for the Andersen–Gill model comparing rates of recurrent events. Statistics in Medicine, 41(20):4079, September 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2022:RMS

[989] Chengfeng Zhang, Baoyi Huang, Hongji Wu, Hao Yuan, Yawen Hou, and Zheng Chen. Restricted mean survival time regression model with timedependent covariates. Statistics in Medicine, 41(21):4081–4090, September 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Milienos:2022:RFF

[990] Fotios S. Milienos. On a reparameterization of a flexible family of cure models. Statistics in Medicine, 41(21):4091–4111, September 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

He:2022:TLH

[991] Yong He, Qiushi Li, Qinqin Hu, and Lei Liu. Transfer learning in highdimensional semiparametric graphical models with application to brain connectivity analysis. Statistics in Medicine, 41(21):4112–4129, September 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Takahashi:2022:STT

[992] Kenichi Takahashi, Ryota Ishii, Kazushi Maruo, and Masahiko Gosho. Statistical tests for two-stage adaptive seamless design using short- and long-term binary outcomes. Statistics in Medicine, 41(21):4130–4142, September 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tai:2022:IRE

[993] An-Shun Tai and Sheng-Hsuan Lin. Identification and robust estimation of swapped direct and indirect effects: Mediation analysis with unmeasured mediator-outcome confounding and intermediate confounding. Statistics in Medicine, 41(21):4143–4158, September 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Paul:2022:DDD

[994] Rajib Paul, Dan Han, Elise DeDoncker, and Diana Prieto. Dynamic downscaling and daily nowcasting from influenza surveillance data. Statistics in Medicine, 41(21):4159–4175, September 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Parra:2022:TES

[995] Camila Olarte Parra, Ingeborg Waernbaum, Staffan Schön, and Els Goetghebeur. Trial emulation and survival analysis for disease incidence registers: a case study on the causal effect of pre-emptive kidney transplantation. Statistics in Medicine, 41(21):4176–4199, September 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hirakawa:2022:ELT

[996] Akihiro Hirakawa, Hiroyuki Sato, Ryoichi Hanazawa, Keisuke Suzuki, and Japanese Alzheimer’s Disease Neuroimaging Initiative. Estimating the longitudinal trajectory of cognitive function measurement using short-term data with different disease stages: Application in Alzheimer’s disease. Statistics in Medicine, 41(21):4200–4214, September 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2022:UME

[997] Guanqing Chen, Valerie A. Lewis, Daniel J. Gottlieb, and A. James O’Malley. Using a mixed-effect model with a parameter-space of heterogeneous dimension to evaluate whether accountable care organizations are associated with greater uniformity across constituent practices. Statistics in Medicine, 41(21):4215–4226, September 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cai:2022:COS

[998] Hengrui Cai, Wenbin Lu, Rachel Marceau West, Devan V. Mehrotra, and Lingkang Huang. CAPITAL: Optimal subgroup identification via constrained policy tree search. Statistics in Medicine, 41(21):4227–4244, September 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2022:MBB

[999] Junni L. Zhang and Per Johansson. Model-based Bayesian inference under computer assisted balance-improving designs. Statistics in Medicine, 41(21):4245–4265, September 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mundo:2022:GAM

[1000] Ariel I. Mundo, John R. Tipton, and Timothy J. Muldoon. Generalized additive models to analyze nonlinear trends in biomedical longitudinal data using R: Beyond repeated measures ANOVA and linear mixed models. Statistics in Medicine, 41(21):4266–4283, September 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ma:2022:AWE

[1001] Huijuan Ma, Weicai Pang, Liuquan Sun, and Wei Xu. Augmented weighting estimators for the additive rates model under multivariate recurrent event data with missing event type. Statistics in Medicine, 41 (22):4285–4298, September 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kahan:2022:EFT

[1002] Brennan C. Kahan, Tim P. Morris, Beatriz Goulão, and James Carpenter. Estimands for factorial trials. Statistics in Medicine, 41(22): 4299–4310, September 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kenny:2022:ASW

[1003] Avi Kenny, Emily C. Voldal, Fan Xia, Patrick J. Heagerty, and James P. Hughes. Analysis of stepped wedge cluster randomized trials in the presence of a time-varying treatment effect. Statistics in Medicine, 41(22): 4311–4339, September 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fu:2022:CVS

[1004] Han Fu, Deedra Nicolet, Krzysztof Mrózek, Richard M. Stone, Ann-Kathrin Eisfeld, John C. Byrd, and Kellie J. Archer. Controlled variable selection in Weibull mixture cure models for high-dimensional data. Statistics in Medicine, 41(22):4340–4366, September 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2022:BLE

[1005] Yilin Liu, Michael Kane, Denise Esserman, Ondrej Blaha, Daniel Zelterman, and Wei Wei. Bayesian local exchangeability design for phase II basket trials. Statistics in Medicine, 41(22):4367–4384, September 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wijesuriya:2022:MIA

[1006] Rushani Wijesuriya, Margarita Moreno-Betancur, John Carlin, Anurika Priyanjali De Silva, and Katherine Jane Lee. Multiple imputation approaches for handling incomplete three-level data with time-varying clustermemberships. Statistics in Medicine, 41(22):4385–4402, September 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yang:2022:ARD

[1007] Ce Yang, Liqun Diao, and Richard J. Cook. Adaptive responsedependent two-phase designs: Some results on robustness and efficiency. Statistics in Medicine, 41(22):4403–4425, September 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Austin:2022:BVA

[1008] Peter C. Austin. Bootstrap vs asymptotic variance estimation when using propensity score weighting with continuous and binary outcomes. Statistics in Medicine, 41(22):4426–4443, September 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wigle:2022:BUA

[1009] Augustine Wigle and Audrey Béliveau. Bayesian unanchored additive models for component network meta-analysis. Statistics in Medicine, 41 (22):4444–4466, September 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2022:DSM

[1010] Xingruo Zhang and Donald Hedeker. Defining R-squared measures for mixed-effects location scale models. Statistics in Medicine, 41(22):4467– 4483, September 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2022:MMA

[1011] Hwiyoung Lee, Chixiang Chen, Peter Kochunov, Liyi Elliot Hong, and Shuo Chen. Modeling multivariate age-related imaging variables with dependencies. Statistics in Medicine, 41(22):4484–4500, September 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Senn:2022:NRA

[1012] Stephen Senn, Susanne Schmitz, Anna Schritz, and Artur Araujo. A note regarding alternative explanations for heterogeneity in meta-analysis. Statistics in Medicine, 41(22):4501–4509, September 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Josey:2022:CAT

[1013] Kevin P. Josey, Fan Yang, Debashis Ghosh, and Sridharan Raghavan. A calibration approach to transportability and data-fusion with observational data. Statistics in Medicine, 41(23):4511–4531, October 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Eliseussen:2022:RBB

[1014] Emilie Eliseussen, Thomas Fleischer, and Valeria Vitelli. Rank-based Bayesian variable selection for genome-wide transcriptomic analyses. Statistics in Medicine, 41(23):4532–4553, October 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zivich:2022:TML

[1015] Paul N. Zivich, Michael G. Hudgens, Maurice A. Brookhart, James Moody, David J. Weber, and Allison E. Aiello. Targeted maximum likelihood estimation of causal effects with interference: a simulation study.

Statistics in Medicine, 41(23):4554–4577, October 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Park:2022:EBP

[1016] Yeonhee Park, Zhihua Su, and Dongjun Chung. Envelope-based partial partial least squares with application to cytokine-based biomarker analysis for COVID-19. Statistics in Medicine, 41(23):4578–4592, October 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Boyle:2022:ECS

[1017] Joseph Boyle, Mary H. Ward, Stella Koutros, Margaret R. Karagas, Molly Schwenn, Debra Silverman, and David C. Wheeler. Estimating cumulative spatial risk over time with low-rank kriging multiple membership models. Statistics in Medicine, 41(23):4593–4606, October 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2022:BCB

[1018] Yiming Zhang, Ming-Hui Chen, and Feng Guo. Bayesian criterion-based assessments of recurrent event models with applications to commercial truck driver behavior studies. Statistics in Medicine, 41(23):4607–4628, October 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gillam:2022:MFH

[1019] Jess Gillam, Rebecca Killick, Jack Heal, and Ben Norwood. Modeling and forecasting of at home activity in older adults using passive sensor technology. Statistics in Medicine, 41(23):4629–4646, October 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jiang:2022:BMI

[1020] Xi Jiang, Guanghua Xiao, and Qiwei Li. A Bayesian modified Ising model for identifying spatially variable genes from spatial transcriptomics data. Statistics in Medicine, 41(23):4647–4665, October 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pan:2022:BAP

[1021] Anqi Pan, Xiao Song, and Hanwen Huang. Bayesian analysis for partly linear Cox model with measurement error and time-varying covariate effect. Statistics in Medicine, 41(23):4666–4681, October 15, 2022. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Self:2022:CPD

[1022] Stella Self, Christopher McMahan, and Stefani Mokalled. Capturing the pool dilution effect in group testing regression: a Bayesian approach. Statistics in Medicine, 41(23):4682–4696, October 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jiang:2022:CDR

[1023] Huijun Jiang, Quefeng Li, Jessica T. Lin, and Feng-Chang Lin. Classification of disease recurrence using transition likelihoods with expectationmaximization algorithm. Statistics in Medicine, 41(23):4697–4715, October 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Witte:2022:MIT

[1024] Janine Witte, Ronja Foraita, and Vanessa Didelez. Multiple imputation and test-wise deletion for causal discovery with incomplete cohort data. Statistics in Medicine, 41(23):4716–4743, October 15, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schoenfeld:2022:DLC

[1025] David A. Schoenfeld, Ritesh Ramchandani, and Dianne M. Finkelstein. Designing a longitudinal clinical trial based on a composite endpoint: Sample size, monitoring, and adaptation. Statistics in Medicine, 41 (24):4745–4755, October 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pfeiffer:2022:APD

[1026] Ruth M. Pfeiffer, Yiyao Chen, Mitchell H. Gail, and Donna P. Ankerst. Accommodating population differences when validating risk prediction models. Statistics in Medicine, 41(24):4756–4780, October 30, 2022. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thiessen:2022:UEC

[1027] David Luke Thiessen, Yang Zhao, and Dongsheng Tu. Unified estimation for Cox regression model with nonmonotone missing at random covariates. Statistics in Medicine, 41(24):4781–4790, October 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2022:SAG

[1028] Ling-Wan Chen, Jason P. Fine, Eric Bair, Victor S. Ritter, Thomas F. McElrath, David E. Cantonwine, John D. Meeker, Kelly K. Ferguson, and Shanshan Zhao. Semiparametric analysis of a generalized linear model with multiple covariates subject to detection limits. Statistics in Medicine, 41(24):4791–4808, October 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2022:PSL

[1029] Xiudi Li, Susanne May, Ilana M. Trumble, Nancie M. Archin, and Michael G. Hudgens. Paired serial limiting dilution assays. Statistics in Medicine, 41(24):4809–4821, October 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Riley:2022:CPE

[1030] Richard D. Riley, Miriam Hattle, Gary S. Collins, Rebecca Whittle, and Joie Ensor. Calculating the power to examine treatment-covariate interactions when planning an individual participant data meta-analysis of randomized trials with a binary outcome. Statistics in Medicine, 41 (24):4822–4837, October 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wu:2022:WGS

[1031] Yougui Wu. Weighted generalized score test for comparing predictive values in the presence of verification bias. Statistics in Medicine, 41 (24):4838–4859, October 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Blaha:2022:DAC

[1032] Ondrej Blaha, Denise Esserman, and Fan Li. Design and analysis of cluster randomized trials with time-to-event outcomes under the additive hazards mixed model. Statistics in Medicine, 41(24):4860–4885, October 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jadhav:2022:FBA

[1033] Sneha Jadhav, Carmen D. Tekwe, and Yuanyuan Luan. A function-based approach to model the measurement error in wearable devices. Statistics in Medicine, 41(24):4886–4902, October 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2022:DRE

[1034] Zhang Zhang, Danhui Yi, and Yiwei Fan. Doubly robust estimation of optimal dynamic treatment regimes with multicategory treatments and survival outcomes. Statistics in Medicine, 41(24):4903–4923, October 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cai:2022:CDH

[1035] Zhanrui Cai, Dong Xi, Xuan Zhu, and Runze Li. Causal discoveries for high dimensional mixed data. Statistics in Medicine, 41(24):4924– 4940, October 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2022:FLT

[1036] Hong Wang, Zhenyuan Shen, Zhelun Tan, Zhuan Zhang, and Gang Li. Fast Lasso-type safe screening for Fine–Gray competing risks model with ultrahigh dimensional covariates. Statistics in Medicine, 41(24):4941– 4960, October 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Papanikos:2022:UCM

[1037] Tasos Papanikos, John R. Thompson, Keith R. Abrams, and Sylwia Bujkiewicz. Use of copula to model within-study association in bivariate meta-analysis of binomial data at the aggregate level: a Bayesian approach and application to surrogate endpoint evaluation. Statistics in Medicine, 41(25):4961–4981, November 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hu:2022:FAC

[1038] Liangyuan Hu, Jiayi Ji, Ronald D. Ennis, and Joseph W. Hogan. A flexible approach for causal inference with multiple treatments and clustered survival outcomes. Statistics in Medicine, 41(25):4982–4999, November 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Quartagno:2022:SMC

[1039] Matteo Quartagno and James R. Carpenter. Substantive model compatible multilevel multiple imputation: a joint modeling approach. Statistics in Medicine, 41(25):5000–5015, November 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chang:2022:FPS

[1040] Ting-Hsuan Chang, Trang Quynh Nguyen, Youjin Lee, John W. Jackson, and Elizabeth A. Stuart. Flexible propensity score estimation strategies for clustered data in observational studies. Statistics in Medicine, 41 (25):5016–5032, November 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brannath:2022:SST

[1041] Werner Brannath, Martin Scharpenberg, and Sylvia Schmidt. Singlestage, three-arm, adaptive test strategies for non-inferiority trials with an unstable reference. Statistics in Medicine, 41(25):5033–5045, November 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Niu:2022:CMW

[1042] Xin Niu, Jiangtao Gou, Hansoo Chang, Michael Lowe, and Fengqing (Zoe) Zhang. Classification model with weighted regularization to improve the reproducibility of neuroimaging signature selection. Statistics in Medicine, 41(25):5046–5060, November 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yu:2022:CTO

[1043] Yian Yu, Cong Xu, Junjiang Zhong, and Siu Hung Cheung. Comparison of treatments with ordinal responses in trials with sequential monitoring and response-adaptive randomization. Statistics in Medicine, 41(25): 5061–5083, November 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Di:2022:CEE

[1044] Fengrui Di, Lei Wang, and Heng Lian. Communication-efficient estimation and inference for high-dimensional quantile regression based on smoothed decorrelated score. Statistics in Medicine, 41(25):5084–5101, November 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gallagher:2022:NTC

[1045] Shannon K. Gallagher, Jing Wang, Keith Lumbard, Lori E. Dodd, and Michael Proschan. Noninferiority testing with censoring when the event rate is low. Statistics in Medicine, 41(25):5102–5112, November 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Luo:2022:OTW

[1046] Lan Luo and Lexin Li. Online two-way estimation and inference via linear mixed-effects models. Statistics in Medicine, 41(25):5113–5133, November 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Diao:2022:SSC

[1047] Guoqing Diao, Haijun Ma, Donglin Zeng, Chunlei Ke, and Joseph G. Ibrahim. Synthesizing studies for comparing different treatment sequences in clinical trials. Statistics in Medicine, 41(25):5134–5149, November 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shutta:2022:GGM

[1048] Katherine H. Shutta, Roberta De Vito, Denise M. Scholtens, and Raji Balasubramanian. Gaussian graphical models with applications to omics analyses. Statistics in Medicine, 41(25):5150–5187, November 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Aliverti:2022:DMI

[1049] Emanuele Aliverti and Massimiliano Russo. Dynamic modeling of the Italians’ attitude towards Covid-19. Statistics in Medicine, 41(26):5189– 5202, November 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Evrenoglou:2022:NMA

[1050] Theodoros Evrenoglou, Ian R. White, Sivem Afach, Dimitris Mavridis, and Anna Chaimani. Network meta-analysis of rare events using penalized likelihood regression. Statistics in Medicine, 41(26):5203–5219, November 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2022:LBR

[1051] Jiachen Chen, Ruofan Bie, Yichen Qin, Yang Li, and Shuangge Ma. Lqbased robust analytics on ultrahigh and high dimensional data. Statistics in Medicine, 41(26):5220–5241, November 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Huang:2022:PSD

[1052] Xin Huang, Lu Tian, Yan Sun, Saptarshi Chatterjee, and Viswanath Devanarayan. Predictive signature development based on maximizing the area between receiver operating characteristic curves. Statistics in Medicine, 41(26):5242–5257, November 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Morzywolek:2022:ECV

[1053] Paweł Morzywołek, Johan Steen, Wim Van Biesen, Johan Decruyenaere, and Stijn Vansteelandt. On estimation and cross-validation of dynamic treatment regimes with competing risks. Statistics in Medicine, 41(26): 5258–5275, November 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Toyoizumi:2022:BCB

[1054] Kiichiro Toyoizumi and Shigeyuki Matsui. Bias correction based on weighted likelihood for conditional estimation of subgroup effects in randomized clinical trials. Statistics in Medicine, 41(26):5276–5289, November 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Han:2022:ISM

[1055] Larry Han, Xuan Wang, and Tianxi Cai. Identifying surrogate markers in real-world comparative effectiveness research. Statistics in Medicine, 41(26):5290–5304, November 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2022:SPW

[1056] Tuo Wang and Lu Mao. Stratified proportional win-fractions regression analysis. Statistics in Medicine, 41(26):5305–5318, November 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chi:2022:BBO

[1057] Xiaohan Chi, Zhangsheng Yu, and Ruitao Lin. BOB: Bayesian optimal design for biosimilar trials with co-primary endpoints. Statistics in Medicine, 41(26):5319–5334, November 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sheng:2022:AMC

[1058] Yanghui Sheng, Ce Yang, Sharon Curhan, Gary Curhan, and Molin Wang. Analytical methods for correlated data arising from multicenter hearing studies. Statistics in Medicine, 41(26):5335–5348, November 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

He:2022:MST

[1059] Yizeng He, Soyoung Kim, Lu Mao, and Kwang Woo Ahn. Marginal semiparametric transformation models for clustered multivariate competing risks data. Statistics in Medicine, 41(26):5349–5364, November 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhan:2022:DBI

[1060] Tianyu Zhan. DL 101: Basic introduction to deep learning with its application in biomedical related fields. Statistics in Medicine, 41(26): 5365–5378, November 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Guo:2022:CWL

[1061] Xiaohan Guo and Ai Ni. Contrast weighted learning for robust optimal treatment rule estimation. Statistics in Medicine, 41(27):5379–5394, November 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chakraborty:2022:ULR

[1062] Saptarshi Chakraborty, Anran Liu, Robert Ball, and Marianthi Markatou. On the use of the likelihood ratio test methodology in pharmacovigilance. Statistics in Medicine, 41(27):5395–5420, November 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yang:2022:OSP

[1063] Zehan Yang, HaiYing Wang, and Jun Yan. Optimal subsampling for parametric accelerated failure time models with massive survival data. Statistics in Medicine, 41(27):5421–5431, November 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jin:2022:DST

[1064] Jin Jin, Xinyuan Song, and Liuquan Sun. Dynamic semiparametric transformation models for recurrent event data with a terminal event. Statistics in Medicine, 41(27):5432–5447, November 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cheng:2022:RAC

[1065] Chao Cheng, Xingdong Feng, Xiaoguang Li, and Mengyun Wu. Robust analysis of cancer heterogeneity for high-dimensional data. Statistics in Medicine, 41(27):5448–5462, November 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2022:IPB

[1066] Hongtao Zhang, Alan Y. Chiang, and Jixian Wang. Improving the performance of Bayesian logistic regression model with overdose control in oncology dose-finding studies. Statistics in Medicine, 41(27):5463–5483, November 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See commentaries [1067, 1069] and rejoinder [1070].

Yuan:2022:CIP

[1067] Ying Yuan and Yixuan Zhao. Commentary on “Improving the performance of Bayesian logistic regression model with overdose control in oncology dose-finding studies”. Statistics in Medicine, 41(27):5484–5490, November 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1066] and rejoinder [1070].

Frankel:2022:EUC

[1068] Paul H. Frankel, Elizabeth Garrett-Mayer, and Mark D. Krailo. Explicit underdose control based on toxicity: Four points to consider. Statistics in Medicine, 41(27):5491–5493, November 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Iasonos:2022:IPB

[1069] Alexia Iasonos and Graham M. Wheeler. “Improving the performance of Bayesian logistic regression model with overdose control in oncology dose-finding studies” by Hongtao Zhang, Alan Chiang, and Jixian Wang. Statistics in Medicine, 41(27):5494–5496, November 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1066] and rejoinder [1070].

Zhang:2022:RIP

[1070] Hongtao Zhang, Alan Y. Chiang, and Jixian Wang. Rejoinder: Improving the performance of Bayesian logistic regression model with overdose control in oncology dose-finding studies. Statistics in Medicine, 41(27): 5497–5500, November 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1066, 1067, 1069].

Beesley:2022:CSB

[1071] Lauren J. Beesley and Bhramar Mukherjee. Case studies in bias reduction and inference for electronic health record data with selection bias and phenotype misclassification. Statistics in Medicine, 41(28):5501– 5516, December 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tsiatis:2022:GSM

[1072] Anastasios A. Tsiatis and Marie Davidian. Group sequential methods for interim monitoring of randomized clinical trials with time-lagged outcome. Statistics in Medicine, 41(28):5517–5536, December 10, 2022. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Raket:2022:PMR

[1073] Lars Lau Raket. Progression models for repeated measures: Estimating novel treatment effects in progressive diseases. Statistics in Medicine, 41 (28):5537–5557, December 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Remiro-Azocar:2022:TEP

[1074] Antonio Remiro-Azócar. Target estimands for population-adjusted indirect comparisons. Statistics in Medicine, 41(28):5558–5569, December 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See commentary [1075, 1076].

Schiel:2022:CTE

[1075] Anja Schiel. Commentary on “Target estimands for population-adjusted indirect comparisons”. Statistics in Medicine, 41(28):5570–5572, December 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1074].

Russek-Cohen:2022:DTE

[1076] Estelle Russek-Cohen. Discussion of “Target estimands for populationadjusted indirect comparisons” by Antonio Remiro-Azocar. Statistics in Medicine, 41(28):5573–5576, December 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1074].

VanLancker:2022:EHT

[1077] Kelly Van Lancker, Tat-Thang Vo, and Mouna Akacha. Estimands in health technology assessment: a causal inference perspective. Statistics in Medicine, 41(28):5577–5585, December 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Senn:2022:CSM

[1078] Stephen Senn. Conditions for success and margins of error: Estimation in clinical trials. Statistics in Medicine, 41(28):5586–5588, December 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Spieker:2022:CDB

[1079] Andrew J. Spieker. Comments on the debate between marginal and conditional estimands. Statistics in Medicine, 41(28):5589–5591, December 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Remiro-Azocar:2022:SCT

[1080] Antonio Remiro-Azócar. Some considerations on target estimands for health technology assessment. Statistics in Medicine, 41(28):5592–5596, December 10, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kurum:2022:BMT

[1081] Esra Kürüm, Danh V. Nguyen, Sudipto Banerjee, Yihao Li, Connie M. Rhee, and Damla Sentürk. A Bayesian multilevel time-varying framework for joint modeling of hospitalization and survival in patients on dialysis. Statistics in Medicine, 41(29):5597–5611, December 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Postmus:2022:SML

[1082] Douwe Postmus, Francesco Pignatti, Hans L. Hillege, and Tommi Tervonen. A simulated maximum likelihood procedure for analyzing imprecise trade-off thresholds between the benefits and harms of medicines. Statistics in Medicine, 41(29):5612–5621, December 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Igeta:2022:BSS

[1083] Masataka Igeta and Shigeyuki Matsui. Blinded sample size re-estimation for comparing over-dispersed count data incorporating follow-up lengths. Statistics in Medicine, 41(29):5622–5644, December 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ma:2022:RAC

[1084] Wei Ma, Fuyi Tu, and Hanzhong Liu. Regression analysis for covariateadaptive randomization: a robust and efficient inference perspective. Statistics in Medicine, 41(29):5645–5661, December 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Utazi:2022:CPR

[1085] Chigozie Edson Utazi, Justice Moses K. Aheto, Ho Man Theophilus Chan, Andrew J. Tatem, and Sujit K. Sahu. Conditional probability and ratio-based approaches for mapping the coverage of multi-dose vaccines. Statistics in Medicine, 41(29):5662–5678, December 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Boyle:2022:EME

[1086] Joseph Boyle, Mary H. Ward, James R. Cerhan, Nat Rothman, and David C. Wheeler. Estimating mixture effects and cumulative spatial risk over time simultaneously using a Bayesian index low-rank kriging multiple membership model. Statistics in Medicine, 41(29):5679–5697, December 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xie:2022:FHC

[1087] Can Xie, Xuelin Huang, Ruosha Li, and Peter W. T. Pisters. A flexiblehazards cure model with application to patients with soft tissue sarcoma. Statistics in Medicine, 41(29):5698–5714, December 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2022:MTP

[1088] Bing Wang, Jialiang Li, and Xiaoguang Wang. Multi-threshold proportional hazards model and subgroup identification. Statistics in Medicine, 41(29):5715–5737, December 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhu:2022:PSI

[1089] Jian Zhu and Rui (Sammi) Tang. A proper statistical inference framework to compare clinical trial and real-world progression-free survival data. Statistics in Medicine, 41(29):5738–5752, December 20, 2022. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Andrillon:2022:SCB

[1090] Anaïs Andrillon, Sylvie Chevret, Shing M. Lee, and Lucie Biard. Surv-CRM-12: a Bayesian phase I/II survival CRM for right-censored toxicity endpoints with competing disease progression. Statistics in Medicine, 41 (29):5753–5766, December 20, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Barnett:2022:DFS

[1091] Helen Barnett, Oliver Boix, Dimitris Kontos, and Thomas Jaki. Dose finding studies for therapies with late-onset toxicities: a comparison study of designs. Statistics in Medicine, 41(30):5767–5788, December 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mozgunov:2022:PIP

[1092] Pavel Mozgunov, Thomas Jaki, Ioannis Gounaris, Thomas Goddemeier, Anja Victor, and Marianna Grinberg. Practical implementation of the partial ordering continual reassessment method in a Phase I combinationschedule dose-finding trial. Statistics in Medicine, 41(30):5789–5809, December 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

LeClair:2022:SSD

[1093] Jessica LeClair, Joseph Massaro, Oleksandr Sverdlov, Leslie Gordon, and Yorghos Tripodis. Sample size determination for the association between longitudinal and time-to-event outcomes using the joint modeling timedependent slopes parameterization. Statistics in Medicine, 41(30):5810– 5829, December 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2022:OAS

[1094] Woojoo Lee, Donghwan Lee, and Yudi Pawitan. Overall assessment for selected markers from high-throughput data. Statistics in Medicine, 41 (30):5830–5843, December 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cao:2022:REI

[1095] Yi Cao, Heather Allore, Brent Vander Wyk, and Roee Gutman. Review and evaluation of imputation methods for multivariate longitudinal data with mixed-type incomplete variables. Statistics in Medicine, 41(30): 5844–5876, December 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bujkiewicz:2022:BMA

[1096] Sylwia Bujkiewicz. Bayesian meta-analytical methods to incorporate multiple surrogate endpoints in drug development process. Statistics in Medicine, 41(30):5877–5878, December 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Alt:2023:STP

[1097] Ethan M. Alt, Brady Nifong, Xinxin Chen, Matthew A. Psioda, and Joseph G. Ibrahim. The scale transformed power prior for use with historical data from a different outcome model. Statistics in Medicine, 42 (1):1–14, January 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Castelletti:2023:BGM

[1098] Federico Castelletti and Guido Consonni. Bayesian graphical modeling for heterogeneous causal effects. Statistics in Medicine, 42(1):15–32, January 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhu:2023:APV

[1099] Angela Yaqian Zhu, Nandita Mitra, and Jason Roy. Addressing positivity violations in causal effect estimation using Gaussian process priors. Statistics in Medicine, 42(1):33–51, January 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2023:DIC

[1100] Zheng Wang, Zi Wang, Lingyun Lyu, Yu Cheng, Eric C. Seaberg, Samantha A. Molsberry, Ann Ragin, and James T. Becker. Dynamic impairment classification through arrayed comparisons. Statistics in Medicine, 42(1):52–67, January 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Parast:2023:USH

[1101] Layla Parast, Tianxi Cai, and Lu Tian. Using a surrogate with heterogeneous utility to test for a treatment effect. Statistics in Medicine, 42(1): 68–88, January 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhai:2023:TST

[1102] Jingyi Zhai and Hui Jiang. Two-sample test with g-modeling and its applications. Statistics in Medicine, 42(1):89–104, January 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Payares-Garcia:2023:SIB

[1103] David Payares-Garcia, Jorge Mateu, and Wiebke Schick. Spatially informed Bayesian neural network for neurodegenerative diseases classification. Statistics in Medicine, 42(2):105–121, January 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Robertson:2023:PEAa

[1104] David S. Robertson, Babak Choodari-Oskooei, Munya Dimairo, Laura Flight, Philip Pallmann, and Thomas Jaki. Point estimation for adaptive trial designs I: a methodological review. Statistics in Medicine, 42(2): 122–145, January 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Khan:2023:ATS

[1105] Josephine N. Khan, Peter K. Kimani, Ekkehard Glimm, and Nigel Stallard. Adjusting for treatment selection in phase II/III clinical trials with time to event data. Statistics in Medicine, 42(2):146–163, January 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Huang:2023:CCD

[1106] Ying Huang and Yunda Huang. Comparing and combining data from immune assays based on left-censored multivariate normal model assuming common assay differences across settings. Statistics in Medicine, 42(2): 164–177, January 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Talbot:2023:DRE

[1107] Denis Talbot, Erica E. M. Moodie, and Caroline Diorio. Double robust estimation of optimal partially adaptive treatment strategies: an application to breast cancer treatment using hormonal therapy. Statistics in Medicine, 42(2):178–192, January 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Saha:2023:VSD

[1108] Abhisek Saha and Rajeshwari Sundaram. Variable selection for discrete survival model with frailty in presence of left truncation and right censoring: Studying association of environmental toxicants on time-to-pregnancy. Statistics in Medicine, 42(2):193–208, January 30, 2023. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Selukar:2023:RRE

[1109] Subodh Selukar and Megan Othus. RECeUS: Ratio estimation of censored uncured subjects, a different approach for assessing cure model appropriateness in studies with long-term survivors. Statistics in Medicine, 42(3):209–227, February 10, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Peng:2023:MEV

[1110] Yingwei Peng, Yuyao Wang, and Ronghui Xu. Measures of explained variation under the mixture cure model for survival data. Statistics in Medicine, 42(3):228–245, February 10, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Um:2023:BAR

[1111] Seungha Um, Antonio R. Linero, Debajyoti Sinha, and Dipankar Bandyopadhyay. Bayesian additive regression trees for multivariate skewed responses. Statistics in Medicine, 42(3):246–263, February 10, 2023. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2023:GMR

[1112] Wenwen Li, Huijuan Ma, David Faraggi, and Gregg E. Dinse. Generalized mean residual life models for survival data with missing censoring indicators. Statistics in Medicine, 42(3):264–280, February 10, 2023. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2023:OCI

[1113] Weizhen Wang, Shuiyun Lu, and Tianfa Xie. Optimal confidence intervals for the relative risk and odds ratio. Statistics in Medicine, 42 (3):281–296, February 10, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Teng:2023:NSC

[1114] Wen Teng, Wenyu Jiang, and Bingshu E. Chen. A nonparametric simultaneous confidence band for biomarker effect on the restricted mean survival time. Statistics in Medicine, 42(3):297–315, February 10, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tu:2023:GVA

[1115] Jieqi Tu and Jiehuan Sun. Gaussian variational approximate inference for joint models of longitudinal biomarkers and a survival outcome. Statistics in Medicine, 42(3):316–330, February 10, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Buch:2023:SRE

[1116] Gregor Buch, Andreas Schulz, Irene Schmidtmann, Konstantin Strauch, and Philipp S. Wild. A systematic review and evaluation of statistical methods for group variable selection. Statistics in Medicine, 42(3):331– 352, February 10, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Samoilenko:2023:ERB

[1117] Mariia Samoilenko and Geneviève Lefebvre. An exact regression-based approach for the estimation of natural direct and indirect effects with a binary outcome and a continuous mediator. Statistics in Medicine, 42 (3):353–387, February 10, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2023:EDR

[1118] Li Chen, Mark Burkard, Jianrong Wu, Jill M. Kolesar, and Chi Wang. Estimating the distribution of ratio of paired event times in phase II oncology trials. Statistics in Medicine, 42(3):388–406, February 10, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Geng:2023:GFT

[1119] Ziqi Geng, Jialiang Li, Yi Niu, and Xiaoguang Wang. Goodness-of-fit test for a parametric mixture cure model with partly interval-censored data. Statistics in Medicine, 42(4):407–421, February 20, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wickramarachchi:2023:MAM

[1120] Deshanee Senevirathne Wickramarachchi, Laura Huey Mien Lim, and Baoluo Sun. Mediation analysis with multiple mediators under unmeasured mediator-outcome confounding. Statistics in Medicine, 42(4):422– 432, February 20, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Drouin:2023:SSC

[1121] Pierre Drouin, Aymeric Stamm, Laurent Chevreuil, Vincent Graillot, Laetitia Barbin, Pierre-Antoine Gourraud, David-Axel Laplaud, and Lise Bellanger. Semi-supervised clustering of quaternion time series: Application to gait analysis in multiple sclerosis using motion sensor data. Statistics in Medicine, 42(4):433–456, February 20, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kojima:2023:ACF

[1122] Masahiro Kojima. Adjusted closed-form confidence interval formulas for network meta-analysis with a small number of studies. Statistics in Medicine, 42(4):457–469, February 20, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Su:2023:RMA

[1123] Xiaogang Su, Youngjoo Cho, Liqiang Ni, Lei Liu, and Elise Dusseldorp. Refined moderation analysis with categorical outcomes in precision medicine. Statistics in Medicine, 42(4):470–486, February 20, 2023. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Varga:2023:DCO

[1124] Anita Natalia Varga, Alejandra Elizabeth Guevara Morel, Joran Lokkerbol, Johanna Maria van Dongen, Maurits Willem van Tulder, and Judith Ekkina Bosmans. Dealing with confounding in observational studies: a scoping review of methods evaluated in simulation studies with singlepoint exposure. Statistics in Medicine, 42(4):487–516, February 20, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yan:2023:CIP

[1125] Ying Yan and Mingchen Ren. Consistent inverse probability of treatment weighted estimation of the average treatment effect with mismeasured time-dependent confounders. Statistics in Medicine, 42(4):517–535, February 20, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Meis:2023:OTS

[1126] Jan Meis, Maximilian Pilz, Carolin Herrmann, Björn Bokelmann, Geraldine Rauch, and Meinhard Kieser. Optimization of the two-stage group sequential three-arm gold-standard design for non-inferiority trials. Statistics in Medicine, 42(4):536–558, February 20, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Davis-Plourde:2023:PAS

[1127] Kendra Davis-Plourde, Monica Taljaard, and Fan Li. Power analyses for stepped wedge designs with multivariate continuous outcomes. Statistics in Medicine, 42(4):559–578, February 20, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Geraci:2023:JRM

[1128] Marco Geraci. Joint regression modelling of intensity and timing of accelerometer counts. Statistics in Medicine, 42(4):579–595, February 20, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Farewell:2023:OAL

[1129] Vern Farewell. Obituary: Anthony Leonard Johnson (1943–2022). Statistics in Medicine, 42(5):597–599, February 28, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Freedman:2023:TJM

[1130] Laurence S. Freedman and Theodore Colton. Tony Johnson — memories and a tribute. Statistics in Medicine, 42(5):600–602, February 28, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Starkopf:2023:MSM

[1131] Liis Starkopf, Shahzleen Rajan, Theis Lange, and Thomas Alexander Gerds. Marginal structural models with monotonicity constraints: a case study in out-of-hospital cardiac arrest patients. Statistics in Medicine, 42(5):603–618, February 28, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schuemie:2023:ABS

[1132] Martijn J. Schuemie, Fan Bu, Akihiko Nishimura, and Marc A. Suchard. Adjusting for both sequential testing and systematic error in safety surveillance using observational data: Empirical calibration and MaxSPRT. Statistics in Medicine, 42(5):619–631, February 28, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Butera:2023:MLC

[1133] Nicole M. Butera, Donglin Zeng, Gerardo Heiss, and Jianwen Cai. Modeling longitudinal change in biomarkers using data from a complex survey sampling design: an application to the Hispanic Community Health Study/Study of Latinos. Statistics in Medicine, 42(5):632–655, February 28, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2023:NRE

[1134] Shaoxin Wang, Chaoping Xie, and Xiaoning Kang. A novel robust estimation for high-dimensional precision matrices. Statistics in Medicine, 42(5):656–675, February 28, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schauberger:2023:TBM

[1135] Gunther Schauberger, Luana Fiengo Tanaka, and Moritz Berger. A treebased modeling approach for matched case-control studies. Statistics in Medicine, 42(5):676–692, February 28, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yuki:2023:EVH

[1136] Shintaro Yuki, Kensuke Tanioka, and Hiroshi Yadohisa. Estimation and visualization of heterogeneous treatment effects for multiple outcomes. Statistics in Medicine, 42(5):693–715, February 28, 2023. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yan:2023:CIA

[1137] Qinling Yan, Robert A. Cheke, and Sanyi Tang. Coupling an individual adaptive-decision model with a SIRV model of influenza vaccination reveals new insights for epidemic control. Statistics in Medicine, 42(5): 716–729, February 28, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Breskin:2023:CFD

[1138] Alexander Breskin, Stephen R. Cole, Jessie K. Edwards, Ron Brookmeyer, Joseph J. Eron, and Adimora A. Adimora. Corrigendum to: Fusion designs and estimators for treatment effects. Statistics in Medicine, 42(5):730, February 28, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [515].

Delaigle:2023:GTR

[1139] Aurore Delaigle and Ruoxu Tan. Group testing regression analysis with covariates and specimens subject to missingness. Statistics in Medicine, 42(6):731–744, March 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kang:2023:ADF

[1140] Tong Kang, Jeremy Gaskins, Steven Levy, and Somnath Datta. Analyzing dental fluorosis data using a novel Bayesian model for clustered longitudinal ordinal outcomes with an inflated category. Statistics in Medicine, 42(6):745–760, March 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yu:2023:VIE

[1141] Duo Yu and Hulin Wu. Variable importance evaluation with personalized odds ratio for machine learning model interpretability with applications to electronic health records-based mortality prediction. Statistics in Medicine, 42(6):761–780, March 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhou:2023:LBS

[1142] Yi Zhou, Ao Huang, and Satoshi Hattori. A likelihood-based sensitivity analysis for publication bias on the summary receiver operating characteristic in meta-analysis of diagnostic test accuracy. Statistics in Medicine, 42(6):781–798, March 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rava:2023:DRE

[1143] Denise Rava and Ronghui Xu. Doubly robust estimation of the hazard difference for competing risks data. Statistics in Medicine, 42(6):799–814, March 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Webb:2023:CMT

[1144] Annabel Webb and Jun Ma. Cox models with time-varying covariates and partly-interval censoring — a maximum penalised likelihood approach. Statistics in Medicine, 42(6):815–833, March 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Feng:2023:NAC

[1145] Qunqiang Feng, Pan Liu, Pei-Fen Kuan, Fei Zou, Jianan Chen, and Jialiang Li. A network approach to compute hypervolume under receiver operating characteristic manifold for multi-class biomarkers. Statistics in Medicine, 42(6):834–859, March 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ye:2023:RAL

[1146] Yuan Ye, Zhongchun Liu, Deng Pan, and Yuanshan Wu. Regression analysis of logistic model with latent variables. Statistics in Medicine, 42(6):860–877, March 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Giai:2023:INB

[1147] Joris Giai, Julien Péron, Matthieu Roustit, Jean-Luc Cracowski, Pascal Roy, Brice Ozenne, Marc Buyse, and Delphine Maucort-Boulch. Individualized Net Benefit estimation and meta-analysis using generalized pairwise comparisons in N-of-1 trials. Statistics in Medicine, 42(6):878–893, March 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ding:2023:FAR

[1148] Jie Ding, Jialiang Li, Yang Han, Ian W. McKeague, and Xiaoguang Wang. Fitting additive risk models using auxiliary information. Statistics in Medicine, 42(6):894–916, March 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sauer:2023:PSO

[1149] Sara Sauer, Bethany Hedt-Gauthier, and Sebastien Haneuse. Practical strategies for operationalizing optimal allocation in stratified clusterbased outcome-dependent sampling designs. Statistics in Medicine, 42 (7):917–935, March 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Uno:2023:RDA

[1150] Hajime Uno and Miki Horiguchi. Ratio and difference of average hazard with survival weight: New measures to quantify survival benefit of new therapy. Statistics in Medicine, 42(7):936–952, March 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2023:CSE

[1151] Ziyi Li, Yijian Huang, Dattatraya Patil, Mark Rubin, and Martin G. Sanda. Covariate-specific evaluation of continuous biomarker. Statistics in Medicine, 42(7):953–969, March 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shu:2023:RCI

[1152] Di Shu, Peisong Han, Sean Hennessy, and Todd A. Miano. Robust causal inference of drug-drug interactions. Statistics in Medicine, 42(7):970–992, March 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Collazos:2023:MED

[1153] Julian A. A. Collazos, Ronaldo Dias, and Marcelo C. Medeiros. Modeling the evolution of deaths from infectious diseases with functional data models: The case of COVID-19 in Brazil. Statistics in Medicine, 42(7): 993–1012, March 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Malenica:2023:POE

[1154] Ivana Malenica, Rachael V. Phillips, Antoine Chambaz, Alan E. Hubbard, Romain Pirracchio, and Mark J. van der Laan. Personalized online ensemble machine learning with applications for dynamic data streams. Statistics in Medicine, 42(7):1013–1044, March 30, 2023. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Katahira:2023:EPP

[1155] Kentaro Katahira. Evaluating the predictive performance of subtyping: a criterion for cluster mean-based prediction. Statistics in Medicine, 42 (7):1045–1065, March 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rubio:2023:IFE

[1156] Francisco J. Rubio, Hein Putter, and Aurélien Belot. Individual frailty excess hazard models in cancer epidemiology. Statistics in Medicine, 42 (7):1066–1081, March 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Segalas:2023:PSM

[1157] Corentin Ségalas, Clémence Leyrat, James R. Carpenter, and Elizabeth Williamson. Propensity score matching after multiple imputation when a confounder has missing data. Statistics in Medicine, 42(7):1082–1095, March 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ghosh:2023:NAA

[1158] Palash Ghosh, Xiaoxi Yan, and Bibhas Chakraborty. A novel approach to assess dynamic treatment regimes embedded in a SMART with an ordinal outcome. Statistics in Medicine, 42(7):1096–1111, March 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xu:2023:MAN

[1159] Wenfu Xu, Yuli Hou, and Tong-Yu Lu. Multiple assessments of noninferiority trials with ordinal endpoints. Statistics in Medicine, 42(8): 1113–1126, April 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Turner:2023:PAB

[1160] Rebecca M. Turner, Michelle N. Clements, Matteo Quartagno, Victoria Cornelius, Suzie Cro, Deborah Ford, Conor D. Tweed, A. Sarah Walker, and Ian R. White. Practical approaches to Bayesian sample size determination in non-inferiority trials with binary outcomes. Statistics in Medicine, 42(8):1127–1138, April 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Horiguchi:2023:ASB

[1161] Miki Horiguchi, Lu Tian, and Hajime Uno. On assessing survival benefit of immunotherapy using long-term restricted mean survival time. Statistics in Medicine, 42(8):1139–1155, April 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2023:PRC

[1162] Kim May Lee, Rebecca M. Turner, Guy E. Thwaites, A. Sarah Walker, and Ian R. White. The Personalised Randomized Controlled Trial: Evaluation of a new trial design. Statistics in Medicine, 42(8):1156–1170, April 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bather:2023:ICE

[1163] Jemar R. Bather, Nicholas J. Horton, Brent A. Coull, and Paige L. Williams. The impact of correlated exposures and missing data on multiple informant models used to identify critical exposure windows. Statistics in Medicine, 42(8):1171–1187, April 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Efthimiou:2023:MPP

[1164] Orestis Efthimiou, Jeroen Hoogland, Thomas P. A. Debray, Michael Seo, Toshiaki A. Furukawa, Matthias Egger, and Ian R. White. Measuring the performance of prediction models to personalize treatment choice. Statistics in Medicine, 42(8):1188–1206, April 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mao:2023:TPD

[1165] Fangya Mao and Richard J. Cook. Two-phase designs with current status data. Statistics in Medicine, 42(8):1207–1232, April 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chauvet:2023:FCG

[1166] Jocelyn Chauvet and Virginie Rondeau. A flexible class of generalized joint frailty models for the analysis of survival endpoints. Statistics in Medicine, 42(8):1233–1262, April 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2023:ARR

[1167] Danping Liu, Emily Wu, Joanna H. Shih, Cari M. Kitahara, and Li C. Cheung. Absolute and relative risk estimation in the presence of outcome ascertainment gaps and competing risks. Statistics in Medicine, 42(8): 1263–1276, April 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bottomley:2023:PPW

[1168] C. Bottomley, M. Ooko, A. Gasparrini, and Rh Keogh. In praise of Prais–Winsten: an evaluation of methods used to account for autocorrelation in interrupted time series. Statistics in Medicine, 42(8):1277–1288, April 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Song:2023:STD

[1169] Minsun Song, Efstathia Bura, Roman Parzer, and Ruth M. Pfeiffer. Structured time-dependent inverse regression (STIR). Statistics in Medicine, 42(9):1289–1307, April 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sheikh:2023:NCI

[1170] Md. Tuhin Sheikh, Ming-Hui Chen, Jonathan A. Gelfond, Wei Sun, and Joseph G. Ibrahim. New C-indices for assessing importance of longitudinal biomarkers in fitting competing risks survival data in the presence of partially masked causes. Statistics in Medicine, 42(9):1308–1322, April 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Antognini:2023:SAB

[1171] Alessandro Baldi Antognini, Marco Novelli, and Maroussa Zagoraiou. Simulated annealing for balancing covariates. Statistics in Medicine, 42 (9):1323–1337, April 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Maronge:2023:MMR

[1172] Jacob M. Maronge, Jonathan S. Schildcrout, and Paul J. Rathouz. Model misspecification and robust analysis for outcome-dependent sampling designs under generalized linear models. Statistics in Medicine, 42(9):1338– 1352, April 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Huang:2023:DMD

[1173] Hengzhen Huang, Min-Qian Liu, Ming T. Tan, and Hong-Bin Fang. Design and modeling for drug combination experiments with order effects. Statistics in Medicine, 42(9):1353–1367, April 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Buhler:2023:MMF

[1174] Alexandra Bühler, Richard J. Cook, and Jerald F. Lawless. Multistate models as a framework for estimand specification in clinical trials of complex processes. Statistics in Medicine, 42(9):1368–1397, April 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2023:CCA

[1175] Wen Li, Ruosha Li, Qingxiang Yan, Ziding Feng, and Jing Ning. Conditional concordance-assisted learning under matched case-control design for combining biomarkers for population screening. Statistics in Medicine, 42(9):1398–1411, April 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Miller:2023:FSI

[1176] Ryan Miller and Patrick Breheny. Feature-specific inference for penalized regression using local false discovery rates. Statistics in Medicine, 42(9): 1412–1429, April 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gill:2023:FEM

[1177] Nathan Gill and Donald Hedeker. Fast estimation of mixed-effects location-scale regression models. Statistics in Medicine, 42(9):1430–1444, April 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Berube:2023:BHM

[1178] Sophie Bérubé, Tamaki Kobayashi, Amy Wesolowski, Douglas E. Norris, Ingo Ruczinski, William J. Moss, and Thomas A. Louis. A Bayesian hierarchical model for signal extraction from protein microarrays. Statistics in Medicine, 42(9):1445–1460, April 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Denz:2023:CDM

[1179] Robin Denz, Renate Klaaßen-Mielke, and Nina Timmesfeld. A comparison of different methods to adjust survival curves for confounders. Statistics in Medicine, 42(10):1461–1479, May 10, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wu:2023:GSM

[1180] Jianrong Wu, Yimei Li, and Liang Zhu. Group sequential multi-arm multi-stage trial design with treatment selection. Statistics in Medicine, 42(10):1480–1491, May 10, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zou:2023:MFM

[1181] Haotian Zou, Luo Xiao, Donglin Zeng, Sheng Luo, and for the Alzheimer’s Disease Neuroimaging Initiative. Multivariate functional mixed model with MRI data: an application to Alzheimer’s disease. Statistics in Medicine, 42(10):1492–1511, May 10, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhao:2023:MAE

[1182] Bangyao Zhao and Lili Zhao. Mining adverse events in large frequency tables with ontology, with an application to the vaccine adverse event reporting system. Statistics in Medicine, 42(10):1512–1524, May 10, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Austin:2023:ITE

[1183] Peter C. Austin, Daniele Giardiello, and Stef van Buuren. Impute-then-exclude versus exclude-then-impute: Lessons when imputing a variable used both in cohort creation and as an independent variable in the analysis model. Statistics in Medicine, 42(10):1525–1541, May 10, 2023. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rytgaard:2023:RAT

[1184] Helene C. W. Rytgaard, Claus T. Ekstrøm, Lars V. Kessing, and Thomas A. Gerds. Ranking of average treatment effects with generalized random forests for time-to-event outcomes. Statistics in Medicine, 42(10):1542–1564, May 10, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ren:2023:MFD

[1185] Rui Ren, Kuangnan Fang, Qingzhao Zhang, and Xiaofeng Wang. Multivariate functional data clustering using adaptive density peak detection. Statistics in Medicine, 42(10):1565–1582, May 10, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xiong:2023:EMF

[1186] Wei Xiong, Han Pan, Jianrong Wang, and Maozai Tian. An efficient model-free approach to interaction screening for high dimensional data. Statistics in Medicine, 42(10):1583–1605, May 10, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fuyama:2023:ICB

[1187] Kanako Fuyama, Mitsunori Ogawa, Junki Mizusawa, Yukihide Kanemitsu, Shin Fujita, Takuya Kawahara, Kentaro Sakamaki, and Koji Oba. Impact of correlations between prioritized outcomes on the net benefit and its estimate by generalized pairwise comparisons. Statistics in Medicine, 42(10):1606–1624, May 10, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wu:2023:IIG

[1188] Peitao Wu, Josée Dupuis, and Ching-Ti Liu. Identifying important gene signatures of BMI using network structure-aided nonparametric quantile regression. Statistics in Medicine, 42(10):1625–1639, May 10, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cai:2023:CSM

[1189] Jianwen Cai, Donglin Zeng, Haolin Li, Nicole M. Butera, Pedro L. Baldoni, Poulami Maitra, and Li Dong. Comparisons of statistical methods for handling attrition in a follow-up visit with complex survey sampling. Statistics in Medicine, 42(11):1641–1668, May 20, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cheng:2023:MAP

[1190] Chao Cheng, Donna Spiegelman, and Fan Li. Mediation analysis in the presence of continuous exposure measurement error. Statistics in Medicine, 42(11):1669–1686, May 20, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McKenzie:2023:SEA

[1191] Katelyn A. McKenzie and Jonathan D. Mahnken. Simulating and estimating agreement in the presence of multiple raters and covariates. Statistics in Medicine, 42(11):1687–1698, May 20, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2023:BET

[1192] Ming Zhang, Jackson Barth, Johan Lim, and Xinlei Wang. Bayesian estimation and testing in random-effects meta-analysis of rare binary events allowing for flexible group variability. Statistics in Medicine, 42 (11):1699–1721, May 20, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Alt:2023:BAS

[1193] Ethan M. Alt, Matthew A. Psioda, and Joseph G. Ibrahim. A Bayesian approach to study design and analysis with type I error rate control for response variables of mixed types. Statistics in Medicine, 42(11): 1722–1740, May 20, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thomassen:2023:BMR

[1194] Doranne Thomassen, Ewout Steyerberg, and Saskia le Cessie. A Bayesian (meta-)regression model for treatment effects on the risk difference scale. Statistics in Medicine, 42(11):1741–1759, May 20, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhao:2023:FTM

[1195] Ruochen Zhao and Bo Lu. Flexible template matching for observational study design. Statistics in Medicine, 42(11):1760–1778, May 20, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stromer:2023:BMS

[1196] Annika Strömer, Nadja Klein, Christian Staerk, Hannah Klinkhammer, and Andreas Mayr. Boosting multivariate structured additive distributional regression models. Statistics in Medicine, 42(11):1779–1801, May 20, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Harrison:2023:SSC

[1197] Linda J. Harrison and Rui Wang. Sample size calculation for randomized trials via inverse probability of response weighting when outcome data are missing at random. Statistics in Medicine, 42(11):1802–1821, May 20, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bayer:2023:CIP

[1198] Damon M. Bayer, Michael P. Fay, and Barry I. Graubard. Confidence intervals for prevalence estimates from complex surveys with imperfect assays. Statistics in Medicine, 42(11):1822–1867, May 20, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Han:2023:DEC

[1199] Shasha Han and Xiao-Hua Zhou. Defining estimands in clinical trials: a unified procedure. Statistics in Medicine, 42(12):1869–1887, May 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gascoigne:2023:PSS

[1200] Connor Gascoigne and Theresa Smith. Penalized smoothing splines resolve the curvature identifiability problem in age-period-cohort models with unequal intervals. Statistics in Medicine, 42(12):1888–1908, May 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lu:2023:TLC

[1201] Xiaoming Lu, Thierry Chekouo, Hua Shen, and Alexander R. de Leon. A two-level copula joint model for joint analysis of longitudinal and competing risks data. Statistics in Medicine, 42(12):1909–1930, May 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Denti:2023:MHS

[1202] Francesco Denti, Stefano Peluso, Michele Guindani, and Antonietta Mira. Multiple hypothesis screening using mixtures of non-local distributions with applications to genomic studies. Statistics in Medicine, 42(12):1931–1945, May 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ciocanea-Teodorescu:2023:CIS

[1203] Iuliana Ciocanea-Teodorescu, Els Goetghebeur, Ingeborg Waernbaum, Staffan Schön, and Erin E. Gabriel. Causal inference in survival analysis under deterministic missingness of confounders in register data. Statistics in Medicine, 42(12):1946–1964, May 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chang:2023:NHC

[1204] Joonha Chang, Hei Kit Chan, Jeffrey Lin, and Wenyaw Chan. Nonhomogeneous continuous-time Markov chain with covariates: Applications to ambulatory hypertension monitoring. Statistics in Medicine, 42 (12):1965–1980, May 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Boher:2023:NLE

[1205] Jean Marie Boher, Thomas Filleron, Pierre Bunouf, and Richard J. Cook. New late-emphasis and combination tests based on infimum and supremum logrank statistics with application in oncology trials. Statistics in Medicine, 42(12):1981–1994, May 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sanusi:2023:NEM

[1206] Busola Sanusi, Jianwen Cai, and Michael G. Hudgens. Nonparametric estimation of marked survival data in the presence of dependent censoring. Statistics in Medicine, 42(12):1995–2008, May 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hauser:2023:BGL

[1207] Paloma Hauser, Xianming Tan, Fang Chen, and Joseph G. Ibrahim. Bayesian generalized linear low rank regression models for the detection of vaccine-adverse event associations. Statistics in Medicine, 42(12): 2009–2026, May 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2023:CCP

[1208] Anonymous. Correction to “Controlled pattern imputation for sensitivity analysis of longitudinal binary and ordinal outcomes with nonignorable dropout”. Statistics in Medicine, 42(12):2027–2028, May 30, 2023. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [2].

Dahabreh:2023:SAU

[1209] Issa J. Dahabreh, James M. Robins, Sebastien J.-P. A. Haneuse, Iman Saeed, Sarah E. Robertson, Elizabeth A. Stuart, and Miguel A. Hernán. Sensitivity analysis using bias functions for studies extending inferences from a randomized trial to a target population. Statistics in Medicine, 42(13):2029–2043, June 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liao:2023:ATU

[1210] Ziwei Liao, Min Qian, Ian M. Kronish, and Ying Kuen Cheung. Analysis of N-of-1 trials using Bayesian distributed lag model with autocorrelated errors. Statistics in Medicine, 42(13):2044–2060, June 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jiang:2023:NCM

[1211] Meilin Jiang, Seonjoo Lee, A. James O’Malley, Yaakov Stern, and Zhigang Li. A novel causal mediation analysis approach for zero-inflated mediators. Statistics in Medicine, 42(13):2061–2081, June 15, 2023. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Peng:2023:DPB

[1212] Mengjiao Peng and Liming Xiang. Disease progression based feature screening for ultrahigh-dimensional survival-associated biomarkers. Statistics in Medicine, 42(13):2082–2100, June 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2023:CTA

[1213] Wenhao Li, Liang Li, and Brad C. Astor. A comparison of two approaches to dynamic prediction: Joint modeling and landmark modeling. Statistics in Medicine, 42(13):2101–2115, June 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shutta:2023:SEA

[1214] Katherine H. Shutta, Laura B. Balzer, Denise M. Scholtens, and Raji Balasubramanian. SpiderLearner: an ensemble approach to Gaussian graphical model estimation. Statistics in Medicine, 42(13):2116–2133, June 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xu:2023:PML

[1215] Changchang Xu and Shelley B. Bull. Penalized maximum likelihood inference under the mixture cure model in sparse data. Statistics in Medicine, 42(13):2134–2161, June 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anyaso-Samuel:2023:AIC

[1216] Samuel Anyaso-Samuel and Somnath Datta. Adjusting for informative cluster size in pseudo-value-based regression approaches with clustered time to event data. Statistics in Medicine, 42(13):2162–2178, June 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hartman:2023:PCI

[1217] Nicholas Hartman, Sehee Kim, Kevin He, and John D. Kalbfleisch. Pitfalls of the concordance index for survival outcomes. Statistics in Medicine, 42(13):2179–2190, June 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Keogh:2023:CIS

[1218] Ruth H. Keogh, Jon Michael Gran, Shaun R. Seaman, Gwyneth Davies, and Stijn Vansteelandt. Causal inference in survival analysis using longitudinal observational data: Sequential trials and marginal structural models. Statistics in Medicine, 42(13):2191–2225, June 15, 2023. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ditzhaus:2023:SPM

[1219] Marc Ditzhaus, Menggang Yu, and Jin Xu. Studentized permutation method for comparing two restricted mean survival times with small sample from randomized trials. Statistics in Medicine, 42(13):2226–2240, June 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Deng:2023:BAT

[1220] Yangqing Deng, Dongsheng Tu, Chris J. O’Callaghan, Derek J. Jonker, Christos S. Karapetis, Jeremy Shapiro, Geoffrey Liu, and Wei Xu. A Bayesian approach for two-stage multivariate Mendelian randomization with mixed outcomes. Statistics in Medicine, 42(13):2241–2256, June 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vimalajeewa:2023:EDO

[1221] Dixon Vimalajeewa, Scott Alan Bruce, and Brani Vidakovic. Early detection of ovarian cancer by wavelet analysis of protein mass spectra. Statistics in Medicine, 42(13):2257–2273, June 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hsu:2023:MIB

[1222] Chiu-Hsieh Hsu, Yulei He, Chengcheng Hu, and Wei Zhou. A multiple imputation-based sensitivity analysis approach for regression analysis with a missing not at random covariate. Statistics in Medicine, 42(14): 2275–2292, June 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Feng:2023:SRA

[1223] Fan Feng, Shuwei Li, Peijie Wang, Jianguo Sun, and Chaofu Ke. Semiparametric regression analysis of length-biased and partly intervalcensored data with application to an AIDS cohort study. Statistics in Medicine, 42(14):2293–2310, June 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Andreella:2023:PBT

[1224] Angela Andreella, Jesse Hemerik, Livio Finos, Wouter Weeda, and Jelle Goeman. Permutation-based true discovery proportions for functional magnetic resonance imaging cluster analysis. Statistics in Medicine, 42 (14):2311–2340, June 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Arntzen:2023:TRA

[1225] Vera H. Arntzen, Marta Fiocco, Nils Leitzinger, and Ronald B. Geskus. Towards robust and accurate estimates of the incubation time distribution, with focus on upper tail probabilities and SARS-CoV-2 infection. Statistics in Medicine, 42(14):2341–2360, June 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pullenayegum:2023:CIL

[1226] Eleanor M. Pullenayegum, Catherine Birken, Jonathon Maguire, and The Target Kids! Collaboration. Causal inference with longitudinal data subject to irregular assessment times. Statistics in Medicine, 42 (14):2361–2393, June 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wen:2023:SHT

[1227] Jiyang Wen, Mei-Cheng Wang, and Chen Hu. Simultaneous hypothesis testing for multiple competing risks in comparative clinical trials. Statistics in Medicine, 42(14):2394–2408, June 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Saha:2023:CPO

[1228] Pooja T. Saha, Jason P. Fine, and Anastasia Ivanova. CRM and partial order CRM with adaptive rescaling for dose-finding in immunotherapy trials with a continuous outcome. Statistics in Medicine, 42(14):2409– 2419, June 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Neely:2023:EMM

[1229] Megan L. Neely, Carl F. Pieper, Bida Gu, Natalia O. Dmitrieva, and Jane F. Pendergast. Exploration of model misspecification in latent class methods for longitudinal data: Correlation structure matters. Statistics in Medicine, 42(14):2420–2438, June 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rover:2023:SEI

[1230] Christian Röver, Sibylle Sturtz, Jona Lilienthal, Ralf Bender, and Tim Friede. Summarizing empirical information on between-study heterogeneity for Bayesian random-effects meta-analysis. Statistics in Medicine, 42(14):2439–2454, June 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See correction [1372].

Wu:2023:MML

[1231] Jing Wu, Lijiang Geng, Angela Starkweather, and Ming-Hui Chen. Modeling and maximum likelihood based inference of interval-censored data with unknown upper limits and time-dependent covariates. Statistics in Medicine, 42(14):2455–2474, June 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Robertson:2023:OER

[1232] David S. Robertson, James M. S. Wason, Franz König, Martin Posch, and Thomas Jaki. Online error rate control for platform trials. Statistics in Medicine, 42(14):2475–2495, June 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Robertson:2023:PEAb

[1233] David S. Robertson, Babak Choodari-Oskooei, Munya Dimairo, Laura Flight, Philip Pallmann, and Thomas Jaki. Point estimation for adaptive trial designs II: Practical considerations and guidance. Statistics in Medicine, 42(14):2496–2520, June 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Aghayerashti:2023:PLB

[1234] Maryam Aghayerashti, Ehsan Bahrami Samani, and Ahmad Pour-Rashidi. Partially linear Bayesian modeling of longitudinal rank and time-to-event data using accelerated failure time model with application to brain tumor data. Statistics in Medicine, 42(14):2521–2556, June 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2023:MRL

[1235] Chyong-Mei Chen, Shuo-Chun Weng, Jia-Ren Tsai, and Pao sheng Shen. The mean residual life model for the right-censored data in the presence of covariate measurement errors. Statistics in Medicine, 42(15):2557– 2572, July 10, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2023:SSI

[1236] Tong Wang, Wenlu Tang, Yuanyuan Lin, and Wen Su. Semi-supervised inference for nonparametric logistic regression. Statistics in Medicine, 42(15):2573–2589, July 10, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Paukner:2023:DST

[1237] Mitchell Paukner and Richard Chappell. Designing superiority trials with window mean survival time as a primary endpoint. Statistics in Medicine, 42(15):2590–2599, July 10, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pal:2023:PEB

[1238] Suvra Pal and Souvik Roy. On the parameter estimation of Box–Cox transformation cure model. Statistics in Medicine, 42(15):2600–2618, July 10, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Huling:2023:DSS

[1239] Jared D. Huling, Jennifer P. Lundine, and Julie C. Leonard. Doubly structured sparsity for grouped multivariate responses with application to functional outcome score modeling. Statistics in Medicine, 42(15): 2619–2636, July 10, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2023:PSA

[1240] Yan Li and Liang Li. Propensity score analysis with local balance. Statistics in Medicine, 42(15):2637–2660, July 10, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bhattacharya:2023:NPB

[1241] Indrabati Bhattacharya, Brent A. Johnson, William J. Artman, Andrew Wilson, Kevin G. Lynch, James R. McKay, and Ashkan Ertefaie. A non-parametric Bayesian approach for adjusting partial compliance in sequential decision making. Statistics in Medicine, 42(15):2661–2691, July 10, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ozturk:2023:MCR

[1242] Omer Ozturk, Olena Kravchuk, and Richard Jarrett. Models for cluster randomized designs using ranked set sampling. Statistics in Medicine, 42(15):2692–2710, July 10, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gosho:2023:CBA

[1243] Masahiko Gosho, Ryota Ishii, Hisashi Noma, and Kazushi Maruo. A comparison of bias-adjusted generalized estimating equations for sparse binary data in small-sample longitudinal studies. Statistics in Medicine, 42(15):2711–2727, July 10, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Maidman:2023:QPL

[1244] Adam Maidman, Lan Wang, Xiao-Hua Zhou, and Ben Sherwood. Quantile partially linear additive model for data with dropouts and an application to modeling cognitive decline. Statistics in Medicine, 42(16): 2729–2745, July 20, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Johnsen:2023:SAS

[1245] Pål V. Johnsen, Øyvind Bakke, Thea Bjørnland, Andrew Thomas De-Wan, and Mette Langaas. Saddlepoint approximations to score test statistics in logistic regression for analyzing genome-wide association studies. Statistics in Medicine, 42(16):2746–2759, July 20, 2023. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

He:2023:RFT

[1246] Liang He, Ian Philipp, Stephanie Webster, Jacob v. B. Hjelmborg, and Alexander M. Kulminski. A robust and fast two-sample test of equal correlations with an application to differential co-expression. Statistics in Medicine, 42(16):2760–2776, July 20, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cohn:2023:SSC

[1247] Eric R. Cohn, Tianchen Qian, and Susan A. Murphy. Sample size considerations for micro-randomized trials with binary proximal outcomes. Statistics in Medicine, 42(16):2777–2796, July 20, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sugitani:2023:BBB

[1248] Yasuo Sugitani, Satoshi Morita, Akiyoshi Nakakura, and Hideharu Yamamoto. Biomarker-based Bayesian randomized clinical trial design for identifying a target population. Statistics in Medicine, 42(16):2797–2810, July 20, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Viele:2023:APT

[1249] Kert Viele. Allocation in platform trials to maintain comparability across time and eligibility. Statistics in Medicine, 42(16):2811–2818, July 20, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Whitehead:2023:BBB

[1250] Lou E. Whitehead, Oliver Sailer, Miles D. Witham, and James M. S. Wason. Bayesian borrowing for basket trials with longitudinal outcomes. Statistics in Medicine, 42(16):2819–2840, July 20, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Serra:2023:BMA

[1251] Alessandra Serra, Pavel Mozgunov, and Thomas Jaki. A Bayesian multiarm multi-stage clinical trial design incorporating information about treatment ordering. Statistics in Medicine, 42(16):2841–2854, July 20, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2023:CAA

[1252] Liang Li and Thomas Jemielita. Confounding adjustment in the analysis of augmented randomized controlled trial with hybrid control arm. Statistics in Medicine, 42(16):2855–2872, July 20, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thomadakis:2023:IEI

[1253] Christos Thomadakis, Nikos Pantazis, and Giota Touloumi. Issues with the expected information matrix of linear mixed models provided by popular statistical packages under missingness at random dropout. Statistics in Medicine, 42(16):2873–2885, July 20, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lang:2023:RRB

[1254] Yanda Lang, Joseph W. McKean, and Omer Ozturk. Robust rank-based meta-analyses for two-sample designs with application to platelet counts of malaria infection data. Statistics in Medicine, 42(17):2887–2913, July 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhou:2023:MJM

[1255] Grace Chen Zhou, Seongho Song, and Rhonda D. Szczesniak. Multilevel joint model of longitudinal continuous and binary outcomes for hierarchically structured data. Statistics in Medicine, 42(17):2914–2927, July 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ge:2023:TCR

[1256] Lin Ge, Yuzi Zhang, Kevin C. Ward, Timothy L. Lash, Lance A. Waller, and Robert H. Lyles. Tailoring capture-recapture methods to estimate registry-based case counts based on error-prone diagnostic signals. Statistics in Medicine, 42(17):2928–2943, July 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vutov:2023:MMS

[1257] Vladimir Vutov and Thorsten Dickhaus. Multiple multi-sample testing under arbitrary covariance dependency. Statistics in Medicine, 42(17): 2944–2961, July 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Devogel:2023:ADS

[1258] Nicholas Devogel, Paul L. Auer, Regina Manansala, and Tao Wang. On asymptotic distributions of several test statistics for familial relatedness in linear mixed models. Statistics in Medicine, 42(17):2962–2981, July 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tian:2023:MSI

[1259] Zibo Tian and Peihua Qiu. Multivariate single index modeling of longitudinal data with multiple responses. Statistics in Medicine, 42(17): 2982–2998, July 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fu:2023:BJM

[1260] Jingyan Fu, Matthew D. Koslovsky, Andreas M. Neophytou, and Marina Vannucci. A Bayesian joint model for compositional mediation effect selection in microbiome data. Statistics in Medicine, 42(17):2999–3015, July 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McGee:2023:IBK

[1261] Glen McGee, Ander Wilson, Brent A. Coull, and Thomas F. Webster. Incorporating biological knowledge in analyses of environmental mixtures and health. Statistics in Medicine, 42(17):3016–3031, July 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhou:2023:LSL

[1262] Yiwang Zhou and Peter X. K. Song. Longitudinal self-learning of individualized treatment rules in a nutrient supplementation trial with missing data. Statistics in Medicine, 42(17):3032–3049, July 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kanapka:2023:FOR

[1263] Lauren Kanapka and Anastasia Ivanova. Fully order restricted multi-arm multi-stage clinical trial design. Statistics in Medicine, 42(17):3050–3066, July 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Weisenthal:2023:RSM

[1264] Samuel J. Weisenthal, Sally W. Thurston, and Ashkan Ertefaie. Relative sparsity for medical decision problems. Statistics in Medicine, 42 (18):3067–3092, August 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bondi:2023:ABC

[1265] Laura Bondi, Marco Bonetti, Denitsa Grigorova, and Antonio Russo. Approximate Bayesian computation for the natural history of breast cancer, with application to data from a Milan cohort study. Statistics in Medicine, 42(18):3093–3113, August 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Begun:2023:DCM

[1266] Alexander Begun, Elena Kulinskaya, and Njabulo Ncube. A double-Cox model for non-proportional hazards survival analysis with frailty. Statistics in Medicine, 42(18):3114–3127, August 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Orindi:2023:CMF

[1267] Benedict Orindi, Adrian Quintero, Luk Bruyneel, Baoyue Li, and Emmanuel Lesaffre. A combined multilevel factor analysis and covariance regression model with mixed effects in the mean and variance structure. Statistics in Medicine, 42(18):3128–3144, August 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2023:SSL

[1268] Wei Liu, Huazhen Lin, Li Liu, Yanyuan Ma, Ying Wei, and Yi Li. Supervised structural learning of semiparametric regression on highdimensional correlated covariates with applications to eQTL studies. Statistics in Medicine, 42(18):3145–3163, August 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Poulet:2023:MDP

[1269] Pierre-Emmanuel Poulet and Stanley Durrleman. Multivariate disease progression modeling with longitudinal ordinal data. Statistics in Medicine, 42(18):3164–3183, August 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pate:2023:DPM

[1270] Alexander Pate, Matthew Sperrin, Richard D. Riley, Jamie C. Sergeant, Tjeerd Van Staa, Niels Peek, Mamas A. Mamas, Gregory Y. H. Lip, Martin O’Flaherty, Iain Buchan, and Glen P. Martin. Developing prediction models to estimate the risk of two survival outcomes both occurring: a comparison of techniques. Statistics in Medicine, 42(18):3184–3207, August 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2023:CIR

[1271] Wenling Zhang and Cecilia A. Cotton. Causal inference for recurrent events via aggregated marginal odds ratio. Statistics in Medicine, 42 (18):3208–3235, August 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zong:2023:EDP

[1272] Wei Zong, Marianne L. Seney, Kyle D. Ketchesin, Michael T. Gorczyca, Andrew C. Liu, Karyn A. Esser, George C. Tseng, Colleen A. McClung, and Zhiguang Huo. Experimental design and power calculation in omics circadian rhythmicity detection using the cosinor model. Statistics in Medicine, 42(18):3236–3258, August 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ounajim:2023:MLF

[1273] Amine Ounajim, Yousri Slaoui, Pierre-Yves Louis, Maxime Billot, Denis Frasca, and Philippe Rigoard. Mixture of longitudinal factor analyzers and their application to the assessment of chronic pain. Statistics in Medicine, 42(18):3259–3282, August 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Silva:2023:PTR

[1274] Ivair R. Silva and Joselito Montalban. The person-time ratio distribution for the exact monitoring of adverse events: Historical vs surveillance Poisson data. Statistics in Medicine, 42(18):3283–3301, August 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kim:2023:UCR

[1275] Sungduk Kim, Neil E. Caporaso, Fangyi Gu, Elizabeth B. Klerman, and Paul S. Albert. Uncovering circadian rhythms in metabolic longitudinal data: a Bayesian latent class modeling approach. Statistics in Medicine, 42(18):3302–3315, August 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2023:CCE

[1276] Anonymous. Correction to “Causal estimands and confidence intervals associated with Wilcoxon–Mann–Whitney tests in randomized experiments”. Statistics in Medicine, 42(18):3316, August 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [3].

Ma:2023:SPS

[1277] Wanying Ma, Mengya Liu, Jian Zhu, Qing Li, Elaine Hoffman, and Jianchang Lin. Semiparametric pseudo-score and pseudo-likelihood for evaluating correlate of protection in vaccine trials. Statistics in Medicine, 42 (19):3317–3332, August 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hansen:2023:EIF

[1278] Spencer Hansen and Kenneth Rice. Exact inference for fixed effects meta-analysis of 2 × 2 tables. Statistics in Medicine, 42(19):3333–3352, August 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kundu:2023:BHS

[1279] Debamita Kundu, Ritendranath Mitra, Paul S. Albert, and Jeremy T. Gaskins. A Bayesian hierarchical sparse factor model for estimating simultaneous covariance matrices for gestational outcomes in consecutive pregnancies. Statistics in Medicine, 42(19):3353–3370, August 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jansen:2023:MSN

[1280] Jeroen P. Jansen, Devin Incerti, and Thomas A. Trikalinos. Multi-state network meta-analysis of progression and survival data. Statistics in Medicine, 42(19):3371–3391, August 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tong:2023:SSC

[1281] Guangyu Tong, Monica Taljaard, and Fan Li. Sample size considerations for assessing treatment effect heterogeneity in randomized trials with heterogeneous intracluster correlations and variances. Statistics in Medicine, 42(19):3392–3412, August 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wan:2023:REM

[1282] Ke Wan, Kensuke Tanioka, and Toshio Shimokawa. Rule ensemble method with adaptive group lasso for heterogeneous treatment effect estimation. Statistics in Medicine, 42(19):3413–3442, August 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Benitez:2023:DEE

[1283] Alejandra Benitez, Maya L. Petersen, Mark J. van der Laan, Nicole Santos, Elizabeth Butrick, Dilys Walker, Rakesh Ghosh, Phelgona Otieno, Peter Waiswa, and Laura B. Balzer. Defining and estimating effects in cluster randomized trials: a methods comparison. Statistics in Medicine, 42(19):3443–3466, August 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ribaud:2023:IPS

[1284] Mélina Ribaud, Edith Gabriel, Joseph Hughes, and Samuel Soubeyrand. Identifying potential significant factors impacting zero-inflated proportion data. Statistics in Medicine, 42(19):3467–3486, August 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liang:2023:TDP

[1285] Menglu Liang, Zheng Li, Liang Li, Vernon M. Chinchilli, Lijun Zhang, and Ming Wang. Tackling dynamic prediction of death in patients with recurrent cardiovascular events. Statistics in Medicine, 42(19):3487– 3507, August 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

deJong:2023:PBS

[1286] Valentijn M. T. de Jong, Jeroen Hoogland, Karel G. M. Moons, Richard D. Riley, Tri-Long Nguyen, and Thomas P. A. Debray. Propensity-based standardization to enhance the validation and interpretation of prediction model discrimination for a target population. Statistics in Medicine, 42(19):3508–3528, August 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yelland:2023:HMS

[1287] Lisa N. Yelland, Jennie Louise, Brennan C. Kahan, Tim P. Morris, Katherine J. Lee, and Thomas R. Sullivan. Handling misclassified stratification variables in the analysis of randomised trials with continuous outcomes. Statistics in Medicine, 42(19):3529–3546, August 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Seng:2023:IVM

[1288] Loraine Liping Seng, Ching-Ti Liu, Jingli Wang, and Jialiang Li. Instrumental variable model average with applications in Mendelian randomization. Statistics in Medicine, 42(19):3547–3567, August 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Teerenstra:2023:SSP

[1289] Steven Teerenstra, Jessica Kasza, Ruslan Leontjevas, and Andrew B. Forbes. Sample size for partially nested designs and other nested or crossed designs with a continuous outcome when adjusted for baseline. Statistics in Medicine, 42(19):3568–3592, August 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Goyal:2023:ECN

[1290] Ravi Goyal, Nicole Carnegie, Sally Slipher, Philip Turk, Susan J. Little, and Victor De Gruttola. Estimating contact network properties by integrating multiple data sources associated with infectious diseases. Statistics in Medicine, 42(20):3593–3615, September 10, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2023:LRL

[1291] Yuzhe Zhang, Xu Zhang, Hong Zhang, Aiyi Liu, and Catherine C. Liu. Low-rank latent matrix-factor prediction modeling for generalized highdimensional matrix-variate regression. Statistics in Medicine, 42(20): 3616–3635, September 10, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lin:2023:DMS

[1292] Pei-Sheng Lin, Yih-Jeng Yu, and Jun Zhu. Disease mapping for spatially semi-continuous data by estimating equations with application to dengue control. Statistics in Medicine, 42(20):3636–3648, September 10, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hu:2023:SIT

[1293] Dingding Hu, Meng Yuan, Tao Yu, and Pengfei Li. Statistical inference for the two-sample problem under likelihood ratio ordering, with application to the ROC curve estimation. Statistics in Medicine, 42(20): 3649–3664, September 10, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yao:2023:DCF

[1294] Yuchen Yao, Dipnil Charkraborty, Lin Zhang, Xiaotong Shen, Alzheimer’s Disease Neuroimaging Initiative, and Wei Pan. Deep causal feature extraction and inference with neuroimaging genetic data. Statistics in Medicine, 42(20):3665–3684, September 10, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2023:HCC

[1295] Lili Liu, Kevin He, Di Wang, Shujie Ma, Annie Qu, Lu Lin, J. Philip Miller, and Lei Liu. Healthcare center clustering for Cox’s proportional hazards model by fusion penalty. Statistics in Medicine, 42(20):3685– 3698, September 10, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yu:2023:IPW

[1296] Youfei Yu, Min Zhang, and Bhramar Mukherjee. An inverse probability weighted regression method that accounts for right-censoring for causal inference with multiple treatments and a binary outcome. Statistics in Medicine, 42(20):3699–3715, September 10, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wei:2023:SAL

[1297] Kecheng Wei, Guoyou Qin, and Zhongyi Zhu. Subgroup analysis for longitudinal data based on a partial linear varying coefficient model with a change plane. Statistics in Medicine, 42(20):3716–3731, September 10, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

deUna-Alvarez:2023:TIS

[1298] Jacobo de Uña-Álvarez. Testing for an ignorable sampling bias under random double truncation. Statistics in Medicine, 42(20):3732–3744, September 10, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tawiah:2023:MJF

[1299] Richard Tawiah and Howard Bondell. Multilevel joint frailty model for hierarchically clustered binary and survival data. Statistics in Medicine, 42(21):3745–3763, September 20, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ryan:2023:MOC

[1300] Mary M. Ryan, Denise Esserman, and Fan Li. Maximin optimal cluster randomized designs for assessing treatment effect heterogeneity. Statistics in Medicine, 42(21):3764–3785, September 20, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Watson:2023:PBM

[1301] Samuel I. Watson, Joshua O. Akinyemi, and Karla Hemming. Permutation-based multiple testing corrections for P -values and confidence intervals for cluster randomized trials. Statistics in Medicine, 42 (21):3786–3803, September 20, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mews:2023:MMM

[1302] Sina Mews, Bastian Surmann, Lena Hasemann, and Svenja Elkenkamp. Markov-modulated marked Poisson processes for modeling disease dynamics based on medical claims data. Statistics in Medicine, 42(21): 3804–3815, September 20, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Strandberg:2023:NPI

[1303] Rickard Strandberg, Kamila Czene, Per Hall, and Keith Humphreys. Novel predictions of invasive breast cancer risk in mammography screening cohorts. Statistics in Medicine, 42(21):3816–3837, September 20, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Roy:2023:GST

[1304] Samrat Roy, Ting Ye, Ashkan Ertefaie, Tat-Thang Vo, James Flory, Sean Hennessy, and Dylan Small. Group sequential testing under instrumented difference-in-differences approach. Statistics in Medicine, 42(21):3838– 3859, September 20, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pabon-Rodriguez:2023:BML

[1305] Felix M. Pabon-Rodriguez, Grant D. Brown, Breanna M. Scorza, and Christine A. Petersen. Bayesian multivariate longitudinal model for immune responses to Leishmania: a tick-borne co-infection study. Statistics in Medicine, 42(21):3860–3876, September 20, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tanaka:2023:ESN

[1306] Shiro Tanaka, M. Alan Brookhart, and Jason Fine. G-estimation of structural nested mean models for interval-censored data using pseudoobservations. Statistics in Medicine, 42(21):3877–3891, September 20, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ocampo:2023:SWI

[1307] Alex Ocampo and Jemar R. Bather. Single-world intervention graphs for defining, identifying, and communicating estimands in clinical trials. Statistics in Medicine, 42(21):3892–3902, September 20, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Park:2023:VCR

[1308] Seyoung Park, Eun Ryung Lee, and Hyokyoung G. Hong. Varyingcoefficients for regional quantile via KNN-based LASSO with applications to health outcome study. Statistics in Medicine, 42(22):3903–3918, September 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chang:2023:CAR

[1309] Chia-Rui Chang, Yue Song, Fan Li, and Rui Wang. Covariate adjustment in randomized clinical trials with missing covariate and outcome data. Statistics in Medicine, 42(22):3919–3935, September 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Arsham:2023:CEA

[1310] Aryana Arsham, Ionut Bebu, and Thomas Mathew. Cost-effectiveness analysis under multiple effectiveness outcomes: a probabilistic approach. Statistics in Medicine, 42(22):3936–3955, September 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Egbon:2023:PEG

[1311] Osafu Augustine Egbon, Diego Nascimento, and Francisco Louzada. Prior elicitation for Gaussian spatial process: an application to TMS brain mapping. Statistics in Medicine, 42(22):3956–3980, September 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chipman:2023:RFS

[1312] Jonathan J. Chipman, Lindsay Mayberry, and Robert A. Greevy, Jr. Rematching on-the-fly: Sequential matched randomization and a case for covariate-adjusted randomization. Statistics in Medicine, 42(22):3981– 3995, September 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Proust-Lima:2023:DCD

[1313] Cécile Proust-Lima, Tiphaine Saulnier, Viviane Philipps, Anne Pavy-Le Traon, Patrice Péran, Olivier Rascol, Wassilios G. Meissner, and Alexandra Foubert-Samier. Describing complex disease progression using joint latent class models for multivariate longitudinal markers and clinical endpoints. Statistics in Medicine, 42(22):3996–4014, September 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wu:2023:STC

[1314] Mixia Wu, Xian Sun, Aiyi Liu, Chenchen Peng, and Zhaohai Li. Significance tests for covariates in the diagnostic accuracy index of a biomarker against a continuous gold standard. Statistics in Medicine, 42(22):4015– 4027, September 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Niu:2023:ECM

[1315] Fang Niu, Cheng Zheng, and Lei Liu. Exploring causal mechanisms and quantifying direct and indirect effects using a joint modeling approach for recurrent and terminal events. Statistics in Medicine, 42(22):4028– 4042, September 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Song:2023:CSS

[1316] Xiao Song. A corrected smoothed score approach for semiparametric accelerated failure time model with error-contaminated covariates. Statistics in Medicine, 42(22):4043–4055, September 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schneider:2023:CCS

[1317] Silvana Schneider, Rodrigo Citton P. dos Reis, Maicon M. F. Gottselig, Patrícia Fisch, Daniela Riva Knauth, and Álvaro Vigo. Clayton copula for survival data with dependent censoring: an application to a tuberculosis treatment adherence data. Statistics in Medicine, 42(23):4057–4081, October 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dey:2023:ICA

[1318] Rajib Dey, J. A. Hanley, and P. Saha-Chaudhuri. Inference for covariateadjusted time-dependent prognostic accuracy measures. Statistics in Medicine, 42(23):4082–4110, October 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Aselisewine:2023:IDT

[1319] Wisdom Aselisewine and Suvra Pal. On the integration of decision trees with mixture cure model. Statistics in Medicine, 42(23):4111–4127, October 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Konstantinou:2023:SMD

[1320] Konstantinos Konstantinou, Farnaz Ghorbanpour, Umberto Picchini, Adam Loavenbruck, and Aila Särkkä. Statistical modeling of diabetic neuropathy: Exploring the dynamics of nerve mortality. Statistics in Medicine, 42(23):4128–4146, October 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Salditt:2023:PNP

[1321] Marie Salditt and Steffen Nestler. Parametric and nonparametric propensity score estimation in multilevel observational studies. Statistics in Medicine, 42(23):4147–4176, October 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Takahashi:2023:HTP

[1322] Kanae Takahashi, Kouji Yamamoto, Aya Kuchiba, Ayumi Shintani, and Tatsuki Koyama. Hypothesis testing procedure for binary and multi-class F_{1} -scores in the paired design. Statistics in Medicine, 42(23):4177–4192, October 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Turchetta:2023:TDP

[1323] Armando Turchetta, Nicolas Savy, David A. Stephens, Erica E. M. Moodie, and Marina B. Klein. A time-dependent Poisson–Gamma model for recruitment forecasting in multicenter studies. Statistics in Medicine, 42(23):4193–4206, October 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brathovde:2023:LAF

[1324] Mari Brathovde, Tron A. Moger, Odd O. Aalen, Tom Grotmol, Marit B. Veierød, and Morten Valberg. A lean additive frailty model: With an application to clustering of melanoma in Norwegian families. Statistics in Medicine, 42(23):4207–4235, October 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shen:2023:EEI

[1325] Jenny Shen, Rebecca A. Hubbard, and Kristin A. Linn. Estimation and evaluation of individualized treatment rules following multiple imputation. Statistics in Medicine, 42(23):4236–4256, October 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vancak:2023:SAE

[1326] Valentin Vancak and Arvid Sjölander. Sensitivity analysis of Gestimators to invalid instrumental variables. Statistics in Medicine, 42 (23):4257–4281, October 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ross:2023:ANM

[1327] Rachael K. Ross, Stephen R. Cole, Jessie K. Edwards, Daniel Westreich, Julie L. Daniels, and Jeffrey S. A. Stringer. Accounting for nonmonotone missing data using inverse probability weighting. Statistics in Medicine, 42(23):4282–4298, October 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2023:CET

[1328] Anonymous. Correction to “estimation of treatment effect in 2-in-1 adaptive design and some of its extensions”. Statistics in Medicine, 42(23): 4299, October 15, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yu:2023:USR

[1329] Jih-Chang Yu and Yen-Tsung Huang. Unified semicompeting risks analysis of hepatitis natural history through mediation modeling. Statistics in Medicine, 42(24):4301–4318, October 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2023:ECP

[1330] Xin Li, Godwin Yung, Jianchang Lin, and Jian Zhu. Estimation of conditional power in the presence of auxiliary data. Statistics in Medicine, 42 (24):4319–4332, October 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tu:2023:RIC

[1331] Shengxin Tu, Chun Li, Donglin Zeng, and Bryan E. Shepherd. Rank intraclass correlation for clustered data. Statistics in Medicine, 42(24):4333– 4348, October 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gao:2023:IVA

[1332] Dexiang Gao, Junxiao Hu, Cathy J. Bradley, and Fan Yang. Instrumental variable analysis for cost outcome: Application to the effect of primary care visit on medical cost among low-income adults. Statistics in Medicine, 42(24):4349–4376, October 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kaplan:2023:SBM

[1333] Adam Kaplan and David Nelson. Simple Bayesian models for missing binary outcomes in randomized controlled trials. Statistics in Medicine, 42(24):4377–4391, October 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Daniells:2023:CBI

[1334] Libby Daniells, Pavel Mozgunov, Alun Bedding, and Thomas Jaki. A comparison of Bayesian information borrowing methods in basket trials and a novel proposal of modified exchangeability-nonexchangeability method. Statistics in Medicine, 42(24):4392–4417, October 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xiong:2023:FCI

[1335] Ruoxuan Xiong, Allison Koenecke, Michael Powell, Zhu Shen, Joshua T. Vogelstein, and Susan Athey. Federated causal inference in heterogeneous observational data. Statistics in Medicine, 42(24):4418–4439, October 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fang:2023:SPR

[1336] Lijun Fang, Shuwei Li, Liuquan Sun, and Xinyuan Song. Semiparametric probit regression model with misclassified current status data. Statistics in Medicine, 42(24):4440–4457, October 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Trinh:2023:FED

[1337] Kelly Trinh, Andrew Staib, and Anton Pak. Forecasting emergency department waiting time using a state space representation. Statistics in Medicine, 42(24):4458–4483, October 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cattaneo:2023:GRD

[1338] Matias D. Cattaneo, Luke Keele, and Rocío Titiunik. A guide to regression discontinuity designs in medical applications. Statistics in Medicine, 42(24):4484–4513, October 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

vanKlaveren:2023:MPP

[1339] David van Klaveren, Carolien C. H. M. Maas, and David M. Kent. Measuring the performance of prediction models to personalize treatment choice: Defining observed and predicted pairwise treatment effects. Statistics in Medicine, 42(24):4514–4515, October 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wilson:2023:BDA

[1340] Kevin J. Wilson. Bayesian design and analysis of two-arm cluster randomized trials using assurance. Statistics in Medicine, 42(25):4517–4531, November 10, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sun:2023:CIC

[1341] Min Woo Sun and Robert Tibshirani. Confidence intervals for the Cox model test error from cross-validation. Statistics in Medicine, 42(25): 4532–4541, November 10, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yoneoka:2023:ETE

[1342] Daisuke Yoneoka, Takayuki Kawashima, Yuta Tanoue, Shuhei Nomura, and Akifumi Eguchi. Estimation of the time of exposure based on interval and censored data using the ϵ-accelerated EM algorithm. Statistics in Medicine, 42(25):4542–4555, November 10, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hazelton:2023:SES

[1343] Martin L. Hazelton. Shrinkage estimators of the spatial relative risk function. Statistics in Medicine, 42(25):4556–4569, November 10, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dugourd:2023:CTM

[1344] Charlotte Dugourd, Amna Abichou-Klich, René Ecochard, and Fabien Subtil. Clustering of trajectories with mixed effects classification model: Inference taking into account classification uncertainties. Statistics in Medicine, 42(25):4570–4581, November 10, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2023:SSA

[1345] Yu Wang, Sharon D. Yeatts, Renee’ H. Martin, Robert Silbergleit, Gaylan L. Rockswold, William G. Barsan, Frederick K. Korley, Sarah Rockswold, and Byron J. Gajewski. Selection of a statistical analysis method for the Glasgow Outcome Scale-Extended endpoint for estimating the probability of favorable outcome in future severe TBI clinical trials. Statistics in Medicine, 42(25):4582–4601, November 10, 2023. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Guolo:2023:ALP

[1346] Annamaria Guolo. Approximate likelihood and pseudo-likelihood inference in meta-analysis of diagnostic accuracy studies accounting for disease prevalence and study design. Statistics in Medicine, 42(25):4602– 4617, November 10, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2023:TLT

[1347] Zhanfeng Wang, Tao Li, Liqun Xiao, and Dongsheng Tu. A threshold longitudinal tobit quantile regression model for identification of treatmentsensitive subgroups based on interval-bounded longitudinal measurements and a continuous covariate. Statistics in Medicine, 42(25):4618– 4631, November 10, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shi:2023:FQL

[1348] Yiming Shi, Huilin Li, Chan Wang, Jun Chen, Hongmei Jiang, Ya-Chen T. Shih, Haixiang Zhang, Yizhe Song, Yang Feng, and Lei Liu. A flexible quasi-likelihood model for microbiome abundance count data. Statistics in Medicine, 42(25):4632–4643, November 10, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2023:CPD

[1349] Jinjuan Wang, Na Li, Zhen Meng, and Qizhai Li. Change point detection for high dimensional data via kernel measure with application to human aging brain data. Statistics in Medicine, 42(25):4644–4663, November 10, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ghosal:2023:DSD

[1350] Nairita Ghosal, Sanjb Basu, and Dulal Bhaumik. Detection of sparse differential dependent functional brain connectivity. Statistics in Medicine, 42(25):4664–4680, November 10, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2023:SAU

[1351] Wei Li, Shanshan Luo, Yangbo He, and Zhi Geng. Subgroup analysis using Bernoulli-gated hierarchical mixtures of experts models. Statistics in Medicine, 42(26):4681–4695, November 20, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wattanachit:2023:CCM

[1352] Nutcha Wattanachit, Evan L. Ray, Thomas C. McAndrew, and Nicholas G. Reich. Comparison of combination methods to create calibrated ensemble forecasts for seasonal influenza in the U.S. Statistics in Medicine, 42(26):4696–4712, November 20, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhou:2023:CPE

[1353] Lili Zhou, Daniel Andrés Díaz-Pachón, Chen Zhao, J. Sunil Rao, and Ola Hössjer. Correcting prevalence estimation for biased sampling with testing errors. Statistics in Medicine, 42(26):4713–4737, November 20, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Collier:2023:HMW

[1354] Willem Collier, Benjamin Haaland, Lesley Inker, and Tom Greene. Handling missing within-study correlations in the evaluation of surrogate endpoints. Statistics in Medicine, 42(26):4738–4762, November 20, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhong:2023:ASF

[1355] Yujie Zhong, Richard J. Cook, and Aiai Yu. Analysis of secondary failure time responses in studies with response-dependent sampling schemes. Statistics in Medicine, 42(26):4763–4775, November 20, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kim:2023:LFA

[1356] Sungduk Kim, Laura E. Beane Freeman, and Paul S. Albert. A latent functional approach for modeling the effects of multidimensional exposures on disease risk. Statistics in Medicine, 42(26):4776–4793, November 20, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ma:2023:BMS

[1357] Shaopei Ma, Keming Yu, Man lai Tang, Jianxin Pan, Wolfgang Karl Härdle, and Maozai Tian. A Bayesian multistage spatio-temporally dependent model for spatial clustering and variable selection. Statistics in Medicine, 42(26):4794–4823, November 20, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hattori:2023:FAB

[1358] Satoshi Hattori and Satoshi Morita. Frequentist analysis of basket trials with one-sample Mantel–Haenszel procedures. Statistics in Medicine, 42 (26):4824–4849, November 20, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Seitidis:2023:IIN

[1359] Georgios Seitidis, Stavros Nikolakopoulos, Ioannis Ntzoufras, and Dimitris Mavridis. Inconsistency identification in network meta-analysis via stochastic search variable selection. Statistics in Medicine, 42(26):4850– 4866, November 20, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Majumdar:2023:BME

[1360] Arunabha Majumdar and Bogdan Pasaniuc. A Bayesian method for estimating gene-level polygenicity under the framework of transcriptomewide association study. Statistics in Medicine, 42(26):4867–4885, November 20, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Guan:2023:MAL

[1361] Zhong Guan. Maximum approximate likelihood estimation in accelerated failure time model for interval-censored data. Statistics in Medicine, 42 (26):4886–4896, November 20, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vera:2023:TSL

[1362] J. Fernando Vera and José A. Roldán-Nofuentes. A two-step log-linear procedure for graphical representation and inference of associations in cross-classified data for disease diagnosis. Statistics in Medicine, 42(27): 4897–4916, November 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Turner:2023:NAE

[1363] Rebecca M. Turner, Tim Band, Tim P. Morris, David J. Fisher, Julian P. T. Higgins, James R. Carpenter, and Ian R. White. A new approach to evaluating loop inconsistency in network meta-analysis. Statistics in Medicine, 42(27):4917–4930, November 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kamat:2023:BRL

[1364] Gauri Kamat, Mingyang Shan, and Roee Gutman. Bayesian record linkage with variables in one file. Statistics in Medicine, 42(27):4931–4951, November 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Marquez:2023:CLP

[1365] Maritza Márquez, Cristian Meza, Dae-Jin Lee, and Rolando De la Cruz. Classification of longitudinal profiles using semi-parametric nonlinear mixed models with P-Splines and the SAEM algorithm. Statistics in Medicine, 42(27):4952–4971, November 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2023:SMJ

[1366] Jiaqing Chen, Yangxin Huang, and Qing Wang. Semiparametric multivariate joint model for skewed-longitudinal and survival data: a Bayesian approach. Statistics in Medicine, 42(27):4972–4989, November 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yao:2023:BPM

[1367] Zixuan Yao, Satoshi Morita, Sumiyuki Nishida, and Haruo Sugiyama. Bayesian predictive model averaging approach to joint longitudinalsurvival modeling: Application to an immuno-oncology clinical trial. Statistics in Medicine, 42(27):4990–5006, November 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Booth:2023:UTR

[1368] Sarah Booth, Sarwar I. Mozumder, Lucinda Archer, Joie Ensor, Richard D. Riley, Paul C. Lambert, and Mark J. Rutherford. Using temporal recalibration to improve the calibration of risk prediction models in competing risk settings when there are trends in survival over time. Statistics in Medicine, 42(27):5007–5024, November 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Segura-Buisan:2023:AMD

[1369] Juan Segura-Buisan, Clemence Leyrat, and Manuel Gomes. Addressing missing data in the estimation of time-varying treatments in comparative effectiveness research. Statistics in Medicine, 42(27):5025–5038, November 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Luo:2023:ITS

[1370] Yuanhui Luo and Xinzhou Guo. Inference on tree-structured subgroups with subgroup size and subgroup effect relationship in clinical trials. Statistics in Medicine, 42(27):5039–5053, November 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Maleyeff:2023:SSR

[1371] Lara Maleyeff, Rui Wang, Sebastien Haneuse, and Fan Li. Sample size requirements for testing treatment effect heterogeneity in cluster randomized trials with binary outcomes. Statistics in Medicine, 42(27): 5054–5083, November 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2023:CSE

[1372] Anonymous. Correction to “Summarizing empirical information on between-study heterogeneity for Bayesian random-effects meta-analysis”. Statistics in Medicine, 42(27):5084, November 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1230].

Wang:2023:BHM

[1373] Zheng Wang, Thomas A. Murray, Mengli Xiao, Lifeng Lin, Demissie Alemayehu, and Haitao Chu. Bayesian hierarchical models incorporating study-level covariates for multivariate meta-analysis of diagnostic tests without a gold standard with application to COVID-19. Statistics in Medicine, 42(28):5085–5099, December 10, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Siddique:2023:JMF

[1374] Juned Siddique, Michael J. Daniels, Gül Inan, Samuel Battalio, Bonnie Spring, and Donald Hedeker. Joint modeling the frequency and duration of accelerometer-measured physical activity from a lifestyle intervention trial. Statistics in Medicine, 42(28):5100–5112, December 10, 2023. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Treszoks:2023:EIC

[1375] Jodi Treszoks and Suvra Pal. On the estimation of interval censored destructive negative binomial cure model. Statistics in Medicine, 42(28): 5113–5134, December 10, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Samawi:2023:PTD

[1376] Hani Samawi, Marwan Alsharman, Mario Keko, and Jing Kersey. Posttest diagnostic accuracy measures under tree ordering of disease classes. Statistics in Medicine, 42(28):5135–5159, December 10, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Teunis:2023:ESR

[1377] Peter F. M. Teunis, Yuke Wang, Kristen Aiemjoy, Mirjam Kretzschmar, and Marc Aerts. Estimating seroconversion rates accounting for repeated infections by approximate Bayesian computation. Statistics in Medicine, 42(28):5160–5188, December 10, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Awasthi:2023:FSM

[1378] Achal Awasthi, Volodymyr M. Minin, Jenny Huang, Daniel Chow, and Jason Xu. Fitting a stochastic model of intensive care occupancy to noisy hospitalization time series during the COVID-19 pandemic. Statistics in Medicine, 42(28):5189–5206, December 10, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nan:2023:NAM

[1379] Nan Nan and Lili Tian. A new accuracy metric under three classes when subclasses are involved and its confidence interval estimation. Statistics in Medicine, 42(28):5207–5228, December 10, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2023:SSO

[1380] Fengqing Zhang and Jiangtao Gou. Sample size optimization for clinical trials using graphical approaches for multiplicity adjustment. Statistics in Medicine, 42(28):5229–5246, December 10, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kim:2023:JSK

[1381] Byung-Jun Kim and Inyoung Kim. Joint semiparametric kernel network regression. Statistics in Medicine, 42(28):5247–5265, December 10, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sun:2023:BFP

[1382] Zequn Sun, Dongjun Chung, Brian Neelon, Andrew Millar-Wilson, Stephen P. Ethier, Feifei Xiao, Yinan Zheng, Kristin Wallace, and Gary Hardiman. A Bayesian framework for pathway-guided identification of cancer subgroups by integrating multiple types of genomic data.

Statistics in Medicine, 42(28):5266–5284, December 10, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Coart:2023:MRC

[1383] Elisabeth Coart, Perrine Bamps, Emmanuel Quinaux, Geneviève Sturbois, Everardo D. Saad, Tomasz Burzykowski, and Marc Buyse. Minimization in randomized clinical trials. Statistics in Medicine, 42(28): 5285–5311, December 10, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lindner:2023:DLR

[1384] Hanna Lindner, Phyllis A. Gimotty, and Warren B. Bilker. The diagnostic likelihood ratio function and modified test for trend: Identifying, evaluating, and validating nontraditional biomarkers in case-control studies. Statistics in Medicine, 42(29):5313–5337, December 20, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jin:2023:BAD

[1385] Huaqing Jin, Mi-Ok Kim, Aaron Scheffler, and Fei Jiang. Bayesian adaptive design for covariate-adaptive historical control information borrowing. Statistics in Medicine, 42(29):5338–5352, December 20, 2023. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

You:2023:BDE

[1386] Na You, Xueyi He, Hongsheng Dai, and Xueqin Wang. Ball divergence for the equality test of crossing survival curves. Statistics in Medicine, 42 (29):5353–5368, December 20, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mei:2023:NCA

[1387] Hao Mei, Jiaxin Xie, Yichen Qin, and Yang Li. Network and covariate adjusted response-adaptive design for binary response. Statistics in Medicine, 42(29):5369–5388, December 20, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bai:2023:SER

[1388] Fangfang Bai and Xiaoran Yang. Semiparametric estimation of restricted mean survival time as a function of restriction time. Statistics in Medicine, 42(29):5389–5404, December 20, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mamouris:2023:LTI

[1389] Pavlos Mamouris, Vahid Nassiri, Geert Verbeke, Arne Janssens, Bert Vaes, and Geert Molenberghs. A longitudinal transition imputation model for categorical data applied to a large registry dataset. Statistics in Medicine, 42(29):5405–5418, December 20, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zamanian:2023:AIS

[1390] Alireza Zamanian, Narges Ahmidi, and Mathias Drton. Assessable and interpretable sensitivity analysis in the pattern graph framework for nonignorable missingness mechanisms. Statistics in Medicine, 42(29):5419– 5450, December 20, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ojeda:2023:CML

[1391] Francisco M. Ojeda, Max L. Jansen, Alexandre Thiéry, Stefan Blankenberg, Christian Weimar, Matthias Schmid, and Andreas Ziegler. Calibrating machine learning approaches for probability estimation: a comprehensive comparison. Statistics in Medicine, 42(29):5451–5478, December 20, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xiang:2023:SIM

[1392] Qingyan Xiang, Ronald J. Bosch, and Judith J. Lok. The survivalincorporated median vs the median in the survivors or in the alwayssurvivors: What are we measuring? and why? Statistics in Medicine, 42 (29):5479–5490, December 20, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tang:2023:SNT

[1393] An-Ming Tang, Cheng Peng, and Niansheng Tang. Semiparametric normal transformation joint model of multivariate longitudinal and bivariate time-to-event data. Statistics in Medicine, 42(29):5491–5512, December 20, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lu:2023:JCM

[1394] Zihang Lu, Mojtaba Ahmadiankalati, and Zhiwen Tan. Joint clustering multiple longitudinal features: a comparison of methods and software packages with practical guidance. Statistics in Medicine, 42(29):5513– 5540, December 20, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Neijzen:2023:ULM

[1395] Dorien Neijzen and Gerton Lunter. Unsupervised learning for medical data: a review of probabilistic factorization methods. Statistics in Medicine, 42(30):5541–5554, December 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Park:2023:STD

[1396] Jaewoo Park, Seorim Yi, Won Chang, and Jorge Mateu. A spatiotemporal Dirichlet process mixture model for coronavirus disease-19. Statistics in Medicine, 42(30):5555–5576, December 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ma:2023:SAF

[1397] Ding Ma, Jun Ma, and Petra L. Graham. On semiparametric accelerated failure time models with time-varying covariates: a maximum penalised likelihood estimation. Statistics in Medicine, 42(30):5577–5595, December 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2023:SET

[1398] Li Liu, Wen Su, and Xingqiu Zhao. Semiparametric estimation and testing for panel count data with informative interval-censored failure event. Statistics in Medicine, 42(30):5596–5615, December 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2023:GWS

[1399] Wenda Zhang, Zichen Ma, Lianming Wang, Daping Fan, and Yen-Yi Ho. Genome-wide search algorithms for identifying dynamic gene coexpression via Bayesian variable selection. Statistics in Medicine, 42 (30):5616–5629, December 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Huang:2023:SSI

[1400] Yunxiang Huang, Chiung-Yu Huang, and Mi-Ok Kim. Simultaneous selection and incorporation of consistent external aggregate information. Statistics in Medicine, 42(30):5630–5645, December 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ngan:2023:ENM

[1401] Esther Ngan, Wenyaw Chan, Luis Leon-Novelo, and Valory Pavlik. Estimation of non-monotonic transition rates in a semi-Markov process with covariates adjustments and application to caregivers’ stress data.

Statistics in Medicine, 42(30):5646–5656, December 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Infante:2023:SSP

[1402] Gabriele Infante, Rosalba Miceli, and Federico Ambrogi. Sample size and predictive performance of machine learning methods with survival data: a simulation study. Statistics in Medicine, 42(30):5657–5675, December 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cai:2023:RSN

[1403] Yun Cai, Hong Gu, and Toby Kenney. Rank selection for non-negative matrix factorization. Statistics in Medicine, 42(30):5676–5693, December 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

PhD:2023:DPI

[1404] Virginia B. Shipes PhD, Caitlyn Meinzer PhD, Bethany J. Wolf PhD, Hong Li PhD, Mathew J. Carpenter PhD, Hooman Kamel Md, and Renee H. Martin PhD. Designing a phase-III time-to-event clinical trial using a modified sample size formula and Poisson-gamma model for subject accrual that accounts for the lag in site initiation using the PERT distribution. Statistics in Medicine, 42(30):5694–5707, December 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bi:2023:BBH

[1405] Dehua Bi, Meizi Liu, Jianchang Lin, and Rachael Liu. BEATS: Bayesian hybrid design with flexible sample size adaptation for time-to-event endpoints. Statistics in Medicine, 42(30):5708–5722, December 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Parner:2023:WLP

[1406] Erik T. Parner and Morten Overgaard. Win-loss parameters for rightcensored event data, with application to recurrent events. Statistics in Medicine, 42(30):5723–5735, December 30, 2023. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rong:2024:KCP

[1407] Yaohua Rong, Sihai Dave Zhao, Xia Zheng, and Yi Li. Kernel Cox partially linear regression: Building predictive models for cancer patients’ survival. Statistics in Medicine, 43(1):1–15, January 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2024:MMA

[1408] Wei Wang, Guangyu Tong, Shashivadan P. Hirani, Stanton P. Newman, Scott D. Halpern, Dylan S. Small, Fan Li, and Michael O. Harhay. A mixed model approach to estimate the survivor average causal effect in cluster-randomized trials. Statistics in Medicine, 43(1):16–33, January 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sun:2024:PSQ

[1409] Shuo Sun, Johanna G. Neslehová, and Erica E. M. Moodie. Principal stratification for quantile causal effects under partial compliance. Statistics in Medicine, 43(1):34–48, January 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2024:CRT

[1410] Kenneth Menglin Lee and Yin Bun Cheung. Cluster randomized trial designs for modeling time-varying intervention effects. Statistics in Medicine, 43(1):49–60, January 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sun:2024:LFD

[1411] Xiaoya Sun and Yan Fu. Local false discovery rate estimation with competition-based procedures for variable selection. Statistics in Medicine, 43(1):61–88, January 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nestler:2024:LVM

[1412] Steffen Nestler and Shelley A. Blozis. A latent variable mixed-effects location scale model that also considers between-person differences in the autocorrelation. Statistics in Medicine, 43(1):89–101, January 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Qiao:2024:TLC

[1413] Xinhui Qiao, Hua He, Liuquan Sun, Shuo Bai, and Peng Ye. Testing latent classes in gut microbiome data using generalized Poisson regression models. Statistics in Medicine, 43(1):102–124, January 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wen:2024:BZI

[1414] Chun-Che Wen, Nathaniel Baker, Rajib Paul, Elizabeth Hill, Kelly Hunt, Hong Li, Kevin Gray, and Brian Neelon. A Bayesian zero-inflated beta-binomial model for longitudinal data with group-specific changepoints. Statistics in Medicine, 43(1):125–140, January 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2024:BCG

[1415] Li Zhang, Xinyan Zhang, and Nengjun Yi. Bayesian compositional generalized linear models for analyzing microbiome data. Statistics in Medicine, 43(1):141–155, January 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kanapka:2024:FDB

[1416] Lauren Kanapka and Anastasia Ivanova. A frequentist design for basket trials using adaptive lasso. Statistics in Medicine, 43(1):156–172, January 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2024:BPI

[1417] Wen Zhang, Glen Laird, Josh Chen, and Ying Yuan. A Bayesian phase II proof-of-concept design for clinical trials with longitudinal endpoints. Statistics in Medicine, 43(1):173–183, January 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Broomfield:2024:MMS

[1418] Jonathan Broomfield, Keith R. Abrams, Suzanne Freeman, Nicholas Latimer, Mark J. Rutherford, Michael J. Crowther, Project Hercules, the Cooperative International Neuromuscular Research Group investigators, and Duchenne Regulatory Science Consortium members. Modeling the multi-state natural history of rare diseases with heterogeneous individual patient data: a simulation study. Statistics in Medicine, 43(1):184–200, January 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Qiu:2024:ETC

[1419] Hongxiang Qiu, Andrea J. Cook, and Jennifer F. Bobb. Evaluating tests for cluster-randomized trials with few clusters under generalized linear mixed models with covariate adjustment: a simulation study. Statistics in Medicine, 43(2):201–215, January 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2024:SIT

[1420] Tuo Wang, Lu Mao, Aldo Cocco, and KyungMann Kim. Statistical inference for time-to-event data in non-randomized cohorts with selective attrition. Statistics in Medicine, 43(2):216–232, January 30, 2024. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ganguly:2024:FMB

[1421] Ayon Ganguly, Debanjan Mitra, Narayanaswamy Balakrishnan, and Debasis Kundu. A flexible model based on piecewise linear approximation for the analysis of left truncated right censored data with covariates, and applications to Worcester Heart Attack Study data and Channing House data. Statistics in Medicine, 43(2):233–255, January 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2024:EMR

[1422] Katherine Li and Yajuan Si. Embedded multilevel regression and poststratification: Model-based inference with incomplete auxiliary information. Statistics in Medicine, 43(2):256–278, January 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2024:MCM

[1423] Yunxiao Li, Yi-Juan Hu, and Glen A. Satten. MERIT: Controlling Monte-Carlo error rate in large-scale Monte-Carlo hypothesis testing. Statistics in Medicine, 43(2):279–295, January 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vo:2024:CRL

[1424] Thanh Huan Vo, Valérie Garès, Li-Chun Zhang, André Happe, Emmanuel Oger, Stéphane Paquelet, and Guillaume Chauvet. Cox regression with linked data. Statistics in Medicine, 43(2):296–314, January 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2024:GST

[1425] Ruyi Liu, Fan Li, Denise Esserman, and Mary M. Ryan. Group sequential two-stage preference designs. Statistics in Medicine, 43(2):315–341, January 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Siddique:2024:TST

[1426] Arman Alam Siddique, Mireille E. Schnitzer, Narayanaswamy Balakrishnan, Giovanni Sotgiu, Mario H. Vargas, Dick Menzies, and Andrea Benedetti. Two-stage targeted maximum likelihood estimation for mixed aggregate and individual participant data analysis with an application to multidrug resistant tuberculosis. Statistics in Medicine, 43(2):342– 357, January 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2024:DIR

[1427] Xueqi Wang, Elizabeth L. Turner, and Fan Li. Designing individually randomized group treatment trials with repeated outcome measurements using generalized estimating equations. Statistics in Medicine, 43(2): 358–378, January 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Amorim:2024:TPG

[1428] Gustavo Amorim, Ran Tao, Sarah Lotspeich, Pamela A. Shaw, Thomas Lumley, Rena C. Patel, and Bryan E. Shepherd. Three-phase generalized raking and multiple imputation estimators to address error-prone data. Statistics in Medicine, 43(2):379–394, January 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bu:2024:BSS

[1429] Fan Bu, Martijn J. Schuemie, Akihiko Nishimura, Louisa H. Smith, Kristin Kostka, Thomas Falconer, Jody-Ann McLeggon, Patrick B. Ryan, George Hripcsak, and Marc A. Suchard. Bayesian safety surveillance with adaptive bias correction. Statistics in Medicine, 43(2):395– 418, January 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2024:REM

[1430] Mushan Li and Yanyuan Ma. Robust estimation of mean-variance relation. Statistics in Medicine, 43(2):419–434, January 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Davila-Payan:2024:AYT

[1431] C. S. Davila-Payan, A. Hill, L. Kayembe, J. P. Alexander, M. Lynch, and S. W. Pallas. Analysis of the yearly transition function in measles disease modeling. Statistics in Medicine, 43(3):435–451, February 10, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schochet:2024:EAT

[1432] Peter Z. Schochet. Estimating average treatment effects for clustered RCTs with recruitment bias. Statistics in Medicine, 43(3):452–474, February 10, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xi:2024:OWB

[1433] Dong Xi and Yao Chen. Optimal weighted Bonferroni tests and their graphical extensions. Statistics in Medicine, 43(3):475–500, February 10, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dufault:2024:FMM

[1434] Suzanne M. Dufault, Angela M. Crook, Katie Rolfe, and Patrick P. J. Phillips. A flexible multi-metric Bayesian framework for decision-making in phase II multi-arm multi-stage studies. Statistics in Medicine, 43(3): 501–513, February 10, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Munoz:2024:MII

[1435] Johanna Muñoz, Orestis Efthimiou, Vincent Audigier, Valentijn M. T. de Jong, and Thomas P. A. Debray. Multiple imputation of incomplete multilevel data using Heckman selection models. Statistics in Medicine, 43(3):514–533, February 10, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gabriel:2024:IPT

[1436] Erin E. Gabriel, Michael C. Sachs, Torben Martinussen, Ingeborg Waernbaum, Els Goetghebeur, Stijn Vansteelandt, and Arvid Sjölander. Inverse probability of treatment weighting with generalized linear outcome models for doubly robust estimation. Statistics in Medicine, 43(3):534–547, February 10, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2024:BHM

[1437] Hao Li, Dooti Roy, and Qiqi Deng. Bayesian hierarchical modeling in interim futility analysis for two parallel clinical trials. Statistics in Medicine, 43(3):548–559, February 10, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lu:2024:BIP

[1438] Yuelin Lu, Bradley P. Carlin, and John W. Seaman, Jr. Bayesian inference for prediction of survival probability in prime-boost vaccination regimes. Statistics in Medicine, 43(3):560–577, February 10, 2024. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Charvadeh:2024:AME

[1439] Yasin Khadem Charvadeh and Grace Y. Yi. Accommodating misclassification effects on optimizing dynamic treatment regimes with Q-learning. Statistics in Medicine, 43(3):578–605, February 10, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brewer:2024:CEC

[1440] Benjamin C. Brewer and Leonidas E. Bantis. Cutoff estimation and construction of their confidence intervals for continuous biomarkers under ternary umbrella and tree stochastic ordering settings. Statistics in Medicine, 43(3):606–623, February 10, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ren:2024:SIL

[1441] Jingchen Ren and Wei Pan. Statistical inference with large-scale trait imputation. Statistics in Medicine, 43(4):625–641, February 20, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mirzaei:2024:AAS

[1442] Sedigheh Mirzaei, José M. Martínez, Eric J. Chow, Kirsten K. Ness, Melissa M. Hudson, Gregory T. Armstrong, and Yutaka Yasui. Association analysis of self-reported outcomes with a validated subset. Statistics in Medicine, 43(4):642–655, February 20, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tai:2024:MRE

[1443] An-Shun Tai and Sheng-Hsuan Lin. Multiply robust estimation of natural indirect effects with multiple ordered mediators. Statistics in Medicine, 43(4):656–673, February 20, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ye:2024:JMA

[1444] Peng Ye, Shuo Bai, Wan Tang, Han Feng, Xinhua Qiao, Shengjia Tu, and Hua He. Joint modeling approaches for censored predictors due to detection limits with applications to metabolites data. Statistics in Medicine, 43(4):674–688, February 20, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See erratum [1626].

Yang:2024:EBS

[1445] Chao Yang and Yisheng Li. An extended Bayesian semi-mechanistic dose-finding design for phase I oncology trials using pharmacokinetic and pharmacodynamic information. Statistics in Medicine, 43(4):689– 705, February 20, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sangnawakij:2024:NER

[1446] Patarawan Sangnawakij, Dankmar Böhning, Heinz Holling, and Katrin Jansen. Nonparametric estimation of the random effects distribution for the risk or rate ratio in rare events meta-analysis with the arm-based and contrast-based approaches. Statistics in Medicine, 43(4):706–730, February 20, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sjolander:2024:CMR

[1447] Arvid Sjölander, Thomas Frisell, Sara Öberg, Yunzhang Wang, and Sara Hägg. Combining Mendelian randomization with the sibling comparison design. Statistics in Medicine, 43(4):731–755, February 20, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jackson:2024:MEB

[1448] Dan Jackson, Wolfgang Viechtbauer, and Robbie C. M. van Aert. Multistep estimators of the between-study covariance matrix under the multivariate random-effects model for meta-analysis. Statistics in Medicine, 43(4):756–773, February 20, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Duan:2024:FES

[1449] Yunshan Duan and Layla Parast. Flexible evaluation of surrogate markers with Bayesian model averaging. Statistics in Medicine, 43(4):774– 792, February 20, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shook-Sa:2024:FTD

[1450] Bonnie E. Shook-Sa, Paul N. Zivich, Samuel P. Rosin, Jessie K. Edwards, Adaora A. Adimora, Michael G. Hudgens, and Stephen R. Cole. Fusing trial data for treatment comparisons: Single vs multi-span bridging. Statistics in Medicine, 43(4):793–815, February 20, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chase:2024:MBB

[1451] Elizabeth C. Chase, Jeremy M. G. Taylor, and Philip S. Boonstra. Modeling basal body temperature data using horseshoe process regression. Statistics in Medicine, 43(5):817–832, February 28, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Boulaguiem:2024:FSC

[1452] Younes Boulaguiem, Julie Quartier, Maria Lapteva, Yogeshvar N. Kalia, Maria-Pia Victoria-Feser, Stéphane Guerrier, and Dominique-Laurent Couturier. Finite sample corrections for average equivalence testing. Statistics in Medicine, 43(5):833–854, February 28, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

vanZwet:2024:ESE

[1453] Erik W. van Zwet, Lu Tian, and Robert Tibshirani. Evaluating a shrinkage estimator for the treatment effect in clinical trials. Statistics in Medicine, 43(5):855–868, February 28, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2024:EJC

[1454] Jia Wang, Jingjing Yin, and Lili Tian. Evaluating joint confidence region of hypervolume under ROC manifold and generalized Youden index. Statistics in Medicine, 43(5):869–889, February 28, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2024:PSW

[1455] Fan Li, Xinyuan Chen, Zizhong Tian, Rui Wang, and Patrick J. Heagerty. Planning stepped wedge cluster randomized trials to detect treatment effect heterogeneity. Statistics in Medicine, 43(5):890–911, February 28, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Steen:2024:DTD

[1456] Johan Steen, Paweł Morzywołek, Wim Van Biesen, Johan Decruyenaere, and Stijn Vansteelandt. Dealing with time-dependent exposures and confounding when defining and estimating attributable fractions-revisiting estimands and estimators. Statistics in Medicine, 43(5):912–934, February 28, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nyberg:2024:PFR

[1457] Joakim Nyberg, E. Niclas Jonsson, Mats O. Karlsson, and Jonas Häggström. Properties of the full random-effect modeling approach with missing covariate data. Statistics in Medicine, 43(5):935–952, February 28, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kuh:2024:ULO

[1458] Swen Kuh, Lauren Kennedy, Qixuan Chen, and Andrew Gelman. Using leave-one-out cross validation (LOO) in a multilevel regression and poststratification (MRP) workflow: a cautionary tale. Statistics in Medicine, 43(5):953–982, February 28, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mallick:2024:IBF

[1459] Himel Mallick, Anupreet Porwal, Satabdi Saha, Piyali Basak, Vladimir Svetnik, and Erina Paul. An integrated Bayesian framework for multiomics prediction and classification. Statistics in Medicine, 43(5):983– 1002, February 28, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shen:2024:ARF

[1460] Biyi Shen, Haoyu Ren, Michelle Shardell, Jason Falvey, and Chixiang Chen. Analyzing risk factors for post-acute recovery in older adults with Alzheimer’s disease and related dementia: a new semi-parametric model for large-scale Medicare claims. Statistics in Medicine, 43(5):1003–1018, February 28, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Devisri:2024:FGA

[1461] B. Devisri and M. Kavitha. Fetal growth analysis from ultrasound videos based on different biometrics using optimal segmentation and hybrid classifier. Statistics in Medicine, 43(5):1019–1047, February 28, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lovblom:2024:MMC

[1462] Leif Erik Lovblom, Laurent Briollais, Bruce A. Perkins, and George Tomlinson. Modeling multiple correlated end-organ disease trajectories: a tutorial for multistate and joint models with applications in diabetes complications. Statistics in Medicine, 43(5):1048–1082, February 28, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Abad:2024:ITA

[1463] Ariel Alonso Abad, Fenny Ong, Florian Stijven, Wim Van der Elst, Geert Molenberghs, Ingrid Van Keilegom, Geert Verbeke, and Andrea Callegaro. An information-theoretic approach for the assessment of a continuous outcome as a surrogate for a binary true endpoint based on causal inference: Application to vaccine evaluation. Statistics in Medicine, 43 (6):1083–1102, March 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wan:2024:CSL

[1464] Fang Wan, Wei Liu, and Frank Bretz. Confidence sets for a level set in linear regression. Statistics in Medicine, 43(6):1103–1118, March 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dunias:2024:CHT

[1465] Zoë S. Dunias, Ben Van Calster, Dirk Timmerman, Anne-Laure Boulesteix, and Maarten van Smeden. A comparison of hyperparameter tuning procedures for clinical prediction models: a simulation study. Statistics in Medicine, 43(6):1119–1134, March 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cremaschi:2024:JMA

[1466] Andrea Cremaschi, Maria De Iorio, Narasimhan Kothandaraman, Fabian Yap, Mya Thway Tint, and Johan Eriksson. Joint modeling of association networks and longitudinal biomarkers: an application to childhood obesity. Statistics in Medicine, 43(6):1135–1152, March 15, 2024. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dai:2024:BFM

[1467] Xiaotian Dai, Nicole Acosta, Xuewen Lu, Casey R. J. Hubert, Jangwoo Lee, Kevin Frankowski, Maria A. Bautista, Barbara J. Waddell, Kristine Du, Janine McCalder, Jon Meddings, Norma Ruecker, Tyler Williamson, Danielle A. Southern, Jordan Hollman, Gopal Achari, M. Cathryn Ryan, Steve E. Hrudey, Bonita E. Lee, Xiaoli Pang, Rhonda G. Clark, Michael D. Parkins, and Thierry Chekouo. A Bayesian framework for modeling COVID-19 case numbers through longitudinal monitoring of SARS-CoV-2 RNA in wastewater. Statistics in Medicine, 43(6):1153– 1169, March 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2024:IBR

[1468] Yizhuo Wang and Susan Murray. -inflated beta regression model for censored recurrent events. Statistics in Medicine, 43(6):1170–1193, March 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhao:2024:SSS

[1469] Wenle Zhao, Kerstine Carter, Oleksandr Sverdlov, Annika Scheffold, Yevgen Ryeznik, Christy Cassarly, and Vance W. Berger. Steady-state statistical properties and implementation of randomization designs with maximum tolerated imbalance restriction for two-arm equal allocation clinical trials. Statistics in Medicine, 43(6):1194–1212, March 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2024:EEC

[1470] Xuerong Chen, Yalu Ping, and Jianguo Sun. Efficient estimation of Cox model with random change point. Statistics in Medicine, 43(6):1213– 1226, March 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Perumalraja:2024:SPR

[1471] Rengaraju Perumalraja, B. Felcia Logan’s Deshna, and N. Swetha. Statistical performance review on diagnosis of leukemia, glaucoma and diabetes mellitus using AI. Statistics in Medicine, 43(6):1227–1237, March 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Curnow:2024:MIS

[1472] Elinor Curnow, Rachael A. Hughes, Kate Birnie, Kate Tilling, and Michael J. Crowther. Multiple imputation strategies for missing event times in a multi-state model analysis. Statistics in Medicine, 43(6):1238– 1255, March 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xu:2024:BRA

[1473] Zhongying Xu, Tianzhou Ma, Lu Tang, Victor B. Talisa, and Chung-Chou H. Chang. Bayesian response adaptive randomization design with a composite endpoint of mortality and morbidity. Statistics in Medicine, 43(6):1256–1270, March 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Marschner:2024:CDT

[1474] Ian C. Marschner. Confidence distributions for treatment effects in clinical trials: Posteriors without priors. Statistics in Medicine, 43(6):1271– 1289, March 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brantner:2024:CMC

[1475] Carly Lupton Brantner, Trang Quynh Nguyen, Tengjie Tang, Congwen Zhao, Hwanhee Hong, and Elizabeth A. Stuart. Comparison of methods that combine multiple randomized trials to estimate heterogeneous treatment effects. Statistics in Medicine, 43(7):1291–1314, March 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rizopoulos:2024:ODP

[1476] Dimitris Rizopoulos and Jeremy M. G. Taylor. Optimizing dynamic predictions from joint models using super learning. Statistics in Medicine, 43(7):1315–1328, March 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2024:CFM

[1477] Xiaoyan Liu and Jiming Jiang. Classified functional mixed effects model prediction. Statistics in Medicine, 43(7):1329–1340, March 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jiang:2024:ASH

[1478] Xinyang Jiang, Wen Li, Ruosha Li, and Jing Ning. Addressing subject heterogeneity in time-dependent discrimination for biomarker evaluation. Statistics in Medicine, 43(7):1341–1353, March 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2024:NEL

[1479] Yaliang Zhang and Yunro Chung. Nonparametric estimation of linear personalized diagnostics rules via efficient grid algorithm. Statistics in Medicine, 43(7):1354–1371, March 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Du:2024:LRC

[1480] Zhiyuan Du, Pang Du, and Aiyi Liu. Likelihood ratio combination of multiple biomarkers via smoothing spline estimated densities. Statistics in Medicine, 43(7):1372–1383, March 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thomassen:2024:ESS

[1481] Doranne Thomassen, Saskia le Cessie, Hans C. van Houwelingen, and Ewout W. Steyerberg. Effective sample size: a measure of individual uncertainty in predictions. Statistics in Medicine, 43(7):1384–1396, March 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tan:2024:SSD

[1482] Xianming Tan, William Wang, Donglin Zeng, Guanghan F. Liu, Guoqing Diao, Niusha Jafari, Ethan M. Alt, and Joseph G. Ibrahim. Safety signal detection with control of latent factors. Statistics in Medicine, 43(7): 1397–1418, March 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hudson:2024:SIQ

[1483] Aaron Hudson and Ali Shojaie. Statistical inference on qualitative differences in the magnitude of an effect. Statistics in Medicine, 43(7): 1419–1440, March 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Boyle:2024:MVM

[1484] Joseph Boyle, Mary H. Ward, James R. Cerhan, Nat Rothman, and David C. Wheeler. Modeling variation in mixture effects over space with a Bayesian spatially varying mixture model. Statistics in Medicine, 43 (7):1441–1457, March 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rabideau:2024:MRG

[1485] Dustin J. Rabideau, Fan Li, and Rui Wang. Multiply robust generalized estimating equations for cluster randomized trials with missing outcomes. Statistics in Medicine, 43(7):1458–1474, March 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Munoz:2024:SAB

[1486] Joel Muñoz, Jordi Ocaña, Rolando Suárez, and Carolina Millapán. Scaled average bioequivalence methods for highly variable drugs: Leveling-off soft limits and the EMA’s 2010 guideline (some ways to improve its type I error control). Statistics in Medicine, 43(7):1475–1488, March 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Poulos:2024:TLO

[1487] Jason Poulos, Marcela Horvitz-Lennon, Katya Zelevinsky, Tudor Cristea-Platon, Thomas Huijskens, Pooja Tyagi, Jiaju Yan, Jordi Diaz, and Sharon-Lise Normand. Targeted learning in observational studies with multi-valued treatments: an evaluation of antipsychotic drug treatment safety. Statistics in Medicine, 43(8):1489–1508, April 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2024:SVS

[1488] Zili Liu and Hong Wang. Simultaneous variable selection and estimation for survival data via the Gaussian seamless-penalty. Statistics in Medicine, 43(8):1509–1526, April 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2024:MPL

[1489] Kuo-Jung Lee, Chanmin Kim, Jae Keun Yoo, and Keunbaik Lee. Multivariate probit linear mixed models for multivariate longitudinal binary data. Statistics in Medicine, 43(8):1527–1548, April 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Meng:2024:SAI

[1490] Zhuo Meng, Jingshen Wang, Lifeng Lin, and Chong Wu. Sensitivity analysis with iterative outlier detection for systematic reviews and metaanalyses. Statistics in Medicine, 43(8):1549–1563, April 15, 2024. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sun:2024:GFL

[1491] Jiehuan Sun and Kuang-Yao Lee. Generalized functional linear model with a point process predictor. Statistics in Medicine, 43(8):1564–1576, April 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Meis:2024:PEC

[1492] Jan Meis, Maximilian Pilz, Björn Bokelmann, Carolin Herrmann, Geraldine Rauch, and Meinhard Kieser. Point estimation, confidence intervals, and P -values for optimal adaptive two-stage designs with normal endpoints. Statistics in Medicine, 43(8):1577–1603, April 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2024:DFH

[1493] Wei Liu, Frank Bretz, and Mario Cortina-Borja. Distribution-free hyperrectangular tolerance regions for setting multivariate reference regions in laboratory medicine. Statistics in Medicine, 43(8):1604–1614, April 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Warren:2024:SKD

[1494] Joshua L. Warren, Qi Wang, and Maria M. Ciarleglio. A scaled kernel density estimation prior for dynamic borrowing of historical information with application to clinical trial design. Statistics in Medicine, 43(8): 1615–1626, April 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2024:MED

[1495] Rui Wang, Mengqi Cen, Yunda Huang, George Qian, Natalie E. Dean, Susan S. Ellenberg, Thomas R. Fleming, Wenbin Lu, and Ira M. Longini. Methods for the estimation of direct and indirect vaccination effects by combining data from individual- and cluster-randomized trials. Statistics in Medicine, 43(8):1627–1639, April 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Seo:2024:ULL

[1496] Jaehyun Seo and Chanmin Kim. Utilizing local likelihood in regression discontinuity design: Investigating the impact of antiretroviral therapy eligibility on retention in clinical HIV care in south Africa. Statistics in Medicine, 43(8):1640–1659, April 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jiang:2024:MCP

[1497] Shu Jiang and Graham A. Colditz. Modeling correlated pairs of mammogram images. Statistics in Medicine, 43(8):1660–1668, April 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lipkovich:2024:CUP

[1498] Ilya Lipkovich, Bohdana Ratitch, Yongming Qu, Xiang Zhang, Mingyang Shan, and Craig Mallinckrodt. Correction to “Using principal stratification in analysis of clinical trials”. Statistics in Medicine, 43(8):1669, April 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [976].

Chung:2024:SRA

[1499] Yunro Chung, Anastasia Ivanova, and Jason P. Fine. Shape restricted additive hazards models: Monotone, unimodal, and U-shape hazard functions. Statistics in Medicine, 43(9):1671–1687, April 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2024:TOU

[1500] Peiran Liu, Ming-Hui Chen, Susie Sinks, and Peng Sun. Are the tests overpowered or underpowered? A unified solution to correctly specify type I errors in design of clinical trials for two sample proportions. Statistics in Medicine, 43(9):1688–1707, April 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thao:2024:HMD

[1501] Le Thi Phuong Thao, Rory Wolfe, Stephane Heritier, and Ronald Geskus. Handling missing disease information due to death in diseases that need two visits to diagnose. Statistics in Medicine, 43(9):1708–1725, April 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2024:MRB

[1502] Chang Wang and Mingyue Du. Martingale-residual-based greedy model averaging for high-dimensional current status data. Statistics in Medicine, 43(9):1726–1742, April 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cao:2024:DAB

[1503] Wenhao Cao, Hongjian Zhu, Li Wang, Lixin Zhang, and Jun Yu. Doubly adaptive biased coin design to improve Bayesian clinical trials with time-to-event endpoints. Statistics in Medicine, 43(9):1743–1758, April 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cheung:2024:ETP

[1504] Yin Bun Cheung, Xiangmei Ma, K. F. Lam, Chee Fu Yung, and Paul Milligan. Estimation of trajectory of protective efficacy in infectious disease prevention trials using recurrent event times. Statistics in Medicine, 43(9):1759–1773, April 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Loewinger:2024:OEC

[1505] Gabriel Loewinger, Rolando Acosta Nunez, Rahul Mazumder, and Giovanni Parmigiani. Optimal ensemble construction for multistudy prediction with applications to mortality estimation. Statistics in Medicine, 43(9):1774–1789, April 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2024:RBL

[1506] Ching-Yun Wang, Li Hsu, and Tabitha Harrison. Robust best linear weighted estimator with missing covariates in survival analysis. Statistics in Medicine, 43(9):1790–1803, April 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schreck:2024:SPS

[1507] Nicholas Schreck, Alla Slynko, Maral Saadati, and Axel Benner. Statistical plasmode simulations-potentials, challenges and recommendations. Statistics in Medicine, 43(9):1804–1825, April 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chowell:2024:PEF

[1508] Gerardo Chowell, Amanda Bleichrodt, and Ruiyan Luo. Parameter estimation and forecasting with quantified uncertainty for ordinary differential equation models using QuantDiffForecast: a MATLAB toolbox and tutorial. Statistics in Medicine, 43(9):1826–1848, April 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Munko:2024:RBM

[1509] Merle Munko, Marc Ditzhaus, Dennis Dobler, and Jon Genuneit. RMSTbased multiple contrast tests in general factorial designs. Statistics in Medicine, 43(10):1849–1866, May 10, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Huang:2024:DCT

[1510] Jenny Huang, Raphaël Morsomme, David Dunson, and Jason Xu. Detecting changes in the transmission rate of a stochastic epidemic model. Statistics in Medicine, 43(10):1867–1882, May 10, 2024. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2024:TSS

[1511] Yanping Chen, Yong Lin, Shou-En Lu, Weichung J. Shih, and Hui Quan. Two-stage stratified designs with survival outcomes and adjustment for misclassification in predictive biomarkers. Statistics in Medicine, 43(10): 1883–1904, May 10, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cao:2024:BNM

[1512] Wenhao Cao, Haitao Chu, Timothy Hanson, and Lianne Siegel. A Bayesian nonparametric meta-analysis model for estimating the reference interval. Statistics in Medicine, 43(10):1905–1919, May 10, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Troendle:2024:UWT

[1513] James F. Troendle, Eric S. Leifer, Song Yang, Neal Jeffries, Dong-Yun Kim, Jungnam Joo, and Christopher M. O’Connor. Use of win time for ordered composite endpoints in clinical trials. Statistics in Medicine, 43(10):1920–1932, May 10, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schmidt:2024:IUS

[1514] Patrick W. Schmidt. Inference under superspreading: Determinants of SARS-CoV-2 transmission in Germany. Statistics in Medicine, 43(10): 1933–1954, May 10, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2024:ERB

[1515] Kenneth Menglin Lee and Yin Bun Cheung. Estimation and reduction of bias in self-controlled case series with non-rare event dependent outcomes and heterogeneous populations. Statistics in Medicine, 43(10):1955–1972, May 10, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

DeSantis:2024:DPF

[1516] Fulvio De Santis and Stefania Gubbiotti. On the distribution of the power function for the scale parameter of exponential families. Statistics in Medicine, 43(10):1973–1992, May 10, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhan:2024:CCM

[1517] Tianyu Zhan. A class of computational methods to reduce selection bias when designing Phase 3 clinical trials. Statistics in Medicine, 43(10): 1993–2006, May 10, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Song:2024:PWS

[1518] Yanan Song, Haohui Han, Liya Fu, and Ting Wang. Penalized weighted smoothed quantile regression for high-dimensional longitudinal data. Statistics in Medicine, 43(10):2007–2042, May 10, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bouvier:2024:DML

[1519] Florie Bouvier, Etienne Peyrot, Alan Balendran, Corentin Ségalas, Ian Roberts, François Petit, and Raphaël Porcher. Do machine learning methods lead to similar individualized treatment rules? A comparison study on real data. Statistics in Medicine, 43(11):2043–2061, May 20, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lou:2024:NUM

[1520] Yichen Lou, Yuqing Ma, and Mingyue Du. A new and unified method for regression analysis of interval-censored failure time data under semiparametric transformation models with missing covariates. Statistics in Medicine, 43(11):2062–2082, May 20, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sullivan:2024:CCC

[1521] Thomas R. Sullivan, Tim P. Morris, Brennan C. Kahan, Alana R. Cuthbert, and Lisa N. Yelland. Categorisation of continuous covariates for stratified randomisation: How should we adjust? Statistics in Medicine, 43(11):2083–2095, May 20, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2024:PMZ

[1522] Yin Liu and Ziwen Gao. Predicting the multivariate zero-inflated counts: a novel model averaging method under Pearson loss. Statistics in Medicine, 43(11):2096–2121, May 20, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chukhrova:2024:MEP

[1523] Nataliya Chukhrova, Oskar Plate, and Arne Johannssen. Monitoring epidemic processes under political measures. Statistics in Medicine, 43 (11):2122–2160, May 20, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Suresh:2024:MAE

[1524] Krithika Suresh, Carsten Görg, and Debashis Ghosh. Model-agnostic explanations for survival prediction models. Statistics in Medicine, 43 (11):2161–2182, May 20, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McCaffrey:2024:EGP

[1525] Daniel F. McCaffrey, Beth Ann Griffin, Michael Robbins, Yajnaseni Chakraborti, Donna L. Coffman, and Brian Vegetabile. Estimating generalized propensity scores with survey and attrition weighted data. Statistics in Medicine, 43(11):2183–2202, May 20, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cui:2024:SSA

[1526] Lu Cui. Sample size adaptation designs and efficiency comparison with group sequential designs. Statistics in Medicine, 43(11):2203–2215, May 20, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dobler:2024:NRT

[1527] Dennis Dobler and Kathrin Möllenhoff. A nonparametric relative treatment effect for direct comparisons of censored paired survival outcomes. Statistics in Medicine, 43(11):2216–2238, May 20, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Heath:2024:HCP

[1528] Bethany Heath, Sofía S. Villar, and David S. Robertson. How could a pooled testing policy have performed in managing the early stages of the COVID-19 pandemic? Results from a simulation study. Statistics in Medicine, 43(11):2239–2262, May 20, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hu:2024:CIT

[1529] Mengtong Hu, Xu Shi, and Peter X.-K. Song. Collaborative inference for treatment effect with distributed data-sharing management in multicenter studies. Statistics in Medicine, 43(11):2263–2279, May 20, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Han:2024:IIS

[1530] Wei Han, Sanguo Zhang, Shuangge Ma, and Mingyang Ren. Informationincorporated sparse hierarchical cancer heterogeneity analysis. Statistics in Medicine, 43(11):2280–2297, May 20, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chernofsky:2024:CMA

[1531] Ariel Chernofsky, Ronald J. Bosch, and Judith J. Lok. Causal mediation analysis with mediator values below an assay limit. Statistics in Medicine, 43(12):2299–2313, May 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dodd:2024:AEN

[1532] Matthew Dodd, James Carpenter, Jennifer A. Thompson, Elizabeth Williamson, Katherine Fielding, and Diana Elbourne. Assessing efficacy in non-inferiority trials with non-adherence to interventions: Are intention-to-treat and per-protocol analyses fit for purpose? Statistics in Medicine, 43(12):2314–2331, May 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ben-Michael:2024:ABW

[1533] Eli Ben-Michael, Lindsay Page, and Luke Keele. Approximate balancing weights for clustered observational study designs. Statistics in Medicine, 43(12):2332–2358, May 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cao:2024:CIO

[1534] Shiwei Cao and Sin-Ho Jung. Confidence intervals for odds ratio from multistage randomized phase II trials. Statistics in Medicine, 43(12): 2359–2367, May 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schuurhuis:2024:TSG

[1535] Stephen Schüürhuis, Frank Konietschke, and Cornelia Ursula Kunz. A two-stage group-sequential design for delayed treatment responses with the possibility of trial restart. Statistics in Medicine, 43(12):2368–2388, May 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Etzkorn:2024:JFM

[1536] Lacey H. Etzkorn, Quentin Le Coënt, Mark van den Boogaard, Virginie Rondeau, and Elizabeth Colantuoni. A joint frailty model for recurrent and competing terminal events: Application to delirium in the ICU. Statistics in Medicine, 43(12):2389–2402, May 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hartman:2024:IEN

[1537] Nicholas Hartman and Kevin He. Individualized empirical null estimation for exact tests of healthcare quality. Statistics in Medicine, 43(12): 2403–2420, May 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jonker:2024:BFI

[1538] Marianne A. Jonker, Hassan Pazira, and Anthony Cc Coolen. Bayesian federated inference for estimating statistical models based on non-shared multicenter data sets. Statistics in Medicine, 43(12):2421–2438, May 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wei:2024:BPT

[1539] Wei Wei, Ondrej Blaha, Denise Esserman, Daniel Zelterman, Michael Kane, Rachael Liu, and Jianchang Lin. A Bayesian platform trial design with hybrid control based on multisource exchangeability modelling. Statistics in Medicine, 43(12):2439–2451, May 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

You:2024:DAM

[1540] Lu You, Xiang Liu, and Jeffrey Krischer. A discrete approximation method for modeling interval-censored multistate data. Statistics in Medicine, 43(12):2452–2471, May 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yan:2024:SCM

[1541] Zhiwu Yan and Min Yang. Statistical considerations in model-based dose finding for binary responses under model uncertainty. Statistics in Medicine, 43(12):2472–2485, May 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhao:2024:FBS

[1542] Beibo Zhao, Jason Fine, and Anastasia Ivanova. Finding the best subgroup with differential treatment effect with multiple outcomes. Statistics in Medicine, 43(13):2487–2500, June 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zou:2024:OSH

[1543] Yudan Zou, Xinyuan Song, and Qian Zhao. Order selection for heterogeneous semiparametric hidden Markov models. Statistics in Medicine, 43(13):2501–2526, June 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sarkar:2024:MPL

[1544] Shoumi Sarkar, Samuel Anyaso-Samuel, Peihua Qiu, and Somnath Datta. Multiblock partial least squares and rank aggregation: Applications to detection of bacteriophages associated with antimicrobial resistance in the presence of potential confounding factors. Statistics in Medicine, 43 (13):2527–2546, June 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sims:2024:PAM

[1545] Andrew Sims, D. Leann Long, Hemant K. Tiwari, Jinhong Cui, Dustin M. Long, Todd M. Brown, Melissa J. Smith, and Emily B. Levitan. Population-average mediation analysis for zero-inflated count outcomes. Statistics in Medicine, 43(13):2547–2559, June 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Choi:2024:VCT

[1546] Jaihee Choi, Zhichao Xu, and Ryan Sun. Variance-components tests for genetic association with multiple interval-censored outcomes. Statistics in Medicine, 43(13):2560–2574, June 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2024:WEC

[1547] Jooyoung Lee, Shuji Ogino, and Molin Wang. Weighting estimation in the cause-specific Cox regression with partially missing causes of failure. Statistics in Medicine, 43(13):2575–2591, June 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gregorich:2024:FPG

[1548] Mariella Gregorich, Sean L. Simpson, and Georg Heinze. Flexible parametrization of graph-theoretical features from individual-specific networks for prediction. Statistics in Medicine, 43(13):2592–2606, June 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2024:ELT

[1549] Wen Li, Qian Wang, Jing Ning, Jing Zhang, Zhouxuan Li, Sean I. Savitz, Amirali Tahanan, and Mohammad H. Rahbar. Enhancing long-term survival prediction with two short-term events: Landmarking with a flexible varying coefficient model. Statistics in Medicine, 43(13):2607–2621, June 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fairfax:2024:DIA

[1550] Sarah R. Fairfax and Shu Yang. Distributional imputation for the analysis of censored recurrent events. Statistics in Medicine, 43(13):2622–2640, June 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Farrington:2024:CIC

[1551] Paddy Farrington and Solène Lellinger. Cumulative incidence of cardiac surgery associated with exposure to benfluorex: a retrospective analysis based on compensation claims data. Statistics in Medicine, 43(13):2641– 2654, June 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2024:DHS

[1552] Ziyue Liu and Wensheng Guo. Dynamic hierarchical state space forecasting. Statistics in Medicine, 43(13):2655–2671, June 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kostouraki:2024:VEI

[1553] Andriana Kostouraki, David Hajage, Bernard Rachet, Elizabeth J. Williamson, Guillaume Chauvet, Aurélien Belot, and Clémence Leyrat. On variance estimation of the inverse probability-of-treatment weighting estimator: a tutorial for different types of propensity score weights. Statistics in Medicine, 43(13):2672–2694, June 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Derkach:2024:MAU

[1554] Andriy Derkach, Elizabeth D. Kantor, Joshua N. Sampson, and Ruth M. Pfeiffer. Mediation analysis using incomplete information from publicly available data sources. Statistics in Medicine, 43(14):2695–2712, June 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sun:2024:TVD

[1555] Lizhe Sun, Aiying Zhang, and Faming Liang. Time-varying dynamic Bayesian network learning for an fMRI study of emotion processing. Statistics in Medicine, 43(14):2713–2733, June 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Luo:2024:OCI

[1556] Lan Luo, Malcolm Risk, and Xu Shi. Online causal inference with application to near real-time post-market vaccine safety surveillance. Statistics in Medicine, 43(14):2734–2746, June 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bridge:2024:NVC

[1557] Helen Bridge, Katy E. Morgan, and Chris Frost. Negative variance components and intercept-slope correlations greater than one in magnitude: How do such “non-regular” random intercept and slope models arise, and what should be done when they do? Statistics in Medicine, 43(14): 2747–2764, June 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yang:2024:LOB

[1558] Bin Yang, Xingche Guo, Ji Meng Loh, Qinxia Wang, and Yuanjia Wang. Learning optimal biomarker-guided treatment policy for chronic disorders. Statistics in Medicine, 43(14):2765–2782, June 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2024:PSM

[1559] Tongrong Wang, Honghe Zhao, Shu Yang, Shuhan Tang, Zhanglin Cui, Li Li, and Douglas E. Faries. Propensity score matching for estimating a marginal hazard ratio. Statistics in Medicine, 43(14):2783–2810, June 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Harari:2024:ADP

[1560] Ofir Harari, Jay J. H. Park, Prince Kumar Lat, and Edward J. Mills. Adaptive designs in public health: Vaccine and cluster randomized trials go Bayesian. Statistics in Medicine, 43(14):2811–2829, June 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pate:2024:CPM

[1561] Alexander Pate, Matthew Sperrin, Richard D. Riley, Niels Peek, Tjeerd Van Staa, Jamie C. Sergeant, Mamas A. Mamas, Gregory Y. H. Lip, Martin O’Flaherty, Michael Barrowman, Iain Buchan, and Glen P. Martin. Calibration plots for multistate risk prediction models. Statistics in Medicine, 43(14):2830–2852, June 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kilpatrick:2024:FOS

[1562] Kayla W. Kilpatrick, Chanhwa Lee, and Michael G. Hudgens. G-formula for observational studies under stratified interference, with application to bed net use on malaria. Statistics in Medicine, 43(15):2853–2868, July 10, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2024:CMA

[1563] Minghao Chen and Yingchun Zhou. Causal mediation analysis with a three-dimensional image mediator. Statistics in Medicine, 43(15):2869– 2893, July 10, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kosko:2024:FBA

[1564] Matthew Kosko, Lin Wang, and Michele Santacatterina. A fast bootstrap algorithm for causal inference with large data. Statistics in Medicine, 43 (15):2894–2927, July 10, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Du:2024:UBF

[1565] Yu Du, Jianghao Li, Sohini Raha, and Yongming Qu. A unified Bayesian framework for bias adjustment in multiple comparisons from clinical trials. Statistics in Medicine, 43(15):2928–2943, July 10, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wu:2024:SSC

[1566] Yougui Wu. Sample size calculation for comparing two screening tests when the gold standard is missing at random. Statistics in Medicine, 43(15):2944–2956, July 10, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Orsini:2024:REM

[1567] Letizia Orsini, Kamila Czene, and Keith Humphreys. Random effects models of tumour growth for investigating interval breast cancer. Statistics in Medicine, 43(15):2957–2971, July 10, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yang:2024:DSS

[1568] Peng Yang, Daniel Li, Ruitao Lin, Bo Huang, and Ying Yuan. Design and sample size determination for multiple-dose randomized phase II trials for dose optimization. Statistics in Medicine, 43(15):2972–2986, July 10, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2024:NNL

[1569] Guoqiao Wang, Whedy Wang, Brian Mangal, Yijie Liao, Lon Schneider, Yan Li, Chengjie Xiong, Eric McDade, Richard Kennedy, Randall Bateman, and Gary Cutter. Novel non-linear models for clinical trial analysis with longitudinal data: a tutorial using SAS for both frequentist and Bayesian methods. Statistics in Medicine, 43(15):2987–3004, July 10, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Adegboye:2024:HEU

[1570] Oyelola A. Adegboye, Tomoki Fujii, Denis Heng-Yan Leung, and Li Siyu. HIV estimation using population-based surveys with non-response: a partial identification approach. Statistics in Medicine, 43(16):3005–3019, July 20, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2024:RIC

[1571] Yan-Lin Chen, Yan-Hong Chen, Pei-Fang Su, Huang-Tz Ou, and An-Shun Tai. Robust inference for causal mediation analysis of recurrent event data. Statistics in Medicine, 43(16):3020–3035, July 20, 2024. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2024:PSB

[1572] Youjin Lee, Peter P. Reese, Amelia H. Tran, and Douglas E. Schaubel. Prognostic score-based methods for estimating center effects based on survival probability: Application to post-kidney transplant survival. Statistics in Medicine, 43(16):3036–3050, July 20, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yang:2024:EMP

[1573] James J. Yang and Anne Buu. Efficient matrix profile computation with Euclidean distance using Eigen transformation: Performance evaluation based on beat-to-beat interval (BBI) data. Statistics in Medicine, 43(16): 3051–3061, July 20, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dayimu:2024:SSD

[1574] Alimu Dayimu, Nikola Simidjievski, Nikolaos Demiris, and Jean Abraham. Sample size determination for prediction models via learning-type curves. Statistics in Medicine, 43(16):3062–3072, July 20, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Porter:2024:FCP

[1575] Erica M. Porter, Christopher T. Franck, and Stephen Adams. Flexible cost-penalized Bayesian model selection: Developing inclusion paths with an application to diagnosis of heart disease. Statistics in Medicine, 43 (16):3073–3091, July 20, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hopkins:2024:SOA

[1576] Will G. Hopkins and David S. Rowlands. Standardization and other approaches to meta-analyze differences in means. Statistics in Medicine, 43(16):3092–3108, July 20, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Korn:2024:DRC

[1577] Edward L. Korn and Boris Freidlin. Design of randomized clinical trials with a binary endpoint: Conditional versus unconditional analyses of a two-by-two table. Statistics in Medicine, 43(16):3109–3123, July 20, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bannick:2024:ECS

[1578] Marlena Bannick, Deborah Donnell, Richard Hayes, Oliver Laeyendecker, and Fei Gao. An enhanced cross-sectional HIV incidence estimator that incorporates prior HIV test results. Statistics in Medicine, 43(17):3125–3139, July 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Filla:2024:BVM

[1579] Tim Filla, Holger Schwender, and Oliver Kuss. Balancing versus modelling in weighted analysis of non-randomised studies with survival outcomes: a simulation study. Statistics in Medicine, 43(17):3140–3163, July 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2024:SLT

[1580] Guanbo Wang, Yi Lian, Archer Y. Yang, Robert W. Platt, Rui Wang, Sylvie Perreault, Marc Dorais, and Mireille E. Schnitzer. Structured learning in time-dependent Cox models. Statistics in Medicine, 43(17): 3164–3183, July 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Parast:2024:AHS

[1581] Layla Parast, Lu Tian, and Tianxi Cai. Assessing heterogeneity in surrogacy using censored data. Statistics in Medicine, 43(17):3184–3209, July 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fang:2024:FAC

[1582] Jialu Fang and Guosheng Yin. Fractional accumulative calibration-free odds (f-aCFO) design for delayed toxicity in phase I clinical trials. Statistics in Medicine, 43(17):3210–3226, July 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yang:2024:SMP

[1583] James J. Yang, Megan E. Piper, Premananda Indic, and Anne Buu. Statistical methods for predicting e-cigarette use events based on beatto-beat interval (BBI) data collected from wearable devices. Statistics in Medicine, 43(17):3227–3238, July 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dong:2024:MII

[1584] Mingfei Dong, Donatello Telesca, Michele Guindani, Catherine Sugar, Sara J. Webb, Shafali Jeste, Abigail Dickinson, April R. Levin, Frederick Shic, Adam Naples, Susan Faja, Geraldine Dawson, James C. Mc-Partland, and Damla Sentürk. Modeling intra-individual inter-trial EEG response variability in autism. Statistics in Medicine, 43(17):3239–3263, July 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Austin:2024:ENC

[1585] Peter C. Austin. The effect of number of clusters and magnitude of within-cluster homogeneity in outcomes on the performance of four variance estimators for a marginal multivariable Cox regression model fit to clustered data in the context of observational research. Statistics in Medicine, 43(17):3264–3279, July 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Alvares:2024:BGB

[1586] Danilo Alvares and François Mercier. Bridging the gap between twostage and joint models: The case of tumor growth inhibition and overall survival models. Statistics in Medicine, 43(17):3280–3293, July 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Huang:2024:DRN

[1587] Xianzheng Huang and Hongmei Zhang. Detecting responsible nodes in differential Bayesian networks. Statistics in Medicine, 43(17):3294–3312, July 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Krisam:2024:UIP

[1588] Johannes Krisam, Yevgen Ryeznik, Kerstine Carter, Olga Kuznetsova, and Oleksandr Sverdlov. Understanding an impact of patient enrollment pattern on predictability of central (unstratified) randomization in a multi-center clinical trial. Statistics in Medicine, 43(17):3313–3325, July 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Voldal:2024:RAS

[1589] Emily C. Voldal, Avi Kenny, Fan Xia, Patrick Heagerty, and James P. Hughes. Robust analysis of stepped wedge trials using composite likelihood models. Statistics in Medicine, 43(17):3326–3352, July 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tada:2024:BMD

[1590] Keisuke Tada, Kazushi Maruo, and Masahiko Gosho. A Bayesian method to detect drug-drug interaction using external information for spontaneous reporting system. Statistics in Medicine, 43(18):3353–3363, August 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Boumendil:2024:TSR

[1591] Luana Boumendil, Sylvie Chevret, Vincent Lévy, and Lucie Biard. Twostage randomized clinical trials with a right-censored endpoint: Comparison of frequentist and Bayesian adaptive designs. Statistics in Medicine, 43(18):3364–3382, August 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2024:SPI

[1592] Yuhan Li, Yiding Zhang, Gu Mi, and Ji Lin. A seamless phase II/III design with dose optimization for oncology drug development. Statistics in Medicine, 43(18):3383–3402, August 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Philipp:2024:MBB

[1593] Morgane Philipp, Adrien Tessier, Mark Donnelly, Lanyan Fang, Kairui Feng, Liang Zhao, Stella Grosser, Guoying Sun, Wanjie Sun, France Mentré, and Julie Bertrand. Model-based bioequivalence approach for sparse pharmacokinetic bioequivalence studies: Model selection or model averaging? Statistics in Medicine, 43(18):3403–3416, August 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thomsen:2024:FEM

[1594] Henrik F. Thomsen, Nanna L. Lausvig, Christian B. Pipper, Søren Andersen, Lars H. Damgaard, Scott S. Emerson, and Henrik Ravn. Familywise error for multiple time-to-event endpoints in a group sequential design. Statistics in Medicine, 43(18):3417–3431, August 15, 2024. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Beall:2024:BHP

[1595] Jonathan Beall, Hong Li, Bonnie Martin-Harris, Brian Neelon, Jordan Elm, Evan Graboyes, and Elizabeth Hill. Bayesian hierarchical profile regression for binary covariates. Statistics in Medicine, 43(18):3432–3446, August 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Burnett:2024:AET

[1596] Thomas Burnett, Franz König, and Thomas Jaki. Adding experimental treatment arms to multi-arm multi-stage platform trials in progress. Statistics in Medicine, 43(18):3447–3462, August 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2024:CLD

[1597] Li-Chun Zhang and Tiziana Tuoto. Categorical linkage-data analysis. Statistics in Medicine, 43(18):3463–3483, August 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yang:2024:RBR

[1598] Cheng-Han Yang, Evan Kwiatkowski, J. Jack Lee, and Ruitao Lin. RE-DOMA: Bayesian random-effects dose-optimization meta-analysis using spike-and-slab priors. Statistics in Medicine, 43(18):3484–3502, August 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wen:2024:CSA

[1599] Chi-Chung Wen and Yi-Hau Chen. Conditional score approaches to errors-in-variables competing risks data in discrete time. Statistics in Medicine, 43(18):3503–3523, August 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sadatsafavi:2024:NPI

[1600] Mohsen Sadatsafavi and John Petkau. Non-parametric inference on calibration of predicted risks. Statistics in Medicine, 43(18):3524–3538, August 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rohde:2024:BTM

[1601] Maximilian D. Rohde, Benjamin French, Thomas G. Stewart, and Frank E. Harrell, Jr. Bayesian transition models for ordinal longitudinal outcomes. Statistics in Medicine, 43(18):3539–3561, August 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2024:CTH

[1602] Xiaoxi Zhang, Somnath Datta, and Peihua Qiu. Comparing two hazard curves when there is a treatment time-lag effect. Statistics in Medicine, 43(19):3563–3577, August 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tian:2024:GSI

[1603] Zibo Tian and Peihua Qiu. Generalized single index modeling of longitudinal data with multiple binary responses. Statistics in Medicine, 43 (19):3578–3594, August 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Salsbury:2024:AMD

[1604] James A. Salsbury, Jeremy E. Oakley, Steven A. Julious, and Lisa V. Hampson. Assurance methods for designing a clinical trial with a delayed treatment effect. Statistics in Medicine, 43(19):3595–3612, August 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Greenstreet:2024:MAM

[1605] Peter Greenstreet, Thomas Jaki, Alun Bedding, Chris Harbron, and Pavel Mozgunov. A multi-arm multi-stage platform design that allows preplanned addition of arms while still controlling the family-wise error. Statistics in Medicine, 43(19):3613–3632, August 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lu:2024:SFM

[1606] Zihang Lu and Noirrit Kiran Chandra. A sparse factor model for clustering high-dimensional longitudinal data. Statistics in Medicine, 43 (19):3633–3648, August 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shi:2024:IMA

[1607] Xiaosong Shi, Jo A. Wick, Renee’ L. Martin, Jonathan Beall, Robert Silbergleit, Gaylan L. Rockswold, William G. Barsan, Frederick K. Korley, Sarah Rockswold, and Byron J. Gajewski. Improved mortality analysis in early-phase dose-ranging clinical trials for emergency medical diseases using Bayesian time-to-event models with active comparators. Statistics in Medicine, 43(19):3649–3663, August 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nguyen:2024:SAP

[1608] Trang Quynh Nguyen, Elizabeth A. Stuart, Daniel O. Scharfstein, and Elizabeth L. Ogburn. Sensitivity analysis for principal ignorability violation in estimating complier and noncomplier average causal effects. Statistics in Medicine, 43(19):3664–3688, August 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Norman:2024:DNA

[1609] Patrick A. Norman, Wanlu Li, Wenyu Jiang, and Bingshu E. Chen. deep-AFT: a nonlinear accelerated failure time model with artificial neural network. Statistics in Medicine, 43(19):3689–3701, August 30, 2024. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Robbins:2024:DFP

[1610] Michael W. Robbins, Sebastian Bauhoff, and Lane Burgette. Data fusion for predicting long-term program impacts. Statistics in Medicine, 43 (19):3702–3722, August 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Alvandi:2024:BMM

[1611] Amirhossein Alvandi, Sedigheh Omidvar, Armin Hatefi, Mohammad Jafari Jozani, Omer Ozturk, and Nader Nematollahi. Bayesian mixture modelling with ranked set samples. Statistics in Medicine, 43(19):3723– 3741, August 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Barbehenn:2024:NEB

[1612] Alton Barbehenn and Sihai Dave Zhao. Nonparametric empirical Bayes biomarker imputation and estimation. Statistics in Medicine, 43(19): 3742–3758, August 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zeng:2024:CPS

[1613] Shuxi Zeng, Fan Li, Rui Wang, and Fan Li. Correction to “Propensity score weighting for covariate adjustment in randomized clinical trials”. Statistics in Medicine, 43(19):3759, August 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [376].

Ding:2024:RRA

[1614] Jie Ding, Jialiang Li, and Xiaoguang Wang. Renewable risk assessment of heterogeneous streaming time-to-event cohorts. Statistics in Medicine, 43(20):3761–3777, September 10, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Noma:2024:RIM

[1615] Hisashi Noma, Shonosuke Sugasawa, and Toshi A. Furukawa. Robust inference methods for meta-analysis involving influential outlying studies. Statistics in Medicine, 43(20):3778–3791, September 10, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2024:IDL

[1616] Feifei Wang, Ke Jia, and Yang Li. Integrative deep learning with prior assisted feature selection. Statistics in Medicine, 43(20):3792–3814, September 10, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wei:2024:PSW

[1617] Wei Wei, Yunxuan Zhang, Satrajit Roychoudhury, and the Alzheimer’s Disease Neuroimaging Initiative. Propensity score weighted multi-source exchangeability models for incorporating external control data in randomized clinical trials. Statistics in Medicine, 43(20):3815–3829, September 10, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ji:2024:DNK

[1618] Jiadong Ji, Zhendong Hou, Yong He, Lei Liu, Fuzhong Xue, Hao Chen, and Zhongshang Yuan. Differential network knockoff filter with application to brain connectivity analysis. Statistics in Medicine, 43(20):3830– 3861, September 10, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wu:2024:MMA

[1619] Qiong Wu, Yuan Zhang, Xiaoqi Huang, Tianzhou Ma, L. Elliot Hong, Peter Kochunov, and Shuo Chen. A multivariate to multivariate approach for voxel-wise genome-wide association analysis. Statistics in Medicine, 43(20):3862–3880, September 10, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fan:2024:RWE

[1620] Yefeng Fan and Simon R. White. Review of weighted exponential random graph models frameworks applied to neuroimaging. Statistics in Medicine, 43(20):3881–3898, September 10, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhao:2024:ADM

[1621] Kaiqiong Zhao, Karim Oualkacha, Yixiao Zeng, Cathy Shen, Kathleen Klein, Lajmi Lakhal-Chaieb, Aurélie Labbe, Tomi Pastinen, Marie Hudson, Inés Colmegna, Sasha Bernatsky, and Celia M. T. Greenwood. Addressing dispersion in mis-measured multivariate binomial outcomes: a novel statistical approach for detecting differentially methylated regions in bisulfite sequencing data. Statistics in Medicine, 43(20):3899–3920, September 10, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Leon:2024:ESI

[1622] Larry F. León, Thomas Jemielita, Zifang Guo, Rachel Marceau West, and Keaven M. Anderson. Exploratory subgroup identification in the heterogeneous Cox model: a relatively simple procedure. Statistics in Medicine, 43(20):3921–3942, September 10, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kuo:2024:LCM

[1623] Jacky C. Kuo, Wenyaw Chan, Luis Leon-Novelo, David R. Lairson, Armand Brown, and Kayo Fujimoto. Latent classification model for censored longitudinal binary outcome. Statistics in Medicine, 43(20):3943– 3957, September 10, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Omar:2024:BNS

[1624] Zayd Omar, David A. Stephens, Alexandra M. Schmidt, and David L. Buckeridge. A Bayesian non-stationary heteroskedastic time series model for multivariate critical care data. Statistics in Medicine, 43(20):3958– 3974, September 10, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Alvares:2024:BSA

[1625] Danilo Alvares, Janet van Niekerk, Elias Teixeira Krainski, Håvard Rue, and Denis Rustand. Bayesian survival analysis with INLA. Statistics in Medicine, 43(20):3975–4010, September 10, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2024:EJM

[1626] Anonymous. Erratum to “Joint modeling for censored predictors due to detection limits with applications to metabolites data”. Statistics in Medicine, 43(20):4011, September 10, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1444].

Sun:2024:BBH

[1627] Na Sun, Jiadong Chu, Qida He, Yu Wang, Qiang Han, Nengjun Yi, Ruyang Zhang, and Yueping Shen. BHAFT: Bayesian heredityconstrained accelerated failure time models for detecting gene-environment interactions in survival analysis. Statistics in Medicine, 43(21):4013– 4026, September 20, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jaspers:2024:CAG

[1628] Stijn Jaspers, Johan Verbeeck, and Olivier Thas. Covariate-adjusted generalized pairwise comparisons in small samples. Statistics in Medicine, 43(21):4027–4042, September 20, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zoh:2024:BSP

[1629] Roger S. Zoh, Yuanyuan Luan, Lan Xue, David B. Allison, and Carmen D. Tekwe. A Bayesian semi-parametric scalar-on-function regression with measurement error using instrumental variables. Statistics in Medicine, 43(21):4043–4054, September 20, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hartman:2024:SMA

[1630] Holly Hartman, Matthew Schipper, and Kelley Kidwell. A sequential, multiple assignment, randomized trial design with a tailoring function. Statistics in Medicine, 43(21):4055–4072, September 20, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ran:2024:ERD

[1631] Xin Ran, Nancy E. Morden, Ellen Meara, Erika L. Moen, Daniel N. Rockmore, and A. James O’Malley. Exploiting relationship directionality to enhance statistical modeling of peer-influence across social networks. Statistics in Medicine, 43(21):4073–4097, September 20, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Turner:2024:DSS

[1632] Rebecca M. Turner, Kim May Lee, A. Sarah Walker, Sally Ellis, Michael Sharland, Julia A. Bielicki, Wolfgang Stöhr, and Ian R. White. Determining sample size in a personalized randomized controlled (PRACTical) trial. Statistics in Medicine, 43(21):4098–4112, September 20, 2024. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

French:2024:PCB

[1633] Joshua P. French, Mohammad Meysami, and Ettie M. Lipner. Prefiltered component-based greedy (PreCoG) scan method. Statistics in Medicine, 43(21):4113–4130, September 20, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McDonnell:2024:DUG

[1634] Erin I. McDonnell, Shanghong Xie, Karen Marder, Fanyu Cui, and Yuanjia Wang. Dynamic undirected graphical models for time-varying clinical symptom and neuroimaging networks. Statistics in Medicine, 43(21): 4131–4147, September 20, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zou:2024:EDA

[1635] Lily Zou, Yun-Hee Choi, Leonardo Guizzetti, Di Shu, Joshua Zou, and Guangyong Zou. Extending the DeLong algorithm for comparing areas under correlated receiver operating characteristic curves with missing data. Statistics in Medicine, 43(21):4148–4162, September 20, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Abbott:2024:LVA

[1636] Madeline R. Abbott, Inbal Nahum-Shani, Cho Y. Lam, Lindsey N. Potter, David W. Wetter, and Walter H. Dempsey. A latent variable approach to jointly modeling longitudinal and cumulative event data using a weighted two-stage method. Statistics in Medicine, 43(21):4163–4177, September 20, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Beltran-Sanchez:2024:BMS

[1637] Miguel Ángel Beltrán-Sánchez, Miguel-Angel Martinez-Beneito, and Ana Corberán-Vallet. Bayesian modeling of spatial ordinal data from health surveys. Statistics in Medicine, 43(21):4178–4193, September 20, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Reeder:2024:AID

[1638] Harrison T. Reeder, Kyu Ha Lee, Stefania I. Papatheodorou, and Sebastien Haneuse. An augmented illness-death model for semi-competing risks with clinically immediate terminal events. Statistics in Medicine, 43(21):4194–4211, September 20, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ojeda:2024:CML

[1639] Francisco M. Ojeda, Stuart G. Baker, and Andreas Ziegler. Calibrating machine learning approaches for probability estimation: a short expansion. Statistics in Medicine, 43(21):4212–4215, September 20, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Remiro-Azocar:2024:TMB

[1640] Antonio Remiro-Azócar. Transportability of model-based estimands in evidence synthesis. Statistics in Medicine, 43(22):4217–4249, September 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gkatzionis:2024:UIS

[1641] Apostolos Gkatzionis, Eric J. Tchetgen Tchetgen, Jon Heron, Kate Northstone, and Kate Tilling. Using instruments for selection to adjust for selection bias in Mendelian randomization. Statistics in Medicine, 43 (22):4250–4271, September 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2024:UMD

[1642] Chunlin Wang, Zheyu Liu, and Xinyu Wang. Using a monotonic density ratio model to increase the power of the goodness-of-fit test for logistic regression models with case-control data. Statistics in Medicine, 43(22): 4272–4286, September 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rucker:2024:SPR

[1643] Gerta Rücker, Theodoros Papakonstantinou, Adriani Nikolakopoulou, Guido Schwarzer, Tobias Galla, and Annabel L. Davies. Shortest path or random walks? A framework for path weights in network meta-analysis. Statistics in Medicine, 43(22):4287–4304, September 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Maltzahn:2024:SCD

[1644] Niklas N. Maltzahn, Ingrid Sivesind Mehlum, and Jon Michael Gran. Separable and controlled direct effects for competing events: Estimation of component specific effects on sickness absence. Statistics in Medicine, 43(22):4305–4327, September 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liang:2024:OMA

[1645] Zhongqi Liang, Suojin Wang, and Li Cai. Optimal model averaging for partially linear models with missing response variables and error-prone covariates. Statistics in Medicine, 43(22):4328–4348, September 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kuhlemeier:2024:VPI

[1646] Alena Kuhlemeier, Thomas Jaki, Katie Witkiewitz, Elizabeth A. Stuart, and M. Lee Van Horn. Validation of predicted individual treatment effects in out of sample respondents. Statistics in Medicine, 43(22):4349– 4360, September 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2024:DAD

[1647] Yunfei Wang. Dual-agent dose-finding in Phase I clinical trial-an extension of rapid enrollment design. Statistics in Medicine, 43(22):4361–4371, September 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Luo:2024:IEC

[1648] Shanshan Luo, Wei Li, Wang Miao, and Yangbo He. Identification and estimation of causal effects in the presence of confounded principal strata. Statistics in Medicine, 43(22):4372–4387, September 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lipkovich:2024:MAE

[1649] Ilya Lipkovich, David Svensson, Bohdana Ratitch, and Alex Dmitrienko. Modern approaches for evaluating treatment effect heterogeneity from clinical trials and observational data. Statistics in Medicine, 43(22): 4388–4436, September 30, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Adenyo:2024:ECC

[1650] David Adenyo, Jason R. Guertin, Bernard Candas, Caroline Sirois, and Denis Talbot. Evaluation and comparison of covariate balance metrics in studies with time-dependent confounding. Statistics in Medicine, 43 (23):4437–4455, October 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lok:2024:HEN

[1651] Judith J. Lok. How estimating nuisance parameters can reduce the variance (with consistent variance estimation). Statistics in Medicine, 43 (23):4456–4480, October 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hoogland:2024:EIT

[1652] J. Hoogland, O. Efthimiou, T. L. Nguyen, and T. P. A. Debray. Evaluating individualized treatment effect predictions: a model-based perspective on discrimination and calibration assessment. Statistics in Medicine, 43(23):4481–4498, October 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shao:2024:ETE

[1653] Yongwu Shao, Zhishen Ye, and Zhiwei Zhang. Exact test and exact confidence interval for the Cox model. Statistics in Medicine, 43(23):4499– 4518, October 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Smith:2024:PSC

[1654] Melissa J. Smith, Leslie A. McClure, and D. Leann Long. Path-specific causal decomposition analysis with multiple correlated mediator variables. Statistics in Medicine, 43(23):4519–4541, October 15, 2024. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chatzilena:2024:MFA

[1655] Anastasia Chatzilena, Nikolaos Demiris, and Konstantinos Kalogeropoulos. A modeling framework for the analysis of the SARS-CoV2 transmission dynamics. Statistics in Medicine, 43(23):4542–4558, October 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Smirnova:2024:SFR

[1656] Ekaterina Smirnova, Erjia Cui, Lucia Tabacu, and Andrew Leroux. Scalar-on-function regression: Estimation and inference under complex survey designs. Statistics in Medicine, 43(23):4559–4574, October 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nieser:2024:CMA

[1657] Kenneth J. Nieser and Alex H. S. Harris. Comparing methods for assessing the reliability of health care quality measures. Statistics in Medicine, 43(23):4575–4594, October 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2024:NBC

[1658] Shuyan Chen, Zhiqing Fang, Zhong Li, and Xin Liu. A novel blockcoordinate gradient descent algorithm for simultaneous grouped selection of fixed and random effects in joint modeling. Statistics in Medicine, 43 (23):4595–4613, October 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kormaksson:2024:DPA

[1659] Matthias Kormaksson, Markus Reiner Lange, David Demanse, Susanne Strohmaier, Jiawei Duan, Qing Xie, Mariana Carbini, Claudia Bossen, Achim Guettner, and Antonella Maniero. Dynamic path analysis for exploring treatment effect mediation processes in clinical trials with timeto-event endpoints. Statistics in Medicine, 43(23):4614–4634, October 15, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).